L 1

A Bibliography of Papers in Lecture Notes in Computer Science (1997), Part 2 of 2

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.25

Title word cross-reference

(log n) [1539, 1534]. (q) [2233]. 0 − 1 [910]. 1 [1839, 330]. 2 [1461, 2205, 1896, 670, 2147, 2001, 124, 2164, 1839, 55, 1855, 2121, 26, 2099, 25, 2131]. 2n [1235]. 3 [1986, 2205, 2181, 110, 1993, 1763, 580, 2004, 487, 2008, 1368, 1029, 2011, 2012, 1857, 2164, 2021, 2137, 1434, 1839, 1852, 1757, 1501, 712, 2041, 522, 1848, 2122, 1149, 399, 2199, 1574, 566, 511, 1507, 562, 2062, 527, 2065, 518, 565, 520, 2066, 2110, 2131, 2130, 2140, 485, 870, 2078, 519, 471, 2141]. 4 [1839, 584]. ^{2} [1810]. β [1979]. d [27]. f [1602]. h [2211]. k [1481, 25]. LT_{2} [1192]. N [2229, 962, 232, 1117, 394]. O(n log n) [1235]. Π [1810]. T [2229]. v [2211].

-ary [25, 27]. -block [1235]. -Calculus [1810]. -CCSP [1461]. -Connectivity [55, 1501]. -D [2181, 2001, 2137, 2065, 26, 2099, 870, 2078]. -Dimensional [124, 1839, 25, 394, 2008, 1855, 1763]. -Grams [962]. -Path [1481]. -skeletons [1979]. -Sparse [1602]. -theories [1896]. -tuple [1117]. -WAY [1434].

0103C [676]. 0p2 [735].

10 [173].

2 [259, 415, 1749]. 2.0 [1344]. 2.1 [128].

3 [804]. 3.4 [1326, 1327].

4 [1335, 1408, 400]. 4-GL [1408].

5 [1605].

95 [2272]. ’96 [337, 783].

= [1423, 954].

A-Method [1162]. A-NETL [1513]. A02 [1423]. A1 [998]. Abduction [877, 876]. Abductive [939, 1825]. Abilities [1084]. Abridged [1186]. Abstract [202, 191, 1903, 207, 1558, 208, 209, 304, 1533, 206, 205, 1176, 85, 1824, 204, 83, 187, 1909]. Abstraction [715, 1009, 1767, 1781]. Abstractions [157]. AC [1254]. AC-Unification [1254]. Accelerated [1082]. accelerators [279]. Accents [764]. Access [434, 423, 2245, 2229]. accidentalness [702]. Accommodating [1108]. According [352]. Accounting [757]. Accounts [1043]. Accumulation [1213]. Accurate [1242]. Achieve [800]. Achieving [1503, 403]. Acoustical [622]. Acquiring [1786]. Acquisition [805, 1707, 1996, 915, 611, 813, 814, 831, 1153, 335, 820, 230, 821, 835, 823]. Across [414, 715]. Action [1660, 698, 1037, 716, 1054, 1548, 239]. Actions [1163, 830, 752, 939, 1860, 1867]. Activation [235]. Active [1987, 1674, 1671, 636, 323, 633, 580, 2005, 346, 489, 637, 1676, 599, 354, 2034, 2107, 2040, 562, 2066, 1672, 2068, 635, 634, 1594, 2268, 2077, 1675, 2102, 1685, 638]. ACTIVE-DESIGN [638]. Actively [2032]. Activities [319, 1289, 1330]. Activity [2114, 1147, 348, 679]. ActivityFlow [1289]. Actor [928]. Acyclic [1465, 745]. Ada [2272]. ADAMS [1774]. Adaptable [437]. Adaptation [1060, 435, 176, 2166]. Adapted [513, 889]. Adapter [811]. Adapting [1805]. Adaptive [1988, 1083, 2229, 1028, 492, 418, 1502, 1755, 446, 570, 1109, 909, 1148, 2118, 617, 213, 2241, 2242, 138, 2244, 496, 2043, 148, 524, 509, 1807, 424, 1718, 1698, 2061, 1564, 123, 569, 473, 1053, 2096, 1150, 1403, 239, 1087, 83, 2144, 1948]. Adaptors [1756]. Adder [2225, 2245]. adders [297]. Adding [542, 1665, 1260, 282]. Addition [2248, 1946, 1435]. addition-shift-sequence [1946]. Additional [247]. Adjacency [445, 1848]. Adjectives [231]. adjoining [1178]. Adjusting [1529]. ADL [1880]. Adler [749]. Administration [1300]. Admissibility [1116, 925]. Adroit [1736]. Advanced [340, 1326, 433, 384]. Advances [1627, 2113, 1626]. Advantage [160]. Aerial [2113]. Aeronautics [113]. Aesthetic [2196]. Affine [499, 40, 1423, 2127, 2128]. after [1043]. against [1436, 1452, 1435]. Age [799]. Agent [864, 1229, 919, 825, 922, 1163, 1725, 1585, 937, 319, 921, 1699, 1386, 1148, 1393, 1392, 528, 320, 868, 1304, 1936, 883, 1457]. Agent-Assisted [320]. Agent-Based [1229, 1386]. Agentry [837]. Agents [1660, 825, 869, 244, 2017, 1391, 1035, 923, 1150, 1960]. Agreement [2226, 1827]. Ahead [1259]. AI [1881]. Aided [581, 679]. Aiding [348]. Air [1362]. Airline [1942]. Airports [1201]. Alexius [749]. ALFRED [632]. Algebra [564, 1354, 1198, 1901, 1902, 1914]. Algebraic [379, 1395, 182, 707, 1800, 2024, 1547, 383, 1398, 1927, 696, 1206, 1438, 1903, 1939]. algebraic-code [1438]. Algebraically [703]. Algebras [1221, 701]. Algorithm [305, 61, 301, 1089, 2158, 2202, 877, 535, 1997, 490, 1080, 845, 487, 1116, 840, 514, 523, 843, 1506, 1681, 2189, 2026, 1168, 2037, 842, 1501, 314, 2248, 1106, 1848, 2150, 2053, 1266, 130, 839, 2183, 1237, 451, 1247, 2214, 1054, 2099, 408, 1040, 911, 2127, 498, 1730, 1139, 2148, 2078, 1948, 1972, 277, 981, 1974, 276, 1426, 1969, 737, 1963, 283, 742].

Algorithmic [1955, 1828, 297]. Algorithms [1349, 1654, 2125, 1470, 88, 1995, 1973, 513, 1364, 1472, 122, 1473, 1551, 925, 1575, 2186, 2184, 921, 1240, 1101, 464, 367, 1547, 1608, 1577, 55, 113, 47, 1354, 2248, 385, 901, 1647, 1357, 2252, 5, 90, 1635, 316, 1618, 317, 1951, 614, 1576, 1649, 1651, 309, 1260, 534, 275, 1964, 284, 1947, 1955, 1979, 686]. All-cavity [1978]. All-to-All [1619]. Allocation [1338, 937, 367, 368, 369, 1484]. Alphanumeric [382]. Alphas [1605]. Alternative [796]. ALV [2055]. ALWAN [1355]. AMnet [2268]. Among [1701, 1706, 1977]. Amongst [1057]. Analog [1165]. Analogue [1044]. Analogy [1793, 218]. Analyses [196]. Analysing [760, 1936]. Analysis [155, 1988, 456, 624, 1779, 443, 1631, 2226, 486, 199, 878, 381, 1719, 167, 1998, 453, 156, 182, 828, 207, 1550, 1046, 186, 1474, 391, 1634, 608, 185, 2232, 611, 1126, 631, 2010, 1131, 1799, 1734, 190, 193, 464, 2138, 1031, 1056, 10, 192, 1488, 660, 1787, 194, 615, 147, 175, 1798, 1123, 2168, 620, 134, 1698, 882, 1784, 2056, 2059, 1245, 210, 1635, 1128, 517, 183, 553, 179, 2064, 127, 465, 621, 473, 188, 2256, 536, 635, 677, 590, 20, 1403, 2130, 1540, 1581, 758, 651, 1898, 1431, 1955, 987, 1930]. Analytical [104]. Analyzer [201]. Analyzing [42, 681, 2169, 1315, 1260]. Anatomical [1069]. Angiography [576]. Angles [2148]. Animal [1154]. Animated [2206]. ANN [1401]. Annealing [2056]. Annotated [1029, 1615]. Annotation [419]. Annotations [2030, 959]. ANNs [1169]. Anomaly [592]. Anonymity [774]. Anonymous [775, 776, 1459, 1961, 1455]. answering [985, 984]. Anti [791]. Anti-Trust [791]. Aperiodic [1537]. Aperture [473]. API [431]. Appearance [2101, 2165, 571, 2133]. Appearance-Based [2165, 571]. Applicability [43]. Application [2223, 1301, 1461, 502, 380, 419, 430, 1340, 892, 1843, 1324, 2007, 1126, 111, 2236, 2038, 832, 1122, 2248, 13, 613, 1696, 1041, 1164, 170, 1119, 1405, 1643, 534, 1421, 695, 1454]. Application-Oriented [1301]. Applications [2115, 1338, 1493, 1717, 1654, 1339, 427, 142, 1746, 439, 1352, 1323, 611, 1742, 1240, 107, 375, 414, 1722, 849, 154, 428, 1842, 1532, 784, 2241, 701, 1262, 1378, 433, 384, 953, 620, 622, 589, 168, 1191, 1573, 1411, 706, 400, 1524, 350, 1333, 83, 1415, 723, 264]. Applied [588, 113, 2052, 312, 1702, 527, 2256, 1695, 678]. Applying [676, 791, 147]. Approach [1301, 1223, 356, 318, 476, 1230, 1213, 1754, 806, 2212, 919, 807, 54, 600, 2229, 888, 930, 340, 960, 182, 394, 1670, 1407, 955, 1599, 2005, 346, 1543, 162, 226, 1536, 956, 1136, 52, 1204, 2235, 1309, 370, 1508, 1391, 1406, 2175, 71, 39, 2173, 1668, 1155, 449, 571, 1122, 1295, 546, 2088, 1565, 1480, 392, 87, 860, 424, 1149, 497, 1121, 1574, 397, 153, 1850, 2057, 2154, 511, 1097, 2139, 565, 912, 2066, 6, 1377, 1040, 559, 702, 1299, 2130, 2102, 2144, 2151, 1863, 1905, 1873, 991, 1943, 687, 258, 1980]. Approaches [748, 2045, 1033, 716, 767, 1958, 297, 1003]. Approaching [1888]. Approximate [1706, 205, 924]. Approximately [1600]. Approximating [1962, 1816, 1579, 56]. Approximation [55, 2248, 1618, 1460, 1233, 1981]. Approximations [879, 1917]. Arabic [605, 1623]. Arbitrarily [507]. Arbitrary [2045]. Arbitration [1568]. Arc [1258, 1170]. ARCH [1321]. Archaeology [2201]. ArchE [2201]. Architectural [322]. Architecture [1760, 903, 869, 2080, 1293, 439, 1774, 1335, 591, 423, 832, 540, 146, 1291, 528, 98, 1753, 1001, 252]. Architectures [100, 177, 538, 139, 1250, 1389, 147, 175, 5, 1401, 1376, 1868, 253]. Archive [326]. Area [2211, 2186, 523, 1408, 2193, 1065]. Area-Based [523]. Arguing [923]. argument [1873]. Arithmetic [535, 1823, 2244, 1969, 298]. Arrangement [1555]. arrangements [1956]. Array [1353, 1813, 1756, 285, 284, 296]. Array-Comprehension [1813]. Arrays [84, 1375, 1125, 38]. Art [443, 1621]. Articulated [2065]. Artificial [1152]. ARTMAP [1132]. ARTMAP-DS [1132]. ary [372, 25, 27]. Asia [2112]. ASICs [292]. asking [734]. Aspect [1773, 566]. Aspect-Based [566]. Aspect-Oriented [1773]. Aspects [380, 1173, 920, 1841, 1531, 628, 696, 1968]. Aspnes [839]. Assemblies [1038]. Assembling [374]. Assembly [1160]. Assessment [513, 1208, 1726]. Assignment [46, 1567, 1418]. Assistant [904, 832]. Assistants [425, 736]. Assisted [1719, 320]. Associated [1138, 466]. Associations [1296, 1021]. Associative [542, 22, 1104, 1108, 558]. ASSOM [1163]. Assuming [2146]. Assumption [1826]. Assumption-Commitment [1826]. Asteroidal [1478]. Asymmetric [1447]. Asymptotic [1818, 1138]. Asynchronous [1321, 100, 2225, 648, 843, 630, 849, 29, 60, 97, 65, 1552]. Athapascan [76]. ATM [429, 1597, 430, 431, 841, 414, 415, 1923, 326, 1487, 1584]. ATM-Routings [1487]. Atomic [63, 847]. Attack [2244, 1441, 1435]. Attacks [782, 1356, 1452]. Attention [1146, 1008]. Attorney [795]. Attractors [558]. Attribute [203, 1667, 976, 1016, 1922, 989]. Attribute-Oriented [1667, 976]. Attributes [1282, 454, 352, 980, 679]. Audio [430, 965, 2264, 544]. Audio-visual [544]. Audio/Video [430]. Audiovisual [965]. Auditable [786]. Auditory [751, 1050]. Augmentation [55]. Augmented [2125]. Augmenting [245, 1952]. Authenticated [854, 1412]. Authentication [783, 1424]. Authenticity [228]. Authoring [419, 1034, 355]. Authority [946]. authorization [1415].

Auto [558, 2017, 296]. Auto-associative [558]. Auto-configurable [296]. Autocorrelation [2090]. Autoepistemic [880]. Automata [186, 1239, 242, 1536, 863, 1791, 1792, 1826, 1613, 330, 1429, 1965, 744]. Automated [155, 1228, 956, 1795, 1279, 885, 1520, 1635, 2064]. Automatic [1227, 275, 605, 1804, 199, 267, 2273, 1743, 3, 597, 1506, 470, 578, 2191, 1186, 1645, 35, 45, 700, 1758, 652, 1588, 2046, 335, 1644, 33, 1573, 2214, 2068, 1505, 590, 556, 1001, 2062]. Automatically [575, 2106, 200, 1010]. Automating [946, 986]. Automation [616, 1735, 1277]. Autonomous [1157, 1163, 869, 319, 592, 1148, 2143, 1165, 996, 1153, 885, 2163, 868, 1387, 821, 1690, 1150, 1875]. autonomy [975]. Autoregressive [612]. Auxiliary [1346]. Availability [1601]. average [734]. AVL [54]. Avoidance [599, 1148, 1149]. Aware [438]. Awareness [952]. AXE [173]. Axial [516]. Axiomatic [226]. Axis [2190].

B [1938]. Back [1259]. Background [2218, 759, 1093]. Backjumping [886]. Backtracking [1693, 1694, 1245]. Bad [758]. Balancing [117, 122, 1365, 121, 843, 116, 1167, 15]. Ball [553]. Balls [2209]. BAMBI [958]. Banach [740]. Bandwidth [2241, 416]. Bankcheck [1646]. Bankchecks [1645]. Banking [619]. Bar [2093]. Bar-graph [2093]. Bargains [1290]. Barrel [1055]. Base [895, 2096, 1010]. Based [1683, 1700, 452, 1223, 2205, 356, 881, 1989, 502, 1287, 318, 582, 1229, 899, 1231, 2227, 1395, 233, 1993, 576, 2101, 1994, 535, 2147, 552, 554, 1697, 809, 117, 207, 1772, 1705, 810, 607, 357, 186, 487, 391, 760, 1088, 162, 956, 803, 523, 631, 1755, 481, 2165, 1386, 94, 1684, 1131, 587, 909, 2012, 396, 3, 501, 2235, 1810, 190, 1587, 2160, 788, 1735, 2126, 2135, 2240, 2022, 581, 2258, 2026, 1645, 591, 479, 332, 512, 2037, 1841, 1811, 1739, 2107, 329, 217, 852, 1775, 496, 2040, 571, 1752, 1122, 1404, 714, 1381, 526, 610, 75]. Based [132, 2249, 2045, 2163, 1513, 615, 497, 2133, 1798, 2169, 2168, 69, 620, 1125, 358, 407, 1718, 1315, 817, 1617, 1304, 833, 566, 2252, 1850, 586, 326, 1198, 1058, 467, 1572, 1409, 170, 595, 2063, 1292, 1054, 553, 1274, 1085, 696, 15, 221, 2099, 2110, 2067, 2134, 2259, 1666, 1510, 2162, 702, 20, 584, 1034, 1302, 2076, 2093, 2102, 2078, 519, 2123, 330, 1021, 1012, 1899, 1017, 1421, 1001, 968, 1000, 1008, 1873, 273, 531, 280, 274, 1429, 675, 1883, 727, 686, 1890, 995, 1892, 1025, 1321, 840, 1624, 272, 289, 989]. Bases [827, 836, 973]. Basic [1804, 1806, 51, 1977]. Basis [1069, 1, 762, 125, 1676, 1094, 1048, 154, 1037, 1164, 705]. battlefield [1878]. Bayes [2040]. Bayesian [2115, 2195, 931, 882, 2104]. BCC [1857]. Be [1748]. Beat [1721]. become [208]. behaved [980]. behaving [1278]. Behavior [1301, 73, 1340, 1568, 1148, 1149, 1288]. Behaviors [1782]. Behaviour [398, 2015, 1389, 351, 696, 972, 1878]. Behavioural [892, 348, 1795, 70]. Behaviours [301, 715, 1610]. behind [2230, 226]. Belief [1713, 1187, 1710]. Belief-Centered [1187]. Beltrami [2027]. Benchmarking [1730]. Benchmarks [127]. Bends [2203, 2185]. Bergen [2244]. Bergen-Hogan [2244]. Best [2182, 232, 2102, 742]. Between [829, 95, 1063, 1539, 1056, 206, 789, 1233, 2261, 324, 1554]. Beyond [1592, 1327, 67, 441]. Bézier [508]. Bharucha [758]. BI [1893]. BI-POMDP [1893]. Bias [835]. BIBD [1423]. Bibliothèque [965]. Bicliques [1481]. Biconnected [1487]. Bigram [2094]. Biham [1434]. Biham-DES [1434]. Bimodal [470, 595]. Binary [507, 120, 1629, 1844, 457, 2211, 1298, 1244, 470, 1839, 1806, 1916, 737]. binate [270]. Binding [1147, 1758, 1329]. Binocular [2052]. Biological [1771, 1036, 679].

Biologically [302, 1067]. Biomedical [611, 575]. Biometric [2074]. Biometrics [1991]. Biomolecular [1361]. Bipartite [1818, 2177, 1977, 1975, 1951]. Bipartization [1618]. Bisimulation [1835, 1913, 1926]. bisimulations [1912]. Black [1830]. Black-box [1830]. Bleak [1305]. Blind [1127, 1121, 2253, 1119, 1449]. Blink [1053]. Block [2158, 541, 1353, 1497, 1742, 2248, 549, 2252, 83, 1359, 2257, 1235, 1437]. Block-Cyclic [1353]. Block-Irregular [83]. Block-Matching [2158]. Block-Structured [1742]. Blur [2130]. Blurred [1985]. BMP [2048]. Board [885]. Body [400]. Boltzmann [1082]. Bona [1202]. Bony [574]. Book [561]. Boolean [728, 878, 670, 1428, 1426, 977, 1732, 724, 1427]. Booster [1522]. boosters [294]. Boosting [1089, 1722]. BOOT [1409]. both [975]. Bottleneck [1345]. Bottom [486, 715, 683]. Bottom-Up [486, 715, 683]. Bound [1349, 1467, 1606, 510]. Boundaries [764, 1202]. Boundary [104, 509]. Bounded [1831, 303, 841, 1877, 30, 1483, 1893, 740]. Bounding [1265, 135]. Bounds [484, 1597, 1818, 2186, 2184, 851, 2254]. box [1830]. Boxes [1447]. BP [2230]. Bracketings [1179]. Brain [559, 331]. Brain-Like [559]. Branch [1349, 138, 1491, 1980]. branch-and-cut [1980]. Branches [198, 1970]. Breaking [2230, 1124, 1881, 1336]. Breast [931]. Bridge [2261]. Brier [738]. Bright [1305]. Brightnesses [518]. Broadband [422]. Broadcast [63, 22, 1801]. Broadcasting [31, 1957]. Broader [1528]. BROKERAGE [1959]. Brouwer [729]. Brouwer-Heyting-Kolmogorov [729]. browsable [1935]. Browse [955]. Browsers [2205]. Browsing [775, 1728, 438]. BSP [52]. Build [1986, 1328, 831]. Builders [317]. Building [955, 1497, 2032, 170, 824, 1904]. Buildings [2105, 2113]. Bullet [1290]. Bus [132]. Bus-Based [132]. Buses [28]. Business [964, 1292]. Byzantine [1827].

C [1325, 1749, 167, 1743, 1747, 1756, 1752, 1748, 1745]. C. [758]. Cache [34, 1499, 102, 103, 1516, 1491]. Caching [405]. CAD [522, 251]. CAD-oriented [251]. Calculation [114]. Calculi [1589, 918]. Calculus [1796, 1182, 1802, 1810, 84, 1710, 69, 70, 1921, 1889]. Calendars [1542]. Calibration [1168, 2034, 494, 519]. Call [1476, 1982]. Call-Scheduling [1476]. CAME [1311]. Camera [2055, 494, 2162, 2152, 519, 2128]. Cameraman [2017]. Cameras [525, 2066]. CAMILA [1931]. Can [1133, 208, 1065, 1582]. Cancellation [1171]. Cancer [581, 931, 586]. Canonical [339]. CAP [1354]. CAP-Specified [1354]. Capabilities [1226, 1326, 945]. Capacity [2048, 1107]. Capitalizing [822]. caption [287]. Capture [2132, 1734]. capturing [972]. Car [715]. Caricaturing [2163]. Carl [754]. Carlo [687]. Carmichael [2251]. cartesian [569]. Cartographic [1228]. Cascade [1116]. Cascade-Correlation [1116]. Cascaded [711]. Case [1360, 895, 672, 809, 640, 921, 946, 908, 235, 2190, 1920, 839, 1201, 127, 668, 128, 534, 1906, 265, 1884, 283, 653, 1010, 1311, 1312]. cases [1010, 1884]. Cash [781, 774, 778, 1961, 1458, 789]. CASL [1900]. CasPer [1730]. CAST [1434]. Casual [955]. Casualty [594]. Cat [1063]. Catalogs [947]. Catalogues [417]. Categorial [1188]. Categorical [1677, 1914, 1922]. Categorization [2092]. Causal [700, 1709, 1025, 1005]. Causality [830]. Causation [222]. cavity [1978]. CBET [895]. CBL [1071]. CBR [896]. CC [1755, 2266]. CCDs [451]. CCMA [2225]. CCSP [1461]. CDMA [2249]. Cell [1059, 2265, 1038, 1053, 1139]. Cells [1056, 1064]. Cellular [330, 1429]. Center [2025, 1068]. Centered [1187, 2122]. Centipede [479]. Central [1046]. Centric [529]. Certainty [2134]. certificates [1453]. CFD [1763, 95]. CFD-Problems [95]. Chain [2001, 77, 329, 2119, 688]. Chaining [2257]. Challenges [441, 253]. Chamfer [1854]. Change [1196, 1197, 150, 927, 2142, 860, 589, 544, 1025]. Changes [1400, 444]. Changing [307]. Channel [2225, 2229, 2241, 2242, 2245, 2081, 1455]. Channels [2236, 20, 1443]. Character [1999, 1630, 1638, 610, 1086, 604, 2094]. Characteristic [1851]. Characteristics [1634, 608, 322]. Characterization [493, 1535, 1835, 1614, 1571, 1956]. Characterizations [1674]. Characters [1623, 2098, 606]. Charging [964]. Charts [71]. Check [2227, 658]. Checker [661, 1934]. Checking [2221, 2220, 1819, 157, 1791, 1897, 1869]. chemical [1862]. Chen [1438]. CHEOPS [418]. Chess [1575, 678]. Chest [573]. children [978]. Chile [593]. Chinese [1999, 1557, 610, 2094, 1720]. Chip [102, 103, 784, 250]. Choice [866]. Choices [1652]. Cholesky [89]. Chord [757, 758]. Chordal [1471]. chosen [1441]. chosen-message [1441]. Christian [749]. Chrominance [500]. chromosomes [1703]. Chronologies [1542]. Cipher [2230, 2243, 779]. Ciphers [2252, 1437]. ciphertext [1444]. Circal [1901, 1933]. Circle [756]. Circles [2188, 2001]. circuit [253]. circuitry [267]. Circuits [1370, 1823, 273, 265, 266]. Circular [510]. Circulation [60]. circumscription [971]. Cities [1216]. City [1229, 1199, 2105]. Claims [166]. Clam [934]. Class [342, 1395, 1282, 1633, 1294, 310, 1756, 643, 1768, 2214, 1753, 1669, 2094, 993]. Classes [1239, 1770, 1538, 2121]. Classical [1361, 1802, 530, 729]. Classically [1053]. Classification [1129, 447, 1630, 1204, 1094, 1096, 2023, 585, 1075, 626, 714, 2049, 574, 515, 2108, 2064, 926, 482, 2095, 2097, 692]. Classifications [1723]. Classificatory [1186]. Classifier [605, 1117, 1112, 1137, 1124, 246, 321, 556, 447, 992]. Classifiers [446, 1098, 593, 467]. classify [682]. Classifying [146]. Clauses [861, 224, 1024]. Client [439, 1774, 435, 11, 425]. Client-Controlled [435]. Client/Server [439, 1774]. Climbing [1260]. Clipped [1059]. Cliques [2188]. Clock [857, 415]. Clocks [850]. Closed [1535, 287]. Closure [1662, 1538]. Cloth [2008, 2138]. Clothing [2033]. CLP [929]. Cluster [1315, 1517]. Clustered [2187, 91]. Clustering [41, 2188, 344, 2200, 1065, 352]. Clusters [1742, 95, 1500]. Clutter [2157]. Co [825, 956, 101]. Co-occurrence [956]. Co-operation [825]. Co-processor [101]. Coalescing [198]. coalgebraic [1912]. Coarse [1043, 2097]. COAST [1673]. COATIS [830]. COCOMO [176]. Code [1361, 202, 1763, 167, 1743, 409, 1773, 160, 329, 2248, 172, 106, 1438, 1423, 1452]. Code-Based [329]. Codebook [1607]. Coded [2229, 2238, 2241]. Codes [2224, 2225, 2227, 2234, 2240, 1488, 2250, 2255]. codesign [259]. Coding [535, 1046, 1043, 1140, 2242, 2244, 2249, 549, 1047, 1092, 1396, 1828, 543, 1967]. Coefficients [1854]. coevolutionary [991]. Coexistence [406, 1846]. Cogito [1940]. Cognition [1558, 1144, 1214]. Cognitive [1217, 922, 1559, 1208, 747, 748, 1216, 1146, 1731, 940]. Coherence [2031, 1516, 2154]. Coherent [2242]. COLBERT [212]. Collaboration [829, 944]. Collaborative [432, 2017, 1034, 353]. Collaterals [1044]. collected [994]. Collection [955]. Collections [362, 979]. Collective [1519]. Collision [599]. Colloquium [966]. Color [1985, 456, 1990, 483, 554, 2003, 551, 2111, 481, 501, 2023, 448, 2028, 2029, 455, 500, 449, 530, 2043, 2048, 2053, 549, 454, 488, 586, 2058, 2060, 2061, 553, 2112, 2071, 482, 401, 2166, 1949]. Color-Based [554]. Color-Image [2061]. Colored [1479]. Colossus [2230]. Colour [451, 539]. colouring [262]. Combination [599, 1101]. Combinations [1622]. Combinatorial [1243, 1579, 2097, 1941]. Combined [1609]. Combining [1782, 1162, 1876, 426, 609, 209, 2055, 1090, 158]. Combustion [114]. Command [2266]. comments [1438]. Commerce [801, 2260, 1457]. Commitment [1826, 372]. Commodity [1508, 1752]. Common [953]. Commonsense [933]. Communicate [1157]. Communication [846, 864, 854, 1701, 431, 190, 1506, 2261, 1504, 40, 1501, 18, 1519, 2267, 1317, 1318, 1619, 1320, 1333]. Communications [2229, 118, 804, 436, 65, 286]. Community [944]. Commutative [1255, 1178]. Compact [1732]. compaction [1972, 263]. Comparative [882, 1040]. COMPARES [624]. Comparing [1780, 146, 158]. Comparison [1129, 1744, 826, 117, 2007, 1677, 95, 311, 1314, 1408, 613, 2056, 1316, 976]. Compartment [2247]. CompAss [897]. Compensated [2023]. Competition [1056, 1057]. Competitive [825, 617, 1982, 1950]. Compilation [897, 1821, 39, 69, 292]. Compile [1747, 210, 1753]. Compile-Time [210, 1753]. Compiler [208, 1504, 1505, 1554]. Compilers [1588, 33]. Compiling [887, 1271]. Complementation [1538]. Complete [2211, 391, 1532, 1534, 2119, 25]. Completeness [1909, 1023, 1563, 1925, 1468]. Completion [135]. Completions [766]. Complex [1361, 236, 1246, 1161, 650, 1199, 1805, 1400, 378, 1389, 1207, 1306, 2104, 331, 1886]. Complexity [1683, 966, 1679, 1255, 1477, 1555, 1538, 30, 1479, 1111, 1968, 1824, 1583, 733, 1977, 1413, 1946, 1970, 1870].

Complies [1228]. Component [443, 1126, 1124, 1123, 1128, 2259]. Components [322, 1061, 13, 1118, 987]. Composite [1634, 351]. Composition [1566, 2125, 2258, 148, 1927, 402]. Compositional [1833, 825, 156, 809, 665, 1827, 1834]. Compositionality [1549, 1384]. Comprehension [1218, 1027, 43, 1813, 945]. Compressed [1992, 2016, 463]. Compression [535, 2000, 2010, 479, 1396, 539, 534]. Compton [583]. Computable [1540, 1581]. Computation [1760, 1989, 751, 1144, 205, 526, 546, 2154, 1815, 49, 127, 494, 1515, 683, 296, 741, 1020]. Computational [1394, 864, 240, 966, 1173, 958, 878, 302, 1477, 1414, 449, 1521, 1580, 1194, 1403, 1583]. Computations [1348, 96, 3, 77, 43, 1655, 38]. Compute [707, 2034, 133]. Computer [1493, 708, 1719, 1575, 1489, 616, 2126, 581, 755, 441, 1541, 1528, 98, 289, 1930]. Computerized [1410]. Computers [1762, 1497, 131, 1514, 1524]. Computing [1493, 1360, 1322, 1469, 1492, 2203, 107, 1838, 1776, 529, 11, 2036, 1767, 1807, 87, 2047, 1647, 1768, 1777, 838, 1319, 277, 1456, 280, 278, 259, 1979]. Concatenated [2255]. concealing [1444]. Concept [909, 849, 1496, 311, 901, 1790, 833, 918, 1299, 991]. Concepts [872, 1284, 1652, 750, 1572, 230, 1015]. Conceptual [1300, 1286, 1282, 1293, 245, 814, 1031, 1276, 1305, 860, 320, 349, 1281, 819, 1019, 988, 984]. Conceptualizing [231]. Concerning [749]. Concurrency [191, 1833, 665, 1599, 1834]. Concurrent [1796, 54, 1238, 1383, 190, 1049, 48, 72, 1798, 69, 64, 70, 642]. Concurrent-Read [48]. Conditional [1664, 198]. Conditioned [1053]. Conditions [2082, 2100]. Conference [422].

Conferencing [430, 426, 2263]. Confidence [2038]. Configurable [1262, 281, 296, 292]. Configuration [1587, 150, 606, 1894]. Configuration-Based [1587]. Confliction [1725]. Conflicts [1307]. Confluence [1253, 633, 634]. Congestion [841]. Congruence [196]. Conic [2099]. Conjecture [658]. Conjugate [331]. Conjunctive [1662]. Connected [1039, 830, 1188]. Connecting [337]. Connection [1523, 243, 2262]. COnnectionist [624]. Connections [1062, 1420]. connectivities [1952]. Connectivity [55, 1962, 1501]. Connectors [171]. Connotation [1205]. Conquer [561, 1220, 1286]. ConQuer-II [1286]. Conquering [82]. Consensus [839]. Consequence [1188]. Considerations [1735]. Considering [2134]. Consistency [672, 1257, 1489, 1258, 1786, 1244, 851, 72]. Consistent [1670, 645, 1783, 1288, 62, 2081]. Consonance [754]. conspiracy [1452]. Constancy [1990]. Constrained [305, 1662, 408]. Constrainedness [1258]. Constraining [230]. Constraint [1253, 1242, 1263, 1631, 1814, 1246, 1238, 1265, 910, 1249, 1789, 1271, 1693, 3, 1251, 1269, 1244, 1307, 1739, 1241, 1694, 1698, 1191, 1273, 1274, 1248, 1890, 1194, 1695, 1234, 1008, 680]. Constraint-based [1890, 1008]. Constraint-Directed [1694]. Constraints [1084, 1267, 1665, 1814, 1236, 1270, 1472, 1239, 360, 185, 1222, 687, 1174, 1261, 1256, 2049, 1275, 1237, 1164, 1785, 1260, 1025, 2151, 1901, 982, 1886]. Construct [1297, 706, 1010]. Constructing [1179, 341, 428, 815, 1848]. Construction [61, 1428, 946, 1298, 311, 397, 2254, 371, 870, 1440, 1917, 978, 679]. Constructions [1184, 1416]. constructive [1931]. constructs [968]. Consultation [965]. Consumption [49]. Contact [2210]. Container [1198]. Container-Surface [1198]. Containing [1814, 1711, 961, 1012].

Content [529, 1037, 358, 545, 330, 1001]. Content-Based [358, 330, 1001]. Content-Centric [529]. Contents [356, 363]. Contest [2216]. Context [1859, 1536, 1226, 1728, 2126, 1483, 598]. Context-Based [2126]. Context-Free [1536, 1483]. Context-Sensitive [1728]. Contextual [1073, 1209, 927, 1142, 560, 153]. Contextualization [883]. Contingency [1887]. Continuations [766]. Continuity [1164]. Continuous [1196, 862, 1787, 1553, 2172, 1580, 404, 1583]. Continuum [789]. Contour [2003, 1630, 2011, 455, 503]. Contours [1987, 603, 580, 2077]. Contraction [243, 52]. Contradiction [1526]. Contrast [1069]. Contribution [1554, 228, 451]. Control [1987, 1171, 636, 1407, 1161, 774, 1169, 2232, 2143, 1406, 1096, 212, 10, 646, 45, 1668, 1153, 1330, 1143, 2266, 1160, 2259, 556, 1087, 17, 969, 1982]. Controlled [1683, 435, 1049, 1787]. Controller [1378, 420]. Controllers [639, 299, 298]. Controlling [932, 907, 1740, 649, 1690]. Convergence [843, 512, 1106, 97]. Convergences [1136]. Conversion [1628]. Converting [2168]. Convex [2035, 1969]. Convolutional [2233, 2250]. Cooperating [236, 339]. Cooperation [922, 1699]. Cooperative [387, 933, 1323, 1328, 901, 2106, 528]. Coordinate [1210, 1734]. Coordination [1700, 1289]. COP [475]. Coping [445]. Coplanar [1845]. coprocessing [282]. coprocessor [289]. Copyright [363]. CORBA [1421, 1601, 426, 2261, 1616]. Core [77, 1509]. Cores [626]. Corners [475]. Corporate [824]. Corporations [776]. Corpus [874, 1615]. Corpus-Driven [874]. Correct [2123]. Corrected [655]. Correcting [2234, 2240, 526, 9]. Correction [582, 1727, 718]. Correctness [1711, 1778, 1912]. CORRELATE [1751].

Correlated [2224, 1147]. Correlation [1116, 524, 521, 1047, 2254, 598, 482, 1428]. Correlation-Immune [2254]. Correlations [1039]. correlator [288]. Correspondence [2075]. Correspondences [2044, 494, 2129]. Cortex [1069, 1055, 1063, 1062]. Cortical [1049, 1038]. Cortical-Type [1049]. Cost [964, 878, 921, 396, 176, 466, 1516, 184, 258]. Cost-Effective [184]. Costing [84]. Could [1053]. Count [841]. Counter [235]. Counting [1244]. Countings [34]. Coupled [1104]. Cover [621]. Coverage [865, 166]. Covering [1297, 270]. Covers [873, 1479]. CP [1266]. CPLDs [282]. CPU95 [127]. Cracks [508]. Crash [376, 626]. Cray [1364, 93]. Create [1629]. Created [2131]. Creating [1736]. Creation [1088]. Creative [766]. Credible [1723]. Creep [469]. Creep-and-Merge [469]. Crew [1369, 1942]. crew-scheduling [1942]. Criteria [1228]. Critical [155, 1431, 650, 783, 1377, 36]. Criticality [159]. CROMATICA [623]. Cross [483, 1039, 770, 957]. Cross-Correlations [1039]. Cross-Language [957]. Cross-Media [483]. Cross-Synthesis [770]. Crossing [640, 2175, 2177, 2192]. Crucial [1622]. cryptanalysis [1434, 1414]. Cryptographic [657, 166, 1356, 1409, 1454, 1427]. Cryptography [803, 793, 785, 792, 1417, 1447]. Cryptosystem [2228, 784, 1438, 1445]. Cryptosystems [2235, 1441, 1414]. CSCW [428]. CSP [1814, 642, 655]. CSPs [1245, 1260]. CT [581, 577]. Cubes [19]. Cubic [2183]. Cues [2141]. Cultural [1528]. Culture [307]. Cumulative [1263]. Curriculum [1530]. Curse [1336]. Cursive [603, 2019, 1639]. Curvature [2020, 1852, 546, 2050, 1718, 460]. Curvature-Based [1718]. Curve [2021, 1439]. Curved [2041]. Curves [1653, 2235, 2120]. custom [277, 280].

Customer [422, 1550]. Customized [1030]. Customizing [1292, 534]. Cut [1467, 2200, 1460, 1980, 1579]. Cuts [62]. Cutting [243]. Cutwidth [27]. CV [2122]. Cycles [1462, 2227, 1288]. Cyclic [1353]. Cyclically [307]. Cylindrical [2069]. Czech [1615].

D [522, 1986, 2205, 2181, 110, 1993, 670, 2147, 2001, 580, 2004, 487, 1368, 1029, 2011, 2012, 1857, 2164, 2021, 2137, 1852, 1757, 712, 2041, 1848, 2122, 1149, 399, 2121, 1574, 566, 511, 1507, 562, 2062, 527, 2065, 518, 565, 26, 520, 2066, 2099, 2110, 101, 584, 2131, 2130, 2140, 485, 870, 2078, 519, 471, 330, 2141]. D-Point-Grid [519]. D-Scene [520]. D-Visualization [1574]. Dark [2004]. DARPA [836]. Data [612, 452, 1988, 42, 624, 898, 37, 41, 1754, 164, 583, 339, 2101, 1661, 167, 2002, 1268, 80, 1285, 783, 94, 1684, 1131, 279, 1734, 370, 1765, 1761, 8, 1681, 325, 2138, 367, 618, 12, 410, 10, 383, 368, 862, 2041, 1291, 1652, 399, 1656, 1768, 1522, 970, 488, 2056, 326, 1498, 1292, 371, 133, 1211, 1200, 411, 420, 204, 545, 350, 1651, 758, 537, 83, 728, 284, 1006, 1456, 994, 981, 731, 694, 300, 1922, 1015, 1013]. data-localised [284]. Data-Parallel [1765, 1761, 8]. Data-Validation [898]. Database [943, 1287, 388, 1671, 1363, 359, 1367, 1774, 1309, 435, 1677, 154, 367, 382, 365, 2033, 547, 2173, 2105, 368, 1578, 1345, 1678, 629, 1593, 2061, 627, 1277, 1672, 352, 1299, 390, 2152, 1685, 1021, 996]. Databases [63, 1674, 342, 899, 554, 633, 1598, 1599, 360, 391, 337, 1657, 338, 344, 1676, 354, 1279, 1303, 628, 328, 1565, 1667, 392, 364, 369, 533, 1636, 62, 366, 634, 1724, 1594, 1675, 355, 685, 1012, 1000, 1005]. Dataflow [1508]. Datalog [633, 1679, 688, 1676]. Datarol [1496]. Datarol-II [1496]. Daubechies [401]. DAVIC [406]. DB [338, 1292]. DBMS [1592, 347]. DCE [2261]. DCube [2199]. De-noising [2083]. Dead [167]. deadlines [1886]. Deadlock [1465, 1898, 21, 660]. dealing [1011]. Debates [1205]. Debugging [1347, 1865]. Decidable [1184]. Decision [1380, 1661, 1725, 1960, 311, 354, 1806, 224, 1564, 1013, 1078, 1010, 1869, 1413, 1877, 1016, 737, 742]. Decision-making [1960]. Declarative [1556, 636, 346, 1682, 333]. Decoding [2225, 2236, 2245, 2248]. Decomposition [507, 1463, 1263, 1350, 1298, 94, 111, 320, 1874, 1013, 1416]. Decompositions [1409]. Decorrelation [539]. Decoupled [36]. Decoupling [802]. Decremental [1954]. Dedicated [537, 251]. Deduction [1186, 1189, 1520, 1656, 1594]. Deductive [155, 628, 1678, 629, 1667, 627, 1675, 1000]. Deep [1575, 969]. deep-space [969]. Default [1715, 240, 985]. Defeasible [1716]. Defect [160, 626, 158, 1975]. Deferred [1668]. Defining [1527, 812, 466, 224]. Definitions [1667]. Deflection [20]. Deflection-Based [20]. Deforestation [203]. Deformable [580, 509, 503, 471]. Deformation [1986, 112, 2031, 521, 1108]. Deformations [1642, 504, 485]. Deforming [1368]. Degraded [962, 625]. Degree [2024]. Degrees [32]. Delay [1060, 408, 274]. Delay-Constrained [408]. Delegation [884, 1456]. Deletion [1839]. Delft [99]. Delft-Java [99]. Delta [1625]. Delta-Lognormal [1625]. Deltas [1672]. Demand [421, 439, 913]. demodulator [291]. Demonstration [2208]. Dems [466]. Denationalise [781]. Denotation [1205]. Denotational [1588, 187]. Dense [1469, 1384]. Dense-Time [1384]. Density [2227]. Dependence [41, 137, 1498]. Dependence-Free [41]. Dependences [42, 204]. Dependencies [1706]. Dependency [1798]. Dependent [1550, 114, 1859]. Deposit [963]. Depth [2006, 139, 2165, 855, 525, 60, 2128]. Depth-First [855, 60]. Derandomized [724]. Derivation [715, 1052, 385].

Derivational [1177]. Derivations [1297, 346]. Derive [670]. Derived [2073]. Deriving [638, 85]. DES-X [1434]. DESAM [1615]. Described [527, 1233]. Describing [704, 1713, 1769]. Description [899, 1665, 125, 146, 763, 2113, 620, 973]. Descriptions [1718]. Descriptor [2267]. Descriptors [1629, 956]. Design [1301, 2205, 1379, 1589, 808, 345, 1421, 1523, 1293, 122, 14, 602, 171, 1312, 1215, 664, 1827, 1114, 104, 646, 1673, 1279, 832, 1578, 115, 113, 148, 1652, 1768, 320, 1591, 1280, 326, 163, 1401, 1517, 1411, 74, 1749, 1277, 151, 1377, 101, 1745, 353, 336, 1906, 683, 252, 1882, 675, 1937, 1923, 638]. Designer [1298]. Designing [235, 1101, 1501, 815, 1375, 696]. Designs [160, 263, 264]. Desirable [206]. Desired [1114]. Deskewing [607]. Detecting [2098, 2249, 2057, 850]. Detection [588, 495, 2082, 600, 601, 167, 1998, 2085, 506, 557, 1599, 1342, 570, 592, 587, 1506, 2171, 464, 618, 2029, 160, 1124, 2146, 2159, 626, 2157, 2113, 560, 589, 1573, 544, 573, 2063, 594, 444, 2068, 1040, 614, 158, 2077, 2144, 718, 678]. detection/correction [718]. Detector [846, 458, 1051, 310, 286]. Detectors [456]. Deteriorating [97]. Determination [1109, 2050, 749]. determine [1014]. Determines [1072]. Determining [764, 1016, 497, 565]. determinism [654]. Deterministic [1679, 1091, 744, 9, 995]. Develop [650]. Developing [142, 1363, 1364, 811, 1306]. Development [1068, 1341, 159, 1312, 173, 1408, 1062, 1404, 384, 1281, 1387, 1528, 221, 594, 1672, 1064, 1066, 2075, 256, 1940]. Developmental [1088, 1150]. Developments [1388]. Device [110, 111, 250]. Devices [1370, 281]. DEXA [337]. DFA [725]. DFS [840]. DFS-Based [840]. DFT [281]. Diagnosing [1719]. Diagnosis [236, 581, 882, 586, 682]. Diagnostic [236, 809, 938]. Diagonal [1502]. Diagonalization [241, 934].

Diagram [1806]. Diagrams [639, 1209, 644, 1378, 2214, 737]. DIALOGS [684]. Dialogue [239, 967, 986]. Diameter [1482]. Dictionary [1995, 1600]. DICUT [1461]. Diff [1786]. Difference [1770]. Differences [1111]. Different [872, 1360, 1179, 583, 915, 1550, 715, 1313, 95, 2263, 1641, 1167, 2149, 1316, 2256, 918]. Differential [1764, 1757, 1436, 1435]. Difficult [2082]. Diffusion [117, 123]. digit [987]. Digital [943, 374, 373, 1849, 462, 966, 1445, 958, 964, 576, 2000, 952, 2232, 946, 794, 2010, 951, 2240, 1842, 425, 1855, 795, 965, 790, 1850, 2089, 2266, 511, 573, 902, 949, 942, 2256, 947, 1649, 1846, 291, 298]. Digitization [1853]. Digitizing [2062]. Digraphs [466, 1481]. Dijkstra [893]. Dilation [30]. Dimension [1117, 1477, 1401, 740]. Dimensional [612, 477, 703, 2180, 1994, 1996, 124, 2182, 1048, 1734, 1839, 2247, 2199, 25, 1972, 394, 2008, 1945, 252, 1855, 1969, 2178, 1763]. Dimensionality [1129, 1130, 49]. Dimensions [2179, 940]. Diophantine [911]. Direct [324, 1214, 566, 93, 220]. Directed [715, 193, 1604, 1479, 1694]. Directing [315, 2072]. Directional [610]. Directions [1218]. Disaggregate [1204]. Disassembly [807]. Discarding [2265]. Discipline [1527]. DisCo [669]. Disconcepts [1731]. Discontinuity [473]. discordant [1960]. Discounting [1132]. Discourse [889]. Discovering [994, 979]. Discovery [827, 1793, 378, 1565, 685, 970, 1005]. Discrepancy [1817]. Discrete [1847, 1385, 2210, 1550, 1851, 1852, 1398, 1710, 1440]. Discretely [1787]. Discretization [1084, 993]. Discriminant [871, 1624]. Discriminating [1637]. Discrimination [890, 1132, 1043]. Discs [411]. disjoint [1973]. disjunctions [746, 743]. Disjunctive [1263, 1270, 366, 1825]. Disk [410]. disorder [996]. Disorders [1719].

Disparity [2052]. Displacement [1350, 610]. Displacements [505]. Display [549, 488]. Distance [1213, 474, 1560, 1530, 1854, 1214, 1976, 1561, 1974]. distance-hereditary [1976]. Distances [225]. Distinctness [1606]. Distortion [710]. Distributed [1700, 1338, 61, 1322, 419, 402, 673, 1351, 1365, 1551, 1323, 1586, 1490, 856, 1373, 1776, 1402, 1693, 432, 1251, 1328, 645, 1284, 1406, 853, 367, 145, 363, 1547, 644, 412, 1577, 1501, 192, 47, 314, 438, 75, 1500, 1516, 433, 953, 1519, 2262, 130, 312, 1454, 369, 326, 838, 1337, 1517, 923, 404, 1514, 850, 408, 420, 1567, 86, 1302, 1948, 1911]. Distributed-Memory [1351, 1519]. Distributing [1303]. Distribution [800, 94, 46, 12, 1422]. Distributions [2147, 1138, 1133, 1243, 480, 1403, 726]. Disturbance [1172]. Disunification [1255]. diterpene [676]. Divergence [655]. Divide [561]. Divide-and-Conquer [561]. Dividing [82]. Diving [1522]. Divisibility [1799]. Division [1823]. Do [143]. Document [1628, 955, 1633, 1634, 979, 2092, 1680, 2091, 1636, 1635, 1632, 1026, 1422]. Documents [374, 373, 960, 955, 357, 434, 900, 961, 1632, 822, 249, 437, 1018, 973, 1020, 1019]. does [1861]. Doesn’t [1290]. Domain [892, 94, 1279, 815, 957, 372, 1034, 1021, 1863, 1909, 269]. domain-specific [269]. Domains [887, 1589, 1816, 1789, 2246, 2149, 835, 534, 739]. Dominance [1065]. Dominant [1702]. DOOD [1680, 377]. Door [2039]. Dopamine [1036]. Double [1080]. DOVE [1937]. Down [486, 862, 689]. Drape [2138]. Drawable [2174]. Drawing [2205, 2215, 2181, 2212, 2202, 2187, 2216, 2206, 2184, 2201, 2191, 2193, 2179, 2199, 2214, 2204]. Drawings [2211, 2203, 2182, 2186, 2189, 2178, 2183, 2192]. Drive [1065]. Driven [1360, 874, 826, 367, 646, 364, 273]. Driver [2264]. DRS [903]. drug [675]. DS [1132]. DSM [59, 2266]. DSM-CC [2266]. DSP [282, 106]. DSS [1405]. Duality [1427]. due [1084]. Duration [1791]. During [1747, 1279, 1066, 978]. Dutch [963]. Dynamic [1988, 1348, 1338, 2195, 73, 1976, 2080, 920, 1365, 125, 1688, 558, 783, 1172, 1693, 1408, 2034, 2037, 628, 4, 1614, 2172, 1033, 641, 461, 408, 1785, 516, 1643, 250, 259, 253]. Dynamical [1136, 1099, 1040]. dynamically [268]. Dynamics [1361, 1229, 1366, 1052, 235, 1314, 1042, 1336].

E-Cash [774, 789]. E-hyper [217]. E-Money [791]. Earley [867]. Easier [1260]. Easy [494]. ECCC [966]. ECEM [1769]. ECG [681, 875, 246]. Echo [1171]. Ecologies [302]. Economical [1079, 1842]. ECS [1708]. Edge [456, 2082, 458, 2085, 2197, 464, 2029, 2035, 2103, 560, 2089, 444, 1952]. Edge-Preserving [2035]. Edges [475, 2118, 2189, 2192]. EDI [370]. Editing [1818]. EDPEPPS [14]. Education [1529, 1530]. Effect [139, 1625, 2196]. Effective [1117, 458, 197, 180, 184, 1426]. Effectively [118]. Effects [733, 1073, 1771, 765, 179, 1859]. Efficiency [76, 2241, 2246, 1521, 1403, 1876]. Efficient [1964, 469, 1028, 877, 1440, 2003, 1600, 891, 800, 426, 409, 1794, 788, 1689, 1611, 982, 660, 1439, 1656, 1798, 2150, 69, 1458, 60, 19, 691, 1750, 778, 686, 198, 2257, 1981, 297, 737, 1961, 1432]. efficiently [734]. Ehrenfels [749]. Eigenspace [2129]. Eigenvalues [91]. Eight [2256, 2129]. Elastic [2190, 555, 505]. Electric [1170]. Electronic [966, 781, 955, 801, 1370, 800, 370, 417, 780, 963, 175, 778, 798, 2260, 1457]. Electrostatic [1771]. Element [1755, 1606]. Elements [507, 171, 1533, 282]. Elevation [1649]. Elicitation [828]. Eliminate [1504]. Elimination [1471, 678, 1922]. Ellipses [1853]. Elliptic [2235, 1757, 93, 1439].

Elliptical [1094]. Elman [1100, 1103]. elucidation [676]. Embedded [1782, 1474, 1383, 1499]. Embedding [2176, 971, 124, 1334, 2185, 28, 399, 880, 25]. Embeddings [2061]. Embodiment [1145]. Embodiments [940]. Embodying [81]. Emergence [1221]. Empirical [1129, 1264, 813, 1722, 310, 1208, 158, 994]. Empiricism [749]. employing [276]. Emulation [28]. EMV [783]. Enabling [338]. enactment [906]. EnCal [1520]. Encapsulated [955]. Encapsulating [363]. Encoders [2233, 2250]. Encoding [871, 1872, 541, 1842, 691]. Encodings [1612]. Encryption [2263, 2251, 1412]. End [1489, 1050, 2259]. End-To-End [1489]. endgame [678]. endings [1014]. Enemies [1486]. Energy [1379, 450, 49, 482]. Enforcing [1450, 2151]. Engine [2198, 99, 1017]. Engineering [1589, 810, 658, 1386, 393, 154, 417, 365, 1609, 782, 1381, 140, 1740, 1376, 395, 2271, 143]. Engineers [143, 144]. Engines [1273]. English [983]. Enhance [1226, 947]. Enhanced [138]. Enhancement [452, 1397, 1998, 104, 264]. Enhancing [1744, 1027, 282, 944, 2077]. Ensemble [1341]. Ensuring [1928]. entailment [745]. Enterprise [810, 1291, 1310]. entities [1878]. Entity [1307, 1278]. Entity-Relationship [1307, 1278]. Entropy [468, 1119]. Entry [72]. enumerating [1951]. Enumeration [2250]. Environment [1229, 419, 542, 402, 1340, 340, 1311, 579, 1776, 1370, 1386, 109, 1328, 1759, 1406, 1343, 148, 1330, 424, 1033, 1517, 649, 1274, 1672, 1331, 2260, 353, 299, 278, 1422, 1939, 1890]. Environmental [1805, 1214]. Environments [427, 244, 1601, 1029, 432, 307, 7, 1807, 2267, 403, 1319]. Epipolar [1997, 2139, 494]. EPK [417]. EPK-fix [417]. Equality [221]. Equation [1138]. Equations [1764, 1546, 1048, 2237, 40, 1757, 2153, 911, 1359, 994]. Equipment [422]. Equivalence [2221, 1821, 1526, 1977]. Era [1529]. EREW [53, 1963]. EREW-PRAM [53]. Erich [769]. Error [2224, 2232, 2234, 2135, 2240, 526, 1698, 2059, 2064, 2256, 510, 2257, 718]. Error-Correcting [2240, 526]. Error-Propagating [2257]. Errors [1854, 2139]. Escrow [2246, 1451]. ESPRIT [620]. Esterel [1812]. Estimate [1115]. Estimates [595]. Estimating [2006, 521]. Estimation [2158, 1130, 1997, 921, 396, 1499, 2164, 618, 2031, 1852, 500, 496, 1516, 2150, 2052, 454, 2056, 2065, 1853, 400, 2071, 2162, 184, 2140, 2151, 1436, 268]. Estimations [1213]. Estimator [1142, 1085, 727]. Estimators [2155]. Ethernet [1315]. EUBAFES [1097]. Euclidean [474, 2100, 2087]. Euler [1851, 2148]. European [1352]. evaluated [265]. Evaluating [878, 1746, 780, 1319, 1741]. Evaluation [429, 126, 1340, 14, 1495, 1624, 1688, 234, 813, 531, 209, 1689, 311, 103, 10, 2263, 2038, 1641, 2248, 615, 1500, 178, 1318, 635, 1320, 1803, 2074, 980, 1937, 254]. Event [1223, 1385, 631, 646, 1404, 1710, 163, 1889, 1685, 1899]. Event-Based [1223, 1899]. Event-Driven [646]. Events [73, 398, 1788]. Everything [1541]. Evidence [230, 2134, 796, 1004, 998]. Evidence-Based [2134]. Evidential [1598]. Evolution [1287, 388, 319, 1283, 909, 1144, 307, 304, 1149, 312, 1749]. Evolutionary [305, 318, 303, 1088, 2235, 311, 1168, 2052, 320, 315, 1702, 1150, 472, 276]. Evolved [1167]. Evolving [1380, 972]. Exact [736, 1818, 194, 1054, 735, 1969, 737]. Examining [378]. Example [2161]. Examples [1794, 1002, 721, 695, 730, 725]. Exception [1603]. Exceptionally [1248]. Exceptions [248, 184]. Exchange [19]. Exdasy [12]. Executable [171, 1794, 1708]. Executing [427, 1875]. Execution [638, 1780, 1, 1861, 1882]. Executions [631]. Exercise [105, 838]. Existing [822, 1010]. Expansion [2030, 949]. Expect [144]. Expectancies [765]. Expectation [766]. Experience [1300, 177, 1969]. Experiences [42, 1363, 159, 1751, 11]. Experiencing [1387]. Experiment [657, 160, 758, 1014, 675]. Experimental [878, 2163, 81, 1320, 614, 51]. Experiments [507, 108, 1549, 523, 1226, 1336, 93, 123, 1093, 1359]. Expert [236, 1706, 247, 268, 606]. Experts [1081, 1115]. Explaining [1293]. Explanation [1225, 1181, 935]. Explicit [67, 148]. Exploiting [63, 269, 1279, 1668, 835, 1524, 1010]. Exploration [895, 908, 2051, 333]. explorer [979]. Exploring [1029, 1699]. Explosion [649]. Exponential [572, 1910]. Exponentials [220]. exponentiation [1439]. Expressing [1864, 172]. Expression [2045, 2067]. Expressions [830, 1765, 2169, 2168, 1732, 66]. Expressive [1238, 1679, 1822, 1829, 772]. expressivity [1876]. Extendable [12]. Extended [703, 1285, 664, 1824, 923, 1202, 1103]. Extending [388, 445, 972, 1067, 1693, 1294, 1880, 253, 2214, 1730, 1929, 1879]. extensibility [999]. Extensible [340, 1761, 1750]. Extension [1322, 647, 1666]. Extensional [1657, 876]. Extensions [1466, 1378, 728, 1953]. External [1654, 218, 651]. External-Memory [1654]. extortion [1458]. Extracting [475, 930, 2106, 331]. Extraction [2167, 871, 2003, 608, 1645, 455, 2119, 610, 1732, 2131, 1017, 977]. Eye [1053]. Eye-Blink [1053].

F [106]. F-Code [106]. fabrication [252]. Face [2167, 2165, 587, 2173, 2042, 2045, 491, 2077, 1953, 148]. Face-Only [2173]. Faces [2014, 2161]. Facets [870]. Facial [570, 2164, 2171, 2163, 2169, 2168, 2067].

Facilities [1680]. facility [1943]. Factor [1056]. Factorization [89]. facts [1885]. Fading [2241, 2242]. Failure [846, 655]. Failure-Divergence [655]. Fair [758, 1458, 1449]. Fairness [853]. Families [554, 925, 1533]. Family [24, 1184, 85, 816]. Farming [1735]. Fashion [2033]. Fast [1986, 1865, 2158, 1992, 233, 1995, 281, 587, 478, 1852, 526, 1500, 1315, 1979, 2060, 93, 565, 2096, 584, 180, 1437]. Faster [1748, 1972]. Fat [22, 23]. Fat-Trees [22, 23]. Fault [1801, 851, 1114, 782, 23, 1455, 837, 133, 1487, 1948, 1897, 1898, 264, 1911, 1913, 1957]. fault-monotonicity [1913]. fault-tolerance [1913]. Fault-Tolerant [837, 133, 1487, 1948, 1898, 1911, 1957]. Faults [2239]. Faustian [1290]. FCC [1857]. Feasibility [1563]. Feasible [165, 49]. Feature [1213, 896, 506, 1141, 1147, 570, 909, 1184, 1063, 2084, 332, 2119, 1256, 2169, 480, 1572, 1097, 679, 584, 1105, 977]. Feature-Based [332, 1572]. Features [452, 871, 1282, 1624, 2086, 609, 585, 1175, 2090, 2030, 610, 1637, 517, 753, 598, 2095, 743]. FEC [409]. Federated [387, 388, 389, 338, 390]. Federation [1309]. Feed [1704]. Feed-Forward [1704]. Feedback [1298, 412, 1160, 949]. Ferrite [626]. FFT [1514]. Fiat [1202]. Fide [1202]. Field [1069, 457, 1072, 1765, 1082, 574, 2056, 572, 1211]. Fields [1073, 1068, 564, 2023, 497, 1064]. Fifths [756]. Figures [1210]. Files [946]. Fill [1470]. Fill-In [1470]. Filled [1645]. Filled-in [1645]. Filling [1653]. Film [328]. Filmification [328]. Filter [2011, 2012, 2081, 286]. Filtering [881, 1697, 2013, 1728, 714, 1237]. Filters [2086]. Financial [803, 784, 1704, 785, 792]. find [1963]. Finder [2142]. Finders [1734]. Finding [31, 2182, 1269, 2237, 2025, 1692, 711, 1973, 1974, 1426]. Fine [54, 101]. Fine-Grain [101]. Fine-Grained [54].

Finger [1738]. Fingerprint [1994, 2064, 596]. Fingerprint-Based [1994]. Fingerprints [2014]. Fingertips [938]. Finishing [247]. Finite [1789, 1755, 1838, 1840, 1770, 1610, 1965, 744]. finite-memory [744]. Fire [1060, 1051, 1042, 1045]. First [919, 359, 1599, 855, 862, 60, 820, 1485, 745, 974]. First-Order [919, 862, 745]. Fitted [530]. Fitting [522, 2078]. Fitting-and-Splitting [2078]. Five [1264]. fix [417]. Fixating [2066]. Fixed [1614, 926, 282]. Fixed-Point [1614]. Fixpoints [1532, 1709, 1916]. Flexible [1021, 866, 1311, 607, 905, 906, 1289, 817, 787, 259, 1015]. Flies [1148]. flights [1942]. flip [238]. Floating [1910]. Flow [1989, 164, 636, 1697, 1770, 193, 10, 2265, 194, 2153, 2154, 2156, 17, 1928]. Flows [1368, 114]. Flowshop [315]. Fluid [1368]. Fly [955]. FOL [901]. Folders [327]. Folk [955]. Follow [1133]. FOLRE [629]. Forbidden [2174]. Force [1161, 1160]. Forcing [1061]. forest [1963]. Form [1633, 765, 1644, 671, 2096, 1023]. Formal [1674, 1589, 359, 764, 1221, 1181, 650, 662, 658, 1402, 2271, 1604, 71, 1532, 1534, 1382, 1308, 666, 1295, 1787, 663, 641, 1908, 1920, 1923]. Formalism [1632]. formalisms [998]. Formalization [1782, 704]. Formalizing [177, 644, 1562]. format [1935]. Formation [1061, 1074, 1002]. Formed [1727]. Forms [1706, 698, 2025]. Formulae [878, 1954, 1966]. Formulas [1525]. Formulation [1946, 1510]. Fortran [1512, 1748]. Forward [804, 1704, 1520, 26]. Foundation [1540, 1581]. Foundations [1677, 2271]. Founded [1678]. Fourier [532, 1398]. Fovea [588, 492]. Foveated [2051]. FPGA [275, 271, 263, 264, 294, 291, 297, 252, 286, 253, 251, 298, 537]. FPGA-SRAM [275]. FPGAs [257, 255, 290, 273, 272, 274, 262, 254]. FPLD [276]. Fractal [477, 871, 2000, 541, 705, 1856].

Fragmentation [1303]. Frame [945]. Frames [874, 1085]. Framework [379, 240, 825, 1764, 359, 636, 1772, 1224, 348, 1755, 801, 435, 1765, 1766, 332, 2120, 1757, 416, 1381, 1308, 146, 148, 666, 2122, 763, 1518, 69, 1281, 606, 1777, 163, 2266, 510, 2073, 1788, 377, 1302, 1958, 1022, 1430]. Frameworks [1744, 826]. France [965]. Free [846, 41, 1469, 1171, 1536, 243, 857, 858, 505, 1483, 1470]. FreeBSD [2264]. Frequency [600, 1403]. Friendly [12]. Front [2005, 1050, 226]. FSM [273]. Full [1926, 17, 254]. Fully [1311, 853, 92, 1517, 1001]. Function [1338, 1094, 2234, 1164, 527, 1910, 1454, 1924]. Function-Described [527]. Functional [1556, 1706, 203, 341, 450, 84, 1062, 67, 200, 81, 1803, 187]. Functionality [806, 1107]. functionals [740]. Functions [1849, 199, 2007, 1095, 466, 1426, 2252, 2254, 705, 1074, 1428, 980, 746, 720, 724, 1427]. Fundamental [1578, 706]. fundus [682]. Furnace [1170]. Fusion [452, 2006, 599, 754]. Future [1628, 1621, 1575, 1489, 1305, 1494, 797]. Futures [940]. Fuzzy [1717, 495, 429, 445, 345, 468, 234, 1560, 154, 1704, 1637, 1640, 1211, 1731, 463, 1233, 1561, 299].

GA [990, 989]. GA-based [989]. Gabor [2086]. Gadgets [1460]. Gallery [2205]. Game [1551, 1699, 1617, 553, 719, 738]. Gamma [1049]. Ganglion [1056]. Gas [1135, 2136]. Gateway [801, 2263]. Gateways [337]. gathering [1860]. Gauge [96]. Gauss [2194]. Gaussian [1081, 2135, 2050, 595]. GCD [296]. GEM [1381]. General [129, 607, 1836, 2197, 925, 368, 1262, 1509, 2050, 1815, 1777, 768, 2145, 1685]. General-Purpose [2197]. Generalised [640, 2255]. Generalization [1084, 588, 1100, 1209, 79, 1651, 976, 1017]. generalizations [690]. Generalized [582, 607, 1007, 119, 1983, 1829, 1788]. Generating [889, 907, 2008, 2141]. Generation [1683, 888, 2231, 1642, 1743, 697, 1625, 217, 1488, 1189, 1588, 151, 2067, 336, 1944, 1885, 1887, 970]. Generative [1309, 1866]. Generator [656, 170, 1429]. Generic [638, 1001, 154, 1381, 314, 1295, 1266, 1616, 397]. Genes [1702]. Genetic [301, 1997, 1101, 2026, 2037, 1501, 113, 314, 901, 312, 316, 317, 1426, 1883, 309]. GENIUS [1266]. GENIUS-CP [1266]. GenSAT [912, 238]. Genuine [847]. Geocognostics [1224]. Geodesic [2005]. Geographic [1205]. Geographical [339, 1225, 947]. Geometric [717, 2101, 1076, 1268, 707, 1816, 2198, 2010, 1222, 1063, 713, 1841, 1647, 1655, 520, 702, 2151]. Geometric-Based [702]. Geometrical [485]. Geometrically [503, 2123]. Geometrically-Correct [2123]. Geometries [588]. Geometry [582, 1997, 1076, 1072, 710, 706]. Gerhart [1554]. Gestalt [771, 751, 755, 750, 749]. Gestalten [760]. Gestalts [753]. Gesture [2160, 571, 1738]. Gesture-Recognition [571]. Gestures [2126, 2172]. GF [2233]. GIS [1654, 1229, 1230, 1231, 1364, 1226, 1648, 1215]. Given [1471]. Glasses [2171]. glaucoma [682]. Global [896, 1081, 514, 2049, 351, 1237, 850, 1331]. Goal [967, 715, 1798]. Goal-Directed [715]. Goal-oriented [967]. Goals [932]. Gofer [652]. GOLOG [1690]. Good [1442, 1484, 758, 721]. Goodenough [1554]. Gossip [26]. Government [799]. Gradation [1203]. Gradient [1080, 442, 454]. Gradually [2151]. Grain [101]. Grained [54]. Grammar [875, 1191]. Grammars [865, 2116, 203, 1536, 1178]. Grammatical [1632]. Grams [962]. Grand [803]. Granularities [1267]. Granularity [872, 935]. GRAPh [2207, 2205, 2215, 2181, 2212, 2176, 2188, 2195, 2202, 1818, 1471, 2182, 2216, 2206, 1570, 162, 2198, 2184, 1298, 2185, 2201, 2191, 1915, 1914, 1479, 55, 1155, 526, 2179, 2199, 262, 2200, 1483, 1358, 635, 2204, 2192, 355, 1948, 686, 2093]. Graph-Drawing [2216]. Graph-Partitioning [1358]. Graphic [641]. Graphical [1225, 14, 640, 157, 645, 35, 246, 1636]. Graphics [1627]. Graphplan [1876, 1879]. Graphs [1462, 1464, 1463, 1469, 1470, 1818, 2210, 1474, 1475, 2187, 1801, 1477, 2209, 1269, 1478, 2193, 119, 2103, 1480, 2178, 1481, 2183, 1619, 527, 1487, 2213, 1964, 1976, 1977, 1975, 1880, 1956, 1951]. Grappa [2207]. GraVis [2208]. Greatest [2196, 1916]. Greedy [1475, 842]. Grey [457, 591, 2098, 517]. Grey-Level [591, 517]. Grey-Scale [2098]. Grid [233, 24, 2187, 2185, 2178, 2089, 2203, 313, 519]. Grid-Based [233]. Grids [513, 124, 1843, 1857]. Ground [1210, 2034, 2146]. Groundness [183]. Group [2269, 848, 160, 1320, 1830]. Grouping [2118, 711, 702, 1105]. Groupoids [1195]. Groups [701, 1419]. Growing [474, 1139]. Growth [1056, 556, 2095]. Guarantees [434]. Guidance [1165, 395, 1005]. Guided [2009, 2143]. Guidelines [2272, 795]. Guido [749]. GUMP [803]. GWT [2054].

Hamiltonian [44, 1464]. Hand

[2126, 2037, 2094]. Hand-Written [2094]. Handling [34, 489, 693, 1603, 862, 392, 1371, 1401, 101, 678]. Hands [1171, 2018]. Hands-Free [1171]. Handwriting [1621, 609, 1641]. Handwritten [1623, 1999, 602, 610, 1086, 606, 1637, 1640, 987, 1622]. Hard [31, 1944, 886, 1248, 1971, 2185, 1941]. Hardness [2198, 1460]. Hardware [832, 540, 292, 671, 537, 1526, 295, 277, 265, 283, 293]. Hardware/Software [832, 671, 277]. Harmonic [770]. Harmony [760]. Hash [2234, 2252]. Haskell [86]. Having [2050]. Haystacks [1692]. Hazard [1499]. HCC [2234]. HDL [276]. Head [1895, 2165, 577]. Head-tactics [1895]. Heads [494, 2068]. Heartbeat [846]. Heavy [1243]. Heavy-Tailed [1243]. Hebbian [1151, 1060, 1059, 1058, 1064]. Helical [581]. Helpful [769, 726]. hereditary [1976]. Herlihy [839]. Hesitated [2172]. Hestenes [704]. Heterogeneous [339, 1324, 2261, 13, 2267, 230, 997, 272]. Heuristic [1566, 925, 1786, 1272, 315, 1358, 352]. Heuristics [1575, 905]. Heyting [729]. HHD [1470]. HHD-Free [1470]. Hidden [1444, 1838, 1924]. Hiding [1444, 2253, 1431]. Hierarchical [1717, 2117, 1814, 2080, 2001, 169, 2189, 2026, 1792, 1704, 525, 515, 404, 559, 1200, 1419, 1874, 1892]. Hierarchically [2193]. hierarchies [1960]. Hierarchy [486, 1298]. High [1493, 1754, 600, 1994, 1746, 1492, 1836, 1497, 1601, 1490, 107, 609, 836, 104, 2241, 2243, 1767, 175, 33, 5, 1514, 1319, 106, 276, 1923, 1512]. High-Dimensional [1994]. High-Level [1754, 609, 276, 1923]. High-Performance [1493, 107, 1767, 33]. High-Speed [104]. Higher [1254, 2024, 1820, 2090, 200]. Higher-Order [1254, 2090, 200]. Higher-Orders [1820]. Highly [1263, 802, 1356, 1426]. Highway [534]. Hill [1260]. Hill-Climbing [1260]. Hippocampal [1061]. Histogram [576, 554, 595]. Histogram-Based [576]. Histograms [401]. Historical [373, 748]. Histories [347, 631]. HMM [1624, 2160, 1639]. HMM-based [1624]. Hoare [1833, 1834]. Hoare-Style [1833, 1834]. Hodgkin [1048]. Hogan [2244]. HOL [1935, 1910, 655]. Holographic [532]. Homogeneity [467]. Homogeneous [1463, 1736]. homomorphic [722]. Homophonic [1967]. Homotopy [1855]. Hong [2033]. Hop [841]. Hopfield [1107]. Horizontal [342]. Horn [745, 1954, 861, 1953, 224, 1024, 882]. Hostile [782]. Hot [1724]. Hough [2079, 711]. HPC [1494]. HPSG [1184]. HPSG-Style [1184]. HTML [1632]. HTN [1873]. HTN/ [1873]. hull [1969]. Human [2167, 2114, 2018, 2126, 496, 2196, 491, 2068, 2162, 2077]. Humans [1043]. Huxley [1048]. Hwang [1438]. Hybrid [888, 869, 1366, 1162, 1508, 1813, 1791, 1639, 1958, 1873]. Hydrographic [1228]. HyNoDe [421]. hyper [217]. hypercube [1949]. Hypercubes [2177]. hyperelliptic [1440]. hypergraphs [1954]. Hypermedia [1028, 889, 418, 350, 336]. Hypersequential [1510]. Hypertableaux [216]. Hypertext [1027, 888, 1030, 1031, 1032, 1033, 1034]. Hypothesis [2101, 2135, 755].

I.P. [542]. IBM [122]. IBPhM [2123]. ICC [1746]. IceT [1776]. Iconic [699]. IDAF [1936]. IDEA [1658]. Ideals [568, 1399]. Idempotent [1255]. Identifiability [722]. Identification [1993, 1994, 1633, 2014, 1244, 591, 1170, 2037, 2091, 2096, 1211, 267, 1425]. Identifying [2161, 491, 36]. Identity [1183, 2015, 2042]. IDL [902, 954]. Idle [315]. Idle-Time [315]. IFS [1856]. Ignoring [1885]. II [1674, 1286, 1844, 1496, 1481]. Ill [1727, 835, 1012]. Ill-Formed [1727]. ill-known [1012]. Ill-Structured [835]. Illumination [1985, 2133]. Illusion [2163]. Illusion-Experimental [2163]. Illusions [1076, 449]. ILP [44, 676, 693, 687]. Image [612, 452, 1397, 456, 1493, 1779, 871, 2125, 459, 445, 493, 532, 1071, 576, 535, 1995, 1719, 2085, 704, 2000, 554, 551, 357, 579, 468, 611, 105, 616, 446, 2009, 2010, 234, 501, 531, 2018, 2160, 1769, 2118, 2026, 2027, 699, 618, 625, 591, 1842, 2090, 479, 2030, 2033, 547, 2159, 1806, 500, 530, 2092, 1400, 548, 2119, 546, 550, 2045, 615, 2168, 555, 2053, 358, 505, 549, 2121, 454, 2154, 586, 2083, 1201, 326, 1198, 2061, 533, 1636, 1396, 2108, 553, 2065, 518, 465, 473, 400, 98, 2099, 594, 2070, 559, 2071, 2131, 2130, 2073, 537, 2152, 485]. Image [2123, 2144, 471, 2141, 472, 1001, 968, 288, 989]. Image-Based [2045, 2123]. Image-Derived [2073]. Image-Processing [105]. Image-Schemata-Based [1198]. ImageRoadMap [358]. Imagery [1227, 2016]. Images [477, 2167, 356, 443, 582, 476, 1844, 600, 2101, 2003, 541, 2006, 575, 450, 478, 470, 2020, 2171, 464, 581, 2136, 1839, 455, 2098, 577, 1855, 961, 1848, 2045, 2172, 2113, 2048, 620, 2091, 2054, 2057, 2089, 1696, 334, 511, 533, 2062, 2139, 621, 498, 598, 539, 401, 2148, 330, 331, 1070, 722, 682]. Imaging [2082, 613, 622, 614]. Imitation [1157]. Immune [2254, 1428]. Impact [2269]. Impediment [1204]. Imperative [1821, 1797]. Imperfection [392]. implementable [1456]. Implementation [1379, 1754, 1350, 110, 345, 430, 2000, 255, 1755, 1335, 409, 1496, 1813, 784, 425, 1673, 1756, 1500, 349, 1357, 1616, 667, 1824, 1517, 81, 635, 1332, 1803, 281, 683, 294, 291, 219, 1554, 286, 985, 283, 1437, 298]. Implementations [122, 2079, 192]. Implemented [537]. Implementing [1512, 2197, 1789, 1355, 1680, 2266, 1058, 661, 1969]. implication [690]. Implications [1145]. Implicit [2024, 1164]. Importance [1489]. Imposed [2049]. Imprecise [445]. improve [990]. Improved [1990, 489, 450, 1106, 2108, 1437, 1830, 981, 1429]. Improvement [149, 1043, 1590, 1403, 1745]. Improvements [459]. Improving [202, 603, 1568, 1215, 2079, 1120, 1520, 415, 1247, 1097, 1460]. incomplete [689, 1007]. Inconsistent [1245]. Incorporating [1230, 334]. Increase [791, 279]. Increasing [2246, 1107, 975]. Incremental [343, 214, 1663, 909, 173, 1289, 1153, 386, 385, 2179, 2204, 1893]. Independence [1469, 918]. Independent [1124, 1061, 1123, 1118, 1720]. Independently [82]. Index [381].

Indexing [1994, 541, 956, 332, 547, 2053]. Indices [1506]. Indirect [198]. Induced [1055]. Induction [1721, 1794, 931, 782, 862, 995, 689, 982]. Inductive [1657, 684, 1712, 863, 876, 218, 674, 1565, 835, 733, 981, 731, 675, 679, 677, 993]. Industrial [108, 1352, 177, 2022]. Industry [2033, 144]. Inequalities [195, 1222]. Inertial [1374]. Inessential [1175]. Infeasible [164]. Inferability [720]. Inference [863, 907, 713, 2120, 1159, 2106, 1273, 733, 1023, 731]. Inferences [1198]. Inferring [873, 730]. Infinite [1610, 2273]. Infinite-State [1610, 2273]. Influence [1084]. Infomax [1120]. informal [1908]. Informatics [1529, 1530]. Information [387, 2224, 1988, 1223, 2167, 1860, 881, 495, 1027, 164, 339, 1996, 323, 1046, 1199, 938, 391, 1224, 327, 1204, 1529, 1728, 1645, 324, 332, 322, 500, 1037, 1214, 2106, 1281, 1574, 1216, 1696, 1573, 1035, 1145, 957, 366, 1306, 822, 333, 331, 1596, 997, 1007, 1894, 1011, 799]. Information-Theoretic [1046]. InfoSphere [1574]. InfoSphere-V [1574]. Infrastructure [953]. Infrastructures [433]. Inharmonic [759]. Inheritance [343, 1671, 1298, 1687, 364, 917]. Initial [1734, 675]. Initiative [932, 967]. initiator [254]. Injection [1109, 1317]. Inlining [180]. innocuous [1444]. Innovative [2115]. Input [1549, 1212, 1109, 1061, 1110]. Inputs [1068, 1172]. Insertion [1600]. Inspection [1990, 108, 601, 592, 2022]. Inspired [302, 1389]. instance [1944, 1790]. Instant [419]. Instruction [136, 1491]. Instruction-Level [136]. Instructions [100, 1498]. Instrument [826]. Integer [1349, 34, 1703]. Integral [497, 491]. Integrate [1060, 1051, 1042, 1045]. Integrate-and-Fire [1060, 1042, 1045]. Integrated [653, 1379, 476, 14, 150, 382, 629, 1727, 153]. Integrating [1339, 810, 1908, 2014, 628, 72, 221]. Integration [495, 1300, 1556, 1351, 161, 234, 1307, 2045, 1593, 663, 622, 917, 642, 997, 1920]. Integrity [360, 687, 361]. Intelligent [1391, 936, 914, 1020, 1387, 1280, 902, 1405, 619, 1878, 1011, 1457]. Intended [916]. Intensional [1670, 1657, 1688]. Intensity [502, 582, 2101, 1400]. Intensity-Based [502]. Intensive [127, 133]. Intentions [920]. Inter [1985, 1504]. Inter-Processor [1504]. Inter-reflection [1985]. Interaction [1231, 453, 1342, 2126, 853, 1154, 192, 1038]. Interactionally [1586]. Interactive [402, 2202, 1549, 552, 580, 2203, 2184, 414, 440, 436, 531, 833, 1507, 2213, 1008]. InteractiveGiotto [2202]. Interchange [370]. Interconnecting [13]. Interconnection [1465, 129, 2180, 18]. Interdisciplinary [1213, 860]. Interface [1231, 1328, 617, 382, 1344, 643, 246, 1738, 1749, 1808, 1310]. Interfaces [157, 1408, 213]. Interleaving [2256]. Intermediate [193, 1790, 106]. internal [988]. International [2246]. Internet [1453, 163, 1808]. Internet-Scale [163]. Interoperability [403]. Interoperation [1324, 1009]. Interplay [1038]. Interpolation [618, 1396]. Interpretation [202, 191, 207, 208, 209, 206, 205, 772, 2134, 204, 249, 2104, 187, 1903, 1909, 1004, 729, 989]. Interpreted [1382]. Interpreting [1217]. Interprocedural [37]. Interreflections [2076]. Intersection [2210]. Interval [356, 1782, 1818, 760, 840, 30, 1261, 1003]. Interval-Based [356]. Interverted [770]. Intonational [773]. Intracortical [1071]. Intranets [338, 776]. Intrinsic [1130, 2100]. Introducing [1385, 876]. Introduction [1173, 2217, 2230, 630, 747, 1647, 1595]. Intrusion [9]. intrusive [2132]. Invariance [2133]. Invariant [41, 564, 2086, 2090, 1190]. Invariants [707, 1912, 2084, 2024, 1916, 1791, 2087, 491, 2073]. Investigating [1048]. Investigation [1208]. Invisible [1347]. Invited [337]. IQ [291]. IRO [338]. IRO-DB [338]. Iron [1379]. Irregular [76, 1746, 1512, 3, 515, 83, 1942]. irrelevant [1885]. Isabelle [655]. Isabelle/HOL [655]. Island [309]. Iso [1848]. Iso-surfaces [1848]. Isomorphism [755]. Isomorphisms [526]. Isophotes [2049]. Issues [1218, 1991, 121, 145, 1393, 793, 1828]. issuing [1453]. Italian [890, 1300]. Items [1600]. Iterated [1111]. Iteration [1896, 1820]. Iterations [499]. Iterative [1364, 1512, 2248, 2150, 5, 1695]. Iteratively [1609]. ITU [426]. ITU-T.120 [426].

J. [758]. Jalapeño [593]. JAPE [1759]. Java [1760, 2207, 1899, 99, 801, 1761, 1759, 278, 1757, 1758, 1329, 349, 2259, 1302]. Java-Based [1302]. Job [929, 1567]. join [1971]. Joint [618]. Jumps [198]. Junctions [2120]. Just [1290].

KADS [834]. Kahn [1614]. Kalman [1697, 2011, 2012]. KARAT [822]. Karhuenen [105]. Karhuenen-Loeve [105]. Karl [769]. KDD [975]. Kernel [1095, 1128]. Key [2226, 2235, 265, 2039, 2246, 2251, 945, 1445, 1453, 1434, 1451, 1429]. Key-specific [265]. Keys [547, 1442]. KIDS [834]. Kit [1761]. Knapsack [910]. Knitting [2008]. know [1861]. Knowledge [1218, 805, 1707, 933, 808, 827, 1395, 915, 968, 2231, 892, 809, 1705, 810, 1026, 1657, 752, 813, 393, 245, 831, 836, 1645, 1739, 1279, 1714, 815, 1565, 817, 820, 227, 880, 883, 1093, 821, 835, 823, 1724, 824, 1005, 685, 999, 1010, 1413, 1006, 979, 973, 970]. Knowledge-Based [809, 1705, 817, 968]. Knowledgeable [229]. known [1012]. Kohonen [1136, 1142]. Kolmogorov [733, 729, 1374]. KOMET [997]. Kong [2033]. Krausz [1477]. KRK [678].

Krumhansl [758]. KUMP [101]. KUMP/D [101].

L1 [1974, 1474]. l1-Spheres [1474]. Lab [1658]. labeled [742]. Labeling [2189, 460]. Labelled [223, 1189, 463]. Labels [2190]. Laboratory [160]. LABYRINTH [2243]. Lach [749]. lambda [695]. lambda-subsumption [695]. Lambek [1182]. Lamellibranch [556]. Laminar [1038]. Lanczos [91]. Land [1205, 621]. Land-Cover [621]. Land-Use [1205]. Landscapes [308, 235, 1260]. Language [708, 1754, 1210, 922, 1181, 1238, 1684, 2018, 84, 1751, 831, 1185, 382, 212, 383, 1752, 1252, 941, 1798, 200, 2091, 312, 1797, 957, 151, 1234, 106, 2094, 1595, 1415, 732, 1414, 674, 1939]. Languages [379, 639, 190, 193, 1533, 1532, 1534, 893, 1678, 146, 1588, 69, 64, 1193, 1483, 1803, 734, 721, 722, 731, 1929]. Large [149, 1362, 1686, 2165, 951, 1943, 112, 660, 1521, 953, 2048, 1736, 1482, 350, 1724, 2213, 273]. Large-Scale [1362, 1943, 1521, 953]. Large-Strain [112]. Larva [556]. Laser [593, 2055]. LASS [1809]. Latencies [100]. latency [268]. Later [820]. Latest [1388]. Lattice [96, 1840, 232]. Lattices [1532, 1299]. Laundering [797]. Law [792]. Layer [375, 436, 2175]. Layered [1772, 1105]. Layout [1597, 2195, 2180, 1645, 2214, 249, 2093, 1584]. Layouts [841]. Lazy [37, 1786, 1803]. Leaning [744]. Learn [1053]. Learned [1327, 1906]. learners [732]. Learning [745, 872, 874, 1157, 1083, 896, 1733, 568, 1399, 1079, 506, 1138, 734, 829, 1059, 393, 1723, 432, 692, 2015, 1101, 1391, 721, 2021, 1082, 311, 617, 875, 215, 1075, 743, 213, 1638, 1393, 1106, 941, 682, 694, 901, 2133, 335, 725, 1718, 2149, 2121, 2057, 1015, 1113, 1092, 1158, 1058, 1145, 954, 2066, 1146, 1156, 249, 1567, 1087, 377, 2102, 983, 739, 719, 736, 718, 726, 992, 1875, 991, 1014, 746, 981, 735, 681, 1414, 741, 695, 737, 1970, 727, 978, 680, 724].

Learning-by-Showing [2102]. Least [1081, 690, 97, 1989]. Least-Squares [1081]. Leather [601]. Lecture [337]. left [226]. Legacy [1290]. Legal [793, 957, 796, 998]. LEGO [661]. Lenient [1813]. Lesions [585]. less [2129]. Lessons [1906]. Letters [1625, 1622]. Level [136, 1754, 1629, 457, 1746, 2005, 1836, 538, 1810, 609, 591, 1114, 138, 115, 1032, 1411, 517, 1753, 106, 1503, 276, 1291, 1923, 516]. Level-Set [2005]. Levels [872]. Lexical [890, 245, 2204]. Lexicalized [865]. Lexicographically [1485]. Libraries [943, 374, 37, 952, 946, 951, 1758, 954, 942, 1352]. Library [1321, 373, 380, 1631, 966, 958, 964, 1772, 965, 944, 902, 949, 51]. Licence [597]. Lie [564, 701]. Life [1152, 1541, 1288]. LifeCycle [1616]. Lifelong [1391]. Light [1910, 2069, 2070]. Lighting [1849]. Lightweight [786]. Like [1159, 559, 1096]. Likelihood [1121, 595, 2248]. Limitations [750]. Limited [549]. Limited-Color [549]. Limits [1525, 792]. Linda [1325]. Line [603, 1472, 1476, 1209, 609, 455, 1086, 1481, 2266, 1626, 517, 1640, 2131, 1639, 516, 1643, 1999, 1959, 802, 2046, 1961, 1458, 954, 1956, 298, 789]. Linear [1349, 704, 840, 1112, 1222, 448, 1791, 1757, 119, 1122, 1045, 1354, 1848, 1176, 2153, 1355, 1614, 2183, 2087, 220, 90, 97, 2099, 911, 1460, 1064, 2148, 1178, 1905, 740, 1023, 714, 1261, 1883, 1983, 1963]. linearities [496]. Linearly [309]. Lines [1858, 2096, 711]. Linguistic [1186, 916, 894]. Linguistically [1233]. Linguistics [1173, 1194]. Link [2080, 206, 2215]. Linking [349]. Links [830, 1500, 334]. LINUX [1319]. List [52, 1184]. Lists [1563, 1983]. Literals [1716]. Live [2067]. Liveness [156]. Living [304]. LK [1793]. Load [117, 122, 121, 843, 116, 1498, 15]. Load-Balancing [122]. Loads [36]. Local [1461, 896, 764, 1364, 843, 609, 699, 2090, 1137, 2031, 1252, 1153, 1123, 1637, 1192, 491, 987]. localised [284]. Localization [1993]. Localizer [1252]. Locally [2146]. Locate [830]. Location [1394, 1210, 1943]. Locking [1599]. LOCOCODE [1140]. Loeve [105]. Log [2090, 2248, 2095]. Log-likelihood [2248]. Log-Polar [2090]. logarithm [1440]. logarithmic [1967]. Logging [631]. Logic [1716, 919, 1395, 1384, 1221, 182, 186, 751, 2198, 1657, 243, 1240, 570, 234, 3, 1794, 215, 876, 1739, 927, 1176, 1159, 1704, 1565, 392, 1798, 939, 1824, 1191, 1637, 880, 220, 1193, 923, 1194, 1825, 187, 1178, 267, 1900, 971, 972, 1905, 1872, 684, 692, 689, 973, 1918, 268, 674, 675, 1888, 680, 679, 677, 993]. Logic-Algebraic [1395]. Logical [1683, 1631, 1230, 1173, 608, 183, 1669, 1022]. Logics [1715, 223, 899, 1665, 1782, 1184, 1174, 985, 1925]. Logistic [931]. Logistics [1943]. Lognormal [1625]. LogP [119]. LogP-Machines [119]. LogPQ [1515]. Logspace [1538]. Long [1212, 589, 498]. Long-Term [1212]. Look [1575, 1259]. Look-Ahead [1259]. Look-Back [1259]. lookup [273]. lookup-table [273]. Loop [1506, 1518]. Loosely [1321]. Lorenz [2230]. Lossless [535, 539]. LOTOS [672, 1926]. Lottery [798]. Low [2227, 1629, 848, 175, 258]. Low-Overhead [848]. Lower [1597, 1606]. Lowness [1830]. LR [1604]. LTR [620]. LU [111]. Lung [581, 586]. LUT [272, 274]. LUT-based [272].

M [647, 372, 83]. M-ary [372]. M-Net [647]. M-Tree [83]. MAC [2252, 1247]. Macaque [1069]. Machine [829, 1723, 1508, 1496, 213, 1382, 941, 1521, 2057, 954, 1567, 377, 1014]. Machines [1362, 120, 1366, 1098, 169, 1082, 119, 85, 1582, 277, 280]. Magnetic [582, 614]. Magnification [459]. Magnifying [2089]. Magnitude [442]. Mainstream [208]. Maintaining [346, 1662]. Maintenance [1800, 313, 361, 386, 168, 1664, 1954]. Make [1706, 775, 1486, 1971]. Makes [1260]. Making [1217, 148, 1091, 229, 1449, 824, 1959, 1960]. Mammograms [578]. Managed [1516]. Management [387, 16, 1380, 1379, 904, 2002, 1705, 1367, 1324, 375, 150, 4, 1509, 2262, 1616, 838, 420, 152, 1685, 1422]. Manager [651]. Managing [1246, 1283, 351, 1299]. manifold [2109]. Manipulation [2215, 1684, 2143, 2072]. Manipulations [1856]. Manipulative [2126]. Manipulator [1161, 1569]. Manipulators [1169]. Many [1837, 215, 1942]. Many-one [1837]. Many-Sorted [215]. Map [1072, 1805, 2105, 1123, 2054, 1102, 584, 1074, 2248]. mapped [264]. Mapping [673, 324, 525, 2154, 621, 38, 275, 272, 274, 270]. Mappings [812, 135]. Maps [1217, 1068, 453, 1205, 1141, 1136, 1063, 455, 1062, 1718, 1216, 1143, 1620, 1731, 767]. Marine [594]. Marker [2164]. Markets [800]. Markov [457, 1877, 2023, 574, 1078, 1893]. Markov-model [1893]. Markovian [1710, 1571]. Marriage [140]. MASMA [904]. Masses [1998]. Massively [1509, 51]. Matchable [1475]. matched [286]. Matching [483, 2158, 1992, 1236, 1995, 2001, 552, 490, 1475, 2011, 2012, 664, 2021, 2031, 2107, 2103, 524, 555, 517, 520, 2096, 516, 1643]. matchings [1978, 1951]. Materialized [1663, 386, 385]. Mathematical [1056, 651, 1004]. MATLAB [1605]. Matrices [2227, 2249, 1409, 1856]. Matrix [1512, 94, 1773, 43, 92, 39, 38]. Matrix-by-Vector [94]. Max [1257, 1461, 1958, 1579]. Max-Restricted [1257]. Maximal [1843, 1249, 1485, 1951]. Maximum [1467, 1244, 1121, 595, 1978, 1951]. MC [1766]. Mean [1467, 1205, 1082, 1852, 2056]. Meaning [916, 1017, 1019]. Means [495, 468, 1141, 1634, 451]. Measure [1831, 468, 329, 1485]. Measured [583]. Measurement [487, 176]. Measurements [1209]. Measures [457, 531, 466, 918, 996]. Measuring [2124]. MEC [1126]. Mechanical [1381]. Mechanising [219]. Mechanism [429, 388, 362, 363, 412, 1739, 2106]. Mechanisms [804, 2265, 1736]. Mechanizing [1823]. Media [483, 1770, 382, 1553, 2266, 404, 2112, 420, 258]. Medical [1779, 579, 613, 331, 1010]. MeDoc [964]. meet [253]. Meetings [904]. Meets [941]. Meinong [749]. Melanocytic [585]. Melodic [764, 773, 765, 766]. Members [1701]. Membership [848]. Memory [1380, 1654, 1804, 933, 1780, 1351, 34, 1366, 1489, 1373, 908, 857, 851, 102, 1104, 852, 1155, 1159, 1500, 1516, 72, 1519, 1815, 1517, 1514, 6, 1332, 651, 36, 1967, 744]. Mercury [188]. Merge [469, 48]. mergeability [1884]. Mesh [1986, 31, 28, 27, 1968]. Meshes [542, 123, 2177, 50, 25, 1945, 1950]. Message [1362, 648, 1340, 844, 1341, 131, 370, 71, 11, 1329, 1330, 60, 1371, 101, 1347, 1441, 1443]. Message-Passing [844, 1341, 131, 60]. Messages [73, 1808]. Meta [1339, 937, 803, 370, 1381, 663, 1753, 634, 1920, 1312]. Meta-Applications [1339]. Meta-CASE [1312]. Meta-Level [1753]. Meta-Method [663, 1920]. Meta-Protocols [803]. Meta-Rules [634]. Meta-Societies [937]. Metacomputer [1336]. Metadata [373, 948, 1408]. Metadatabase [1310]. Metamodeling [1295]. Metaphor [1199]. Metaprogramming [1603]. Meter [763]. Metering [786]. Method [1986, 2167, 475, 2212, 1350, 583, 532, 1395, 1999, 1523, 607, 760, 811, 1162, 2008, 121, 908, 2160, 2021, 2022, 2090, 2170, 368, 577, 1519, 663, 515, 369, 589, 1938, 926, 2096, 2129, 635, 1077, 1643, 2097, 1006, 1436, 1920]. method-specific [1006]. Methodology [575, 85, 1646, 1724]. Methods [624, 806, 73, 811, 812, 1755, 1648, 658, 2010, 95, 531, 1689, 311, 417, 2263, 1708, 1261, 1262, 328, 1509, 2052, 816, 397, 78, 97, 91, 556, 1359, 2271, 1443]. Metre [764]. Metric [234, 1886]. Metrical [1232]. Micro [2072]. Micro-manipulation [2072]. Microcontroller [1411]. Microdomain [766]. Micropayment [788, 787]. Micropayments [798]. microprocessor [250]. Microprocessors [1780]. Microscopic [1052]. MIDAS [1363]. Might [1525]. Migration [1336, 15]. Mine [1724]. Minimal [873, 1471, 24, 2100, 217, 135, 2250, 880, 1025]. Minimalism [1177]. Minimalistic [1149]. Minimization [2175, 2035, 40]. Minimize [2139]. Minimizing [1412]. Minimum [61, 1468, 1470, 2211, 2203, 2185, 2174, 2265, 510, 1954, 1974, 1980, 1963]. Minimum-Area [2211]. Mining [1681, 1724, 996, 970]. Miracles [1290]. Missing [618, 2056]. Mistrusting [2246]. Misunderstandings [864]. Mixed [1349, 932, 185, 832, 1703, 2067, 250, 1925]. Mixed-Initiative [932]. Mixing [1368]. Mixture [1081, 2043, 2058]. Mixtures [1115, 550, 595]. ML [1934]. MLP [1093]. MLPs [1171]. Mob [1358]. Mobile [1760, 233, 1585, 599, 423, 1807, 424, 172, 1304, 1871, 1982]. Mobility [1616, 172]. Modal [240, 223, 1867, 392, 1022]. Model [1394, 864, 881, 1073, 1069, 343, 142, 129, 339, 964, 1897, 1071, 2220, 764, 167, 552, 1819, 929, 341, 810, 1169, 489, 157, 1147, 514, 327, 1067, 801, 1312, 1283, 1059, 1625, 432, 1677, 1294, 1786, 176, 936, 2135, 2022, 448, 479, 1382, 512, 2037, 1124, 2107, 217, 376, 1668, 548, 526, 1792, 2122, 1159, 2045, 2163, 72, 620, 574, 757, 407, 503, 1815, 1044, 1292, 2065, 949, 473, 520, 1065, 655, 1306, 1515, 2145, 1332, 1034, 350, 1803, 309, 1064, 2094, 2093, 661, 471, 1685, 1869, 1904, 988, 973, 978, 995, 1893, 1025, 1708, 1068]. Model-Based [881, 552, 2037, 2107, 2045, 620, 1025]. MODEL-ECS [1708]. Modeler [1276]. Modeling [1055, 225, 73, 1296, 2147, 1633, 1148, 1499, 1391, 2136, 1841, 1122, 1045, 1252, 509, 2163, 615, 1305, 860, 1387, 488, 819, 1050, 520, 1103, 128, 1150, 1234, 51, 437, 2166, 651, 1595, 2102, 2078, 2123, 1908, 1968, 1923]. Modelizing [910]. Modelling [612, 826, 825, 1410, 340, 1225, 1725, 920, 319, 810, 303, 162, 348, 347, 1383, 1917, 961, 1704, 384, 349, 820, 632, 1248, 366, 508, 1019]. Models [1986, 59, 1362, 638, 964, 892, 1282, 1293, 580, 1052, 1182, 1370, 761, 1030, 2032, 1042, 383, 852, 1032, 1381, 2042, 2043, 522, 1291, 613, 399, 2169, 335, 2121, 882, 2089, 2058, 49, 1278, 1093, 2139, 1274, 78, 2069, 1649, 2141, 1954]. Modes [1797]. Modification [1054]. Modifications [359]. Modula [201]. Modula-Programs [201]. Modular [1904, 2237, 666]. Modulation [2238]. Module [1808]. modules [275]. Molecular [1361, 1366, 1314]. Molecules [1771]. Moment [2036]. Moments [1984]. Monadic [84]. Money [781, 791, 780, 790, 797, 1959]. money-making [1959]. Monitoring [502, 1339, 1344, 1345, 534, 257]. Monocular [2065, 518, 400]. Monogram [2007]. Monotone [728]. Monotonic [1613, 1872, 2172]. monotonicity [1913]. Montage [2033]. Monte [687]. MOPS [1370]. Morphic [1535]. Morphing [2108, 261]. Morphisms [1714]. Morphological [507, 2170, 284]. Morphology [587]. Morphology-Based [587]. Mortem [631]. Mosaick [553]. Motion [2155, 502, 495, 2116, 2158, 2132, 715, 2160, 2023, 618, 500, 496, 2153, 709, 2157, 497, 521, 400, 2127, 2129, 498, 2162, 614, 590, 2130, 2075, 2148, 516, 2077, 2151, 2128]. Mountain [1290]. mouse [1425]. Mouth [509]. Move [1066]. Movement [1394, 1719, 2164]. Moving [1386, 2170, 2159]. MPEG [412, 415, 400, 536, 543, 534]. MPEG-2 [415]. MPEG-4 [400]. MPI [1321, 1348, 1360, 1322, 1364, 1324, 1313, 1335, 1314, 1329, 1315, 1316, 81, 1749, 15]. MPI-2 [1749]. MPI-Based [1321]. Mpixel [288]. Mpixel/sec [288]. MPPs [1322, 1316]. MPSK [2238]. MRF [2026, 2107, 473]. MRI [331]. mTunnel [407]. Multi [452, 149, 919, 1629, 1814, 825, 2229, 1503, 703, 922, 447, 1742, 921, 1949, 1215, 1335, 857, 859, 2175, 1393, 1392, 2247, 1291, 528, 1703, 2089, 606, 883, 238, 543, 352, 1788, 1333, 516, 2140, 1899, 1907, 1945, 1936, 993]. Multi-access [2229]. Multi-Agent [919, 825, 922, 921, 1393, 1392, 528, 883, 1936]. Multi-chromosomes [1703]. multi-class [993]. Multi-classifier [447]. Multi-color [1949]. Multi-Dimensional [703, 1945]. Multi-expert [606]. Multi-flip [238]. Multi-grid [2089]. Multi-Layer [2175]. Multi-level [1503, 1291, 516]. Multi-Node [1335]. Multi-objects [859]. Multi-output [1814]. Multi-Part [2140]. Multi-Physics [1333]. Multi-point [1788]. Multi-processor [857]. Multi-program [543]. Multi-Purpose [1215]. Multi-resolution [1629]. Multi-sensor [452]. Multi-Site [149]. multi-tasking [1907]. multi-threaded [1899]. Multi-tier [1742]. Multi-valued [352]. Multicast [847, 2269, 440, 407, 408]. Multicasting [405, 2268]. Multichannel [28]. Multicomputer [368]. Multicontext [250]. Multidimensional [381, 697, 368, 114, 555]. Multigraphs [1479]. Multilayer [1151, 1114, 1106, 1113, 556]. Multilevel [359, 1407, 1396, 123, 1491]. Multilinear [698]. Multilingual [955, 956, 957]. Multimedia [419, 427, 142, 431, 338, 434, 435, 436, 325, 2258, 363, 2263, 425, 438, 433, 2262, 620, 364, 420, 437, 967, 402]. Multiobjective [316]. Multipart [2137]. Multiple [1267, 2000, 31, 1734, 1294, 2136, 138, 2245, 13, 1400, 2088, 2113, 790, 525, 706, 2070, 2162, 2081, 2141, 275, 1418]. Multiple-Payment [790]. Multiple-view [706]. Multiplexer [2229]. multiplication [1435]. multiplication-addition [1435]. Multiplications [94]. multipliers [255]. Multiprocessor [1357, 1785]. Multiprocessors [102, 103, 35, 132, 1519, 6]. Multiresolution [563, 588, 1844, 2056]. Multiscale [486, 1998, 2085, 442, 2121, 444]. Multisender [1424]. multisignature [1432]. Multispectral [443]. Multistage [129, 559, 2151]. Multithreaded [1495, 1496, 1757, 135, 1955]. Multithreading [132]. Multitolerant [1827]. Multiuser [2249]. Multivalent [959]. multivariate [1446]. Multiversion [360]. Multiway [2200]. Munsell [2013]. Muscle [2164, 2169]. Muscle-Based [2169]. Mushroom [1735]. Music [752, 761, 750, 753, 992, 402]. Musical [764, 771, 752, 765]. Musicology [747, 748]. Must [1066]. Mutation [307]. Mutual [1696]. Mutually [2246].

Nadaraya [727]. Nadaraya-Watson [727]. Naive [1846]. Name [1228]. NAO [1770]. Nara [1507]. narrative [1019]. Narrowing [1235]. Nasality [768]. National [313]. Nationale [965]. Native [430, 431, 1758]. Natural [477, 1071, 1181, 1874, 2028, 941, 312, 2060, 1193, 1070, 674]. Naturals [911]. Nature [747, 1062]. Navigation [1199, 2206, 1219, 1155, 540, 1513, 868, 2055, 2061, 2072, 1156, 2204]. Navigational [912, 333]. NBV [2102]. NC [55, 1817]. Near [411]. Near-VOD [411]. Nearest [381]. nearly [742]. Neat [1290]. NEC [1335]. need [1861]. Needles [1692]. Needs [150]. Negation [916, 894, 1023]. Negative [366, 1011, 995]. Negotiation [921, 883].

Neighborhood [1272]. Neighbour [381]. Neighbourhood [1850, 315, 467]. Neighbourhood-Based [467]. Neighbourhood-Search [315]. NeMo [1274]. Neocortical [1059]. Nephelometric [1737]. Nepi [1810]. Nerve [1056]. Nervous [1389]. Nested [376]. Net [1179, 1838, 1103, 647]. Netherlands [963]. NETL [1513]. Nets [1733, 1410, 1836, 1837, 1543, 154, 1544, 1547, 192, 2040, 1189, 1545, 132, 1390, 756, 1915]. Nets-Based [2040]. Network [605, 1228, 1151, 1071, 930, 1163, 2080, 960, 1365, 804, 1060, 1367, 1126, 1133, 1051, 1810, 761, 2164, 2020, 1104, 416, 1166, 1125, 1698, 560, 326, 716, 1058, 1145, 133, 1107, 1524, 1108, 1105, 2268, 1422, 1892, 1005]. Network-Balancing [1365]. Networked [403]. Networking [776]. Networks [887, 429, 717, 1597, 854, 1465, 483, 129, 1039, 1130, 2180, 1076, 1523, 557, 1472, 405, 319, 1353, 118, 1088, 863, 1162, 1095, 570, 558, 1369, 414, 1094, 593, 1100, 1722, 1048, 235, 2015, 1101, 1096, 855, 29, 1099, 1049, 931, 842, 1501, 660, 1045, 18, 4, 1153, 1614, 1152, 415, 1580, 882, 2054, 1113, 586, 1047, 1041, 1091, 1092, 1160, 1164, 1482, 1097, 1571, 569, 238, 1569, 1090, 543, 556, 2104, 1584, 974, 1916, 1982]. Neural [605, 717, 483, 1733, 930, 1076, 1163, 2080, 1523, 557, 1161, 1088, 1147, 751, 863, 1095, 558, 593, 1722, 1135, 1048, 2015, 761, 1101, 2164, 2020, 1099, 1390, 1152, 1118, 1580, 1166, 1125, 1698, 756, 560, 1113, 586, 716, 1047, 1041, 1160, 1058, 1145, 1571, 1569, 1090, 1524, 1105, 1074, 974, 741, 1639]. Neural-HMM [1639]. Neurocontrol [1172]. Neurocontrollers [1167]. Neuromorphic [1165]. Neuronal [1052, 1042]. Neurones [1049]. Neurons [1039, 1046, 1060, 1051, 1045, 1092, 1036, 1057]. Neurotrophins [1057]. Neutron [1372]. NewDES [1434]. News [421, 2167]. Newton [1359]. NeXeme [75]. Next [1327, 151, 2102]. Next-Generation [151]. Nexus [75]. NFC [1161]. Nice [1384]. NicheWorks [2213]. NLP [830]. NMR [676]. No [2227, 2192]. Nobel [593]. Noble [1290]. NOBLESSE [620]. Node [1335, 1790]. Noise [678, 1109, 2084, 1042, 1050, 732, 692]. Noise-Robust [1050]. noise-tolerant [732]. noising [2083]. Noisy [2025, 739]. Non [1461, 2132, 1068, 1079, 2147, 1364, 504, 2198, 1133, 1812, 1508, 2109, 654, 2242, 496, 714, 1261, 2172, 200, 569, 9, 702, 1178, 1872, 1024]. Non-accidentalness [702]. Non-cartesian [569]. Non-coherent [2242]. non-commutative [1178]. Non-determinism [654]. Non-deterministic [9]. non-Horn [1024]. Non-intrusive [2132]. Non-linear [714, 1261]. Non-linearities [496]. Non-Local [1364]. Non-manifold [2109]. Non-monotonic [2172, 1872]. Non-Oblivious [1461]. Non-oriented [1068]. Non-rigid [2198]. Non-standard [1812]. Non-stationary [1133]. Non-strict [1508, 200]. Non-uniform [2147]. Non-visible [504]. Nonclassical [1072]. Noncombining [26]. Nondeterminism [1549, 68]. Nonlinear [305, 1999, 1172, 1104, 620, 1121, 1571, 1359, 1426]. Nonmonotonic [222, 1713, 1714, 1687, 1675]. Nonoverlapping [94]. Nonplanar [2186]. Nonstationary [1083]. Nonzero [1488]. Norm [2028]. Normal [671, 444, 1023]. Normalization [1999, 2053]. Nose [1985]. Notation [641]. Notations [658, 1908]. Note [2254, 740]. Notification [163]. Notion [1711, 819]. Nova [8]. Novel [2158, 122, 2236, 2019, 2245, 549, 606, 411, 539, 1108, 250]. Novelty [1051]. NP [1468, 2185, 1602, 1941]. NP-Completeness [1468]. NP-hard [2185, 1941]. NPB [128]. NRDR [1707]. NSPACE [1539]. NTSC [287]. Nuanced [894]. Number [1462, 1469, 2203, 2185]. Numbers [196, 2251, 1736, 2177]. Numeral [602, 1637]. Numerical [1744, 861, 1762, 88, 1351, 1352, 93, 980, 982]. Nurture [1062].

O-Unitaire [1763]. Object

[379, 1984, 943, 1321, 484, 588, 1771, 1744, 903, 342, 899, 1671, 1754, 1764, 1410, 636, 340, 564, 1296, 1772, 439, 1592, 499, 347, 327, 1490, 344, 481, 1676, 1283, 109, 1774, 2011, 2012, 1677, 1294, 1851, 1284, 936, 1751, 2135, 104, 383, 1852, 2170, 2107, 1279, 2103, 1775, 1303, 1752, 628, 1678, 2043, 643, 1295, 1588, 1807, 72, 2133, 1518, 1593, 335, 349, 2121, 64, 566, 1591, 1280, 369, 627, 2124, 1278, 1288, 2062, 2214, 527, 2065, 518, 1274, 74, 565, 945, 642, 2110, 1672, 1669, 1299, 1745, 353, 2102, 2078, 2123, 1908]. Object-Based [1772]. Object-Oriented [379, 1771, 1744, 903, 342, 1754, 1764, 636, 1296, 439, 1592, 327, 1676, 109, 1774, 1677, 1294, 1751, 383, 1279, 1775, 1303, 1752, 628, 1678, 643, 1295, 1588, 72, 1518, 1593, 349, 64, 1591, 1280, 627, 1278, 2214, 1274, 1745, 353]. Object-Z [642]. Objectionable [401]. Objective [305, 1624, 1074]. Objects [1986, 2115, 445, 73, 670, 955, 752, 1821, 1222, 1294, 1857, 1769, 2021, 2137, 363, 2109, 1668, 2142, 378, 2088, 351, 515, 70, 229, 1202, 2099, 1211, 352, 1856, 1302, 2140, 2141, 1905, 859]. Obligation [656]. Oblivious [1461]. observable [1893]. Observation [1641, 163, 1211]. Observations [1549]. Observer [709]. Observing [1544]. Obstacle [1148, 1149, 2063, 2144]. Obstacles [2146]. Obtain [1670]. Obtained [758]. Occam [723]. Occluding [1987, 2137]. Occlusion [489]. occurrence [956]. OCR [962, 2038]. Octaves [773]. Ocular [1065, 682]. Ocular-Dominance [1065]. ODB [899]. ODB-Tools [899]. ODEs [1357]. Off [1999, 603, 1476, 1626, 1640, 1639, 789, 1643, 1959, 1961, 1458, 1068, 1065]. OFF-center [1068]. Off-Line [603, 1476, 1626, 1640, 1639, 1643, 1999, 789, 1959, 1961, 1458]. Offering [1367]. Oja [1138]. OLAP [1661, 1684]. OLCCR [2097]. OMIS [1344]. Omnidirectional [1996]. On-Board [885]. ON-center [1068]. On-Chip [102, 250]. On-Line [1472, 1476, 609, 1086, 2266, 1959, 802, 2046, 954, 789, 298]. ON-OFF [1065]. Onboard [821]. One [1055, 1075, 2070, 1972, 1837, 734, 746, 1446]. one-dimensional [1972]. One-Shot [1075]. one-variable [734]. one-way [1446]. Online [1083, 2206, 1553, 494]. Only [2173, 695]. onto [275]. Ontologies [812, 815]. Ontology [807, 1202, 1277]. OO [1594]. OODB [1686]. OODBMS [343, 1513]. Open [1406, 376, 2039, 1832, 1422]. Operating [413, 1337, 1862, 280]. Operation [1613, 825]. Operational [1253, 2230, 319, 1899]. Operations [22, 670, 646, 1806, 53]. Operator [673, 1697, 2258, 2027, 1873, 1874, 1892]. operator-based [1873, 1892]. Operators [493, 699, 815, 1885]. Opportunities [441]. Optic [1989]. Optical [1084, 1697, 611, 28, 2153, 604, 622, 1619, 411, 2156]. Optics [2022]. Optimal [2082, 1701, 1474, 1112, 1135, 857, 1681, 46, 1607, 1608, 547, 1854, 1612, 1957, 2183, 90, 26, 1487, 1441, 1973, 1413]. Optimality [32, 1116, 464, 183]. Optimisation [318, 37, 1144, 274, 276]. Optimisers [315]. Optimising [2119]. Optimistic [1668]. Optimization [379, 305, 899, 1265, 1365, 662, 125, 1688, 1250, 1765, 8, 713, 77, 115, 1579, 1753, 1119, 1405, 556]. Optimize [1080]. Optimized [1488]. Optimizing [1686, 45, 1501, 767]. Optimum [1397, 1265, 113]. OQL [1663]. Oracles [735, 1412]. oral [969]. Order [887, 1466, 919, 1254, 1544, 2090, 862, 745, 1863, 974, 1445, 1884, 200]. Ordered [384, 372, 737]. Ordering [2155, 1471, 1475, 1272, 1256]. Orderings [199, 1799]. Orders [1820]. Organising [1733]. Organization [751, 1075, 765, 1216]. Organizations [387]. Organizing [1141, 1133, 1123, 1143, 1620, 767, 1102, 1805, 1983, 756]. Orientation [1984, 1073, 1069, 1068, 1072, 227]. Orientations [1465]. Oriented [379, 1321, 1301, 1771, 1744, 903, 342, 899, 1754, 1068, 1764, 1410, 636, 1296, 439, 1592, 162, 327, 344, 1676, 109, 1774, 1677, 1294, 1773, 1751, 383, 1279, 1775, 1303, 1752, 628, 1678, 643, 1295, 1787, 1667, 1588, 72, 1518, 1593, 349, 64, 1203, 1591, 1280, 369, 627, 1278, 2214, 1274, 1745, 353, 976, 967, 1908, 1882, 1969, 1937, 1013, 251]. oriented-modeling [1908]. Origin [747, 756]. Orthogonal [2181, 2212, 2202, 2229, 2203, 2187, 2186, 2184, 2029, 2179, 1142, 2183]. Orthographic [2109]. orthonormal [727]. Oscillation [1272]. Oscillations [1087]. Oscillators [1104, 257]. Oscillatory [1805]. Out-of-Core [77]. Outer [2176]. Outer-Planar [2176]. Outliers [489]. Output [419, 1549, 1604, 1814]. Output-Store [1604]. Outputs [1112]. Overhead [1603, 848]. Overlapping [1506]. Overriding [1671]. Overture [1764]. Overview [1688, 2047]. Owner [48]. Owner-Write [48].

P [1602]. P3L [80]. P4 [294]. PAC

[718, 726, 727]. PACCO [542]. PAckage [2207]. Packages [384]. Packet [543]. packing [271]. PADAM [32]. Pair [181, 31, 606]. Pair-Sharing [181]. Pair-Wise [606]. Pairs [2223]. Pairwise [453, 2192]. PAL [287]. PALLAS [1343]. Pamela [1932]. Pan [494, 2152]. Pan/Tilt [494]. Paper [900]. Paradigm [2115, 1589, 2195, 280, 1888]. Parallel [1700, 1321, 1349, 1361, 1744, 37, 1804, 1754, 110, 1363, 1762, 1340, 88, 1763, 1746, 2000, 1351, 1364, 1473, 1772, 14, 118, 1, 105, 96, 1373, 52, 1755, 131, 1369, 1776, 1370, 109, 1774, 111, 2, 1765, 1761, 1759, 84, 8, 1769, 1751, 112, 1605, 92, 1049, 1168, 1343, 1577, 48, 67, 1775, 1345, 115, 113, 47, 1756, 1752, 13, 1241, 1375, 1521, 1513, 114, 1355, 1509, 85, 1768, 1522, 134, 1356, 312, 1824, 1815, 89, 1371, 262, 93, 511, 1318, 90, 56, 81, 1514, 1358, 98, 1374, 91, 1331, 1515, 1524, 51, 1567, 86, 83]. Parallel [1359, 1964, 1962, 281, 734, 279, 1352]. Parallel-Sequential [1049]. Parallelisation [87]. Parallelism [136, 80, 1367, 1531]. Parallelizable [2154]. Parallelization [1511, 1360, 1503, 1523, 3, 95, 35, 1518, 33, 1372, 1507, 1510, 1485]. Parallelizers [45]. Parallelizing [1504, 1588, 238, 1620, 1505]. Paramenters [2100]. Parameter [2164, 1127, 618, 505, 595, 2140, 1877]. Parameter-Free [505]. parameterised [266]. Parameterized [207, 1916]. Parameters [2231, 400, 2071]. Parametric [567, 1923, 1930]. Parametrization [1398]. parent [1983]. Parity [2227, 1977]. Parla [922]. Parlists [79]. Parsing [1604, 867, 823, 232]. Part [1674, 228, 1562, 2140, 1014]. Part-of-Speech-Tagging [228]. Part-Whole [1562]. Partial [887, 582, 199, 1764, 1706, 723, 264, 1251, 2021, 209, 1174, 1544, 1757, 1695, 1110, 1863, 1900, 1918, 1451, 1884]. partial-order [1884]. Partiality [219]. Partially [854, 2063, 372, 265, 1893]. partially-observable [1893]. Participative [1215]. Particle [1511]. Partition [1220]. Partitioning [342, 671, 1358, 277, 273]. Partly [62]. Partly-Consistent [62]. Parts [2022, 1563]. Pascal [1505]. passage [730]. Passing [1362, 648, 1340, 844, 1341, 131, 855, 11, 1329, 1330, 60, 1347]. Password [2039]. Past [797]. Path [1597, 1257, 1570, 165, 1162, 1799, 30, 714, 1481, 1498, 1569, 1584]. Paths [164, 1474, 1973]. Patient [327]. PATON [1159]. Pattern [1717, 1129, 1132, 664, 1075, 1775, 2142, 441, 1041, 170, 2064, 705, 1040, 1108, 463, 734]. Pattern-Based [1775, 170]. Patterns [1254, 1342, 760, 428, 148, 1591, 1280, 1317]. Paxos [845]. Payment [790, 1459]. Payments [802]. PCG [1350]. PCI [289, 254]. PCI-based [289]. PDB [1522]. PDEs [1762]. Pearl [615, 1275]. Pedestrian [590]. Peephole [662]. Peg [1965]. Peg-solitaire [1965]. PEI [66]. Pen [617, 2047]. PENGUIN [406]. People [1205, 2063]. Perceptible [753]. Perception [1147, 698, 496, 712, 1214, 716, 1572, 1640]. Perceptron [1084, 1114, 1122, 1086]. Perceptron-Based [1122]. Perceptual [564, 751, 765, 757, 756, 454]. Perceptually [2081]. Perfect [718, 1416, 1951]. Performance [484, 1493, 380, 429, 1744, 2225, 126, 129, 381, 117, 1492, 1512, 14, 1497, 139, 773, 1490, 107, 131, 2269, 531, 1499, 310, 836, 1314, 1120, 160, 137, 1354, 1767, 132, 1520, 134, 1315, 1316, 33, 254, 1317, 5, 1318, 1514, 1319, 6, 128, 2256, 1745, 1902, 282, 990, 279]. Performance-Enhancing [1744]. Performances [2079, 677]. Performed [1163]. Performing [844]. Periodic [2223]. Periodicities [1805]. Permissive [1900]. Permutation [707]. Permutations [1409, 1446]. Person [591, 2152]. Personal [904, 2014, 425, 1319, 300]. Personalised [421, 1726]. Personalized [775]. Perspective [490, 1185, 782, 1295, 2153, 1152, 792, 2148, 1909]. Perspectives [2222, 3, 785, 1626]. Petri [1410, 1836, 1837, 1543, 1544, 1547, 1545, 132]. Phase [2212, 1134, 1697, 319, 487, 521]. Phenomena [771]. Phenomenon [768]. Philology [958]. Phonological [1179]. Photo [965]. Photographic [362].

Photometric [2006]. Photorealistic [2123]. Physical [1370, 102]. Physics [1766, 1521, 615, 1333, 978]. Physics-Based [615]. piano [677]. PICGAL [650]. PICS [948]. Pictorial [1212, 545]. Picture [1242, 429, 2116, 1590]. Piecewise [1122]. Pilot [1336]. PINEAPL [1352]. Pinwheels [1066]. Pipeline [1499, 261]. Pipelined [1780, 255]. Pipelines [105, 261]. Pitfalls [1264, 2191]. Pixel [1630, 104]. Place [1156, 1915]. place/transition [1915]. Placement [1228, 37, 46, 411, 271]. Plan [1713, 1891, 1865, 1885, 1887]. Plan-refinement [1891]. Planar [1462, 2176, 1468, 2021, 2024, 2193, 2146, 1964, 1956]. Plane [2034, 2183, 1846]. Planes [1858]. Planetary [326]. Planned [313]. Planner [821, 1861, 1884, 1883]. Planning [1862, 887, 933, 1385, 241, 1869, 1000, 932, 1570, 637, 1162, 1875, 214, 1871, 1878, 2107, 1404, 883, 1569, 239, 2102, 1859, 1863, 1866, 1868, 1872, 1873, 1874, 1880, 1881, 1870, 1889, 1892, 1893]. Plans [1683, 897, 866, 889, 1888, 1860, 1864, 1867, 1882]. Plant [1379, 592, 1862]. Plasma [1511]. Plastic [1043]. Plasticity [1055, 1071, 307]. Plate [597]. platform [1421, 294, 260, 259, 985]. Platforms [1360, 1359, 292]. Plausible [1067]. Play [1551]. Playback [440]. Player [412]. Players [553]. Playing [1945]. PNtalk [1410]. Point [2176, 2007, 1641, 1614, 505, 1850, 926, 1666, 2129, 519, 282, 1910, 1788]. Point-Based [1666]. Point-Neighbourhood [1850]. Pointer [3, 179]. Pointer-Based [3]. Pointers [42]. Points [34, 2049, 711]. Polar [2090]. Polder [1336]. Pole [1167]. Pole-Balancing [1167]. Policies [117, 1495, 791, 1292, 1022]. Policing [429]. Policy [34, 1430]. Polling [788]. Pollution [1362]. Polycrystal [112]. Polygon [1848, 510]. Polyhedra [1111]. Polyhedral [2175, 2142, 2076]. Polymorphic [1182]. Polynomial [199, 1837, 195, 1098, 731, 1396, 680, 1903]. Polynomial-time [1837, 680]. Polynomials [2024, 2029, 1446]. POMDP [1893]. Pomset [1193]. Pool [1052]. POOMA [1765]. Populations [1282]. Porous [1770]. Portable [110, 1772, 14, 1352, 1329, 87, 1517]. Pose [499, 2045, 2150, 2161, 2054, 2065, 2140]. Position [233]. Positioning [518]. positive [695, 731, 694, 995]. positive-only [695]. Possibilistic [1870]. Possibilities [750]. Possibility [391]. Possible [940]. Post [631, 761, 1054]. Post-Mortem [631]. Post-Synaptic [1054]. Post-Tonal [761]. Postal [616]. Posteriori [1115]. Postscript [293]. Posture [2170, 2037]. Potential [2147, 137, 1037]. Potential-Based [2147]. Potentials [1054]. Power [189, 1238, 1624, 1679, 859, 1533, 1822, 1580, 1829, 1413]. Powerful [1297]. powerline [1448]. Powerlists [79]. powersets [999]. PQ [2191]. PQ-Trees [2191]. Practical [1706, 650, 1943, 1378, 210, 1808, 498, 1941]. Practice [1586, 206, 963]. Practices [1306]. Pragmatic [1219, 1187]. PRAM [1606, 48, 53, 1963]. Pre [587, 1054, 539, 537, 2152]. Pre- [1054]. Pre-press [539]. Pre-processing [587, 537]. Pre-recorded [2152]. preanalysis [981]. Precedence [849, 852]. Precedence-Based [852]. Precise [2031, 2096]. Precision [1998, 1655, 179]. Precompilation [526]. Preconditioning [96]. Predicate [1599]. Predicates [850]. Predicting [1317, 742]. Prediction [126, 21, 501, 931, 138, 137, 2040, 714, 1704, 1491, 742, 677]. Prediction/Verification [2040]. Predictions [1725, 159, 679]. Preference [1072]. Preferences [317]. Preferential [1709]. Prefetching [132, 1491]. Preliminary [600, 1735, 584]. Premises [422]. Prepartitioning [121]. Preprocessing [624]. Preprocessor [1761]. Prescribed [505, 1974]. Presence [2239].

Present [1494, 797, 733]. Presentation [436, 1020]. Presentations [2258, 945]. Preservation [1919, 1925]. Preserving [1839, 2035, 2083, 511]. President [1276]. press [539]. Presupposition [1187]. Pretty [1442]. Prevailing [757]. Prevention [1465, 1458]. Price [2259]. Price-Based [2259]. Pricing [301]. prime [1445]. Primitives [2072, 51]. Principal [443, 1126, 165, 1118, 770, 1128, 987]. Principle [1142]. Principles [1293, 1088, 769, 1651]. Printable [1539]. Printed [605, 1623]. Prior [1645]. priori [268]. prioritized [971]. Priority [52, 412, 53, 364]. Priority-Driven [364]. Privacy [778, 1442]. Proactive [244]. Probabilistic [1095, 962, 788, 1139, 1865, 1966]. Probabilities [1115]. Probability [828, 738]. probing [250]. Problem [1700, 866, 903, 1597, 806, 929, 2194, 910, 811, 812, 1251, 2175, 1708, 40, 1578, 927, 2049, 816, 1703, 1740, 818, 168, 315, 1618, 1483, 823, 1567, 1260, 1868, 1943, 1981, 1881, 1946, 1942]. Problem-Solving [866, 806, 811, 812, 816]. Problems [305, 1349, 1263, 108, 1083, 76, 1236, 1246, 1079, 1265, 1512, 1536, 910, 2198, 1249, 1250, 1693, 1244, 759, 1259, 1817, 314, 1579, 886, 451, 56, 1248, 97, 1405, 1485, 1830, 309, 1460, 1583, 1440, 1962, 1977, 1872, 1941, 977, 1886, 993, 95]. Procedural [1185]. Procedure [2019, 180, 1825, 1862, 1869]. Process [1324, 150, 1590, 660, 361, 69, 153, 15, 151, 1901, 1902, 1929]. Process-Integrated [153]. Processes [2219, 1559, 67, 395, 1078, 2273, 1877]. Processing [456, 1493, 1629, 493, 703, 340, 1598, 1772, 611, 105, 616, 1048, 1747, 325, 1605, 2027, 2028, 1120, 1049, 1806, 900, 1125, 227, 1401, 1318, 544, 465, 98, 594, 559, 2112, 1102, 1302, 390, 587, 258, 674, 300, 1939, 537]. Processor [630, 29, 1504, 1375, 38, 282, 857, 287, 101]. Processors [16, 844, 2000, 1495, 138, 537, 285]. Produced [1099, 1442]. Produces [1732]. Producing [752]. Product [807, 417, 1619]. Production [1218, 317]. Professional [2272]. Professor [792]. Profiles [491, 913]. Profiling [168]. Profilometry [487]. Profitability [466]. Progol [683]. Program [718, 1341, 405, 1002, 185, 125, 1508, 43, 160, 134, 168, 17, 684, 543]. Programmable [639, 250, 1429]. Programme [836]. Programmers [2272]. Programming [1321, 1348, 1349, 639, 1771, 1556, 1746, 203, 186, 1742, 234, 3, 2, 1810, 190, 1773, 1587, 1343, 876, 67, 1775, 893, 1752, 1241, 1176, 1565, 1588, 72, 2172, 939, 64, 312, 1191, 1273, 1274, 78, 1194, 1331, 1510, 1234, 1332, 1460, 86, 516, 1643, 1976, 674, 675, 1883, 680, 679, 677, 993]. Programs [42, 1804, 673, 182, 1342, 201, 197, 1794, 8, 39, 35, 1795, 45, 135, 643, 85, 200, 1824, 1778, 1888, 1507, 923, 133, 9, 1484, 1347, 184, 1825, 187, 17, 36, 971, 1872, 1907, 692, 689, 1918, 688]. Progression [758]. Progressions [758]. Project [149, 624, 1263, 958, 1763, 1352, 1751, 623, 1673, 832, 626, 1404, 620, 619]. Projection [1271, 2142, 2153, 2127, 614]. Projection-Reconstruction [614]. Projections [607]. Projective [707]. Projects [174]. Prolog [891, 1496, 1795]. PROMELA [658]. Promising [1668]. Proof [656, 1179, 1833, 665, 2194, 905, 214, 658, 1793, 1174, 1382, 879, 1189, 1825, 1834, 691]. Proof-Replay [905]. Proofs [241, 1181, 906, 1802, 1402, 220, 1413, 1935, 1927]. Propagating [2257]. Propagation [1253, 1263, 1631, 1246, 2005, 1141, 109, 134]. Properties [1847, 1782, 1384, 156, 633, 1072, 1712, 196, 1838, 1795, 894, 2124, 1902]. Property [1297, 2050, 942, 634]. Proportional [762]. Proposal [1425]. Propositional [1706, 879]. Prosody [2270].

Protection [363]. Protention [773]. Protocol [857, 659, 60, 372, 294]. Protocols [657, 2226, 803, 1453, 658, 1516, 1778, 1413, 1911, 1881, 1454]. Prototypal [898]. Prototype [333]. Prototypes [1638]. Prototypical [902]. Prototyping [1931, 1708, 299]. provable [1913, 1436]. Prove [633, 668]. Prover [215, 221]. Provers [907]. Providing [155, 1726]. Proving [653, 1716, 2198, 218, 200]. Proximity [467]. Proxy [1433]. Pruning [1272, 742]. Pseudo [1464, 470, 1956]. Pseudo-Binary [470]. Pseudo-Hamiltonian [1464]. pseudo-line [1956]. Pseudorandomness [1831]. Psychological [762, 758]. Psychoneural [755]. Public [1300, 2235, 2251, 1445, 1453]. Public-Key [2235, 1453]. Publications [963]. Publicly [1451]. Pulse [1048]. Pulse-Processing [1048]. Pulses [772]. Pump [316]. Purkinje [1053]. Purpose [826, 2197, 1215, 1509, 1815]. Purposes [915, 1209]. Pursuit [1699]. Pursuits [1290]. Pushdown [2219]. PVaniM [1339]. PVM [16, 1348, 1349, 1350, 1363, 1342, 1323, 1367, 1334, 1325, 1373, 1326, 1327, 1369, 1313, 1328, 1314, 11, 1357, 1317, 1337, 1319, 1346, 1320, 86]. PVM-Implementation [1350]. PVM-RPC [11]. PVM/MPI [1348]. PVMPI [1324]. PVS [1932, 669, 668]. Pyramid [452, 591]. Pyramid-Based [452]. Pyramidal [1059]. Pyramids [1629].

QAM [2241]. QDDs [189]. QoS [438, 2259]. QSM [1815]. Quadratic [910]. quadrilaterals [1968]. Qualification [893]. Qualitative [1230, 1210, 211, 2002, 393, 1560, 1561, 1197, 1004, 1392, 894, 227, 926, 1741]. Qualitatively [715]. Qualities [749]. Quality [1228, 434, 531, 2265, 615, 407, 2083, 230, 74, 2272]. Quality-of-Service [407]. Quantification [1866]. Quantified [223, 1716, 878]. Quantifiers [1829, 693].

Quantitative

[1267, 513, 112, 1191, 1004, 679].

Quantization [1134, 1135, 496]. Quantizer [1397]. Quantum [1361]. Quasi [2245, 1359]. Quasi-Newton [1359]. Quasi-synchronous [2245]. Queries [356, 1286, 1662, 1686, 1682, 1207, 333, 734, 746, 1970]. Query [379, 1683, 899, 1691, 381, 389, 1598, 1688, 367, 382, 2030, 383, 1680, 949, 2204, 390, 1233, 985, 984]. query-answering [985]. Querying [530, 1636, 1021, 1012, 973]. Quest [1345]. Question [228, 1924]. Queue [52, 53]. Quiescent [846].

R [804]. r.e [721, 1535]. R/3 [804]. RaAP [285]. Race [10]. Radar [1772, 2055, 473]. RADD [1301]. Radial [1164]. Radiance [1373, 1371]. Radio [2232, 109, 2236, 423]. Radiographic [1227]. Radiographs [573]. Radiological [579]. radix [285]. Railroad [640]. Ramification [893, 927]. Random [1242, 1653, 457, 2240, 2023, 1480, 574, 237, 330, 1412, 692]. Randomization [1778, 1443]. Randomized [839, 1963]. Range [476, 478, 1734, 464, 2118, 2138, 1314, 2142, 488, 2062, 2144]. Rangefinding [562]. Rank [2155, 1058]. Rank-Based [1058]. RAP [1594]. Rapid [843, 1312, 1708, 2144]. Raster [466]. Rat [1055]. Rate [410, 2241, 2242]. Ratio [2248, 2200]. Rational [196]. Rationality [303]. Rats [1043]. Raw [2030]. Ray [583, 1857, 2095]. Ray-tracing [1857]. Rayleigh [2241, 2242]. Rays [462, 1840]. razor [723]. RBF [1162, 1096, 2054, 1097]. RBF-like [1096]. RBF-Networks [1162]. RC2 [1434]. RCBR [414]. RDBMS [380]. Re [906, 219]. Re-enactment [906]. Re-implementation [219]. Reachability [167, 156, 1838, 1954]. Reactive [1833, 1549, 340, 891, 669, 212, 646, 1811, 2142, 868, 1864, 1834, 1890]. Read [48, 1797]. Read/Write [1797]. Reading [1644]. Real [648, 413, 1772, 290, 647, 1722, 597, 2018, 324, 540, 1106, 1738, 2262, 1784, 1524, 1785, 590, 1745, 537, 1724, 17, 36, 250, 1907, 720, 1911, 289, 298]. Real-Time [648, 1772, 647, 2018, 1106, 1738, 2262, 1784, 1785, 537, 290, 590, 1907, 1911, 298]. real-valued [720]. Real-World [1745]. Realistic [120, 197]. Reality [2125, 2067, 1310, 1595]. Realization [1076, 1577]. Realizations [2210]. Realm [1290]. Reasoner [1741]. Reasoning [240, 495, 1230, 1221, 394, 211, 892, 809, 640, 222, 1713, 154, 365, 1392, 879, 885, 939, 924, 1783, 928, 1564, 1702, 1200, 248, 1788, 1018, 1867, 977, 1003]. Rebalancing [54]. Receptive [1073, 1068, 1072, 1064]. Recessive [1702]. Recipes [803]. recipient [1452]. Reciprocal [2223]. Recognisers [312]. Recognition [1984, 2115, 484, 563, 605, 1623, 1717, 2167, 1631, 564, 1999, 2001, 1621, 602, 1721, 1624, 2007, 481, 2165, 2011, 2012, 597, 609, 2016, 2160, 1713, 578, 2126, 2135, 1627, 2024, 2090, 2170, 2173, 2038, 2107, 2105, 2103, 571, 2041, 2042, 610, 1086, 2172, 604, 1519, 2133, 2168, 441, 566, 606, 1637, 1041, 705, 945, 1622, 2110, 2134, 619, 2145, 590, 463, 1639, 1720, 2093, 987]. Recognition-Based [2168]. Recognizable [1192]. Recognizer [2094]. Recognizing [2021, 1111, 1481]. Recombinative [306]. Recommendations [1726]. Reconciling [648]. reconfigurability [269]. Reconfigurable [1250, 295, 284, 260, 288, 258, 278, 289, 268, 259, 300, 283]. reconfiguration [264, 250, 268]. Reconstruct [2004]. Reconstructing [462]. Reconstruction [2233, 1840, 2100, 1769, 2137, 2109, 625, 1061, 2040, 712, 562, 520, 614]. recorded [2152]. Recording [419, 1935, 440]. Recovering [561]. Recovery [376, 2153, 2069, 2070, 2127, 2129, 498, 2076]. rectangle [1973]. Rectangular [513, 1298]. Recurrence [44, 40]. Recurrences [1348].

Recurrent [1733, 1039, 235, 1805, 1099, 1104, 1106, 560, 1102, 1105, 990]. Recursion [1268, 1820]. Recursive [863, 1502, 1799, 2127, 2078, 720]. Redesign [818]. Redistribution [1353]. Reduce [1524]. Reduces [451]. Reducing [245]. Reduction [1129, 169, 1689, 1498, 671]. Reductions [1837]. Redundancy [245]. Redundant [181, 1504, 404]. Reed [2238]. Reeling [1552]. Reference [1985, 1993, 810]. References [1631]. Refined [354, 635]. Refinement [808, 1836, 1907, 643, 1232, 1288, 2065, 668, 2204, 1548, 1931, 1866, 1934, 1904, 1968, 1927, 1891]. refinement-type [1934]. Refinements [124, 1912]. Refining [164, 974, 861, 460]. Reflectance [2050, 2124]. Reflection [2071, 1985, 1919]. Reflective [915, 2259]. REGAL3.2 [901]. Region [469, 557, 474, 1974, 313]. Regions [2147, 551, 1196, 2118, 2120, 1207, 1206, 463, 1233]. Register [1484]. Registration [576, 550, 505, 1696]. Regression [1089, 1112, 931, 758]. Regression-Analysis [758]. Regular [1192, 38, 1023, 731]. Regularization [583, 1080]. Regularized [1090]. Regularizing [1093]. Reinforcement [1079, 1036]. Related [24, 1434]. Related-key [1434]. Relating [485]. Relation [1298, 1539, 1848, 1554]. Relational [872, 685, 1556, 1599, 154, 39, 365, 1667, 1593, 335, 2149, 366]. Relational-Functional [1556]. Relations [866, 873, 226, 762, 1208, 1562, 1573, 1232, 1233]. Relationship [1307, 1278, 749]. Relationships [1063, 385, 947]. Relative [1075, 2128]. relatives [1979]. Relaxation [460]. Release [1786, 152]. Relevance [357, 1109, 949, 1016, 741]. Reliability [447]. Reliable [846, 854, 1204]. Remark [1546]. Remarks [1418, 783, 749]. Remote [440, 649, 621, 624]. Renaming [1024]. Rendering [1373, 293]. Reorganization [1043]. Repair [1695, 239]. Repartitioning [123]. Repetitive [1827]. Replacement [1611]. Replanning [1892]. Replay [905]. Replicated [63, 158]. Report [2216, 160]. Reporting [1226]. Repositories [960, 948, 1284, 824]. Repository [396]. Represent [1163]. Representation [1986, 1380, 1715, 356, 476, 1844, 703, 564, 2080, 2085, 211, 1141, 1219, 992, 2198, 1197, 1857, 1382, 329, 1400, 1545, 1787, 894, 1203, 227, 371, 527, 1195, 17, 1012, 999, 1009, 1870]. representation-based [1012]. Representation-Oriented [1203]. Representations [532, 1212, 1560, 1099, 1610, 701, 178, 1561]. Represented [1222]. Representing [1130, 1018, 1915, 1291]. Required [654]. Requirements [1217, 1589, 644, 1609, 2193, 395]. rescheduling [1890]. Rescue [594]. Research [1264, 1991, 751, 2111, 1627, 963, 1275, 620, 441, 769, 2112, 271, 259]. Resemblances [816]. Reservation [1777]. Reserved [416]. Residuated [1195]. Resistance [782]. Resisting [2244]. Resolution [567, 1044, 1629]. resolvable [1423]. Resolving [1725, 1307, 1374]. Resonance [582, 614]. Resource [1831, 937, 1174, 1509, 2267, 1777, 239, 1868]. Resource-Adaptive [239]. Resource-Bounded [1831]. Resources [1796, 1246, 921]. Response [866, 1053]. Responses [697, 1036]. Restart [1613]. Restartable [844]. Restoration [1698, 596]. Restoring [2083]. Restricted [1257, 778]. Restrictions [1611]. Restrictive [2253]. Restructuring [688]. Results [1715, 1804, 1468, 925, 1208, 28, 1829, 128, 584, 1830, 51, 1460]. Rether [1906]. Rethorical [866]. Retina [1696]. Retraction [1789]. Retraining [1110]. Retransmission [659]. Retrieval [2167, 1691, 552, 554, 551, 357, 956, 856, 908, 962, 951, 531, 362, 2030, 332, 530, 961, 548, 546, 555, 358, 1280, 1573, 1035, 957, 947, 510, 545, 330, 1001, 968]. Retriever [1522]. Retrospective [1628]. Reusable [171]. Reuse [664, 816]. Reusing [811]. Reverse [154, 365, 1609, 782]. reversible [722]. Review [538]. Reviews [160]. REVINOS [833]. revising [739]. Revision [1713, 833, 1026]. Revisited [2237, 1433, 1606, 1390, 1850, 754, 1552, 1924, 1889]. Revisiting [845]. Reward [1036]. Rewrite [670]. Rewriting [1822, 1613, 934, 1965]. Rhesus [228]. RHODOS [1337]. Rhythm [764, 1721]. Rib [573]. Rich [1586]. Richard [749]. Ridge [2054]. right [226]. Rights [907, 942]. Rigid [216, 400, 2129, 2198]. Rigorous [1908]. Riley [259]. Riley-2 [259]. ring [257]. ring-oscillators [257]. Rings [1474, 843, 2095]. RLP1 [659]. Road [2106]. Roadmap [795]. Roads [2145]. Robert [749]. Robot [1151, 1169, 1147, 2143, 1154, 540, 868, 1387, 1143, 1160, 1146, 1871, 994]. Robot-Animal [1154]. robot-collected [994]. Robotic [1735]. Robotics [1717, 708, 1388]. Robots [1157, 233, 599, 1168, 1153, 1690, 2072, 300]. Robust [2155, 1989, 2125, 1997, 2001, 2238, 2135, 2031, 500, 522, 2088, 1738, 2060, 1050, 91, 596]. Robustness [1114, 1655]. ROCK [1594]. Role [2230, 884, 398, 1392, 1767, 819, 1376]. Roles [1296]. ROLL [1594]. Room [1557]. Root [757]. Rooted [855]. Roots [2237]. Rotated [1095]. Rotation [2086, 2090]. Rotational [2025]. Rotations [2165]. Rothko [252]. Rotorcraft [115]. rough [995]. round [1413]. round-complexity [1413]. Rounds [1552]. Route [1218, 1213]. Routed [19]. Router [2197]. Routing [32, 24, 840, 1502, 29, 30, 18, 1266, 130, 408, 20, 1948, 271, 1949, 1950]. Routings [1487]. RPC [11]. RSA [1441, 2231, 2239, 1432, 1442]. RSA-keys [1442]. RSA-type [1441, 2239]. Rule [1683, 1700, 638, 827, 873, 1684, 931, 1811, 1278, 632, 635, 1017]. Rule-Based [1683, 1700, 1684, 1811]. Rules [1253, 930, 636, 637, 891, 1681, 875, 862, 1781, 224, 1783, 317, 634, 1675, 1907, 1007, 682, 1925, 995, 1025]. Run [263, 1742, 1743, 266, 260]. Run-Time [1742, 1743, 263, 266, 260]. Running [1362]. Runs [1544]. Runtime [37, 76, 125, 1519]. RVBR [414].

s [757, 1447]. S-Boxes [1447]. Sacrificing [1107]. Safe [376]. SAFER [2228]. Safety [1377]. Saliency [2084]. Salient [2120]. same [915]. Sample [1726, 50, 598]. samples [689]. Sampling [306]. SAP [804]. Sapphira [212]. SAR [2057, 598]. SAT [1958, 1944]. Sather [1750]. Satisfaction [1242, 910, 1249, 1693, 1251, 1244, 1248, 1695]. Satisfiability [295, 1267, 1259, 1460]. Satisfying [195, 1233]. Saturation [451]. Save [1290]. SBS [1355]. SBS-Type [1355]. Scalability [1033, 89]. Scalable [1994, 96, 434, 951, 802, 87, 1458]. Scale [1362, 1943, 2090, 700, 2098, 660, 524, 2044, 1521, 546, 953, 163, 461, 2087, 2156, 740]. scale-sensitive [740]. Scale-Space [700, 524, 461]. Scale-Spaces [2087]. Scales [715, 1075, 700, 2088]. Scanned [455]. Scans [2095]. Scattering [583]. Scenario [1510]. Scenario-Based [1510]. Scenarios [161]. Scene [1734, 2098, 2105, 2051, 544, 520, 2134, 2123]. Scene/Object [2123]. Scenes [2060, 2104]. Schaffer [1044]. Schedule [317]. Schedules [119]. Scheduling [100, 1263, 903, 1264, 120, 1246, 413, 118, 1476, 1369, 410, 116, 35, 313, 842, 314, 1553, 886, 1798, 316, 321, 56, 317, 1785, 279, 1971, 1886, 1942]. Schema [1297, 388, 899, 1599, 1283, 755, 1308, 371]. Schemaless [371]. Schemas [1300, 1307, 1292]. Schemata [1201, 1198]. Scheme [1994, 117, 810, 2007, 1801, 434, 2240, 2040, 378, 75, 364, 2253, 606, 787, 411, 1459, 1456, 1418, 1425, 1981, 1961, 1449].

Schemes [2247, 790, 1412, 1983, 1432, 1416]. Schumann [769]. SCI [1334]. Science [1528, 940]. Scientific [814, 1758, 1768, 1777]. Scientist [1541]. Screening [401]. Script [603, 2091]. SDL [647]. Search [1461, 805, 1707, 2158, 242, 925, 907, 1249, 1240, 1243, 1272, 1252, 334, 315, 237, 947, 1008, 1878, 1886, 1891]. search-space [1891]. Searching [1607, 949]. sec [288]. SECCO [2260]. Secret [1419, 1418, 1416]. Secure [804, 775, 856, 2263, 1422, 2039, 837, 619, 1440, 1456, 1457, 1452]. Security [1438, 2226, 786, 801, 145, 780, 1431, 1421, 1022, 1436, 1424, 1881]. Seek [766]. Segment [575, 517, 2110]. Segment-Based [2110]. Segmentation [612, 477, 2155, 502, 2117, 443, 469, 476, 2080, 603, 580, 468, 450, 478, 501, 1734, 2118, 2026, 2170, 577, 1400, 604, 480, 442, 950, 586, 2058, 451, 2060, 1646, 621, 584, 471, 2081, 472]. Segmenting [2088]. Segments [1209, 1842]. Segregation [769, 1065]. Selected [613, 1651]. Selection [506, 2007, 165, 446, 700, 2267, 1097, 1110, 976, 990, 1947, 1887, 989]. Selective [1073, 1922, 1008]. Selectivity [1069]. Self [61, 1733, 1141, 1133, 1568, 855, 2137, 1805, 1443, 1075, 1608, 1123, 60, 756, 1216, 1143, 1620, 767, 1102, 1983, 292]. Self-Improving [1568]. Self-Occluding [2137]. Self-Organising [1733]. Self-Organization [1075, 1216]. Self-Organizing [1141, 1133, 1123, 1143, 1620, 767, 1102, 1805, 756, 1983]. Self-Stabilizing [61, 855, 1608, 60]. Self-synchronized [1443]. self-validating [292]. Semantic [1988, 899, 960, 201, 875, 1668, 1531, 1572, 642, 947]. Semantics [1253, 207, 1238, 647, 1812, 1679, 1682, 1532, 1811, 68, 893, 1678, 1378, 1588, 1709, 667, 1825, 1675, 1899]. Semaphores [1331]. Semi [960, 362, 1127, 491, 2062]. Semi-automatic [2062]. Semi-Blind [1127]. Semi-Local [491]. Semi-structured [960, 362]. Semiconductor [110, 1370, 111]. Semigroups [1255]. Semilinearity [1190]. Semistructured [371]. sense [1941]. Sensing [624, 599, 2017, 621, 1860, 1871]. Sensitive [1728, 1174, 1732, 740]. Sensitivity [1069, 828, 197, 1784]. Sensor [599, 1735, 2107, 1737, 452, 300]. Sensor-Based [1735]. sensor-data [300]. Sensors [492, 1165, 1125, 622, 292]. Sensory [1056]. Sentences [1727, 745]. Separable [1463, 309]. Separable-Homogeneous [1463]. Separation [1126, 1127, 1121, 1119]. Separator [57]. Sequence [234, 2018, 71, 2065, 400, 498, 1156, 1102, 1946]. Sequences [2223, 1733, 1780, 1141, 1600, 501, 618, 625, 2159, 500, 2042, 2168, 549, 2141, 1415, 1947]. Sequencing [929, 1237, 315]. Sequential [1804, 1762, 161, 1367, 1402, 609, 1049, 1382, 1643, 2273, 1899, 273, 1983]. Sequentiality [1611]. Serial [1500, 777, 753]. Serializability [1700]. Series [1533, 1122, 1571, 1103]. seriously [143]. Server [413, 439, 1774, 11, 404]. Servers [411]. ServIce [402, 421, 414, 434, 426, 407, 1616, 2266, 902]. Services [952, 432, 415, 2262, 619]. Sesame [2039]. Set [181, 2176, 670, 955, 2005, 186, 1239, 185, 1184, 1275, 1726, 1856, 1552, 1024, 995]. Set-Based [186]. Sets [462, 445, 1171, 1537, 1602, 1840, 1535, 1539, 1478, 1188, 1784, 1206, 728, 1975]. Setting [932]. Several [1814, 2045]. Severely [625]. SGML [1018]. Shading [561, 514, 2020, 512, 2049]. Shadows [2004, 573]. Shape [1987, 563, 318, 1629, 1221, 1999, 561, 2002, 506, 489, 514, 759, 2136, 512, 329, 113, 546, 550, 613, 2049, 511, 2124, 562, 2069, 2070, 2145, 2130, 2076, 2140]. Shape-from-Shading [514]. Shaped [507]. Shapes [1844, 2001, 2008, 752, 2024]. Share [1800]. Share-Where [1800]. Shared [1804, 1833, 665, 857, 102, 103, 1500, 1516, 1815, 1517, 6, 1332, 36, 1834]. Shared-Memory [857, 1500, 1815, 6]. Sharing [181, 1658, 182, 1419, 1418, 1416]. Sharper [1829]. SHEARSORT [1945]. Shell [936]. Shift [41, 835, 1946]. Shift-Invariant [41]. Shifted [588]. Short [2194, 1314, 1548]. Short-Range [1314]. Shortest [30]. Shot [1075]. Showing [2102]. Shy [1290]. Sighted [2063]. Sign [2018, 2020, 2050]. Sign-Language [2018]. Signal [703, 2000, 1772, 1120, 1401, 1036, 250]. Signals [697, 1403]. Signature [2240, 1641, 795, 2046, 2253, 1626, 1640, 1643, 1445, 1449]. Signatures [1642, 794, 2239, 482, 796, 1433]. Signed [224]. Significance [1226, 1078, 1427]. Silhouette [2105]. Silver [1290]. Similarities [1132]. Similarity [457, 329, 546]. Simple [1084, 1462, 458, 840, 105, 803, 775, 1699, 2041, 1835, 1703, 1736, 1646, 1064, 725, 744, 742]. simplification [1895]. Simulated [1149, 1040]. Simulating [301]. Simulation [1511, 1779, 1780, 1342, 1370, 109, 111, 112, 1314, 755, 1739, 1404, 1521, 134, 130, 632, 170, 1374, 926]. Simulations [1229, 1368]. Simulator [110]. Simulators [1762]. Simultaneous [2153, 945, 1065]. simultaneously [1952]. Sine [2131]. SINERGY [1883]. Single [305, 1681, 103, 784, 1266, 2161, 1053, 2095]. Single-Chip [103, 784]. Single-Vehicle [1266]. Singular [2049]. Singularities [601, 1971]. Site [149]. Situation [1710, 833]. Situs [465]. Size [1394, 1069, 1171, 1246, 2007, 45, 598, 1966, 1967, 1891]. skeletal [1882]. Skeleton [486]. Skeletonization [513, 1630]. Skeletons [603, 80, 81, 1979]. Sketch [1231, 2002]. Sketch-Based [1231]. Skin [1056, 2166]. Skin-Color [2166]. Slicing [165, 169]. Slit [2131]. Sloppy [1183]. Small [2227, 1998, 2237, 1482, 1484, 598]. Smart [1125]. Smear [2130]. SML [184]. Smoothing [2035, 543]. SMP [1742]. SNePS [917]. Snoop [103]. Soccer [1158, 553]. Societies [937]. Society [1530, 1596]. Soft [1134, 413]. Software [1380, 1809, 1589, 1658, 14, 177, 1906, 159, 109, 396, 409, 176, 1499, 145, 1590, 1894, 160, 143, 832, 1404, 1787, 1516, 147, 1647, 175, 140, 168, 1517, 2214, 671, 221, 9, 144, 152, 353, 158, 1526, 1450, 1444, 277, 2271, 1554, 283, 1437]. Software-Oriented [1787]. Solaris [135]. Solid [580]. solitaire [1965]. Solomon [2238]. Solution [1700, 1762, 514, 659, 1617, 74, 2075]. Solutions [40]. Solvable [1269, 1830]. Solve [236, 95, 893, 886, 1703, 1886]. Solver [1814, 93, 1234, 1745]. Solvers [1512, 1262, 1355, 1372]. Solving [866, 806, 1764, 1236, 1239, 910, 811, 812, 1271, 1941, 1708, 1757, 1817, 314, 2153, 1579, 816, 1740, 818, 90, 911]. SOM [1068, 1131, 1142, 1729]. SOM-Model [1068]. Some [1847, 1536, 783, 701, 1817, 1357, 1619, 749, 758]. Sonar [600]. Sort [48, 50]. Sorted [1600, 215, 1947]. Sorting [593, 51]. Sortings [1235]. Sound [752]. Sound-Producing [752]. Sounds [759]. Source [2147, 1126, 1127, 1214, 1121, 1119, 2069, 1950]. Sources [2224, 2070, 997]. South [313]. SP2 [122]. Space [1511, 1653, 492, 1076, 242, 710, 195, 700, 524, 2044, 546, 320, 60, 1201, 461, 2156, 969, 1891]. Space-Efficient [60]. Space-Variant [492]. Spacecraft [821, 969]. Spacecrafts [885]. Spaces [1199, 770, 2087, 740]. Span [1915, 1914]. Spanners [1468]. Spanning [61, 1963]. Spared [1055]. Sparse [2227, 1512, 1602, 111, 1773, 1747, 43, 1539, 39, 178, 89, 38, 1359]. Sparsely [1039]. Spatial [1223, 1230, 1221, 394, 211, 1224, 1219, 226, 1043, 1212, 1559, 1196, 1204, 1558, 1560, 1208, 1155, 385, 1652, 364, 1198, 467, 1702, 1200, 598, 1651, 968, 1561, 1971]. Spatially [1202]. Spatio [356, 495, 715, 1086, 2149, 1041, 2131]. Spatio-Temporal [495, 715, 2149, 1041, 356, 1086, 2131]. Spatiotemporal [496, 1064]. Speaker [1720]. SPEC [127]. Specialization [197, 1687]. Specialized [579]. Specific [638, 1262, 1753, 534, 1006, 269, 265]. Specification [1901, 1809, 1287, 1589, 398, 650, 640, 645, 1609, 654, 1289, 1739, 1714, 666, 659, 817, 642, 1685, 1902, 1906, 1917, 1919]. Specifications [1800, 652, 1588, 641, 141, 1931, 1922]. Specified [783, 1354]. Specifying [649, 1415]. Spectra [2013, 759, 1374, 676]. Spectrometers [583]. spectrum [286]. Speculation [139]. Speculative [1]. Speech [228, 2039, 1041, 1050, 1720, 1014, 2270]. Speed [853, 104, 2243, 254, 779]. Speeding [541, 989]. Spheres [1474]. Spherical [1986, 2069]. Spicules [578]. Spike [1040]. Spiking [1039, 1046, 1049, 1092]. Spiral [142]. Split [2231]. Splitting [2078]. SPMD [67]. SPOCK [1563]. Spoken [831, 986]. Spots [1724]. Spotting [2172]. spread [286]. Sputum [586]. SQL [1285, 384, 1666]. Squares [1989, 1081, 97]. SRAM [275]. SRT [1823]. SSOR [96]. SSS [1509]. SSS-Core [1509]. st [1964]. st-graphs [1964]. Stabilizing [61, 855, 1608, 60]. Stack [1552]. Stage [510]. Staircase [57]. Stamped [355]. Standard [1934, 2248, 2266, 1812]. Standards [803, 1284]. Standards-Based [803]. Star [1801, 1948]. Stars [1476]. State [1550, 1621, 242, 2111, 169, 1610, 1378, 114, 2273]. Statecharts [1792, 667]. Stateless [1420]. STATEMATE [667]. States [1549, 1712]. Static [207, 558, 196, 45, 28, 135, 192, 4, 147]. Stationary [1079, 1133]. Statistical [1096, 713, 1120, 1075, 613, 515, 1245, 128, 753]. Statistics [2094, 1070]. Steady [114].

Steady-State [114]. Steel [1379]. Steganography [2048]. Step [2158, 1171, 505, 536]. Stepping [1596]. Steps [359]. Stereo [1992, 1996, 2006, 523, 524, 2044, 2059, 1696, 517, 2063, 520, 2110, 2068, 2075, 516]. Stereo-Matching [520]. stereochemistry [675]. stereochemistry-based [675]. stereopsis [290]. Stereovision [2005, 870]. Still [429]. Stimuli [1071]. STL [1756, 1745]. Stochastic [1047, 1091, 1890]. Stones [1596]. Storage [856, 45, 1807, 1790, 2266]. Store [804, 1604, 26]. Store-and-Forward [26]. Stored [414]. Story [1548]. Straight [1858, 1842]. Strain [112]. Strategic [799]. Strategical [2142]. Strategies [1744, 915, 1046, 712, 1158, 1564, 2256, 976, 1959, 1970, 1891]. Strategy [1360, 561, 115, 886, 320, 671, 1886]. Stream [260, 2243, 769, 1429]. Streams [410, 522, 1553, 1928]. Striate [1069]. Strict [1192, 1508, 1436, 200]. string [1965]. Strong [853, 1611, 1554]. Strongly [1368]. Structural [1847, 493, 398, 1101, 297, 596, 679]. Structure [1631, 608, 308, 2269, 1244, 814, 103, 699, 1645, 1488, 2153, 709, 2083, 1035, 404, 1206, 1808, 2127, 498, 537, 2151, 2128, 676, 1435]. Structured [1742, 214, 593, 2258, 961, 1680, 835, 1484, 249, 336, 1864, 960, 1935, 362, 973]. Structures [1988, 42, 41, 933, 1993, 492, 381, 1559, 1558, 1652, 1656, 364, 2149, 1736, 1669, 1200, 1139, 2093]. Structuring [507, 1213, 361, 1201]. Studies [1699, 1920]. Study [897, 1360, 1069, 672, 1725, 809, 640, 715, 921, 946, 1722, 235, 761, 2036, 2049, 839, 1201, 1563, 127, 668, 128, 1040, 1567, 158, 1906, 265, 1884, 283, 679]. Studying [766]. Stumpf [754]. Stuttering [1835]. Style [1833, 1184, 749, 2272, 1834]. Sub [925, 770]. Sub-Admissibility [925].

Sub-Spaces [770]. Subcategorization [874]. Subdefinite [1274]. Subgoal [1570]. Subgraph [1485]. subgroup [1445]. Subject [1172]. Subjective [456, 225, 2134]. subliminal [1443]. Subrange [1374]. subset [1981, 1880]. subset-sum [1981]. Subsets [1269]. subsorted [1900]. Subspace [2167, 2160, 1137]. subspaces [722]. subsumption [695, 686]. Subsymbolic [763]. subsystem [254]. Subsystems [1346]. Subtraction [576]. Subtype [195]. Subtypes [141]. Successful [1306]. successive [1917]. Suffering [931]. sufficient [735]. Suffix [242]. Sugiyama [2214]. Suite [1771]. Suited [431]. sum [1981]. Summarization [1227]. Summary [2222]. Supercover [1858]. Superpositions [1843]. Superquadrics [2078]. Superscalar [139, 138]. Supervised [1990, 617]. Supply [316]. Support [943, 155, 388, 1661, 1746, 1285, 1342, 1199, 1742, 1601, 891, 1098, 1184, 1328, 1813, 423, 7, 858, 397, 1280, 1564, 2260, 711, 1077, 239, 1010, 1882]. Supporting [1221, 167, 828, 1599, 319, 346, 87, 1753, 1785]. Supportive [427]. Surface [2101, 764, 561, 2185, 2020, 2022, 1852, 2041, 2050, 488, 1198, 2124, 460]. Surfaces [2004, 567, 2041, 565, 1848]. Surfacess [1847]. Surgical [2009]. Surrogates [945]. Surveillance [2114, 589, 420]. Survey [1538, 627]. Survival [931]. SVP [787]. swap [1983]. swap-with-parent [1983]. swappable [267]. Switch [1495, 1923]. switching [253]. SX [1335]. SX-4 [1335]. syllabification [983]. Symbolic [1780, 2103, 1718, 1926]. Symmetric [92]. Symmetry [2025]. SymSPACE [1534]. Synapse [1065]. Synaptic [1037, 1054]. Synchronisation [133]. Synchronising [65]. Synchronization [1497, 857, 1406, 415, 1750, 1925, 74]. Synchronized [59, 1443]. Synchronizing [1332]. Synchronous [1321, 2225, 842, 2245, 2249].

Synchronous-CDMA [2249]. Syntactic [1174, 1531, 1190, 1480, 823]. Syntactical [1631]. Syntax [875]. Synthesis [1988, 2045, 615, 1280, 770, 1150, 1862, 1002, 273, 684, 260, 276]. Synthesized [2139]. Synthesizing [732]. Synthetic [473]. System [1700, 1321, 16, 897, 1379, 898, 1228, 808, 1363, 1719, 413, 1407, 665, 1705, 1365, 357, 650, 1624, 162, 804, 1367, 538, 327, 1136, 1162, 1325, 830, 1373, 1774, 813, 1793, 597, 2201, 857, 1406, 791, 788, 853, 581, 12, 591, 876, 2208, 2034, 1392, 416, 629, 1375, 965, 1389, 615, 1807, 1500, 1516, 246, 1520, 1166, 358, 407, 1304, 1644, 1337, 1318, 1507, 1517, 1411, 2062, 247, 221, 2068, 1632, 536, 1040, 101, 420, 51, 401, 239, 1639, 336, 390, 1898, 997, 1448, 999, 1010, 1933, 1017, 1444, 1008, 992, 280, 289, 1936, 268, 251, 730, 256, 300]. system [1011, 1940, 975]. Systematic [806, 747, 748, 702]. Systems [155, 1380, 1361, 236, 388, 1671, 1833, 402, 339, 1549, 1395, 1782, 1385, 670, 922, 323, 389, 1351, 1706, 447, 809, 774, 1586, 2232, 829, 1490, 1489, 1383, 2009, 1172, 1684, 849, 435, 1328, 645, 1587, 529, 325, 367, 831, 669, 1099, 1186, 644, 382, 11, 412, 646, 332, 322, 1393, 1811, 1739, 368, 1344, 1714, 1345, 1241, 438, 1787, 2249, 1710, 175, 1656, 1593, 60, 1281, 817, 2251, 1216, 641, 627, 90, 928, 1376, 1278, 1564, 1702, 321, 642, 696, 850, 1377, 2256, 911, 1832, 333, 2074, 789, 331, 36, 1359, 1685, 1903, 1834, 1875, 1023, 269, 1914]. systems [2270, 1007, 1894, 969, 1424, 286, 727, 1965, 986, 1930].

T [1729]. T-SOM [1729]. T.120 [426]. T3D [1364]. T3E [93]. table [273]. Tableaux [217, 1867]. Tables [1664, 1013]. Tabu [1249, 237, 1886]. Tactical [319]. tactics [1895]. Tactile [1737]. Tagging [228]. Tags [1347]. Tailed [1243]. Tailoring [1262]. take [143]. Tale [174]. Tales [955]. Tamper [782]. Tandem [1550]. Tangential [1062]. Tape [2256]. Target [594, 254]. targeting [273]. Targets [2057, 2066]. Task [887, 1863, 120, 884, 800, 802, 119, 1336, 1784, 321, 1892]. tasking [1907]. Tasks [236, 844, 1161, 118, 812, 1094, 799, 127]. Taxonomy [337, 1203]. TEA [1434, 200]. Teach [143]. teaching [736]. Team [719, 1701]. Technical [2074]. Technique [319, 541, 344, 2011, 2012, 501, 1504, 2245, 660, 594, 539, 38]. Techniques [1129, 1263, 2225, 1028, 211, 405, 303, 1120, 1303, 1641, 130, 1636, 1635, 158, 1431, 1878]. Technologies [1752, 1494]. Technology [781, 1290, 1585, 273, 208, 272, 274, 270, 772]. Technology-driven [273]. telecommunications [251]. Teleeducation [1406]. Telephone [1171]. Teleservice [2267]. Teleteaching [433]. TELLTALE [1033]. Template [1995, 1755, 509]. Templates [664]. Tempo [762, 345]. Temporal [495, 1267, 1733, 877, 1782, 1384, 1282, 1472, 1598, 1599, 1046, 1141, 715, 1559, 1283, 222, 234, 501, 700, 1739, 2157, 364, 2149, 1281, 1696, 1041, 1092, 1401, 928, 947, 1666, 1102, 1859, 356, 968, 972, 1905, 1086, 2131]. Temporal-Spatial [364]. Temporally [1147]. Tensor [706]. Tentative [58]. Term [670, 185, 1212, 589, 731]. terminals [1974]. Termination [199, 633, 354, 200, 188, 635]. Terminological [339, 393, 365, 548]. Terms [861, 1711]. Terrain [1718]. Terrains [1576]. Tesseral [394]. Test [166, 165, 649, 1554]. Testing [636, 755, 68, 1635, 1554]. tetris [1945]. Texel [531]. Texel-based [531]. Text [605, 955, 813, 393, 962, 951, 2019, 2030, 950, 1632, 1444]. Textile [2033]. Texts [961]. Textual [1212, 1019]. Textural [452]. Texture [2117, 453, 771, 2086, 585, 480, 562, 1396, 482, 2095, 2081]. Textured [450]. Textures [457]. Thalamocortical [1038]. Their [1525, 2225, 2226, 812, 829, 1840, 1979, 1853]. Theorem [653, 907, 215, 218, 57]. Theorems [1545, 1195]. Theoretic [1046, 1145, 1414]. Theoretical [1406, 1711, 1393, 1958, 985]. Theories [1660, 96, 2, 814, 1186, 1711, 461, 1896, 974, 739]. Theory [2223, 1180, 2230, 391, 1586, 1224, 751, 762, 1240, 752, 176, 304, 1082, 206, 205, 879, 1617, 1528, 221, 1828, 614, 1002, 1004, 1554, 738]. Theory-Based [391]. There [762, 1534]. Thermal [257]. Thesaurus [956, 1573]. theta [686]. theta-subsumption [686]. Think [1582]. Thinning [2019, 511]. third [970]. Thread [1495, 1497]. Thread-Switch [1495]. threaded [1899]. Threads [348]. Three [2212, 2158, 2180, 1996, 2182, 1734, 2178, 2179, 2199, 49, 2148, 252, 1969]. Three-Dimensional [2180, 1996, 2182, 1734, 2178, 1969]. Three-Dimensionality [49]. Three-Phase [2212]. Three-Step [2158]. threshold [746]. Tickets [798]. tier [1742]. Tiles [1537]. Tilt [494, 2152]. Timbre [769, 767]. Time [155, 648, 1384, 413, 1550, 1772, 1742, 647, 1743, 597, 2018, 1747, 58, 848, 1075, 71, 1608, 700, 135, 540, 1122, 1106, 1882, 114, 2172, 1738, 2262, 1580, 1784, 315, 210, 1571, 1103, 1753, 1524, 1785, 1403, 537, 355, 1868, 263, 1837, 290, 250, 1907, 1911, 260, 289, 741, 731, 730, 266, 680, 298, 1929, 590]. Time-Causal [700]. Time-Dependent [114]. Time-oriented [1882]. Time-Stamped [355]. Time-Triggered [848]. Time-Varying [2172]. Timed [1238, 667, 1902]. Timeout [846]. Timeout-Free [846]. Timetabling [913]. Timing [1054, 1053, 1901]. TIN [1649]. TINA [428]. TINPAR [1504]. TINs [1650]. Tiny [1761]. Tissue [574, 584]. Tissues [331]. TLA [658, 659]. TM [293]. Today [794]. Token [1551, 855, 60]. Tolerance [851, 23, 1897, 264, 1913]. Tolerant [1801, 1114, 837, 133, 1487, 1948, 1898, 732, 1911, 1455, 1957]. Tomography [583, 572]. Tonal [761, 754]. Tonality [757]. Tool [653, 895, 899, 1410, 2203, 1312, 35, 2199, 833, 769, 334, 520, 271, 1937]. Toolkit [638, 1938]. Tools [1658, 915, 707, 417, 1408, 7, 776, 153, 1636, 1346, 252, 899]. Toolset [14]. Top [380, 486, 1325, 689, 862, 1680, 81]. Top-Down [486, 862, 689]. Topic [950]. Topographic [1134]. Topography [1718]. Topological [1816, 1208, 1854, 1207, 1850, 1232, 1233, 1972]. Topologies [1850]. Topology [1130, 1839, 1841, 1969, 511]. Topology-Based [1841]. Topology-Preserving [1839]. Tori [24, 26, 25, 1950]. Toroidal [19]. Torus [1502]. Total [1989, 19, 1974]. Total-Exchange [19]. Touching [2209]. Tower [811]. Tower-of-Adapter [811]. TR [1763]. TR/O [1763]. TR/O-Unitaire [1763]. Trace [1546, 644, 1783]. traceability [1450]. Traceable [1417]. traces [1936]. Tracing [9, 1803, 1857]. Track [1738, 2256]. Tracking [233, 1368, 2018, 2159, 2043, 2142, 1166, 2058, 553, 2068, 2152]. Tractable [1270, 1268, 879]. trade [1457]. Traffic [502, 1407]. Trained [930]. Training [1081, 915, 914, 1524]. Trains [1061, 1037, 1040]. Transaction [1287, 375, 376, 1888]. Transaction-Based [1287]. Transactional [1668, 1659]. Transactions [1660, 858, 777, 1905]. Transfer [2099]. Transform [532, 538, 105, 2079, 711]. Transformation [162, 308, 470, 92, 104, 2034, 1308, 595, 331, 1863]. Transformational [1301]. Transformations [1804, 662, 1734, 1415, 1864, 1006]. Transforming [390]. Transforms [601, 704, 474, 1398]. Transient [310, 2239]. Transition [319, 1915, 1914]. Transitions [1134]. Transitive [1662]. Translation [652, 521]. Translation/Deformation [521]. Translational [709]. Translations [1706, 1670]. Translator [1604]. Transmission [414]. Transparent [858, 1486]. Transport [1766, 1372]. Trapdoor [1446]. Trapezoid [591].

Trauma [1010]. Travel [1246]. Treatment [389, 1187, 629]. Tree [1178, 61, 186, 1239, 242, 863, 1483, 1484, 83, 1916, 731, 742, 1729]. Tree-SOM [1729]. Tree-Width [1483, 1484]. Trees [22, 120, 54, 1725, 2211, 1476, 1555, 1174, 311, 1607, 1256, 23, 1617, 25, 27, 232, 463, 1016, 1957, 2191]. Treewidth [1470, 1473]. Trellis [2225, 2241]. Trends [1628, 1621]. Triangle [1469]. Triangle-Free [1469]. Triangles [1858]. Triangular [24]. Triangulated [1576]. Triangulation [2019, 1980]. Triangulations [1816, 2174]. Triconnected [2183]. Tricubes [1845, 1846]. Tridiagonal [90]. Trigger [1671, 630]. Triggered [848]. Triggering [635]. TriggerMan [630]. Triggers [1660, 345]. Trilinear [706]. Trios [1535]. TRN [1067]. True [191]. Truncated [1359]. Truncation [549]. Trunked [2232]. Trust [791, 1430]. TSP [1579]. TSU [1497]. TTM15 [149]. Tuning [1073, 6, 822]. Tunneling [407, 1498]. tuple [1117]. Turbo [2236, 2248]. Turbulence [1374]. Turing [1390]. Turning [718]. Tutoring [936]. TV [2167]. Twilight [1552]. Twist [1535]. Twist-Closed [1535]. Two [612, 477, 513, 1473, 1953, 375, 1048, 1810, 531, 2190, 2109, 138, 115, 1032, 1432, 2149, 505, 174, 536, 767, 510, 614]. Two-Axis [2190]. Two-Dimensional [477, 1048]. Two-face [1953]. Two-Layer [375]. Two-Level [1810, 138, 115, 1032]. Two-Stage [510]. Two-Step [505]. Type [1180, 999, 190, 193, 1049, 378, 1798, 934, 1797, 221, 83, 1934, 1355, 1441, 2239]. Type-Based [190, 1798]. Type-Directed [193]. Typechecking [66]. Typed [193, 1668, 334]. Types [1814, 1550, 1188, 70]. Typical [2079]. Typing [1796]. Typographic [1634, 608].

UCPOP [1876]. Ultra [2243]. Ultrasound [612, 580]. UML [2214]. Unbounded [853, 68]. Uncalibrated [2128]. Uncaught [184]. Uncertain [100, 1285, 1015, 1003]. Uncertainty [918]. unconditional [1424]. Unconsummated [140]. undeniable [1459]. Underlying [226, 1481]. underpinnings [985]. Understanding [1879, 1185, 2092, 2196, 768, 1247, 1632, 1026, 2270]. Underwater [600, 622]. undirected [1949]. Unexpected [932]. unfair [1449]. Unfolding [583]. Unfoldings [1838]. Unification [1255, 1254]. Unified [708, 2212, 348, 803, 2120, 2130, 350, 2073]. Uniform [1543, 1555, 332, 2147]. Unifying [2]. Unique [1099, 1532]. Uniqueness [1665]. Unit [1497, 2209]. Unitaire [1763]. units [267]. Unity [673, 172]. Univariate [2237, 595]. Universal [1344]. Universality [1390]. Universe [1541]. University [143, 1276]. UNIX [413]. Unlinkable [777]. Unmeasured [1172]. Unsupervised [874, 1138, 2013, 1140, 1638, 480, 1092]. unweighted [1962, 1947]. Update [1627, 1668, 385, 1710, 1921, 1025]. Updates [1660, 1663, 629]. Updating [1670, 1669, 1000]. Upon [1381]. Upper [1467, 2254]. Urban [109, 2105, 1070]. Use [890, 966, 1796, 603, 1706, 447, 1205, 650, 1169, 1204, 197, 168, 317, 106, 824, 1412, 321]. User [881, 1231, 2229, 889, 157, 1030, 617, 12, 1408, 213, 246, 407, 1808, 1425]. User-Adapted [889]. User-Friendly [12]. Users [955, 829, 1452]. Uses [792]. Using [1301, 1397, 605, 2167, 2155, 1809, 866, 108, 301, 1286, 164, 483, 1629, 486, 601, 199, 1171, 2229, 703, 2101, 1246, 1996, 1997, 564, 453, 156, 1725, 915, 2000, 2147, 2001, 1364, 633, 490, 580, 2004, 1705, 1474, 541, 157, 303, 1570, 1367, 812, 1226, 1298, 1283, 1369, 450, 1370, 658, 800, 990, 593, 2011, 2234, 962, 1271, 1770, 501, 698, 1051, 1603, 1734, 1328, 1148, 948, 1165, 2164, 2020, 2021, 2191, 2135, 77, 1082, 831, 2023, 2024, 669, 2026, 2136, 1120, 10, 2029, 2090, 160, 1708, 2173, 2037, 1124, 2107, 2105, 1611, 228, 1501, 893, 500, 1488].

Using [1756, 2042, 2043, 1241, 610, 385, 1355, 615, 1500, 433, 2048, 1798, 480, 2168, 1698, 1357, 2267, 130, 868, 1703, 2251, 560, 397, 2161, 2054, 2058, 770, 2200, 1092, 1736, 467, 1571, 2062, 2108, 1664, 913, 1358, 668, 2110, 2096, 2068, 2259, 1040, 1856, 1695, 1200, 947, 510, 2204, 539, 590, 1491, 51, 2095, 401, 1087, 2104, 1347, 2148, 2077, 2081, 330, 331, 2141, 257, 281, 290, 185, 1874, 1425, 291, 687, 675, 1920, 727, 262, 254, 1888, 298, 677, 1892]. Utilizing [1762].

V [285, 1574, 1505]. V-Pascal [1505]. V-RaAP [285]. V.3 [1505]. V.L.P.’s [597]. Vagueness [1180]. Valid [391, 1204]. Validating [637, 160, 292]. Validation [898, 899, 310, 1918, 969]. Value [1467, 137, 1013]. valued [720, 352]. values [1021, 1012, 742]. Valves [649]. Vanishing [711]. Variability [1625, 853]. Variable [1833, 665, 567, 410, 2241, 2242, 967, 1834, 734]. Variable-Rate [410]. Variables [1109]. Variant [492, 867]. Variants [1193]. Variations [2084]. Various [1236, 2093]. Varying [715, 2172, 2161]. VBR [416]. VC [1117]. VDM [653, 656, 652]. VDM-SL [653, 656, 652]. Vector [1134, 564, 1098, 94, 1135, 496, 497, 2154, 572, 1077]. Vectorial [1625]. Vectorizing [1505]. Vectorizing/Parallelizing [1505]. Vectors [1747, 77]. Vegetation [600]. Vehicle [597, 1165, 2159, 1266]. Verb [874]. verifiable [1451]. Verification [657, 1780, 2101, 156, 809, 166, 662, 1911, 2135, 1823, 669, 1547, 1795, 969, 2040, 1641, 1241, 666, 659, 2046, 839, 1113, 1626, 1640, 614, 1832, 1643, 330, 2273, 1902, 1906, 1910, 1554, 1937, 1923]. Verified [664]. Versatile [120, 791]. versatility [975]. Verschmelzung [754]. Version [1361]. Versioning [1599]. Versions [360]. Versus [1820, 1259, 896, 1653, 1959, 414, 1004, 297]. Vertex [1618, 1952]. Very [106, 2213, 734].

VHDL [291]. Via

[2017, 802, 1984, 736, 1589, 191, 1998, 1869, 357, 587, 197, 2027, 1061, 1637, 444]. Video [1988, 1691, 422, 430, 1992, 1993, 439, 414, 501, 2023, 2263, 412, 496, 522, 965, 945, 536, 543, 420, 355, 534]. Video-Based [1993]. Video-on-Demand [439]. Videoconferences [440]. View [388, 1670, 162, 1489, 2136, 1155, 1641, 385, 1667, 2161, 1507, 1541, 1664, 1000, 706]. View-Graph [1155]. View-Oriented [162]. Viewpoint [1987, 672, 2102]. Viewpoints [2182]. Views [1663, 2109, 386, 385]. Violin [773]. VIRSBS [619]. Virtual [387, 1597, 285, 849, 2269, 102, 776, 944, 403, 20, 1310, 1595, 36, 261, 293]. Virtual-Physical [102]. VIRTUE [2032]. Visibility [565]. visible [504]. Vision [1493, 708, 1996, 2111, 616, 1250, 2143, 2034, 2051, 528, 441, 2059, 2063, 569, 2110, 2068, 2112, 2075, 289]. Vision-Based [2063]. Vision-Guided [2143]. Visions [796]. Visual [1394, 108, 568, 1399, 323, 1076, 2114, 1800, 1147, 710, 327, 1063, 529, 325, 713, 324, 332, 322, 1062, 2146, 1155, 2040, 712, 540, 2163, 2133, 1166, 334, 1146, 619, 2072, 1234, 333, 1066, 331, 2141, 1417, 1008, 544]. Visualization [1205, 1650, 357, 579, 1131, 8, 424, 1574, 1507, 2213]. Visualizing [567]. Vivid [923]. VLSI [1165, 1045]. VOD [411, 405]. voice [1431]. Voiced [1050]. Voicing [757]. Volume [582, 1844, 175]. Volumes [1368]. VoR [416]. Voronoi [1648, 1209]. Vowel [1720]. Voxel [318]. VPR [271]. VR [1365, 914, 2122, 2139]. vs [1631, 1081, 1560, 1561, 103, 1062, 1969, 237].

Wait [857, 858]. Wait-Free [857, 858].

Wales [313]. Walk [237]. Walks [1653, 1480]. Walkthroughs [545]. Wallaschek [749]. Walsh [1409]. WAM [202]. WAMM [1339]. Wang [1537]. Warehousing [1661]. Warp [58]. Warping [501, 485]. Water [316]. watermarking [1452]. Watershed [442]. Watson [727]. Wave [1141, 109, 134]. waveforms [681]. Wavelet [601, 538, 2016, 1396, 1085, 482]. Wavelet-Compressed [2016]. Wavelets [563, 401]. Way [1067, 1446, 1403, 975, 1434]. Wayfinding [1218, 1201]. Weak [1466, 1523, 2148, 1442]. Weak-Order [1466]. Weakly [886]. Web [1658, 1705, 2206, 337, 418, 775, 338, 1029, 1682, 423, 334, 949]. Web-Enabling [338]. Webs [423]. Weight [1084, 2174, 1732, 1142, 1980]. Weighted [1236, 1947, 1191, 1695, 545]. Weighting [909]. Weights [896]. Well [980, 766, 1678, 1278, 733, 742]. Well-behaved [980]. Well-behaving [1278]. Well-Founded [1678]. We’ve [1327]. Wheelchair [1151]. Where [1800, 553]. Which [2196, 451, 1228, 1184]. Whisker [1055]. Whole [1590, 1562]. Wide [865, 1705, 418, 1029]. Wide-Coverage [865]. Width [1483, 1484]. Will [781, 1748]. Win [142]. Win-Win [142]. Windows [1319]. Wireless [1492]. WISDOM [900]. Wise [606]. within [1901, 1210]. Without [604, 219, 1107]. Witness [2253]. Wobbly [2198]. Women [931]. Word [1624, 232, 1639, 2094, 1014]. Word-Class [2094]. word-endings [1014]. WordNet [890, 1017]. Words [1218, 983]. Work [1290, 1328, 1963]. Workbench [377]. Workflow [1705, 631, 375, 1289, 838]. Workflows [1659]. Workload [1338, 127]. Works [443]. Workstation [1369, 95, 1500]. Workstations [1353, 118, 1367, 1315, 133]. World [1722, 1745, 1724, 1705, 418, 1029]. Worlds [324]. Worldwide [1405]. Worm [978]. Wormhole [21, 19, 20, 260]. Wormhole-Routed [19]. WPVM [108]. Write [48, 1797]. Written [2094]. WWW [406, 1692].

x [531, 462, 583, 1840, 1434, 2095]. X- [1840]. X-Ray [583, 2095]. X-Rays [462]. XC6200 [267, 266]. XC6200DS [256]. XC6216 [288]. Xilinx [255, 297, 266]. Xputer [280].

Year [441, 168]. yield [264]. YLC [1325].

Z [672, 1782, 642, 668]. Zernike [1984, 2036]. Zero [2224, 1413, 1603, 746, 722]. Zero-Error [2224]. Zero-knowledge [1413]. zero-one [746]. Zero-Overhead [1603]. zero-reversible [722]. Zn0* [1445]. Zoning [602]. Zoom [2152].

References

Feautrier:1997:BPS

[1] P. Feautrier. Basis of parallel speculative execution. Lecture Notes in Computer Science, 1300:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoare:1997:UTP

[2] T. Hoare and J. He. Unifying theories for parallel programming. Lecture Notes in Computer Science, 1300:15– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermenegildo:1997:API

[3] M. Hermenegildo. Automatic parallelization of irregular and pointer-based computations: Perspectives from logic and constraint programming. Lecture Notes in Computer Science, 1300:31– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1997:SDM

[4] F. Meyer auf der Heide and B. Voecking. Static and dynamic management in networks. Lecture Notes in Computer Science, 1300:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruede:1997:IAH

[5] U. Ruede. Iterative algorithms on high performance architectures. Lecture Notes in Computer Science, 1300: 57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stenstroem:1997:PTA

[6] P. Stenstroem and J. Skeppstedt. A performance tuning approach for shared-memory multiprocessors. Lecture Notes in Computer Science, 1300: 72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malony:1997:STE

[7] A. D. Malony and B. Tourancheau. Support tools and environments. Lecture Notes in Computer Science, 1300: 87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeffery:1997:NVO

[8] C. L. Jeffery, S. G. Dykes, X. Zhang, and G. H. Gonzalez. Nova visualization for optimization of data-parallel programs. Lecture Notes in Computer Science, 1300:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teodorescu:1997:CIT

[9] F. Teodorescu and J. Chassin de Kergommeaux. On correcting the intrusion of tracing non-deterministic programs by software. Lecture Notes in Computer Science, 1300:94–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranzlmueller:1997:UCD

[10] D. Kranzlmueller, S. Grabner, and J. J. Volkert. Using control and data flow analysis for race evaluation. Lecture Notes in Computer Science, 1300:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krantz:1997:CSC

[11] A. T. Krantz and V. S. Sunderam. Client server computing on message passing systems: Experiences with PVM-RPC. Lecture Notes in Computer Science, 1300:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koppler:1997:EUF

[12] R. C. Koppler, G. Kůrka, and J. J. Volkert. Exdasy — A user-friendly and extendable data distribution system. Lecture Notes in Computer Science, 1300:118–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Medeiros:1997:IMH

[13] P. D. Medeiros and J. C. Cunha. Interconnecting multiple heterogeneous parallel application components. Lecture Notes in Computer Science, 1300:128– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delaitre:1997:EIG

[14] T. Delaitre, P. Vekariya, R. Bigeard, and G. R. Justo. EDPEPPS: An integrated graphical toolset for the design and performance evaluation of portable parallel software. Lecture Notes in Computer Science, 1300:140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stellner:1997:LBB

[15] G. Stellner and J. Trinitis. Load balancing based on process migration for MPI. Lecture Notes in Computer Science, 1300:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilar:1997:PMS

[16] J. Aguilar and T. Jimenez. A processors management system for PVM. Lecture Notes in Computer Science, 1300:158–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1997:FPC

[17] H. Zhu and I. Watson. A full program control flow representation for real programs. Lecture Notes in Computer Science, 1300:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1997:RCI

[18] E. Mayr and C. Das. Routing and communication in interconnection networks. Lecture Notes in Computer Science, 1300:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrini:1997:ETE

[19] F. Petrini and M. Vanneschi. Efficient total-exchange in wormholerouted toroidal cubes. Lecture Notes in Computer Science, 1300:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varvarigos:1997:ADB

[20] E. A. Varvarigos and J. P. Lang. An analysis of deflection-based wormhole routing with virtual channels. Lecture Notes in Computer Science, 1300:176– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiIanni:1997:WDP

[21] M. Di Ianni. Wormhole deadlock prediction. Lecture Notes in Computer Science, 1300:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bilardi:1997:BAO

[22] G. Bilardi, B. Codenotti, G. Del Corso, and C. Pinotti. Broadcast and associative operations on fat-trees. Lecture Notes in Computer Science, 1300:196– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikoletseas:1997:FTF

[23] S. Nikoletseas, G. Pantziou, P. Psycharis, and P. Spirakis. On the fault tolerance of fat-trees. Lecture Notes in Computer Science, 1300:208–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deserable:1997:MRT

[24] D. Deserable. Minimal routing in the triangular grid and in a family of related tori. Lecture Notes in Computer Science, 1300:218–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trdlicka:1997:ECK

[25] J. Trdlicka and P. Tvrdik. Embedding complete k-ary trees into 2-dimensional meshes and tori. Lecture Notes in Computer Science, 1300:226–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soch:1997:OGS

[26] M. Soch and P. Tvrdik. Optimal gossip in store-and-forward noncombining 2-D tori. Lecture Notes in Computer Science, 1300:234–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vrto:1997:CMD

[27] I. Vrt’o. Cutwidth of the mesh of dary trees. Lecture Notes in Computer Science, 1300:242–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leppaenen:1997:EER

[28] V. Leppaenen. Embedding and emulation results for static multichannel mesh of optical buses. Lecture Notes in Computer Science, 1300:246–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karaivazoglou:1997:RAP

[29] E. Karaivazoglou and F. Meyer auf der Heide. Routing on asynchronous processor networks. Lecture Notes in Computer Science, 1300:250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kralovic:1997:CSP

[30] R. Kralovic, P. Ruzicka, and D. Stefankovic. The complexity of shortest path and dilation bounded interval routing. Lecture Notes in Computer Science, 1300:258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1997:FPM

[31] P. Damaschke. Finding a pair on a mesh with multiple broadcasting is hard. Lecture Notes in Computer Science, 1300:266–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1997:RPD

[32] B. S. Chlebus, A. Czumaj, and J. F. Sibeyn. Routing on the PADAM: Degrees of optimality. Lecture Notes in Computer Science, 1300:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robert:1997:APH

[33] Y. Robert. Automatic parallelization and high-performance compilers. Lecture Notes in Computer Science, 1300: 283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clauss:1997:HMC

[34] P. Clauss. Handling memory cache policy with integer points countings. Lecture Notes in Computer Science, 1300: 285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok:1997:GTA

[35] Y.-K. Kwok, I. Ahmad, M.-Y. Wu, and W. Shu. A graphical tool for automatic parallelization and scheduling of programs on multiprocessors. Lecture Notes in Computer Science, 1300:294– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1997:ICL

[36] H. Zhu and I. Watson. Identifying critical loads in real programs for decoupled virtual shared memory systems. Lecture Notes in Computer Science, 1300:302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckmann:1997:RID

[37] O. Beckmann and P. H. J. Kelly. Runtime interprocedural data placement optimisation for lazy parallel libraries. Lecture Notes in Computer Science, 1300:306–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wyrzykowski:1997:TMS

[38] R. Wyrzykowski and J. Kanevski. A technique for mapping sparse matrix computations into regular processor arrays. Lecture Notes in Computer Science, 1300:310–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kotlyar:1997:RAC

[39] V. Kotlyar, K. Pingali, and P. Stodghill. A relational approach to the compilation of sparse matrix programs. Lecture Notes in Computer Science, 1300: 318–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loechner:1997:SCM

[40] V. Loechner and C. Mongenet. Solutions to the communication minimization problem for affine recurrence equations. Lecture Notes in Computer Science, 1300:328–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besch:1997:DFC

[41] M. Besch and H. W. Pohl. Dependencefree clustering of shift-invariant data structures. Lecture Notes in Computer Science, 1300:338–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amine:1997:EAD

[42] W. Amine and E. M. Zehendner. Experiences in analyzing data dependences for programs with pointers and structures. Lecture Notes in Computer Science, 1300:342–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kessler:1997:APC

[43] C. W. Kessler. Applicability of program comprehension to sparse matrix computations. Lecture Notes in Computer Science, 1300:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrado:1997:HRI

[44] C. Barrado and J. Labarta. Hamiltonian recurrence for ILP. Lecture Notes in Computer Science, 1300:352– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lefebvre:1997:OSS

[45] V. Lefebvre and P. Feautrier. Optimizing storage size for static control programs in automatic parallelizers. Lecture Notes in Computer Science, 1300:356–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knoop:1997:ODA

[46] J. Knoop and E. Mehofer. Optimal distribution assignment placement. Lecture Notes in Computer Science, 1300:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marzullo:1997:PDA

[47] K. Marzullo and C. Phillips. Parallel and distributed algorithms. Lecture Notes in Computer Science, 1300:377– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:PMS

[48] D. C. Lin, P. W. Dymond, and X. Deng. Parallel merge sort on concurrent-read, owner-write PRAM. Lecture Notes in Computer Science, 1300:379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanders:1997:FMC

[49] P. Sanders, R. Vollmar, and T. Worsch. Feasible models of computation:

Three-dimensionality and energy consumption. Lecture Notes in Computer Science, 1300:384–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sibeyn:1997:SSM

[50] J. F. Sibeyn. Sample sort on meshes. Lecture Notes in Computer Science, 1300:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wachsmann:1997:SMP

[51] A. Wachsmann and R. Wanka. Sorting on a massively parallel system using a library of basic primitives — modeling and experimental results. Lecture Notes in Computer Science, 1300:399– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerbessiotis:1997:PPQ

[52] A. V. Gerbessiotis, C. J. Siniolakis, and A. Tiskin. Parallel priority queue and list contraction: The BSP approach. Lecture Notes in Computer Science, 1300:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marin:1997:PQO

[53] M. Marin. Priority queue operations on EREW-PRAM. Lecture Notes in Computer Science, 1300:417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouge:1997:CRA

[54] L. Bouge, J. Gabarro, X. Messeguer, and N. Schabanel. Concurrent rebalancing of AVL trees: A fine-grained approach. Lecture Notes in Computer Science, 1300:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1997:NAA

[55] W. Liang and G. Havas. NC approximation algorithms for 2-connectivity augmentation in a graph. Lecture Notes in Computer Science, 1300:430– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serna:1997:ASP

[56] M. Serna and F. Xhafa. Approximating scheduling problems in parallel. Lecture Notes in Computer Science, 1300:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1997:NSS

[57] Viet Hai Nguyen. A new staircase separator theorem. Lecture Notes in Computer Science, 1300:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalantery:1997:TTW

[58] N. Kalantery. Tentative time warp. Lecture Notes in Computer Science, 1300:458–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bataller:1997:SDM

[59] J. Bataller and J. Bernabeu. Synchronized DSM models. Lecture Notes in Computer Science, 1300:468–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petit:1997:SES

[60] F. Petit and V. Villain. A spaceefficient and self-stabilizing depth-first token circulation protocol for asynchronous message-passing systems. Lecture Notes in Computer Science, 1300:476–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonoiu:1997:DSS

[61] G. Antonoiu and P. K. Srimani. Distributed self-stabilizing algorithm for minimum spanning tree construction. Lecture Notes in Computer Science, 1300:480–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skubiszewski:1997:PCC

[62] M. Skubiszewski and N. Porteix. Partly-consistent cuts of databases. Lecture Notes in Computer Science, 1300:488–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1997:EAB

[63] D. Agrawal, G. Alonso, A. El Abbadi, and I. Stanoi. Exploiting atomic broadcast in replicated databases. Lecture Notes in Computer Science, 1300:496– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrott:1997:PLC

[64] R. Perrott and G. Agha. Programming languages and concurrent objectoriented programming. Lecture Notes in Computer Science, 1300:507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stewart:1997:SAC

[65] A. Stewart and M. Clint. Synchronising asynchronous communications. Lecture Notes in Computer Science, 1300:511– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Violard:1997:TPE

[66] E. Violard. Typechecking of PEI expressions. Lecture Notes in Computer Science, 1300:521–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loulergue:1997:FPP

[67] F. Loulergue and G. Hains. Functional parallel programming with explicit processes: Beyond SPMD. Lecture Notes in Computer Science, 1300: 530–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LlanaDiaz:1997:TSU

[68] L. F. Llana Diaz and M. Nuñez. Testing semantics for unbounded nondeterminism. Lecture Notes in Computer Science, 1300:538–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oyama:1997:ECF

[69] Y. Oyama, K. Taura, and A. Yonezawa. An efficient compilation framework for languages based on a concurrent process calculus. Lecture Notes in Computer Science, 1300:546–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravara:1997:BTC

[70] A. Ravara and V. Vasconcelos. Behavioural types for a calculus of concurrent objects. Lecture Notes in Computer Science, 1300:554–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosiuczenko:1997:TMS

[71] P. Kosiuczenko. Time in message sequence charts: A formal approach. Lecture Notes in Computer Science, 1300: 562–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebro:1997:IEC

[72] A. J. Nebro, E. Pimentel, and J. M. Troya. Integrating an entry consistency memory model and concurrent object-oriented programming. Lecture Notes in Computer Science, 1300:567– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1997:MDB

[73] R. Breu and R. Grosu. Modeling the dynamic behavior of objects on events, messages and methods. Lecture Notes in Computer Science, 1300: 572–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silva:1997:QDS

[74] A. R. Silva. A quality design solution for object synchronization. Lecture Notes in Computer Science, 1300: 576–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1997:NDS

[75] L. Moreau, D. De Route, and I. Foster. NeXeme: A distributed scheme based on Nexus. Lecture Notes in Computer Science, 1300:581–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briat:1997:ARE

[76] J. Briat, I. Ginzburg, M. Pasin, and B. Plateau. Athapascan runtime: Efficiency for irregular problems. Lecture Notes in Computer Science, 1300:591– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandemir:1997:OCC

[77] M. Kandemir, J. Ramanujam, and A. Choudhary. Optimization of outof-core computations using chain vectors. Lecture Notes in Computer Science, 1300:601–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skillicorn:1997:PMM

[78] D. B. Skillicorn. Programming models and methods. Lecture Notes in Computer Science, 1300:611–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kornerup:1997:PGP

[79] J. Kornerup. Parlists — A generalization of powerlists. Lecture Notes in Computer Science, 1300:614–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Danelutto:1997:SDP

[80] M. Danelutto, F. Pasqualetti, and S. Pelagatti. Skeletons for data parallelism in P3L. Lecture Notes in Computer Science, 1300:619–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serot:1997:EPF

[81] J. Serot. Embodying parallel functional skeletons: An experimental implementation on top of MPI. Lecture Notes in Computer Science, 1300:629–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cole:1997:DCI

[82] M. I. Cole. On dividing and conquering independently. Lecture Notes in Computer Science, 1300:634–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:MTP

[83] Q. Wu, A. J. Field, and P. H. J. Kelly. M-tree: A parallel abstract data type for block-irregular adaptive applications. Lecture Notes in Computer Science, 1300:638–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jay:1997:MCP

[84] C. B. Jay, M. I. Cole, M. Sekanina, and P. Steckler. A monadic calculus for parallel costing of a functional language of arrays. Lecture Notes in Computer Science, 1300:650–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ODonnell:1997:MDP

[85] J. O’Donnell and G. Ruenger. A methodology for deriving parallel programs with a family of parallel abstract machines. Lecture Notes in Computer Science, 1300:662–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winstanley:1997:PDP

[86] N. Winstanley and J. O’Donnell. Parallel distributed programming with Haskell+PVM. Lecture Notes in Computer Science, 1300:670–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nash:1997:PAS

[87] J. M. Nash, P. M. Dew, and J. R. Davy. A parallelisation approach for supporting scalable and portable computing. Lecture Notes in Computer Science, 1300:678–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bungartz:1997:PNA

[88] H.-J. Bungartz. Parallel numerical algorithms. Lecture Notes in Computer Science, 1300:685–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1997:SPS

[89] T. Rauber, G. Ruenger, and C. Scholtes. Scalability of parallel sparse Cholesky factorization. Lecture Notes in Computer Science, 1300:688–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1997:OPA

[90] E. E. Santos. Optimal parallel algorithms for solving tridiagonal linear systems. Lecture Notes in Computer Science, 1300:700–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szularz:1997:RPL

[91] M. Szularz, J. Weston, and M. Clint. Robust parallel Lanczos methods for clustered eigenvalues. Lecture Notes in Computer Science, 1300:710–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kiper:1997:FPS

[92] A. Kiper. A fully parallel symmetric matrix transformation. Lecture Notes in Computer Science, 1300:718– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1997:NEP

[93] T. Rossi and J. Toivanen. Numerical experiments with a parallel fast direct elliptic solver on Cray T3E. Lecture Notes in Computer Science, 1300:722– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haase:1997:NMV

[94] G. Haase. New matrix-by-vector multiplications based on a nonoverlapping domain decomposition data distribution. Lecture Notes in Computer Science, 1300:726–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1997:CBD

[95] W. Huber. A comparison between different parallelization methods on workstation clusters to solve CFD-Problems. Lecture Notes in Computer Science, 1300:734–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frommer:1997:SPS

[96] A. Frommer, T. Lippert, and K. Schilling. Scalable parallel SSOR preconditioning for lattice computations in gauge theories. Lecture Notes in Computer Science, 1300:742–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steihaug:1997:DCA

[97] T. Steihaug and Y. Yalcinkaya. Deteriorating convergence for asynchronous methods on linear least squares problems. Lecture Notes in Computer Science, 1300:750–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stenstroem:1997:PCA

[98] P. Stenstroem and P. Quinton. Parallel computer architecture and image processing. Lecture Notes in Computer Science, 1300:763–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glossner:1997:DJE

[99] C. J. Glossner and S. Vassiliadis. The Delft-Java engine. Lecture Notes in Computer Science, 1300:766–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1997:SIU

[100] D. K. Arvind and S. Sotelo-Salazar. Scheduling instructions with uncertain latencies in asynchronous architectures. Lecture Notes in Computer Science, 1300:771–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomiyasu:1997:CPS

[101] H. Tomiyasu, S. Kusakabe, T. Kawano, and M. Amamiya. Co-processor system design for fine-grain message handling in KUMP/D. Lecture Notes in Computer Science, 1300:779–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:VPC

[102] D. Kim and J. Lee. A virtual-physical on-chip cache for shared memory multiprocessors. Lecture Notes in Computer Science, 1300:789–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kisuki:1997:SVS

[103] T. Kisuki, M. Wakabayashi, J. Yamamoto, and K. Inoue. Shared vs. snoop: Evaluation of cache structure for single-chip multiprocessors. Lecture Notes in Computer Science, 1300:793– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1997:ADH

[104] K. P. Lam and A. Furness. An analytical design of high-speed pixel transformation for object boundary enhancement. Lecture Notes in Computer Science, 1300:807–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleury:1997:KLT

[105] M. Fleury, A. C. Downton, and A. F. Clark. Karhuenen-Loeve transform: An exercise in simple image-processing parallel pipelines. Lecture Notes in Computer Science, 1300:815–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willink:1997:UFC

[106] E. D. Willink, A. V. Shafarenko, and V. B. Muchnick. Use of F-code as a very high level intermediate language for DSP. Lecture Notes in Computer Science, 1300:820–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gentzsch:1997:AHP

[107] W. Gentzsch. Applications of highperformance computing. Lecture Notes in Computer Science, 1300:827–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbosa:1997:EUW

[108] J. G. Barbosa, A. J. Padilha, J.-P. Madier, and T. Neubert. Experiments on using WPVM for industrial visual inspection problems. Lecture Notes in Computer Science, 1300:828–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guidec:1997:OOP

[109] F. Guidec, P. Calegari, and P. Kuonen. Object-oriented parallel software for radio wave propagation simulation in urban environment. Lecture Notes in Computer Science, 1300:832–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouaricha:1997:PPI

[110] A. Bouaricha and S. Mueller. A portable parallel implementation of a 3D semiconductor device simulator. Lecture Notes in Computer Science, 1300:840–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahad:1997:PSL

[111] M. Hahad. A parallel sparse LU decomposition with application to semiconductor device simulation. Lecture Notes in Computer Science, 1300:848– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juganaru:1997:PSQ

[112] M. Juganaru, I. Sakho, C. Maurice, and F. Montheillet. A parallel simulation of a quantitative large-strain polycrystal deformation. Lecture Notes in Computer Science, 1300:852–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marco:1997:PGA

[113] N. Marco, S. Lanteri, J.-A. Desideri, and B. Mantel. Parallel genetic algorithms applied to optimum shape design in aeronautics. Lecture Notes in Computer Science, 1300:856–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muzaferija:1997:PMC

[114] S. Muzaferija, V. Seidl, H. Fogt, and A. Kneer. Parallel multidimensional calculation of steady-state and timedependent flows with combustion. Lecture Notes in Computer Science, 1300: 864–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manke:1997:TLP

[115] J. W. Manke, T. M. Wicks, L. Dadone, and J. E. Hirsh. A two-level parallel strategy for rotorcraft optimization and design. Lecture Notes in Computer Science, 1300:872–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kumar:1997:SLB

[116] V. Kumar, R. Lueling, and C. Roucairol. Scheduling and load balancing. Lecture Notes in Computer Science, 1300:879–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradi:1997:PCL

[117] A. Corradi, L. Leonardi, and F. Zambonelli. Performance comparison of load balancing policies based on a diffusion scheme. Lecture Notes in Computer Science, 1300:882–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1997:ESP

[118] X. Du, Y. Dong, and X. Zhang. Effectively scheduling parallel tasks and communications on networks of workstations. Lecture Notes in Computer Science, 1300:887–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewe:1997:LST

[119] W. Loewe, W. Zimmermann, and J. Eisenbiegler. On linear schedules of task graphs for generalized LogPmachines. Lecture Notes in Computer Science, 1300:895–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boeres:1997:VTS

[120] C. Boeres and V. E. F. Rebello. Versatile task scheduling of binary trees for realistic machines. Lecture Notes in Computer Science, 1300:913–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galtier:1997:LBI

[121] J. Galtier. Load balancing issues in the prepartitioning method. Lecture Notes in Computer Science, 1300:922– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Das:1997:DNL

[122] S. K. Das, D. J. Harvey, and R. Biswas. Design of novel load-balancing algorithms with implementations on an IBM SP2. Lecture Notes in Computer Science, 1300:937–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schloegel:1997:RAM

[123] K. Schloegel, G. Karypis, and V. Kumar. Repartitioning of adaptive meshes: Experiments with multilevel diffusion. Lecture Notes in Computer Science, 1300:945–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAmore:1997:ERD

[124] F. D’Amore, L. Becchetti, S. L. Bezrukov, and A. Marchetti-Spaccamela. On the embedding of refinements of 2-dimensional grids. Lecture Notes in Computer Science, 1300:950–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehring:1997:DPD

[125] J. Gehring. Dynamic program description as a basis for runtime optimization. Lecture Notes in Computer Science, 1300:958–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bode:1997:PEP

[126] A. Bode and J. Dongarra. Performance evaluation and prediction. Lecture Notes in Computer Science, 1300:969– 970, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1997:WAC

[127] J. Simon, M. Vieth, and R. Weicker. Workload analysis of computation intensive tasks: Case study on SPEC CPU95 benchmarks. Lecture Notes in Computer Science, 1300:971–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strohmaier:1997:SPM

[128] E. Strohmaier. Statistical performance modeling: Case study of the NPB 2.1 results. Lecture Notes in Computer Science, 1300:985–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouras:1997:GPM

[129] C. J. Bouras, J. D. Garofalakis, P. G. Spirakis, and V. D. Triantafillou. A general performance model for multistage interconnection networks. Lecture Notes in Computer Science, 1300: 993–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pham:1997:SRA

[130] C. D. Pham, J. Essmeyer, and S. Fdida. Simulation of a routing algorithm using distributed simulation techniques. Lecture Notes in Computer Science, 1300:1001–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Getov:1997:MPP

[131] V. Getov, E. Hernandez, and T. Hey. Message-passing performance of parallel computers. Lecture Notes in Computer Science, 1300:1009–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1997:PMP

[132] E. D. Moreno, S. T. Kofuji, and M. H. Cintra. Prefetching and multithreading performance in bus-based multiprocessors with Petri nets. Lecture Notes in Computer Science, 1300:1017–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:SFT

[133] J. Smith. On synchronisation in faulttolerant data and compute intensive programs over a network of workstations. Lecture Notes in Computer Science, 1300:1025–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pahud:1997:PAP

[134] M. Pahud, F. Guidec, and T. Cornu. Performance analysis of a parallel program for wave propagation simulation. Lecture Notes in Computer Science, 1300:1030–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lundberg:1997:BMC

[135] L. Lundberg. Bounding the minimal completion time of static mappings of multithreaded Solaris programs. Lecture Notes in Computer Science, 1300:1034–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1997:ILP

[136] D. K. Arvind. Instruction-level parallelism. Lecture Notes in Computer Science, 1300:1041–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipasti:1997:PPV

[137] M. H. Lipasti and J. P. Shen. The performance potential of value and dependence prediction. Lecture Notes in Computer Science, 1300:1043–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:ETL

[138] J.-B. Lee, W. Sung, and S.-M. Moon. An enhanced two-level adaptive multiple branch prediction for superscalar processors. Lecture Notes in Computer Science, 1300:1053–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filho:1997:ESD

[139] E. M. C. Filho and E. S. T. Fernandes. The effect of the speculation depth on the performance of superscalar architectures. Lecture Notes in Computer Science, 1300:1061–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnas:1997:SEU

[140] D. L. Parnas. Software engineering: An unconsummated marriage. Lecture Notes in Computer Science, 1301:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rushby:1997:SS

[141] J. Rushby. Subtypes for specifications. Lecture Notes in Computer Science, 1301:4–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehm:1997:DMA

[142] B. Boehm, J. Kwan, and R. Madachy. Developing multimedia applications with the win-win spiral model. Lecture Notes in Computer Science, 1301: 20–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maibaum:1997:WWT

[143] T. S. E. Maibaum. What we teach software engineers in the university: Do we take engineering seriously? Lecture Notes in Computer Science, 1301: 40–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uhl:1997:WWE

[144] J. Uhl. What we expect from software engineers in the industry. Lecture Notes in Computer Science, 1301: 51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kemmerer:1997:SID

[145] R. A. Kemmerer. Security issues in distributed software. Lecture Notes in Computer Science, 1301:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Medvidovic:1997:FCC

[146] N. Medvidovic and R. N. Taylor. A framework for classifying and comparing architecture description languages. Lecture Notes in Computer Science, 1301:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naumovich:1997:ASA

[147] G. Naumovich, G. S. Avrunin, L. A. Clarke, and L. J. Osterweil. Applying static analysis to software architectures. Lecture Notes in Computer Science, 1301:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijler:1997:MDP

[148] T. D. Meijler, S. Demeyer, and R. Engel. Making design patterns explicit in FACE — A framework adaptive composition environment. Lecture Notes in Computer Science, 1301:94–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1997:TLM

[149] G. Allen, L. Davies, G. Lindmark, and E.-A. Karlsson. TTM15 — A large multi-site improvement project. Lecture Notes in Computer Science, 1301: 111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joeris:1997:CMN

[150] G. Joeris. Change management needs integrated process and configuration management. Lecture Notes in Computer Science, 1301:125–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutton:1997:DNG

[151] S. M. Sutton and L. J. Osterweil. The design of a next-generation process language. Lecture Notes in Computer Science, 1301:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDerHoek:1997:SRM

[152] A. Van Der Hoek, R. S. Hall, D. Heimbigner, and A. L. Wolf. Software release management. Lecture Notes in Computer Science, 1301:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pohl:1997:CAP

[153] K. Pohl and K. Weidenhaupt. A contextual approach for process-integrated tools. Lecture Notes in Computer Science, 1301:176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jahnke:1997:GFR

[154] J. H. Jahnke, W. Schaefer, and A. Zuendorf. Generic fuzzy reasoning nets as a basis for reverse engineering relational database applications. Lecture Notes in Computer Science, 1301:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alborghetti:1997:PAS

[155] A. Alborghetti, A. Gargantini, and A. Morzenti. Providing automated support to deductive analysis of time critical systems. Lecture Notes in Computer Science, 1301:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1997:VLP

[156] Shing Chi Cheung, D. Giannakopoulou, and J. Kramer. Verification of liveness properties using compositional reachability analysis. Lecture Notes in Computer Science, 1301:227–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwyer:1997:MCG

[157] M. B. Dwyer, V. Carr, and L. Hines. Model checking graphical user interfaces using abstractions. Lecture Notes in Computer Science, 1301:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wood:1997:CCS

[158] M. Wood, M. Roper, A. Brooks, and J. Miller. Comparing and combining software defect detection techniques: A replicated empirical study. Lecture Notes in Computer Science, 1301:262– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1997:ECP

[159] C. Ebert. Experiences with criticality predictions in software development. Lecture Notes in Computer Science, 1301:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Land:1997:VDD

[160] L. Pek Wee Land, C. Sauer, and R. Jeffery. Validating the defect detection performance advantage of group designs for software reviews: Report of a laboratory experiment using program code. Lecture Notes in Computer Science, 1301:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desharnais:1997:ISS

[161] J. Desharnais, R. Khedri, M. Frappier, and A. Mili. Integration of sequential scenarios. Lecture Notes in Computer Science, 1301:310–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1997:VOA

[162] G. Engels, R. Heckel, G. Taentzer, and H. Ehrig. A view-oriented approach to system modelling based on graph transformation. Lecture Notes in Computer Science, 1301:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosenblum:1997:DFI

[163] D. S. Rosenblum and A. L. Wolf. A design framework for Internet-scale event observation and notification. Lecture Notes in Computer Science, 1301:344– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodik:1997:RDF

[164] R. Bodik, R. Gupta, and M. L. Soffa. Refining data flow information using infeasible paths. Lecture Notes in Computer Science, 1301:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forgacs:1997:FTP

[165] R. Forgacs and A. Bertolino. Feasible test path selection by principal slicing. Lecture Notes in Computer Science, 1301:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devanbu:1997:CVT

[166] P. Devanbu and S. G. Stubblebine. Cryptographic verification of test coverage claims. Lecture Notes in Computer Science, 1301:395–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:CDM

[167] Y.-F. R. Chen, E. R. Gansner, and E. Koutsofios. A C++ data model supporting reachability analysis and dead code detection. Lecture Notes in Computer Science, 1301:414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reps:1997:UPP

[168] T. Reps, T. Ball, M. Das, and J. Larus. The use of program profiling for software maintenance with applications to the year 2000 problem. Lecture Notes in Computer Science, 1301:432–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heimdahl:1997:RSH

[169] M. P. E. Heimdahl and M. W. Whalen. Reduction and slicing of hierarchical state machines. Lecture Notes in Computer Science, 1301:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuetze:1997:PBA

[170] M. Schuetze, J. P. Riegel, and G. Zimmermann. A pattern-based application generator for building simulation. Lecture Notes in Computer Science, 1301:468–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ducasse:1997:ECT

[171] S. Ducasse and T. Richner. Executable connectors: Towards reusable design elements. Lecture Notes in Computer Science, 1301:483–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Picco:1997:ECM

[172] G. P. Picco, G.-C. Roman, and P. J. McCann. Expressing code mobility in mobile UNITY. Lecture Notes in Computer Science, 1301:500–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlsson:1997:IDA

[173] E.-A. Karlsson and L. Taxen. Incremental development for AXE 10. Lecture Notes in Computer Science, 1301: 519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1997:TTP

[174] D. E. Perry and L. G. Votta. The tale of two projects. Lecture Notes in Computer Science, 1301:521–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obbink:1997:ASA

[175] J. H. Obbink. Analysis of software architectures in high and low volume electronic systems. Lecture Notes in Computer Science, 1301:523–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Idri:1997:TAC

[176] A. Idri, B. Griech, and A. El Irak. Towards an adaptation of the CO-COMO cost model to the software measurement theory. Lecture Notes in Computer Science, 1301:525–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dini:1997:FSA

[177] P. Dini, A. Belkhelladi, and W. L. Melo. Formalizing software architectures: An industrial experience. Lecture Notes in Computer Science, 1301: 527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramalingam:1997:SER

[178] G. Ramalingam. On sparse evaluation representations. Lecture Notes in Computer Science, 1302:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shapiro:1997:EPP

[179] M. Shapiro and S. Horwitz. The effects of the precision of pointer analysis. Lecture Notes in Computer Science, 1302:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waddell:1997:FEP

[180] O. Waddell and R. K. Dybvig. Fast and effective procedure inlining. Lecture Notes in Computer Science, 1302: 35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagnara:1997:SSR

[181] R. Bagnara, P. M. Hill, and E. Zaffanella. Set sharing is redundant for pair-sharing. Lecture Notes in Computer Science, 1302:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codish:1997:AAS

[182] M. Codish, V. Lagoon, and F. Bueno. An algebraic approach to sharing analysis of logic programs. Lecture Notes in Computer Science, 1302:68–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scozzari:1997:LOG

[183] F. Scozzari. Logical optimality of groundness analysis. Lecture Notes in Computer Science, 1302:83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yi:1997:TCE

[184] K. Yi and S. Ryu. Toward a costeffective estimation of uncaught exceptions in SML programs. Lecture Notes in Computer Science, 1302:98– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faehndrich:1997:PAU

[185] M. Faehndrich and A. Aiken. Program analysis using mixed term and set constraints. Lecture Notes in Computer Science, 1302:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devienne:1997:SBA

[186] P. Devienne, J. M. Talbot, and S. Tison. Set-based analysis for logic programming and tree automata. Lecture Notes in Computer Science, 1302:127– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zartmann:1997:DAI

[187] F. Zartmann. Denotational abstract interpretation of functional logic programs. Lecture Notes in Computer Science, 1302:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Speirs:1997:TAM

[188] C. Speirs, Z. Somogyi, and H. Sondergaard. Termination analysis for mercury. Lecture Notes in Computer Science, 1302:160–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boigelot:1997:PQ

[189] B. Boigelot, P. Godefroid, B. Willems, and P. Wolper. The power of QDDs. Lecture Notes in Computer Science, 1302:172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Igarashi:1997:TBA

[190] A. Igarashi and N. Kobayashi. Typebased analysis of communication for concurrent programming languages. Lecture Notes in Computer Science, 1302:187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodei:1997:TCA

[191] C. Bodei and C. Priami. True concurrency via abstract interpretation. Lecture Notes in Computer Science, 1302: 202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mackie:1997:SAI

[192] I. Mackie. Static analysis of interaction nets for distributed implementations. Lecture Notes in Computer Science, 1302:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1997:TDF

[193] S. Jagannathan, S. Weeks, and A. Wright. Type-directed flow analysis for typed intermediate languages. Lecture Notes in Computer Science, 1302: 232–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossin:1997:EFA

[194] C. Mossin. Exact flow analysis. Lecture Notes in Computer Science, 1302: 250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frey:1997:SSI

[195] A. Frey. Satisfying subtype inequalities in polynomial space. Lecture Notes in Computer Science, 1302:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granger:1997:SAC

[196] P. Granger. Static analyses of congruence properties on rational numbers. Lecture Notes in Computer Science, 1302:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hornof:1997:ESR

[197] L. Hornof, J. Noye, and C. Consel. Effective specialization of realistic programs via use sensitivity. Lecture Notes in Computer Science, 1302:293– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uh:1997:CCB

[198] G.-R. Uh and D. B. Whalley. Coalescing conditional branches into efficient indirect jumps. Lecture Notes in Computer Science, 1302:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauburger:1997:ATA

[199] J. Brauburger. Automatic termination analysis for partial functions using polynomial orderings. Lecture Notes in Computer Science, 1302:330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panitz:1997:TAP

[200] S. E. Panitz and M. Schmidt-Schauss. TEA: Automatically proving termination of programs in a non-strict higherorder functional language. Lecture Notes in Computer Science, 1302:345– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emelianov:1997:SAM

[201] P. G. Emelianov and D. E. Baburin. Semantic analyzer of Modula-programs. Lecture Notes in Computer Science, 1302:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldan:1997:AII

[202] D. Baldan and G. File. Abstract interpretation for improving WAM code. Lecture Notes in Computer Science, 1302:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correnson:1997:AGF

[203] L. Correnson, E. Duris, D. Parigot, and G. Roussel. Attribute grammars and functional programming deforestation. Lecture Notes in Computer Science, 1302:365–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tzolovski:1997:DDA

[204] S. Tzolovski. Data dependences as abstract interpretation. Lecture Notes in Computer Science, 1302:366–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marriott:1997:AIT

[205] K. Marriott. Abstract interpretation: A theory of approximate computation. Lecture Notes in Computer Science, 1302:367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeCharlier:1997:DLB

[206] B. Le Charlier and P. Flener. On the desirable link between theory and practice in abstract interpretation. Lecture Notes in Computer Science, 1302:379– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousot:1997:AIB

[207] P. Cousot. Abstract interpretation based static analysis parameterized by semantics. Lecture Notes in Computer Science, 1302:388–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1997:CAI

[208] L. Harrison. Can abstract interpretation become a mainstream compiler technology? Lecture Notes in Computer Science, 1302:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1997:CAI

[209] N. D. Jones. Combining abstract interpretation and partial evaluation. Lecture Notes in Computer Science, 1302: 396–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryder:1997:PCT

[210] B. G. Ryder. Practical compile-time analysis. Lecture Notes in Computer Science, 1302:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohn:1997:QSR

[211] A. G. Cohn. Qualitative spatial representation and reasoning techniques. Lecture Notes in Computer Science, 1303:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konolige:1997:CLR

[212] K. Konolige. COLBERT: A language for reactive control in sapphira. Lecture Notes in Computer Science, 1303: 31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langley:1997:MLA

[213] P. Langley. Machine learning for adaptive user interfaces. Lecture Notes in Computer Science, 1303:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerberding:1997:SIP

[214] S. Gerberding and B. Pientka. Structured incremental proof planning. Lecture Notes in Computer Science, 1303:

63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolbe:1997:MSL

[215] T. Kolbe and S. Glesner. Manysorted logic in a learning theorem prover. Lecture Notes in Computer Science, 1303:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuehn:1997:RH

[216] M. Kuehn. Rigid hypertableaux. Lecture Notes in Computer Science, 1303: 87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1997:MMG

[217] W. Lu. Minimal model generation based on E-hyper tableaux. Lecture Notes in Computer Science, 1303:99– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melis:1997:EAI

[218] E. Melis and J. Whittle. External analogy in inductive theorem proving. Lecture Notes in Computer Science, 1303:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerber:1997:MPR

[219] M. Kerber and M. Kohlhase. Mechanising partiality without re-implementation. Lecture Notes in Computer Science, 1303:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandner:1997:LPD

[220] E. Sandner. From linear proofs to direct logic with exponentials. Lecture Notes in Computer Science, 1303:135– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strecker:1997:IEP

[221] M. Strecker and M. Sorea. Integrating an equality prover into a software development system based on type theory. Lecture Notes in Computer Science, 1303:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruenwald:1997:CNT

[222] P. Gruenwald. Causation and nonmonotonic temporal reasoning. Lecture Notes in Computer Science, 1303:159– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1997:LQM

[223] D. Basin, S. Matthews, and L. Vigano. Labelled quantified modal logics. Lecture Notes in Computer Science, 1303: 171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messing:1997:DDR

[224] B. Messing. Defining decision rules in signed Horn clauses. Lecture Notes in Computer Science, 1303:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berendt:1997:MSD

[225] B. Berendt. Modeling subjective distances. Lecture Notes in Computer Science, 1303:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eschenbach:1997:AAS

[226] C. Eschenbach and L. Kulik. An axiomatic approach to the spatial relations underlying “left”-“right” and “in front of”-“behind”. Lecture Notes in Computer Science, 1303:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roehrig:1997:RPQ

[227] R. Roehrig. Representation and processing of qualitative orientation knowledge. Lecture Notes in Computer Science, 1303:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1997:CQA

[228] B. Ludwig. A contribution to the question of authenticity of rhesus using part-of-speech-tagging. Lecture Notes in Computer Science, 1303:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schacht:1997:MOM

[229] S. Schacht and U. Hahn. Making objects more knowledgeable. Lecture Notes in Computer Science, 1303:243– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnattinger:1997:CAC

[230] K. Schnattinger and U. Hahn. Constraining the acquisition of concepts by the quality of heterogeneous evidence. Lecture Notes in Computer Science, 1303:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staab:1997:CA

[231] S. Staab and U. Hahn. Conceptualizing adjectives. Lecture Notes in Computer Science, 1303:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1997:PBT

[232] H. Weber, J. Spilker, and G. Goerz. Parsing N best trees from a word lattice. Lecture Notes in Computer Science, 1303:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burgard:1997:FGB

[233] W. Burgard, D. Fox, and D. Henning. Fast grid-based position tracking for mobile robots. Lecture Notes in Computer Science, 1303:289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haag:1997:IIS

[234] M. Haag, W. Theilmann, K. H. Schaefer, and H.-H. Nagel. Integration of image sequence evaluation and fuzzy metric temporal logic programming. Lecture Notes in Computer Science, 1303: 301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoelldobler:1997:DCA

[235] S. Hoelldobler, Y. Kalinke, and H. Lehmann. Designing a counter: Another case study of dynamics and activation landscapes in recurrent networks. Lecture Notes in Computer Science, 1303:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bamberger:1997:CDE

[236] S. K. Bamberger. Cooperating diagnostic expert systems to solve complex diagnosis tasks. Lecture Notes in Computer Science, 1303:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinmann:1997:TSV

[237] O. Steinmann, A. Strohmaier, and T. Stfitzle. Tabu search vs. random walk. Lecture Notes in Computer Science, 1303:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strohmaier:1997:MFN

[238] A. Strohmaier. Multi-flip networks: Parallelizing GenSAT. Lecture Notes in Computer Science, 1303:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weis:1997:RAA

[239] T. Weis. Resource-adaptive action planning in a dialogue system for repair support. Lecture Notes in Computer Science, 1303:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artosi:1997:MCF

[240] A. Artosi, P. Cattabriga, and G. Governatori. A modal computational framework for default reasoning. Lecture Notes in Computer Science, 1303:373– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheikhrouhou:1997:PDP

[241] L. Cheikhrouhou. Planning diagonalization proofs. Lecture Notes in Computer Science, 1303:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edelkamp:1997:STA

[242] S. Edelkamp. Suffix tree automata in state space search. Lecture Notes in Computer Science, 1303:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fronhoefer:1997:CCC

[243] B. Fronhoefer. Connection cutting for contraction free logic. Lecture Notes in Computer Science, 1303:385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1997:APE

[244] D. Gabbay, R. Nossum, and M. Thielscher. Agents in proactive environments. Lecture Notes in Computer Science, 1303:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartrumpf:1997:RLR

[245] S. Hartrumpf and M. Schulz. Reducing lexical redundancy by augmenting conceptual knowledge. Lecture Notes in Computer Science, 1303:393–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nilson:1997:GUI

[246] J. Nilson and G. Koekai. A graphical user interface for an ECG classifier system. Lecture Notes in Computer Science, 1303:397–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stojanovic:1997:ESA

[247] N. Stojanovic, L. Stoiljkovic, D. Milenovic, and V. Stoiljkovic. Expert system in additional finishing. Lecture Notes in Computer Science, 1303:401–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderTorre:1997:RAE

[248] L. W. N. Van der Torre and Y.- H. Tan. Reasoning about exceptions. Lecture Notes in Computer Science, 1303:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walischewski:1997:LIL

[249] H. Walischewski. Learning and interpretation of the layout of structured documents. Lecture Notes in Computer Science, 1303:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faura:1997:MDR

[250] J. Faura, J. M. Moreno, M. A. Aguirre, and P. Van Duong. Multicontext dynamic reconfiguration and real time probing on a novel mixed signal programmable device with on-chip microprocessor. Lecture Notes in Computer Science, 1304:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyazaki:1997:COF

[251] T. Miyazaki, A. Takahara, M. Katayama, and T. Murooka. CAD-oriented FPGA and dedicated CAD system for telecommunications. Lecture Notes in Computer Science, 1304:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leeser:1997:RTD

[252] M. Leeser, W. M. Meleis, M. M. Vai, and P. Zavracky. Rothko: A three dimensional FPGA architecture, its fabrication, and design tools. Lecture Notes in Computer Science, 1304:21– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McGregor:1997:EDC

[253] G. McGregor and P. Lysaght. Extending dynamic circuit switching to meet the challenges of new FPGA architectures. Lecture Notes in Computer Science, 1304:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1997:PEF

[254] D. Robinson, P. Lysaght, G. McGregor, and H. Dick. Performance evaluation of a full speed PCI initiator and target subsystem using FP-GAs. Lecture Notes in Computer Science, 1304:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Do:1997:IPM

[255] T.-T. Do, H. Kropp, M. Schwiegershausen, and P. Pirsch. Implementation of pipelined multipliers on Xilinx FPGAs. Lecture Notes in Computer Science, 1304:51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nisbet:1997:XDS

[256] S. Nisbet and S. A. Guccione. The XC6200DS development system. Lecture Notes in Computer Science, 1304: 61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boemo:1997:TMF

[257] E. Boemo and S. Lopez-Buedo. Thermal monitoring on FPGAs using ringoscillators. Lecture Notes in Computer Science, 1304:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kostarnov:1997:RAL

[258] I. Kostarnov, S. Morley, J. Osmany, and C. Solomon. A reconfigurable approach to low cost media processing. Lecture Notes in Computer Science, 1304:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mackinlay:1997:RFP

[259] P. I. Mackinlay, P. Y. K. Cheung, W. Luk, and R. Sandiford. Riley-2: A flexible platform for codesign and dynamic reconfigurable computing research. Lecture Notes in Computer Science, 1304:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahne:1997:SSW

[260] B. Kahne and P. Athanas. Stream synthesis for a wormhole run-time reconfigurable platform. Lecture Notes in Computer Science, 1304:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luk:1997:PMV

[261] W. Luk, N. Shirazi, S. R. Guo, and P. Y. K. Cheung. Pipeline morphing and virtual pipelines. Lecture Notes in Computer Science, 1304:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rising:1997:PGC

[262] B. Rising, M. Van Daalen, P. Burge, and J. Shawe-Taylor. Parallel graph colouring using FPGAs. Lecture Notes in Computer Science, 1304:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diessel:1997:RTC

[263] O. Diessel and H. ElGindy. Run-time compaction of FPGA designs. Lecture Notes in Computer Science, 1304:131– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmert:1997:PRF

[264] J. M. Emmert and D. Bhatia. Partial reconfiguration of FPGA mapped designs with applications to fault tolerance and yield enhancement. Lecture Notes in Computer Science, 1304:141– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonard:1997:CSP

[265] J. Leonard and W. H. Mangione-Smith. A case study of partially evaluated hardware circuits: Key-specific DES. Lecture Notes in Computer Science, 1304:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Payne:1997:RTP

[266] R. Payne. Run-time parameterised circuits for the Xilinx XC6200. Lecture Notes in Computer Science, 1304:161– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brebner:1997:AIS

[267] G. Brebner. Automatic identification of swappable logic units in XC6200 circuitry. Lecture Notes in Computer Science, 1304:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysaght:1997:TES

[268] P. Lysaght. Towards an expert system for a priori estimation of reconfiguration latency in dynamically reconfigurable logic. Lecture Notes in Computer Science, 1304:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hutchings:1997:ERT

[269] B. L. Hutchings. Exploiting reconfigurability through domain-specific systems. Lecture Notes in Computer Science, 1304:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Servit:1997:TMB

[270] M. Z. Servit and K. Yi. Technology mapping by binate covering. Lecture Notes in Computer Science, 1304:203– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betz:1997:VNP

[271] V. Betz and J. Rose. VPR: A new packing, placement and routing tool for FPGA research. Lecture Notes in Computer Science, 1304:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inuani:1997:TMH

[272] M. K. Inuani and J. Saul. Technology mapping of heterogeneous LUT-based FPGAs. Lecture Notes in Computer Science, 1304:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feske:1997:TDF

[273] K. Feske, S. Mulka, M. Koegst, and G. Elst. Technology-driven FSM partitioning for synthesis of large sequential circuits targeting lookup-table based FPGAs. Lecture Notes in Computer Science, 1304:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:TML

[274] X. Lin, E. Dagless, and A. Lu. Technology mapping of LUT based FPGAs for delay optimisation. Lecture Notes in Computer Science, 1304:245–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acock:1997:AMA

[275] S. J. B. Acock and K. R. Dimond. Automatic mapping of algorithms onto multiple FPGA-SRAM modules. Lecture Notes in Computer Science, 1304: 255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maunder:1997:FHS

[276] R. B. Maunder, Z. A. Salcic, and G. G. Coghill. FPLD HDL synthesis employing high-level evolutionary algorithm optimisation. Lecture Notes in Computer Science, 1304:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chichkov:1997:HSP

[277] A. V. Chichkov and C. B. Almeida. A hardware/software partitioning algorithm for custom computing machines. Lecture Notes in Computer Science, 1304:274–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lechner:1997:JER

[278] E. Lechner and S. A. Guccione. The Java environment for reconfigurable computing. Lecture Notes in Computer Science, 1304:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartenstein:1997:DSI

[279] R. W. Hartenstein, J. Becker, M. Herz, and U. Nageldinger. Data scheduling to increase performance of parallel accelerators. Lecture Notes in Computer Science, 1304:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kress:1997:OSC

[280] R. Kress, R. W. Hartenstein, and U. Nageldinger. An operating system for custom computing machines based on the xputer paradigm. Lecture Notes in Computer Science, 1304: 304–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dandalis:1997:FPI

[281] A. Dandalis and V. K. Prasanna. Fast parallel implementation of DFT using configurable devices. Lecture Notes in Computer Science, 1304:314–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenfield:1997:EFP

[282] D. Greenfield, C. Crome, M. S. Won, and D. Amos. Enhancing fixed point DSP processor performance by adding CPLDs as coprocessing elements. Lecture Notes in Computer Science, 1304: 324–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shand:1997:CSA

[283] M. Shand. A case study of algorithm implementation in reconfigurable hardware and software. Lecture Notes in Computer Science, 1304:333–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1997:RDL

[284] A. S. Chaudhuri, P. Y. K. Cheung, and W. Luk. A reconfigurable datalocalised array for morphological algorithms. Lecture Notes in Computer Science, 1304:344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bramer:1997:VRA

[285] B. Bramer, D. Chauham, M. K. Ibrahim, and A. Aggoun. Virtual radix array processors (V-RaAP). Lecture Notes in Computer Science, 1304:354– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mathews:1997:FIM

[286] T. Mathews, S. C. Gibb, L. E. Turner, and P. J. W. Graumann. An FPGA implementation of a matched filter detector for spread spectrum communications systems. Lecture Notes in Computer Science, 1304:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teerapanyawatt:1997:NPC

[287] S. Teerapanyawatt and K. Athikulwongse. An NTSC and PAL closed caption processor. Lecture Notes in Computer Science, 1304:374–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kean:1997:MSR

[288] T. Kean and A. Duncan. A 800 mpixel/ sec reconfigurable image correlator on XC6216. Lecture Notes in Computer Science, 1304:382–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lisa:1997:RCP

[289] F. Lisa, F. Cuadrado, D. Rexachs, and J. Carrabina. A reconfigurable coprocessor for a PCI-based real time computer vision system. Lecture Notes in Computer Science, 1304:392–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunn:1997:RTS

[290] P. Dunn and P. Corke. Real-time stereopsis using FPGAs. Lecture Notes in Computer Science, 1304:400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jong:1997:FID

[291] C. C. Jong, Y. Y. H. Lam, and L. S. Ng. FPGA implementation of a digital IQ demodulator using VHDL. Lecture Notes in Computer Science, 1304: 410–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Page:1997:HCC

[292] I. Page. Hardware compilation, configurable platforms and ASICs for selfvalidating sensors. Lecture Notes in Computer Science, 1304:418–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1997:PRV

[293] S. Singh, J. Patterson, J. Burns, and M. Dales. Postscript[TM] rendering with virtual hardware. Lecture Notes in Computer Science, 1304:428–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadzic:1997:PPF

[294] I. Hadzic and J. M. Smith. P4: A platform for FPGA implementation of protocol boosters. Lecture Notes in Computer Science, 1304:438–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abramovici:1997:SRH

[295] M. Abramovici and D. Saab. Satisfiability on reconfigurable hardware. Lecture Notes in Computer Science, 1304: 448–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jebelean:1997:ACA

[296] T. Jebelean. Auto-configurable array for GCD computation. Lecture Notes in Computer Science, 1304:457–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurent:1997:SVA

[297] B. Laurent, G. Bosco, and G. Saucier. Structural versus algorithmic approaches for efficient adders on Xilinx 5200 FPGA. Lecture Notes in Computer Science, 1304:462–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tisserand:1997:FIR

[298] A. Tisserand and M. Dimmler. FPGA implementation of real-time digital controllers using on-line arithmetic. Lecture Notes in Computer Science, 1304:472–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hollstein:1997:PEF

[299] T. Hollstein, A. Kirschbaum, and M. Glesner. A prototyping environment for fuzzy controllers. Lecture Notes in Computer Science, 1304:482– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nukata:1997:RSD

[300] K. Nukata, Y. Shibata, H. Amano, and Y. Anzai. A reconfigurable sensordata processing system for personal robots. Lecture Notes in Computer Science, 1304:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bedingfield:1997:SPB

[301] S. Bedingfield, S. Huxford, and Y. Cheung. Simulating pricing behaviours using a genetic algorithm. Lecture Notes in Computer Science, 1305:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devine:1997:BIC

[302] P. Devine and R. Paton. Biologically inspired computational ecologies. Lecture Notes in Computer Science, 1305: 11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edmonds:1997:MBR

[303] B. Edmonds and S. Moss. Modelling bounded rationality using evolutionary techniques. Lecture Notes in Computer Science, 1305:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalmykov:1997:ATE

[304] V. L. Kalmykov. The abstract theory of evolution of the living. Lecture Notes in Computer Science, 1305:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albright:1997:EAS

[305] H. P. Albright and J. P. Ignizio. An evolutionary algorithm for single objective nonlinear constrained optimization problems. Lecture Notes in Computer Science, 1305:55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

East:1997:RS

[306] I. East. On recombinative sampling. Lecture Notes in Computer Science, 1305:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirst:1997:EMP

[307] T. Hirst. The evolution of mutation, plasticity and culture in cyclically changing environments. Lecture Notes in Computer Science, 1305:75– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirst:1997:STL

[308] T. Hirst. On the structure and transformation of landscapes. Lecture Notes in Computer Science, 1305:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whitley:1997:IMG

[309] D. Whitley, S. Rana, and R. Heckendorn. Island model genetic algorithms and linearly separable problems. Lecture Notes in Computer Science, 1305:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacob:1997:EVP

[310] P. Jacob and A. Ball. Empirical validation of the performance of a class of transient detector. Lecture Notes in Computer Science, 1305:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennedy:1997:CED

[311] H. Kennedy, C. Chinniah, P. Bradbeer, and L. Morss. The construction and evaluation of decision trees: a comparison of evolutionary and concept learning methods. Lecture Notes in Computer Science, 1305:147–162, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poli:1997:PDG

[312] R. Poli. Parallel distributed genetic programming applied to the evolution of natural language recognisers. Lecture Notes in Computer Science, 1305: 163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Langdon:1997:SPM

[313] B. Langdon. Scheduling planned maintenance of the South Wales Region of the National Grid. Lecture Notes in Computer Science, 1305:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIlhagga:1997:SGS

[314] M. McIlhagga. Solving generic scheduling problems with a distributed genetic algorithm. Lecture Notes in Computer Science, 1305:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ross:1997:DSE

[315] P. Ross and A. Tuson. Directing the search of evolutionary and neighbourhood-search optimisers for the flowshop sequencing problem with an idle-time heuristic. Lecture Notes in Computer Science, 1305:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Savic:1997:MGA

[316] D. Savic, G. Walters, and M. Schwab. Multiobjective genetic algorithms for pump scheduling in water supply. Lecture Notes in Computer Science, 1305: 227–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shaw:1997:URP

[317] J. Shaw and P. Fleming. Use of rules and preferences for schedule builders in genetic algorithms for production scheduling. Lecture Notes in Computer Science, 1305:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baron:1997:VBA

[318] P. Baron, R. Fisher, A. Sherlock, and F. Mill. A voxel based approach to evolutionary shape optimisation. Lecture Notes in Computer Science, 1305:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deakin:1997:PTN

[319] A. Deakin and D. Yates. Phase transition networks: A modelling technique supporting the evolution of autonomous agents’ tactical and operational activities. Lecture Notes in Computer Science, 1305:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parmee:1997:EAA

[320] I. Parmee and M. Beck. An evolutionary, agent-assisted strategy for conceptual design space decomposition. Lecture Notes in Computer Science, 1305:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seredynski:1997:TSU

[321] F. Seredynski. Task scheduling with use of classifier systems. Lecture Notes in Computer Science, 1305:287–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leung:1997:CAC

[322] C. H. C. Leung and W. W. S. So. Characteristics and architectural components of visual information systems. Lecture Notes in Computer Science, 1306:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:AVI

[323] S.-K. Chang. Active visual information systems. Lecture Notes in Computer Science, 1306:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunii:1997:TDM

[324] T. L. Kunii, J. Ma, and R. Huang. Towards direct mapping between visual information worlds and real worlds. Lecture Notes in Computer Science, 1306:27–??, 1997. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Jin:1997:VDP

[325] J. S. Jin, H. Greenfield, and R. Kurniawati. Visual data processing in multimedia systems. Lecture Notes in Computer Science, 1306:40–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehatschek:1997:DDP

[326] H. Rehatschek. Design of a distributed planetary image data archive based on an ATM network. Lecture Notes in Computer Science, 1306:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferri:1997:OOM

[327] F. Ferri, P. Grifoni, and F. L. Ricci. An object-oriented model for a visual information system of patient folders. Lecture Notes in Computer Science, 1306:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirenkov:1997:FMF

[328] N. Mirenkov and T. Mirenkova. “filmification” of methods and film databases. Lecture Notes in Computer Science, 1306:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1997:CCB

[329] G. Lu. Chain code-based shape representation and similarity measure. Lecture Notes in Computer Science, 1306: 135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1997:CBR

[330] Z. J. Zheng and C. H. C. Leung. Content-based retrieval using random verification of 1D cellular automata images. Lecture Notes in Computer Science, 1306:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1997:ECT

[331] Z. J. Zheng, C. Pantelis, C. H. C. Leung, and M. T. H. Wong. Extracting complex tissues of MRI brain images using conjugate transformation for medical visual information systems. Lecture Notes in Computer Science, 1306:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:UFF

[332] M. C. Lee and D. A. Adjeroh. A uniform framework for feature-based indexing and retrieval in visual information systems. Lecture Notes in Computer Science, 1306:176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wegner:1997:NED

[333] L. Wegner, S. Thelemann, J. Thamm, and D. Wilke. Navigational exploration and declarative queries in a prototype for visual information systems. Lecture Notes in Computer Science, 1306:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rose:1997:ITL

[334] E. A. Rose. Incorporating typed links in a visual search tool for images on the Web. Lecture Notes in Computer Science, 1306:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palhang:1997:AAO

[335] M. Palhang and A. Sowmya. Automatic acquisition of object models by relational learning. Lecture Notes in Computer Science, 1306:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Worring:1997:SDS

[336] M. Worring, C. Van den Berg, L. Hardman, and A. Tam. System design for structured hypermedia generation. Lecture Notes in Computer Science, 1306:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehmayer:1997:CDW

[337] G. Ehmayer, G. Kappel, and S. Reich. Connecting databases to the Web: A taxonomy of gateways (invited lecture at DEXA ’96). Lecture Notes in Computer Science, 1308:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardarin:1997:MFD

[338] G. Gardarin. Multimedia federated databases on intranets: Web-enabling IRO-DB. Lecture Notes in Computer Science, 1308:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branki:1997:TCD

[339] T. Branki and B. Defude. A terminological canonical data model for cooperating heterogeneous geographical information systems. Lecture Notes in Computer Science, 1308:28–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:EAR

[340] L.-C. Chan and Q. Li. An extensible approach to reactive processing in an advanced object modelling environment. Lecture Notes in Computer Science, 1308:38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1997:CFM

[341] J. Debenham. Constructing the functional model. Lecture Notes in Computer Science, 1308:48–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellatreche:1997:HCP

[342] L. Bellatreche, K. Karlapalem, and A. Simonet. Horizontal class partitioning in object-oriented databases. Lecture Notes in Computer Science, 1308: 58–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benattou:1997:IIM

[343] M. Benattou and L. Lakhal. Incremental inheritance model for an OODBMS. Lecture Notes in Computer Science, 1308:68–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gay:1997:CTO

[344] J.-Y. Gay and L. Gruenwald. A clustering technique for object oriented databases. Lecture Notes in Computer Science, 1308:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouaziz:1997:DIT

[345] T. Bouaziz, J. Karvonen, A. Pesonen, and A. Wolski. Design and implementation of TEMPO fuzzy triggers. Lecture Notes in Computer Science, 1308:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desouki:1997:SMD

[346] M. S. Desouki, A. Simonet, and M. Simonet. Supporting and maintaining derivations: A declarative active approach. Lecture Notes in Computer Science, 1308:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fauvet:1997:MHO

[347] M.-C. Fauvet, J.-F. Canavaggio, and P.-C. Scholl. Modelling histories in object DBMS. Lecture Notes in Computer Science, 1308:112–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Faure:1997:ATU

[348] B. Faure, M. Teisseire, and R. Cicchetti. Activity threads: A unified framework for aiding behavioural modelling. Lecture Notes in Computer Science, 1308:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor:1997:LOO

[349] O. Pastor, E. Insfran, V. Pelechano, and S. Ramirez. Linking objectoriented conceptual modelling with object-oriented implementation in Java. Lecture Notes in Computer Science, 1308:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:TUD

[350] B. Wang. Toward a unified data model for large hypermedia applications. Lecture Notes in Computer Science, 1308:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oussalah:1997:MGB

[351] C. Oussalah, M. Magnan, and S. Vauttier. Managing the global behaviour of composite objects. Lecture Notes in Computer Science, 1308:152–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teuhola:1997:HCD

[352] J. Teuhola. Heuristic clustering of database objects according to multivalued attributes. Lecture Notes in Computer Science, 1308:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieczeraycki:1997:OOE

[353] W. Wieczeraycki. Object-oriented environment for collaborative software design. Lecture Notes in Computer Science, 1308:172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:RTD

[354] S. Y. Lee and T. W. Ling. Refined termination decision in active databases. Lecture Notes in Computer Science, 1308:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zettsu:1997:TSA

[355] K. Zettsu, K. Uehara, K. Tanaka, and N. Kimura. A time-stamped authoring graph for video databases. Lecture Notes in Computer Science, 1308:192– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aritsugi:1997:AST

[356] M. Aritsugi, T. Tagashira, T. Amagasa, and Y. Kanamori. An approach to spatio-temporal queries — interval-based contents representation of images. Lecture Notes in Computer Science, 1308:202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denos:1997:IRS

[357] N. Denos, C. Berrut, and M. Mechkour. An image retrieval system based on the visualization of system relevance via documents. Lecture Notes in Computer Science, 1308:214–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1997:INC

[358] Y. Park and F. Golshani. ImageRoadMap: A new content-based image retrieval system. Lecture Notes in Computer Science, 1308:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buddrus:1997:FSF

[359] F. Buddrus, H. Gaertner, and S.-E. Lautemann. First steps to a formal framework for multilevel database modifications. Lecture Notes in Computer Science, 1308:240–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doucet:1997:VIC

[360] A. Doucet and S. Monties. Versions of integrity constraints in multiversion databases. Lecture Notes in Computer Science, 1308:252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayol:1997:SPI

[361] E. Mayol and E. Teniente. Structuring the process of integrity maintenance. Lecture Notes in Computer Science, 1308:262–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jose:1997:RMS

[362] J. M. Jose and D. J. Harper. A retrieval mechanism for semi-structured photographic collections. Lecture Notes in Computer Science, 1308:276–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kidawara:1997:EMC

[363] Y. Kidawara, K. Tanaka, and K. Uehara. Encapsulating multimedia contents and a copyright protection mechanism into distributed objects. Lecture Notes in Computer Science, 1308:293– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1997:PDI

[364] J. Park, Y. Nah, and S. Lee. A priority-driven inheritance scheme for temporal-spatial structures in multimedia databases. Lecture Notes in Computer Science, 1308:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kouloumdjian:1997:RDR

[365] J. Kouloumdjian and F. Toumani. Relational database reverse engineering and terminological reasoning. Lecture Notes in Computer Science, 1308:313– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunderraman:1997:MND

[366] R. Sunderraman. Modelling negative and disjunctive information in relational databases. Lecture Notes in Computer Science, 1308:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlapalem:1997:QDD

[367] K. Karlapalem and N. M. Pun. Query driven data allocation algorithms for distributed database systems. Lecture Notes in Computer Science, 1308:347– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1997:GMD

[368] Y.-L. Lo, K. A. Hua, and H. C. Young. A general multidimensional data allocation method for multicomputer database systems. Lecture Notes in Computer Science, 1308:357–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravat:1997:DOO

[369] F. Ravat, M. De Michiel, and G. Zurfluh. Distributed object oriented databases: An allocation method. Lecture Notes in Computer Science, 1308: 367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huemer:1997:MMA

[370] C. Huemer, G. Quirchmayr, and A. M. Tjoa. A meta message approach for electronic data interchange (EDI). Lecture Notes in Computer Science, 1308: 377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1997:SRS

[371] D.-Y. Seo, D.-H. Lee, K.-S. Moon, and J. Chang. Schemaless representation of semistructured data and schema construction. Lecture Notes in Computer Science, 1308:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimojo:1997:MAC

[372] I. Shimojo, T. Tachikawa, and M. Takizawa. M-ary commitment protocol with partially ordered domain. Lecture Notes in Computer Science, 1308: 397–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AramburuCabo:1997:MDL

[373] M. J. Aramburu Cabo and R. Berlanga Llavori. Metadata for a digital library of historical documents. Lecture Notes in Computer Science, 1308:409– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahonen:1997:ADD

[374] H. Ahonen, B. Heikkinen, O. Heinonen, and P. Kilpelaeinen. Assembling documents from digital libraries. Lecture Notes in Computer Science, 1308:419– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grefen:1997:TLT

[375] P. Grefen, J. Vonk, E. Boertjes, and P. Apers. Two-layer transaction management for workflow management applications. Lecture Notes in Computer Science, 1308:430–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madria:1997:CRO

[376] S. K. Madria, S. N. Maheshwari, B. Chandra, and B. Bhargava. Crash recovery in an open and safe nested transaction model. Lecture Notes in Computer Science, 1308:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winiwarter:1997:MLW

[377] W. Winiwarter and Y. Kambayashi. A machine learning workbench in a DOOD framework. Lecture Notes in Computer Science, 1308:452–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miura:1997:ECO

[378] T. Miura and I. Shioya. Examining complex objects for type scheme discovery. Lecture Notes in Computer Science, 1308:462–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbas:1997:FAO

[379] I. Abbas and O. Boucelma. A framework for algebraic optimization of object-oriented query languages. Lecture Notes in Computer Science, 1308: 478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balownew:1997:LAT

[380] O. Balownew, T. Bode, A. B. Cremers, and J. Kalinski. A library application on top of an RDBMS: Performance aspects. Lecture Notes in Computer Science, 1308:488–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cha:1997:ANN

[381] G.-H. Cha, H.-H. Park, and C.-W. Chung. Analysis of nearest neighbour query performance in multidimensional index structures. Lecture Notes in Computer Science, 1308:498–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koh:1997:QLI

[382] J.-L. Koh, A. L. P. Chen, P. C. M. Chang, and J. C. C. Chen. A query language and interface for integrated media and alphanumeric database systems. Lecture Notes in Computer Science, 1308:508–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leilahi:1997:AQL

[383] K. Leilahi, R. Souah, and N. Spyratos. An algebraic query language for objectoriented data models. Lecture Notes in Computer Science, 1308:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1997:DOS

[384] W. Ng and M. Levene. The development of ordered SQL packages for modelling advanced applications. Lecture Notes in Computer Science, 1308:529– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moon:1997:IUA

[385] S.-H. Moon and B.-H Hong. Incremental update algorithms for materialized spatial views by using view derivation relationships. Lecture Notes in Computer Science, 1308:539–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohania:1997:IMM

[386] M. Mohania, S. Konomi, and Y. Kambayashi. Incremental maintenance of materialized views. Lecture Notes in Computer Science, 1308:551–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Afsarmanesh:1997:FIM

[387] H. Afsarmanesh and L. M. Camarinha-Matos. Federated information management for cooperative virtual organizations. Lecture Notes in Computer Science, 1308:561–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellahsene:1997:EVM

[388] Z. Bellahsene. Extending a view mechanism to support schema evolution in federated database systems. Lecture Notes in Computer Science, 1308:573– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:QTF

[389] Y. Chen and W. Benn. On the query treatment in federated systems. Lecture Notes in Computer Science, 1308: 583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:QPT

[390] X. Wu. Query processing and transforming for a federated database system. Lecture Notes in Computer Science, 1308:593–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1997:VCI

[391] D. Dubois and H. Prade. Valid or complete information in databases — A possibility theory-based analysis. Lecture Notes in Computer Science, 1308:603–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakata:1997:HID

[392] M. Nakata, G. Resconi, and T. Murai. Handling imperfection in databases: A modal logic approach. Lecture Notes in Computer Science, 1308:613–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1997:TKE

[393] U. Hahn and K. Schnattinger. Text knowledge engineering by qualitative terminological learning. Lecture Notes in Computer Science, 1308:623–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coenen:1997:ATD

[394] F. P. Coenen, B. Beattie, T. J. M. Bench-Capon, and B. M. Diaz. Approach to a tesseral n-dimensional spatial reasoning. Lecture Notes in Computer Science, 1308:633–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Si-Said:1997:GRE

[395] S. Si-Said and C. Rolland. Guidance for requirements engineering processes. Lecture Notes in Computer Science, 1308:643–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrich:1997:RBS

[396] A. Henrich. Repository based software cost estimation. Lecture Notes in Computer Science, 1308:653–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plihon:1997:UGA

[397] V. Plihon and C. Rolland. Using a generic approach to support the construction of methods. Lecture Notes in Computer Science, 1308:663–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1997:RSE

[398] D. Costal, M.-R. Sancho, A. Olive, and A. Rosello. The role of structural events in behaviour specification. Lecture Notes in Computer Science, 1308:673–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohbuchi:1997:EDM

[399] R. Ohbuchi, H. Masuda, and M. Aono. Embedding data in 3D models. Lecture Notes in Computer Science, 1309:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolic:1997:EMP

[400] A. Smolic, B. Makai, G. Lin, and T. Sikora. Estimation of motion parameters of a rigid body from a monocular image sequence for MPEG-4 applications. Lecture Notes in Computer Science, 1309:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:SSO

[401] J. Z. Wang, G. Wiederhold, and O. Firschein. System for screening objectionable images using Daubechies’ wavelets and color histograms. Lecture Notes in Computer Science, 1309:20– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosco:1997:MIM

[402] P. Bosco, G. Martini, and G. Reteuna. MUSIC: An interactive MUltimedia ServIce Composition environment for distributed systems. Lecture Notes in Computer Science, 1309:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soto:1997:ANV

[403] M. Soto and H. Le Van Gong. Achieving networked virtual environments interoperability. Lecture Notes in Computer Science, 1309:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahabi:1997:RHS

[404] C. Shahabi, M. H. Alshayeji, and S. Wang. A redundant hierarchical structure for a distributed continuous media server. Lecture Notes in Computer Science, 1309:51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dammicco:1997:PCM

[405] G. Dammicco and U. Mocci. Program caching and multicasting techniques in VoD networks. Lecture Notes in Computer Science, 1309:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baier:1997:PDW

[406] R. Baier, C. Gran, P. Hoepner, and K. Hofrichter. PENGUIN: DAVIC and the WWW in coexistence. Lecture Notes in Computer Science, 1309:77– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parnes:1997:MMT

[407] P. Parnes, K. Synnes, and D. Schefstroem. mTunnel: A multicast tunneling system with a user based qualityof-service model. Lecture Notes in Computer Science, 1309:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1997:DDC

[408] Q. Sun and H. Langendoerfer. A distributed delay-constrained dynamic multicast routing algorithm. Lecture Notes in Computer Science, 1309:97– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrion:1997:ESI

[409] J.-C Henrion. An efficient software implementation of a FEC code. Lecture Notes in Computer Science, 1309:107– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korst:1997:DSV

[410] J. Korst, V. Pronk, and P. Coumans. Disk scheduling for variable-rate data streams. Lecture Notes in Computer Science, 1309:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsao:1997:NDP

[411] S.-L. Tsao, Y.-M. Huang, C.-C. Lin, and S.-C. Liou. A novel data placement scheme on optical discs for Near-VOD servers. Lecture Notes in Computer Science, 1309:133–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lam:1997:PFM

[412] K.-Y. Lam, C. C. H. Ngan, and J. K. Y. Ng. A priority feedback mechanism for distributed MPEG video player systems. Lecture Notes in Computer Science, 1309:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1997:SRT

[413] H.-H. Chu and K. Nahrstedt. A soft real time scheduling server in UNIX operating system. Lecture Notes in Computer Science, 1309:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gumbrich:1997:SVT

[414] S. Gumbrich and J. Heller. Stored video transmission across ATM networks: RVBR versus RCBR service for interactive applications. Lecture Notes in Computer Science, 1309:163– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noro:1997:ICS

[415] R. Noro and J. P. Hubaux. Improving clock synchronization for MPEG-2 services over ATM networks. Lecture Notes in Computer Science, 1309:176– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsui:1997:VNS

[416] Y. Matsui and H. Tokuda. VoR: A network system framework for VBR over reserved bandwidth. Lecture Notes in Computer Science, 1309:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knapp:1997:EFM

[417] A. Knapp, N. Koch, M. Wirsing, and J. Duckeck. EPK-fix: Methods and tools for engineering electronic product catalogues. Lecture Notes in Computer Science, 1309:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrandino:1997:CAH

[418] S. Ferrandino, A. Negro, and V. Scarano. CHEOPS: Adaptive hypermedia on World Wide Web. Lecture Notes in Computer Science, 1309:210–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benz:1997:AOR

[419] H. Benz, S. Fischer, R. Mecklenburg, and A. Wenger. Application output recording for instant authoring in a distributed multimedia annotation environment. Lecture Notes in Computer Science, 1309:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsukada:1997:VDM

[420] A. Tsukada, T. Hata, F. Matsuda, and K. Sato. Video data management in media controller: A distributed multimedia surveillance system. Lecture Notes in Computer Science, 1309:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1997:PND

[421] Anonymous. Personalised news on demand: The HyNoDe service. Lecture Notes in Computer Science, 1309:241– 250, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1997:BVC

[422] H. Brandt, C. Tittel, P. Todorova, and J. J. Tchouto. Broadband video conference customer premises equipment. Lecture Notes in Computer Science, 1309:251–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1997:RWS

[423] R. Ludwig, N. Niebert, and R. Quinet. Radio webs — support architecture for mobile Web access. Lecture Notes in Computer Science, 1309:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1997:AAV

[424] L. Neumann and A. B. Raposo. An approach for an adaptive visualization in a mobile environment. Lecture Notes in Computer Science, 1309:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauff:1997:MCI

[425] M. Lauff and H.-W. Gellersen. Multimedia client implementation on personal digital assistants. Lecture Notes in Computer Science, 1309:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helbig:1997:CCI

[426] T. Helbig, S. Tretter, and D. Trossen. Combining CORBA and ITU-T.120 to an efficient conferencing service. Lecture Notes in Computer Science, 1309: 296–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bochicchio:1997:SEE

[427] M. Bochicchio and P. Paolini. Supportive environments for executing multimedia applications. Lecture Notes in Computer Science, 1309:308–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koerner:1997:PCC

[428] E. Koerner. Patterns for constructing CSCW applications in TINA. Lecture Notes in Computer Science, 1309:320– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barolli:1997:PEF

[429] L. Barolli, K. Tanno, and A. Greca. Performance evaluation of the fuzzy policing mechanism for still picture in ATM networks. Lecture Notes in Computer Science, 1309:330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braun:1997:IAV

[430] T. Braun and A. Reisenauer. Implementation of an audio/video conferencing application over native ATM. Lecture Notes in Computer Science, 1309: 342–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dresler:1997:NAA

[431] S. Dresler, M. Hofmann, C. Schmidt, and H. R. Wiltfang. A native ATM API suited for multimedia communication. Lecture Notes in Computer Science, 1309:352–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilt:1997:MCS

[432] V. Hilt and W. Geyer. A model for collaborative services in distributed learning environments. Lecture Notes in Computer Science, 1309:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1997:UDM

[433] O. Neumann, S. Rennecke, and A. Schill. Using distributed multimedia infrastructures for advanced teleteaching applications. Lecture Notes in Computer Science, 1309:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hafid:1997:SSA

[434] A. Hafid and R. Boutaba. A scalable scheme to access multimedia documents with quality of service guarantees. Lecture Notes in Computer Science, 1309:386–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hollfelder:1997:CCA

[435] S. Hollfelder, A. Kraiss, and T. C. Rakow. A client-controlled adaptation framework for multimedia database systems. Lecture Notes in Computer Science, 1309:397–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:IMC

[436] C.-M. Huang and C. Wang. Interactive multimedia communications at the presentation layer. Lecture Notes in Computer Science, 1309:410–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirag:1997:MAM

[437] S. Wirag. Modeling of adaptable multimedia documents. Lecture Notes in Computer Science, 1309:420–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michelitsch:1997:QAB

[438] G. Michelitsch, M. Ott, D. Reininger, and G. Welling. QoS aware browsing in distributed multimedia systems. Lecture Notes in Computer Science, 1309: 430–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deicke:1997:OOC

[439] J. Deicke, U. Mayer, and M. Glesner. An object-oriented client/server architecture for video-on-demand applications. Lecture Notes in Computer Science, 1309:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holfelder:1997:IRR

[440] W. Holfelder. Interactive remote recording and playback of multicast videoconferences. Lecture Notes in Computer Science, 1309:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petkovic:1997:COP

[441] D. Petkovic. Challenges and opportunities for pattern recognition and computer vision research in year 2000 and beyond. Lecture Notes in Computer Science, 1310:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olsen:1997:MGM

[442] O. F. Olsen and M. Nielsen. Multiscale gradient magnitude watershed segmentation. Lecture Notes in Computer Science, 1310:6–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baronti:1997:SMI

[443] S. Baronti, A. Casini, F. Lotti, and S. Porcinai. Segmentation of multispectral images of works of art through principal component analysis. Lecture Notes in Computer Science, 1310:14– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sze:1997:MED

[444] C.-J. Sze, H.-Y. M. Liao, H.-L. Hung, and K.-C. Fan. Multiscale edge detection via normal changes. Lecture Notes in Computer Science, 1310:22– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloch:1997:EAF

[445] I. Bloch and H. Maitre. Extending adjacency to fuzzy sets for coping with imprecise image objects. Lecture Notes in Computer Science, 1310:30– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacinto:1997:ASI

[446] G. Giacinto and F. Roli. Adaptive selection of image classifiers. Lecture Notes in Computer Science, 1310:38– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordella:1997:CRU

[447] L. P. Cordella, P. Foggia, C. Sansone, and F. Tortorella. Classification reliability and its use in multi-classifier systems. Lecture Notes in Computer Science, 1310:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:CLM

[448] C.-Y. Kim, Y.-S. Seo, and I.-S. Kweon. Color linear model. Lecture Notes in Computer Science, 1310:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marini:1997:CAC

[449] D. Marini and A. Rizzi. A computational approach to color illusions. Lecture Notes in Computer Science, 1310: 62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grau:1997:ITI

[450] A. Grau and J. Saludes. Improved textured images segmentation using an energy functional. Lecture Notes in Computer Science, 1310:70–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RegincosIsern:1997:CCS

[451] J. Regincos Isern and J. Batlle Grabulosa. Contribution to the colour segmentation by means of an algorithm which reduces the CCDs saturation problems. Lecture Notes in Computer Science, 1310:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aiazzi:1997:PBM

[452] B. Aiazzi, L. Alparone, S. Baronti, and V. Cappellini. Pyramid-based multisensor image data fusion with enhancement of textural features. Lecture Notes in Computer Science, 1310:87– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chetverikov:1997:TAU

[453] D. Chetverikov. Texture analysis using pairwise interaction maps. Lecture Notes in Computer Science, 1310:95– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pujas:1997:ECI

[454] P. Pujas and M.-J. Aldon. Estimation of the color image gradient with perceptual attributes. Lecture Notes in Computer Science, 1310:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lalonde:1997:CLE

[455] M. Lalonde and Y. Li. Contour line extraction from color images of scanned maps. Lecture Notes in Computer Science, 1310:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Androutsos:1997:SAE

[456] P. Androutsos, D. Androutsos, K. N. Plataniotis, and A. N. Venetsanopoulos. Subjective analysis of edge detectors in color image processing. Lecture Notes in Computer Science, 1310:119– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carkacioglu:1997:SMB

[457] A. Carkacioğlu and F. T. Yarman-Vural. Similarity measures for binary and grey level Markov random field textures. Lecture Notes in Computer Science, 1310:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cafforio:1997:SEE

[458] C. Cafforio, E. Di Sciascio, C. Guaragnella, and G. Piscitelli. A simple and effective edge detector. Lecture Notes in Computer Science, 1310:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biancardi:1997:IIM

[459] A. Biancardi, L. Lombardi, and V. Pacaccio. Improvements to image magnification. Lecture Notes in Computer Science, 1310:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilson:1997:RSC

[460] R. C. Wilson and E. R. Hancock. Refining surface curvature with relaxation labeling. Lecture Notes in Computer Science, 1310:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salden:1997:DSS

[461] A. H. Salden. Dynamic scale-space theories. Lecture Notes in Computer Science, 1310:158–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barcucci:1997:RDS

[462] E. Barcucci, A. Del Lungo, M. Nivat, and R. Pinzani. Reconstructing digital sets from X-rays. Lecture Notes in Computer Science, 1310:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wendling:1997:PRC

[463] L. Wendling, J. Desachy, and A. Paries. Pattern recognition from compressed labelled trees of fuzzy regions. Lecture Notes in Computer Science, 1310: 174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1997:OAE

[464] X. Jiang. Optimality analysis of edge detection algorithms for range images. Lecture Notes in Computer Science, 1310:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloboda:1997:ASI

[465] F. Sloboda and B. Zat’Ko. Analysis situs and image processing. Lecture Notes in Computer Science, 1310:190– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsakis:1997:DCF

[466] P. Matsakis, J. Gadiou, and J. Desachy. Defining cost functions and profitability measures for digraphs associated with raster dems. Lecture Notes in Computer Science, 1310:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchez:1997:UPS

[467] J. S. Sanchez, F. Pla, and F. J. Ferri. Using proximity and spatial homogeneity in neighbourhood-based classifiers. Lecture Notes in Computer Science, 1310:206–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiRuberto:1997:ISM

[468] C. Di Ruberto, M. Nappi, and S. Vitulano. Image segmentation by means of fuzzy entropy measure. Lecture Notes in Computer Science, 1310:214– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basman:1997:ERS

[469] A. Basman, J. Lasenby, and R. Cipolla. Efficient region segmentation through “creep-and-merge”. Lecture Notes in Computer Science, 1310:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ithesta:1997:ATB

[470] J. M. Ithesta, P. J. Sanz, and A. P. Del Pobil. An automatic transformation from bimodal to pseudo-binary images. Lecture Notes in Computer Science, 1310:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:NDM

[471] Z. Zhang, M. Braun, and P. Abbott. A new deformable model for 3D image segmentation. Lecture Notes in Computer Science, 1310:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingaretti:1997:EIS

[472] P. Zingaretti, A. Carbonaro, and P. Puliti. Evolutionary image segmentation. Lecture Notes in Computer Science, 1310:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smits:1997:DAM

[473] P. C. Smits, S. G. Dellepiane, and G. Vernazza. Discontinuity adaptive MRF model for synthetic aperture radar image analysis. Lecture Notes in Computer Science, 1310:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuisenaire:1997:RGE

[474] O. Cuisenaire. Region growing Euclidean distance transforms. Lecture Notes in Computer Science, 1310:263– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bae:1997:CNM

[475] S. C. Bae and I. S. Kweon. COP: A new method for extracting edges and corners. Lecture Notes in Computer Science, 1310:271–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellon:1997:IAS

[476] O. R. P. Bellon and C. L. Tozzi. An integrated approach for segmentation and representation of range images. Lecture Notes in Computer Science, 1310:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anh:1997:TDF

[477] V. Anh, J. Maeda, T. Ishizaka, and Y. Suzuki. Two-dimensional fractal segmentation of natural images. Lecture Notes in Computer Science, 1310: 287–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haindl:1997:FSR

[478] M. Haindl and P. Zid. Fast segmentation of range images. Lecture Notes in Computer Science, 1310:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurt:1997:ICB

[479] B. Kurt, M. Goekmen, and A. K. Jain. Image compression based on centipede model. Lecture Notes in Computer Science, 1310:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ojala:1997:UTS

[480] T. Ojala and M. Pietikaeinen. Unsupervised texture segmentation using feature distributions. Lecture Notes in Computer Science, 1310:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gevers:1997:CBO

[481] T. Gevers and A. W. M. Smeulders. Color based object recognition. Lecture Notes in Computer Science, 1310:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeWouwer:1997:CTC

[482] G. Van De Wouwer, S. Livens, P. Scheunders, and D. Van Dyck. Color texture classification by wavelet energy correlation signatures. Lecture Notes in Computer Science, 1310:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boldrin:1997:CMC

[483] E. Boldrin and R. Schettini. Crossmedia color matching using neural networks. Lecture Notes in Computer Science, 1310:335–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1997:ORP

[484] J. K. Aggarwal and S. Shah. Object recognition and performance bounds. Lecture Notes in Computer Science, 1310:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuille:1997:RIW

[485] A. L. Yuille, M. Ferraro, and T. Zhang. Relating image warping to 3D geometrical deformations. Lecture Notes in Computer Science, 1310:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgefors:1997:UTB

[486] G. Borgefors, G. Ramella, and G. Sanniti Di Baja. Using top-down and bottom-up analysis for a multiscale skeleton hierarchy. Lecture Notes in Computer Science, 1310:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiStefano:1997:NAP

[487] L. Di Stefano and F. Boland. A new algorithm for 3D profilometry based on phase measurement. Lecture Notes in Computer Science, 1310:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pulli:1997:SMD

[488] K. Pulli, M. Cohen, T. Duchamp, and H. Hoppe. Surface modeling and display from range and color data. Lecture Notes in Computer Science, 1310: 385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duta:1997:IAS

[489] N. Duta and M. Sonka. An improved active shape model: Handling occlusion and outliers. Lecture Notes in Computer Science, 1310:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cross:1997:PMU

[490] A. D. J. Cross and E. R. Hancock. Perspective matching using the EM algorithm. Lecture Notes in Computer Science, 1310:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:IHF

[491] J. Sato and R. Cipolla. Identifying human face profiles with semi-local integral invariants. Lecture Notes in Computer Science, 1310:414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camacho:1997:AFS

[492] P. Camacho, F. Arrebola, and F. Sandoval. Adaptive fovea structures for space-variant sensors. Lecture Notes in Computer Science, 1310:422–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottoni:1997:SCI

[493] P. Bottoni, L. Cinque, S. Levialdi, and P. Mussio. Structural characterization of image processing operators. Lecture Notes in Computer Science, 1310:430–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spiess:1997:ECP

[494] S. Spiess and M. Li. Easy calibration of pan/tilt camera heads and online computation of the epipolar correspondences. Lecture Notes in Computer Science, 1310:438–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barni:1997:IST

[495] M. Barni, F. Bartolini, V. Cappellini, and F. Lambardi. Integration of spatiotemporal information for motion detection by means of fuzzy reasoning. Lecture Notes in Computer Science, 1310:446–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malo:1997:AME

[496] Y. Malo, F. Ferri, J. Albert, and J. M. Artigas. Adaptive motion estimation and video vector quantization based on spatiotemporal non-linearities of human perception. Lecture Notes in Computer Science, 1310:454–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nomura:1997:IBA

[497] A. Nomura. Integral based approach for determining motion vector fields. Lecture Notes in Computer Science, 1310:462–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trajkovic:1997:PAS

[498] M. Trajkovic and M. Hedley. A practical algorithm for structure and motion recovery from long sequence of images. Lecture Notes in Computer Science, 1310:470–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dornaika:1997:OPA

[499] F. Dornaika and C. Garcia. Object pose by affine iterations. Lecture Notes in Computer Science, 1310:478– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magarey:1997:RME

[500] J. Magarey, A. Kokaram, and N. Kingsbury. Robust motion estimation using chrominance information in color image sequences. Lecture Notes in Computer Science, 1310:486–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herodotou:1997:TPV

[501] N. Herodotou and A. N. Venetsanopoulos. Temporal prediction of video sequences using an image warping technique based on color segmentation. Lecture Notes in Computer Science, 1310:494–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badenas:1997:MIB

[502] J. Badenas, M. Bober, and F. Pla. Motion and intensity-based segmentation and its application to traffic monitoring. Lecture Notes in Computer Science, 1310:502–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raji:1997:GDC

[503] A. Raji, E. Petit, J. Lemoine, and S. Djeziri. A geometrically deformable contour model. Lecture Notes in Computer Science, 1310:510–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durou:1997:NVD

[504] J.-D. Durou, L. Mascarilla, and D. Piau. Non-visible deformations. Lecture Notes in Computer Science, 1310: 519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peckar:1997:TSP

[505] W. Peckar, C. Schnoerr, K. Rohr, and H. S. Stiehl. Two-step parameterfree elastic image registration with prescribed point displacements. Lecture Notes in Computer Science, 1310:527– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucchiara:1997:LFS

[506] R. Cucchiara, M. Piccardi, M. Bariani, and P. Mello. Learning for feature selection and shape detection. Lecture Notes in Computer Science, 1310:535– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anelli:1997:EDA

[507] G. Anelli, A. Broggi, and G. Destri. Experiments on the decomposition of arbitrarily shaped binary morphological structuring elements. Lecture Notes in Computer Science, 1310:543–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varley:1997:BMC

[508] A. Varley and P. Rayner. Bézier modelling of cracks. Lecture Notes in Computer Science, 1310:551–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirhosseini:1997:ADT

[509] A. R. Mirhosseini, K.-M. Lam, and H. Yan. An adaptive deformable template for mouth boundary modeling. Lecture Notes in Computer Science, 1310:559–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tung:1997:TSF

[510] L. H. Tung and I. King. A two-stage framework for polygon retrieval using minimum circular error bound. Lecture Notes in Computer Science, 1310: 567–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saha:1997:TSP

[511] P. K. Saha and D. D. Majumder. Topology and shape preserving parallel thinning for 3D digital images — A new approach. Lecture Notes in Computer Science, 1310:575–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lew:1997:CMB

[512] M. S. Lew, M. Chaudron, N. Huijsmans, and A. She. Convergence of model based shape from shading. Lecture Notes in Computer Science, 1310: 582–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciuc:1997:QAT

[513] M. Ciuc, D. Coquin, and P. Bolon. Quantitative assessment of two skeletonization algorithms adapted to rectangular grids. Lecture Notes in Computer Science, 1310:588–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falcone:1997:AGS

[514] M. Falcone and M. Sagona. An algorithm for the global solution of the shape-from-shading model. Lecture Notes in Computer Science, 1310:596– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peura:1997:SCM

[515] M. Peura. A statistical classification method for hierarchical irregular objects. Lecture Notes in Computer Science, 1310:604–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yip:1997:MLD

[516] R. K. K. Yip. Multi-level dynamic programming for axial motion stereo line matching. Lecture Notes in Computer Science, 1310:612–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreer:1997:AGL

[517] O. Schreer, I. Hartmann, and R. Adams. Analysis of grey-level features for line segment stereo matching. Lecture Notes in Computer Science, 1310:620–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shioyama:1997:OPM

[518] T. Shioyama, H. Y. Wu, W. B. Jiang, and S. Terauchi. 3D object positioning from monocular image brightnesses. Lecture Notes in Computer Science, 1310:628–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:CCB

[519] X.-F. Zhang, A. Luo, W. Tao, and H. Burkhardt. Camera calibration based on 3D-Point-Grid. Lecture Notes in Computer Science, 1310:636–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommellier:1997:GMT

[520] L. Sommellier, E. Tosan, and D. Vandorpe. A geometric modeling tool for stereo-matching and reconstruction of a model of 3D-scene. Lecture Notes in Computer Science, 1310:644–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pla:1997:ETD

[521] F. Pla and M. Bober. Estimating translation/deformation motion through phase correlation. Lecture Notes in Computer Science, 1310:653–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meilhac:1997:RFC

[522] C. Meilhac and C. Nastar. Robust fitting of 3D CAD models to video streams. Lecture Notes in Computer Science, 1310:661–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fusiello:1997:ENA

[523] A. Fusiello, V. Roberto, and E. Trucco. Experiments with a new area-based stereo algorithm. Lecture Notes in Computer Science, 1310:669–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menard:1997:ASM

[524] C. Menard and W. G. Kropatsch. Adaptive stereo matching in correlation scale-space. Lecture Notes in Computer Science, 1310:677–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1997:HDM

[525] J.-I. Park and S. Inoue. Hierarchical depth mapping from multiple cameras. Lecture Notes in Computer Science, 1310:685–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Messmer:1997:FCE

[526] B. T. Messmer and H. Bunke. Fast computation of error-correcting graph isomorphisms based on model precompilation. Lecture Notes in Computer Science, 1310:693–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serratosa:1997:FDG

[527] F. Serratosa and A. Sanfeliu. Functiondescribed graphs applied to 3D object representation. Lecture Notes in Computer Science, 1310:701–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oswald:1997:CVM

[528] N. Oswald and P. Levi. Cooperative vision in a multi-agent architecture. Lecture Notes in Computer Science, 1310:709–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1997:CCC

[529] R. Jain. Content-centric computing in visual systems. Lecture Notes in Computer Science, 1311:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez:1997:CIR

[530] J. Martinez and S. Guillaume. Color image retrieval fitted to “classical” querying. Lecture Notes in Computer Science, 1311:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huijsmans:1997:QMI

[531] D. P. Huijsmans, M. S. Lew, and D. Denteneer. Quality measures for interactive image retrieval with a performance evaluation of two 3 x 3 texel-based methods. Lecture Notes in Computer Science, 1311:22–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruckstein:1997:HIR

[532] A. M. Bruckstein, R. J. Holt, and A. N. Netravali. Holographic image representations: The Fourier transform method. Lecture Notes in Computer Science, 1311:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santini:1997:IDD

[533] S. Santini and R. Jain. Image databases are not databases with images. Lecture Notes in Computer Science, 1311: 38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zingirian:1997:CMV

[534] N. Zingirian, P. Baglietto, M. Maresca, and M. Migliardi. Customizing MPEG video compression algorithms to specific application domains: The case of highway monitoring. Lecture Notes in Computer Science, 1311:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpentieri:1997:NLI

[535] B. Carpentieri. A new lossless image compression algorithm based on arithmetic coding. Lecture Notes in Computer Science, 1311:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teixeira:1997:ATS

[536] L. Teixeira. Analysis of a two step MPEG video system. Lecture Notes in Computer Science, 1311:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiatr:1997:DHP

[537] K. Wiatr. Dedicated hardware processors for a real-time image data pre-processing implemented in FPGA structure. Lecture Notes in Computer Science, 1311:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferretti:1997:WTA

[538] M. Ferretti and D. Rizzo. Wavelet transform architectures: A system level review. Lecture Notes in Computer Science, 1311:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanAssche:1997:LCP

[539] S. Van Assche, W. Philips, and I. Lemahieu. Lossless compression of pre-press images using a novel colour decorrelation technique. Lecture Notes in Computer Science, 1311:85– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marino:1997:RTH

[540] F. Marino, E. Stella, N. Veneziani, and A. Distante. Real time hardware architecture for visual robot navigation. Lecture Notes in Computer Science, 1311:93–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Distasi:1997:SFE

[541] R. Distasi, M. Nappi, and S. Vitulano. Speeding up fractal encoding of images using a block indexing technique. Lecture Notes in Computer Science, 1311:101–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biancardi:1997:AAM

[542] A. Biancardi and A. Merigot. Adding associative meshes to the PACCO I.P. environment. Lecture Notes in Computer Science, 1311:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teixeira:1997:SMM

[543] L. Teixeira and T. Andrade. Smoothing of MPEG multi-program video coding for packet networks. Lecture Notes in Computer Science, 1311:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saraceno:1997:AVP

[544] C. Saraceno and R. Leonardi. Audiovisual processing for scene change detection. Lecture Notes in Computer Science, 1311:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vicario:1997:WWR

[545] E. Vicario and W. X. He. Weighted walkthroughs in retrieval by content of pictorial data. Lecture Notes in Computer Science, 1311:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokhtarian:1997:NAC

[546] F. Mokhtarian, S. Abbasi, and J. Kittler. A new approach to computation of curvature scale space image for shape similarity retrieval. Lecture Notes in Computer Science, 1311:140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lew:1997:OKI

[547] M. S. Lew, D. P. Huijsmans, and D. Denteneer. Optimal keys for image database indexing. Lecture Notes in Computer Science, 1311:148–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meghini:1997:TIR

[548] C. Meghini, F. Sebastiani, and U. Straccia. The terminological image retrieval model. Lecture Notes in Computer Science, 1311:156–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pei:1997:NBT

[549] S.-C. Pei and C.-M. Cheng. Novel block truncation coding of image sequences for limited-color display. Lecture Notes in Computer Science, 1311: 164–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moss:1997:IRS

[550] S. Moss and E. R. Hancock. Image registration with shape mixtures. Lecture Notes in Computer Science, 1311: 172–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelBimbo:1997:IRC

[551] A. Del Bimbo, M. Mugnaini, P. Pala, and F. Turco. Image retrieval by color regions. Lecture Notes in Computer Science, 1311:180–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cinque:1997:IMB

[552] L. Cinque, S. Levialdi, A. Malizia, and R. Mancini. Interactive model-based matching retrieval. Lecture Notes in Computer Science, 1311:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seo:1997:WBP

[553] Y. Seo, S. Choi, H. Kim, and K.- S Hong. Where are the ball and players? soccer game analysis with color based tracking and image mosaick. Lecture Notes in Computer Science, 1311:196–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colombo:1997:HFC

[554] C. Colombo, A. Rizzi, and I. Genovesi. Histogram families for color-based retrieval in image databases. Lecture Notes in Computer Science, 1311:204– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pala:1997:IRM

[555] P. Pala and S. Santini. Image retrieval by multidimensional elastic matching. Lecture Notes in Computer Science, 1311:212–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vass:1997:OMM

[556] G. G. Vass, M. Daoudi, and F. Ghorbel. Optimization methods in multilayer classifier networks for automatic control of lamellibranch larva growth. Lecture Notes in Computer Science, 1311:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucurachi:1997:NNR

[557] G. Cucurachi, G. Tascini, and F. Piazza. Neural networks for region detection. Lecture Notes in Computer Science, 1311:228–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorodnichy:1997:SDA

[558] D. O. Gorodnichy and A. M. Reznik. Static and dynamic attractors of autoassociative neural networks. Lecture Notes in Computer Science, 1311:238– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Timchenko:1997:BLA

[559] L. I. Timchenko, Y. F. Kutaev, M. A. Grudin, and S. V. Cepornyuk. A brainlike approach to multistage hierarchical image processing. Lecture Notes in Computer Science, 1311:246–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinho:1997:CED

[560] A. J. Pinho and L. B. Almeida. Contextual edge detection using a recurrent neural network. Lecture Notes in Computer Science, 1311:254–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1997:DCS

[561] S. I. Cho, H. Saito, and S. Ozawa. A divide-and-conquer strategy in recovering shape of book surface from shading. Lecture Notes in Computer Science, 1311:262–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:RST

[562] Y. Sato, T. Ishikawa, and M. Otsuki. Reconstruction of 3D shape and texture by active rangefinding. Lecture Notes in Computer Science, 1311:270– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albanesi:1997:WMS

[563] M. G. Albanesi and L. Lombardi. Wavelets for multiresolution shape recognition. Lecture Notes in Computer Science, 1311:276–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1997:IOR

[564] J. Chao, A. Karasudani, and K. Minowa. Invariant object representation and recognition using lie algebra of perceptual vector fields. Lecture Notes in Computer Science, 1311:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sirakov:1997:FAD

[565] N. M. Sirakov. A fast approach for determining of visibility of 3D object’s surfaces. Lecture Notes in Computer Science, 1311:292–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pontil:1997:DAB

[566] M. Pontil and A. Verri. Direct aspectbased 3D object recognition. Lecture Notes in Computer Science, 1311:300– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFloriani:1997:VPS

[567] L. De Floriani, P. Magillo, and E. Puppo. Visualizing parametric surfaces at variable resolution. Lecture Notes in Computer Science, 1311:308– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burge:1997:LVIa

[568] M. Burge and W. Burger. Learning visual ideals. Lecture Notes in Computer Science, 1311:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serra:1997:ANC

[569] J. R. Serra and J. B. Subirana. Adaptive non-cartesian networks for vision. Lecture Notes in Computer Science, 1311:324–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorodnichy:1997:ALN

[570] D. O. Gorodnichy, W. W. Armstrong, and X. Li. Adaptive logic networks for facial feature detection. Lecture Notes in Computer Science, 1311:332– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:ABA

[571] J. Martin and J. L. Crowley. An appearance-based approach to gesturerecognition. Lecture Notes in Computer Science, 1311:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straahlen:1997:EVF

[572] K. Straahlen. Exponential vector field tomography. Lecture Notes in Computer Science, 1311:348–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sarkar:1997:DRS

[573] S. Sarkar and S. Chaudhuri. Detection of rib shadows in digital chest radiographs. Lecture Notes in Computer Science, 1311:356–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardo:1997:MRF

[574] J. M. Pardo, D. Cabello, and J. Heras. A Markov random field model for bony tissue classification. Lecture Notes in Computer Science, 1311:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrido:1997:NMA

[575] A. Garrido, N. Perez De La Blanca, and M. Garcia-Silvente. A new methodology to automatically segment biomedical images. Lecture Notes in Computer Science, 1311:372–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buzug:1997:HBI

[576] T. M. Buzug, J. Weese, C. Lorenz, and W. Beil. Histogram-based image registration for digital subtraction angiography. Lecture Notes in Computer Science, 1311:380–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loncaric:1997:MSC

[577] S. Loncaric and D. Kovacevic. A method for segmentation of CT head images. Lecture Notes in Computer Science, 1311:388–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1997:ARS

[578] H. Jiang, W. Tiu, S. Yamamoto, and S.-I Iisaku. Automatic recognition of spicules in mammograms. Lecture Notes in Computer Science, 1311:396– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiLecce:1997:SEM

[579] V. Di Lecce and A. Guerriero. Specialized environment for medical radiological image visualization. Lecture Notes in Computer Science, 1311:404– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dance:1997:ISU

[580] C. R. Dance, M. H. Syn, R. W. Prager, and J. P. M. Gosling. Interactive segmentation of 3D ultrasound using deformable solid models and active contours. Lecture Notes in Computer Science, 1311:412–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawata:1997:CAD

[581] Y. Kawata, K. Kanazawa, S. Toshioka, and N. Niki. Computer aided diagnosis system for lung cancer based on helical CT images. Lecture Notes in Computer Science, 1311:420–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bello:1997:GGI

[582] F. Bello, A. C. F. Colchester, and S. A. Roell. A generalized geometry and intensity based partial volume correction for magnetic resonance images. Lecture Notes in Computer Science, 1311:428–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonifazzi:1997:RMU

[583] C. Bonifazzi, G. Maino, and A. Tartari. A regularization method for unfolding the measured data of different X-ray spectrometers in Compton scattering tomography. Lecture Notes in Computer Science, 1311:436–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vinitski:1997:FTS

[584] S. Vinitski, T. Iwanaga, C. Gonzalez, and D. Andrews. Fast tissue segmentation based on a 4D feature map: Preliminary results. Lecture Notes in Computer Science, 1311:445–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kontinen:1997:TFC

[585] J. Kontinen, J. Roening, and R. M. MacKie. Texture features in the classification of melanocytic lesions. Lecture Notes in Computer Science, 1311: 453–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rachid:1997:SSC

[586] S. Rachid, N. Niki, H. Nishitahi, and S. Nakamura. Segmentation of sputum color image for lung cancer diagnosis based on neural networks. Lecture Notes in Computer Science, 1311: 461–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hah:1997:FFD

[587] C.-C. Hah, H.-Y. M. Liao, G.-J. Yu, and L.-H. Chen. Fast face detection via morphology-based pre-processing. Lecture Notes in Computer Science, 1311: 469–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arrebola:1997:GSF

[588] F. Arrebola, P. Camacho, and F. Sandoval. Generalization of shifted fovea multiresolution geometries applied to object detection. Lecture Notes in Computer Science, 1311:477–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regazzoni:1997:LTC

[589] C. S. Regazzoni, A. Teschioni, and E. Stringa. A long term change detection method for surveillance applications. Lecture Notes in Computer Science, 1311:485–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vannoorenberghe:1997:APR

[590] P. Vannoorenberghe, C. Motamed, J.- M. Blosseville, and J.-G. Postaire. Automatic pedestrian recognition using real-time motion analysis. Lecture Notes in Computer Science, 1311:493– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosugi:1997:PIS

[591] M. Kosugi and K. Yamashita. Person identification system based on a trapezoid pyramid architecture of a greylevel image. Lecture Notes in Computer Science, 1311:501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregori:1997:API

[592] M. Gregori, L. Lombardi, M. Savini, and A. Scianna. Autonomous plant inspection and anomaly detection. Lecture Notes in Computer Science, 1311: 509–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1997:NCJ

[593] F. Hahn and R. Mota. Nobel chile jalapeño sorting using structured laser and neural networks classifiers. Lecture Notes in Computer Science, 1311: 517–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sumimoto:1997:DIP

[594] T. Sumimoto, K. Kuramoto, S. Okada, and H. Miyauchi. Development of image processing technique for detection of the rescue target in the marine casualty. Lecture Notes in Computer Science, 1311:524–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schultz:1997:BHT

[595] N. Schultz and J. M. Carstensen. Bimodal histogram transformation based on maximum likelihood parameter estimates in univariate Gaussian mixtures. Lecture Notes in Computer Science, 1311:532–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yazdi:1997:RSF

[596] M. H. Ghassemian Yazdi. A robust structural fingerprint restoration. Lecture Notes in Computer Science, 1311: 544–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermida:1997:SAR

[597] X. F. Hermida, F. M. Rodriguez, J. L. F. Lijo, and F. P. Sande. A system for the automatic and real time recognition of V.L.P.’s (vehicle licence plate). Lecture Notes in Computer Science, 1311:552–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaccaro:1997:SCF

[598] R. Vaccaro and S. Dellepiane. Spatial correlation features for SAR images in a small sample size context. Lecture Notes in Computer Science, 1311:560– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heng:1997:CAS

[599] T. C. H. Heng, Y. Kuno, and Y. Shirai. Combination of active sensing and sensor fusion for collision avoidance in mobile robots. Lecture Notes in Computer Science, 1311:568–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozzano:1997:UVD

[600] R. Bozzano and A. Siccardi. Underwater vegetation detection in high frequency sonar images: A preliminary approach. Lecture Notes in Computer Science, 1311:576–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branca:1997:LIT

[601] A. Branca, M. G. Abbate, F. P. Lovergine, and G. Attolico. Leather inspection through singularities detection using wavelet transforms. Lecture Notes in Computer Science, 1311:584– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimauro:1997:ZDH

[602] G. Dimauro, S. Impedovo, G. Pirlo, and A. Salzo. Zoning design for handwritten numeral recognition. Lecture Notes in Computer Science, 1311:592– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chianese:1997:IUC

[603] A. Chianese, M. De Santo, and A. Picariello. Improving the use of contours and skeletons for off-line cursive script segmentation. Lecture Notes in Computer Science, 1311:600–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oezdil:1997:OCR

[604] M. A. Oezdil, F. T. Yarman-Vural, and N. Arica. Optical character recognition without segmentation. Lecture Notes in Computer Science, 1311:608–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amin:1997:ARP

[605] A. Amin and M. Kavianifar. Automatic recognition of printed Arabic text using neural network classifier. Lecture Notes in Computer Science, 1311:616–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahman:1997:NPW

[606] A. F. R. Rahman and M. C. Fairhurst. A novel pair-wise recognition scheme for handwritten characters in the framework of a multi-expert configuration. Lecture Notes in Computer Science, 1311:624–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DelNinno:1997:GFD

[607] E. Del Ninno, G. Nicchiotti, and E. Ottaviani. A general and flexible deskewing method based on generalized projections. Lecture Notes in Computer Science, 1311:632–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duffy:1997:LSA

[608] L. Duffy, F. Lebourgeois, and H. Emptoz. Logical structure analysis by typographic characteristics extraction. Lecture Notes in Computer Science, 1311: 639–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1997:CHL

[609] J. Hu, A. S. Rosenthal, and M. K. Brown. Combining high-level features with sequential local features for online handwriting recognition. Lecture Notes in Computer Science, 1311:647– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizukami:1997:HCC

[610] Y. Mizukami and K. Koga. Handwritten Chinese character recognition using displacement extraction based on directional features. Lecture Notes in Computer Science, 1311:655–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farkas:1997:OIA

[611] D. L. Farkas, B. T. Ballou, C. Du, and G. W. Fisher. Optical image acquisition, analysis and processing for biomedical applications. Lecture Notes in Computer Science, 1311:663–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abbott:1997:SUI

[612] P. Abbott and M. Braun. Segmentation of ultrasound image data by two dimensional autoregressive modelling. Lecture Notes in Computer Science, 1311:672–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1997:CAS

[613] A. Neumann and C. Lorenz. Comparison and application of selected statistical shape models in medical imaging. Lecture Notes in Computer Science, 1311:680–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeWalle:1997:TMD

[614] R. Van De Walle, I. Lemahieu, and E. Achten. Two motion detection algorithms for projection-reconstruction magnetic resonance imaging: Theory and experimental verification. Lecture Notes in Computer Science, 1311:688– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagata:1997:IAS

[615] N. Nagata, T. Dobashi, Y. Manabe, and T. Usami. Image analysis and synthesis using physics-based modeling for pearl quality evaluation system. Lecture Notes in Computer Science, 1311:697–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garibotto:1997:CVI

[616] G. Garibotto and C. Scagliola. Computer vision and image processing in postal automation. Lecture Notes in Computer Science, 1311:705–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimura:1997:APU

[617] T. D. Kimura. Adaptive pen user interface with supervised competitive learning. Lecture Notes in Computer Science, 1311:713–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokaram:1997:JDI

[618] A. C. Kokaram and S. J. Godsill. Joint detection, interpolation, motion and parameter estimation for image sequences with missing data. Lecture Notes in Computer Science, 1311:719– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tistarelli:1997:VPV

[619] M. Tistarelli, E. Grosso, J. Bigun, and C. Sacerdoti. The VIRSBS project: Visual intelligent recognition for secure banking services. Lecture Notes in Computer Science, 1311:727–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pahor:1997:ELR

[620] V. Pahor, G. Ramponi, and R. Castagno. The ESPRIT LTR research project: “nonlinear model-based analysis and description of images for multimedia applications (NOBLESSE)”. Lecture Notes in Computer Science, 1311:735– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smits:1997:ASR

[621] P. C. Smits and S. B. Serpico. Analysis and segmentation of remote sensing images for land-cover mapping. Lecture Notes in Computer Science, 1311: 743–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pieroni:1997:IOA

[622] G. G. Pieroni, G. L. Foresti, and V. Murino. Integration of optical and acoustical imaging sensors for underwater applications. Lecture Notes in Computer Science, 1311:749–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khoudour:1997:PC

[623] L. Khoudour, J. P. Deparis, J. L. Bruyelle, and F. Cabestaing. Project CROMATICA. Lecture Notes in Computer Science, 1311:757–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Austin:1997:CPC

[624] J. Austin, G. Giacinto, I. Kanellopoulos, and K. Lees. The COM-PARES project: COnnectionist methods for preprocessing and analysis of REmote sensing data. Lecture Notes in Computer Science, 1311:765–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokaram:1997:RSD

[625] A. C. Kokaram. Reconstruction of severely degraded image sequences. Lecture Notes in Computer Science, 1311:773–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mari:1997:CPD

[626] M. Mari, C. Dambra, D. Chetverikov, and J. Verestoy. The CRASH project: Defect detection and classification in ferrite cores. Lecture Notes in Computer Science, 1311:781–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sampaio:1997:DOO

[627] P. R. F. Sampaio and N. W. Paton. Deductive object-oriented database systems: A survey. Lecture Notes in Computer Science, 1312:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1997:IDA

[628] W. May, C. Schlepphorst, and G. Lausen. Integrating dynamic aspects into deductive object-oriented databases. Lecture Notes in Computer Science, 1312:20–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayol:1997:FDD

[629] E. Mayol, J. A. Pastor, E. Teniente, and T. Urpi. FOLRE: A deductive database system for the integrated treatment of updates. Lecture Notes in Computer Science, 1312:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanson:1997:ITA

[630] E. N. Hanson and S. Khosla. An introduction to the TriggerMan asynchronous trigger processor. Lecture Notes in Computer Science, 1312:51– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geppert:1997:LPM

[631] A. Geppert and D. Tombros. Logging and post-mortem analysis of workflow executions based on event histories. Lecture Notes in Computer Science, 1312:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlesinger:1997:RMS

[632] M. Schlesinger and G. Loerincze. Rule modelling and simulation in ALFRED. Lecture Notes in Computer Science, 1312:83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comai:1997:UPD

[633] S. Comai and L. Tanca. Using the properties of datalog to prove termination and confluence in active databases. Lecture Notes in Computer Science, 1312:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:CPA

[634] X. Wang, J.-H. You, and L. Y. Yuan. On confluence property of active databases with meta-rules. Lecture Notes in Computer Science, 1312: 118–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschudi:1997:IER

[635] M. K. Tschudi, S. D. Urban, S. W. Dietrich, and A. P. Karadimce. An implementation and evaluation of the refined triggering graph method for active rule termination analysis. Lecture Notes in Computer Science, 1312:133–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:CFT

[636] H. W. R. Chan, S. W. Dietrich, and S. D. Urban. On control flow testing of active rules in a declarative objectoriented framework. Lecture Notes in Computer Science, 1312:165–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraternali:1997:VAR

[637] P. Fraternali, E. Teniente, and T. Urpi. Validating active rules by planning. Lecture Notes in Computer Science, 1312:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouzeghoub:1997:ADG

[638] M. Bouzeghoub, F. Fabret, F. Llirbat, and M. Matulovic. ACTIVE-DESIGN: A generic toolkit for deriving specific rule execution models. Lecture Notes in Computer Science, 1312:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anderson:1997:DPL

[639] S. Anderson and K. Tourlas. Diagrams and programming languages for programmable controllers. Lecture Notes in Computer Science, 1313:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dierks:1997:GSR

[640] H. Dierks and C. Dietz. Graphical specification and reasoning: Case study generalised railroad crossing. Lecture Notes in Computer Science, 1313:20– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reggio:1997:GNF

[641] G. Reggio and M. Larosa. A graphic notation for formal specifications of dynamic systems. Lecture Notes in Computer Science, 1313:40–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:SIO

[642] G. Smith. A semantic integration of object-Z and CSP for the specification of concurrent systems. Lecture Notes in Computer Science, 1313:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikhajlova:1997:CRI

[643] A. Mikhajlova and E. Sekerinski. Class refinement and interface refinement in object-oriented programs. Lecture Notes in Computer Science, 1313:82– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleuker:1997:FRD

[644] S. Kleuker. Formalizing requirements for distributed systems with trace diagrams. Lecture Notes in Computer Science, 1313:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1997:CGS

[645] F. Huber, B. Schaetz, and G. Einert. Consistent graphical specification of distributed systems. Lecture Notes in Computer Science, 1313:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1997:DRC

[646] K. Lano and A. Sanchez. Design of reactive control systems for event-driven operations. Lecture Notes in Computer Science, 1313:142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fleischhack:1997:MNS

[647] H. Fleischhack and J. Tapken. An M-Net semantics for a real-time extension of SDL. Lecture Notes in Computer Science, 1313:162–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1997:RRT

[648] M. Broy, R. Grosu, and C. Klein. Reconciling real-time with asynchronous message passing. Lecture Notes in Computer Science, 1313:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenhoff:1997:SRC

[649] M. Schoenhoff and M. Kowsari. Specifying the remote controlling of valves in an explosion test environment. Lecture Notes in Computer Science, 1313: 201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devauchelle:1997:PPU

[650] L. Devauchelle, P. G. Larsen, and H. Voss. PICGAL: Practical use of formal specification to develop a complex critical system. Lecture Notes in Computer Science, 1313:221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Young:1997:MMA

[651] W. D. Young and W. R. Bevier. Mathematical modeling and analysis of an external memory manager. Lecture Notes in Computer Science, 1313:237– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukherjee:1997:ATV

[652] P. Mukherjee. Automatic translation of VDM-SL specifications into gofer. Lecture Notes in Computer Science, 1313:258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agerholm:1997:TIC

[653] S. Agerholm and J. Frost. Towards an integrated CASE and theorem proving tool for VDM-SL. Lecture Notes in Computer Science, 1313:278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1997:SRN

[654] K. Lano, J. Bicarregui, J. Fiadeiro, and A. Lopes. Specification of required non-determinism. Lecture Notes in Computer Science, 1313:298–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tej:1997:CFD

[655] H. Tej and B. Wolff. A corrected failure-divergence model for CSP in isabelle/HOL. Lecture Notes in Computer Science, 1313:318–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aichernig:1997:POG

[656] B. K. Aichernig and P. G. Larsen. A proof obligation generator for VDM-SL. Lecture Notes in Computer Science, 1313:338–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayadi:1997:VCP

[657] M. M. Ayadi and D. Bolignano. Verification of cryptographic protocols: An experiment. Lecture Notes in Computer Science, 1313:358–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gregoire:1997:TPC

[658] J.-C Gregoire. TLA + PROMELA: Conjecture, check, proof. engineering new protocols using methods and formal notations. Lecture Notes in Computer Science, 1313:378–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokkedem:1997:TSS

[659] A. Mokkedem, M. J. Ferguson, and R. deB. Johnston. A TLA solution to the specification and verification of the RLP1 retransmission protocol. Lecture Notes in Computer Science, 1313:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:ETD

[660] J. M. R. Martin and S. A. Jassim. An efficient technique for deadlock analysis of large scale process networks. Lecture Notes in Computer Science, 1313:418–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yu:1997:IMC

[661] S. Yu and Z. Luo. Implementing a model checker for LEGO. Lecture Notes in Computer Science, 1313:442– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dold:1997:FVT

[662] A. Dold, F. W. Von Henke, H. Pfeifer, and H. Ruess. Formal verification of transformations for peephole optimization. Lecture Notes in Computer Science, 1313:459–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paige:1997:MMF

[663] R. F. Paige. A meta-method for formal method integration. Lecture Notes in Computer Science, 1313:473–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hemer:1997:RVD

[664] D. Hemer and P. A. Lindsay. Reuse of verified design templates through extended pattern matching. Lecture Notes in Computer Science, 1313:495– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1997:CPS

[665] F. S. De Boer, U. Hannemann, and W.-P. De Roever. A compositional proof system for shared variable concurrency. Lecture Notes in Computer Science, 1313:515–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1997:FMF

[666] P. Michel and V. Wiels. A framework for modular formal specification and verification. Lecture Notes in Computer Science, 1313:533–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersohn:1997:TSS

[667] C. Petersohn and L. Urbina. A timed semantics for the STATEMATE implementation of statecharts. Lecture Notes in Computer Science, 1313:553– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stringer-Calvert:1997:UPP

[668] D. W. J. Stringer-Calvert, S. Stepney, and I. Wand. Using PVS to prove a Z refinement: A case study. Lecture Notes in Computer Science, 1313:573– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kellomaeki:1997:VRS

[669] P. Kellomaeki. Verification of reactive systems using DisCo and PVS. Lecture Notes in Computer Science, 1313: 589–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cazier:1997:TRS

[670] D. Cazier and J.-F. Dufourd. Term rewrite systems to derive set Boolean operations on 2D objects. Lecture Notes in Computer Science, 1313:605– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Silva:1997:NFR

[671] L. Silva, A. Sampaio, and E. Barros. A normal form reduction strategy for hardware/software partitioning. Lecture Notes in Computer Science, 1313:624–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boiten:1997:VCZ

[672] E. Boiten, H. Bowman, J. Derrick, and M. Steen. Viewpoint consistency in Z and LOTOS: A case study. Lecture Notes in Computer Science, 1313:644– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charpentier:1997:UMO

[673] M. Charpentier. A unity mapping operator for distributed programs. Lecture Notes in Computer Science, 1313: 665–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mooney:1997:ILP

[674] R. J. Mooney. Inductive logic programming for natural language processing. Lecture Notes in Computer Science, 1314:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1997:IES

[675] S. Muggleton, D. Page, and A. Srinivasan. An initial experiment into stereochemistry-based drug design using inductive logic programming. Lecture Notes in Computer Science, 1314: 25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1997:AID

[676] S. Dzeroski, S. Schulze-Kremer, K. R. Heidtke, and K. Sierns. Applying ILP to diterpene structure elucidation from 0103C NMR spectra. Lecture Notes in Computer Science, 1314:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBaelen:1997:APP

[677] E. Van Baelen and L. De Raedt. Analysis and prediction of piano performances using inductive logic programming. Lecture Notes in Computer Science, 1314:55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamberger:1997:NDE

[678] D. Gamberger and N. Lavrac. Noise detection and elimination applied to noise handling in a KRK chess endgame. Lecture Notes in Computer Science, 1314: 72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1997:FCI

[679] A. Srinivasan and R. D. King. Feature construction with inductive logic programming: A study of quantitative predictions of biological activity aided by structural attributes. Lecture Notes in Computer Science, 1314:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1997:PTL

[680] M. Sebag and C. Rouveirol. Polynomialtime learning in logic programming and constraint logic programming. Lecture Notes in Computer Science, 1314:105– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokai:1997:ALE

[681] G. Kokai, Z. Alexin, and T. Gyimothy. Analyzing and learning ECG waveforms. Lecture Notes in Computer Science, 1314:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizoguchi:1997:LRC

[682] F. Mizoguchi, H. Ohwada, M. Daidoji, and S. Shivato. Learning rules that classify ocular fundus images for glaucoma diagnosis. Lecture Notes in Computer Science, 1314:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1997:NDI

[683] H. Fujita, N. Yagi, T. Ozaki, and K. Furukawa. A new design and implementation of progol by bottom-up computation. Lecture Notes in Computer Science, 1314:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flener:1997:ILP

[684] P. Flener. Inductive logic program synthesis with DIALOGS. Lecture Notes in Computer Science, 1314:175– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blockeel:1997:RKD

[685] H. Blockeel and L. De Raedt. Relational knowledge discovery in databases. Lecture Notes in Computer Science, 1314:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheffer:1997:ETS

[686] T. Scheffer, R. Herbrich, and F. Wysotzki. Efficient theta-subsumption based on graph algorithms. Lecture Notes in Computer Science, 1314:212–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jorge:1997:ICI

[687] A. Jorge and P. B. Brazdil. Integrity constraints in ILP using a Monte Carlo approach. Lecture Notes in Computer Science, 1314:229–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rieger:1997:RCD

[688] A. Rieger. Restructuring chain datalog programs. Lecture Notes in Computer Science, 1314:245–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inuzuka:1997:TIL

[689] N. Inuzuka, M. Kamo, N. Ishii, and H. Seki. Top-down induction of logic programs from incomplete samples. Lecture Notes in Computer Science, 1314:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1997:LGU

[690] S.-H. Nienhuys-Cheng and R. De Wolf. Least generalizations under implication. Lecture Notes in Computer Science, 1314:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pompe:1997:EPE

[691] U. Pompe. Efficient proof encoding. Lecture Notes in Computer Science, 1314:299–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horvath:1997:LLP

[692] T. Horvath, R. H. Sloan, and G. Turan. Learning logic programs with random classification noise. Lecture Notes in Computer Science, 1314:315–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goncalves:1997:HQI

[693] M.-E. Goncalves. Handling quantifiers in ILP. Lecture Notes in Computer Science, 1314:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muggleton:1997:LPD

[694] S. Muggleton. Learning from positive data. Lecture Notes in Computer Science, 1314:358–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markov:1997:LSA

[695] Z. Markov. lambda-subsumption and its application to learning from positive-only examples. Lecture Notes in Computer Science, 1314:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommer:1997:AAD

[696] G. Sommer. Algebraic aspects of designing behaviour based systems. Lecture Notes in Computer Science, 1315: 1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Granlund:1997:MSG

[697] G. H. Granlund. From multidimensional signals to the generation of responses. Lecture Notes in Computer Science, 1315:29–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyden:1997:PAU

[698] A. Heyden, G. Sparr, and K. Aastroem. Perception and action using multilinear forms. Lecture Notes in Computer Science, 1315:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koenderink:1997:LIO

[699] J. J. Koenderink and A. J. Van Doorn. Local image operators and iconic structure. Lecture Notes in Computer Science, 1315:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindeberg:1997:AST

[700] T. Lindeberg. On automatic selection of temporal scales in time-causal scale-space. Lecture Notes in Computer Science, 1315:94–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenz:1997:SAR

[701] R. Lenz. Some applications of representations of lie algebras and lie groups. Lecture Notes in Computer Science, 1315:114–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGool:1997:SAG

[702] L. Van Gool. A systematic approach to geometric-based grouping and nonaccidentalness. Lecture Notes in Computer Science, 1315:126–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buelow:1997:MDS

[703] T. Buelow and G. Sommer. Multidimensional signal processing using an algebraically extended signal representation. Lecture Notes in Computer Science, 1315:148–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chernov:1997:HFD

[704] A. V. Chernov and V. M. Chernov. On Hestenes’ formalization for describing linear image transforms. Lecture Notes in Computer Science, 1315:164– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1997:FBF

[705] W. Skarbek and K. Ignasiak. Fractal basis functions for pattern recognition. Lecture Notes in Computer Science, 1315:176–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shashua:1997:TTF

[706] A. Shashua. Trilinear tensor: The fundamental construct of multiple-view geometry and its applications. Lecture Notes in Computer Science, 1315:190– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Csurka:1997:AGT

[707] G. Csurka and O. Faugeras. Algebraic and geometric tools to compute projective and permutation invariants. Lecture Notes in Computer Science, 1315:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayro-Corrochano:1997:ULC

[708] E. Bayro-Corrochano and J. Lasenby. A unified language for computer vision and robotics. Lecture Notes in Computer Science, 1315:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naeve:1997:STO

[709] A. Naeve. Structure from translational observer motion. Lecture Notes in Computer Science, 1315:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fermueller:1997:GVS

[710] C. Fermueller, L. F. Cheong, and Y. Aloimonos. The geometry of visual space distortion. Lecture Notes in Computer Science, 1315:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tuytelaars:1997:CHT

[711] T. Tuytelaars, M. Proesmans, and L. Van Gool. The cascaded Hough transform as support for grouping and finding vanishing points and lines. Lecture Notes in Computer Science, 1315: 278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchand:1997:VPS

[712] E. Marchand and F. Chaumette. Visual perception strategies for 3D reconstruction. Lecture Notes in Computer Science, 1315:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanatani:1997:SOG

[713] K. Kanatani. Statistical optimization and geometric visual inference. Lecture Notes in Computer Science, 1315: 306–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maybank:1997:PPC

[714] S. J. Maybank and A. D. Worrall. Path prediction and classification based on non-linear filtering. Lecture Notes in Computer Science, 1315:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dorst:1997:BDQ

[715] L. Dorst. Bottom-up derivation of the qualitatively different behaviours of a car across varying spatio-temporal scales: A study in abstraction of goaldirected motion. Lecture Notes in Computer Science, 1315:344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1997:NNA

[716] H. Ritter. Neural network approaches for perception and action. Lecture Notes in Computer Science, 1315:356– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayro-Corrochano:1997:GNN

[717] E. Bayro-Corrochano and S. Buchholz. Geometric neural networks. Lecture Notes in Computer Science, 1315:379– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1997:PED

[718] M. Blum. Program error detection/ correction: Turning PAC learning into perfect learning. Lecture Notes in Computer Science, 1316:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1997:TLG

[719] A. Ambainis, K. Apsitis, R. Freivalds, and W. Gasarch. Team learning as a game. Lecture Notes in Computer Science, 1316:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirowatari:1997:IRR

[720] E. Hirowatari and S. Arikawa. Inferability of recursive real-valued functions. Lecture Notes in Computer Science, 1316:18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1997:LRL

[721] S. Jain, S. Lange, and J. Nessel. Learning of r.e. languages from good examples. Lecture Notes in Computer Science, 1316:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1997:ISH

[722] S. Kobayashi and T. Yokomori. Identifiability of subspaces and homomorphic images of zero-reversible languages. Lecture Notes in Computer Science, 1316:48–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo:1997:POR

[723] C. Domingo, T. Tsukiji, and O. Watanabe. Partial Occam’s razor and its applications. Lecture Notes in Computer Science, 1316:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sitharam:1997:DLB

[724] M. Sitharam and T. Straney. Derandomized learning of Boolean functions. Lecture Notes in Computer Science, 1316:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parekh:1997:LDS

[725] R. Parekh and V. Honavar. Learning DFA from simple examples. Lecture Notes in Computer Science, 1316:116– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denis:1997:PLU

[726] F. Denis and R. Gilleron. PAC learning under helpful distributions. Lecture Notes in Computer Science, 1316: 132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiao:1997:PLU

[727] H. Qiao, N. S. V. Rao, and V. Protopopescu. PAC learning using Nadaraya-Watson estimator based on orthonormal systems. Lecture Notes in Computer Science, 1316:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boros:1997:MEB

[728] E. Boros, T. Ibaraki, and K. Makino. Monotone extensions of Boolean data sets. Lecture Notes in Computer Science, 1316:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:CBH

[729] M. Sato. Classical Brouwer-Heyting-Kolmogorov interpretation. Lecture Notes in Computer Science, 1316:176– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukouchi:1997:ISE

[730] Y. Mukouchi. Inferring a system from examples with time passage. Lecture Notes in Computer Science, 1316:197– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1997:PTI

[731] S. Matsumoto, Y. Hayashi, and T. Shoudai. Polynomial time inductive inference of regular term tree languages from positive data. Lecture Notes in Computer Science, 1316:212–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1997:SNT

[732] J. Case, S. Jain, and A. Sharma. Synthesizing noise-tolerant language learners. Lecture Notes in Computer Science, 1316:228–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1997:EKC

[733] A. Ambainis, K. Apsitis, C. Calude, and R. Freivalds. Effects of Kolmogorov complexity present in inductive inference as well. Lecture Notes in Computer Science, 1316:244–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erlebach:1997:LOV

[734] T. Erlebach, P. Rossmanith, H. Stadtherr, and A. Steger. Learning one-variable pattern languages very efficiently on average, in parallel, and by asking queries. Lecture Notes in Computer Science, 1316:260–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koebler:1997:OSE

[735] J. Koebler and W. Lindner. Oracles in 0p2 are sufficient for exact learning. Lecture Notes in Computer Science, 1316:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arvind:1997:ELT

[736] V. Arvind and N. V. Vinodchandran. Exact learning via teaching assistants. Lecture Notes in Computer Science, 1316:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakamura:1997:EEL

[737] A. Nakamura. An efficient exact learning algorithm for ordered binary decision diagrams. Lecture Notes in Computer Science, 1316:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vovk:1997:PTB

[738] V. Vovk. Probability theory for the brier game. Lecture Notes in Computer Science, 1316:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:LRT

[739] X. Zhang and M. Numao. Learning and revising theories in noisy domains. Lecture Notes in Computer Science, 1316:339–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurvits:1997:NSS

[740] L. Gurvits. A note on a scale-sensitive dimension of linear bounded functionals in Banach spaces. Lecture Notes in Computer Science, 1316:352–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maass:1997:RTN

[741] W. Maass. On the relevance of time in neural computation and learning. Lecture Notes in Computer Science, 1316:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takimoto:1997:SAP

[742] E. Takimoto, K. Hirai, and A. Maruoka. A simple algorithm for predicting nearly as well as the best pruning labeled with the best prediction values of a decision tree. Lecture Notes in Computer Science, 1316:385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwek:1997:LDF

[743] S. Kwek. Learning disjunctions of features. Lecture Notes in Computer Science, 1316:401–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakamoto:1997:LSD

[744] H. Sakamoto. Leaning simple deterministic finite-memory automata. Lecture Notes in Computer Science, 1316: 416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arimura:1997:LAF

[745] H. Arimura. Learning acyclic firstorder Horn sentences from entailment. Lecture Notes in Computer Science, 1316:432–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hegedues:1997:LDZ

[746] T. Hegedues and P. Indyk. On learning disjunctions of zero-one threshold functions with queries. Lecture Notes in Computer Science, 1316:446–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leman:1997:ONC

[747] M. Leman and A. Schneider. Origin and nature of cognitive and systematic musicology: An introduction. Lecture Notes in Computer Science, 1317:13– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louhivuori:1997:SCH

[748] J. Louhivuori. Systematic, cognitive and historical approaches in musicology. Lecture Notes in Computer Science, 1317:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1997:EGQ

[749] M. Weber. Empiricism, gestalt qualities, and determination of style: Some remarks concerning the relationship of Guido Adler to Richard Wallaschek, Alexius Meinong, Christian von Ehrenfels, and Robert Lach. Lecture Notes in Computer Science, 1317:42–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reybrouck:1997:GCM

[750] M. Reybrouck. Gestalt concepts and music: Limitations and possibilities. Lecture Notes in Computer Science, 1317:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eichert:1997:LGT

[751] R. Eichert, L. Schmidt, and U. Seifert. Logic, gestalt theory, and neural computation in research on auditory perceptual organization. Lecture Notes in Computer Science, 1317:70–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godoey:1997:KMT

[752] R. I. Godoey. Knowledge in music theory by shapes of musical objects and sound-producing actions. Lecture Notes in Computer Science, 1317:89– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ungeheuer:1997:SGP

[753] E. Ungeheuer. Statistical gestalts — perceptible features in serial music. Lecture Notes in Computer Science, 1317:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1997:VTF

[754] A. Schneider. “verschmelzung”, tonal fusion, and consonance: Carl stumpf revisited. Lecture Notes in Computer Science, 1317:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leman:1997:SGT

[755] M. Leman and F. Carreras. Schema and gestalt: Testing the hypothesis of psychoneural isomorphism by computer simulation. Lecture Notes in Computer Science, 1317:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petroni:1997:SON

[756] N. C. Petroni and M. Tricarico. Selforganizing neural nets and the perceptual origin of the circle of fifths. Lecture Notes in Computer Science, 1317: 169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parncutt:1997:MPR

[757] R. Parncutt. A model of the perceptual root(s) of a chord accounting for voicing and prevailing tonality. Lecture Notes in Computer Science, 1317:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Werts:1997:GFB

[758] D. Werts. ‘good’, ‘fair’, and ‘bad’ chord progressions: A regression-analysis of some psychological chord progression data obtained in an experiment by J. Bharucha and C. Krumhansl. Lecture Notes in Computer Science, 1317:200– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keuler:1997:PSB

[759] J. Keuler. Problems of shape and background in sounds with inharmonic spectra. Lecture Notes in Computer Science, 1317:214–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eberlein:1997:MAH

[760] R. Eberlein. A method of analysing harmony, based on interval patterns or “gestalten”. Lecture Notes in Computer Science, 1317:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isaacson:1997:NNM

[761] E. Isaacson. Neural network models for the study of post-tonal music. Lecture Notes in Computer Science, 1317:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franek:1997:TRT

[762] M. Franek and J. Mates. Tempo relations: Is there a psychological basis for proportional tempo theory? Lecture Notes in Computer Science, 1317:

253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moelants:1997:FSD

[763] D. Moelants. A framework for the subsymbolic description of meter. Lecture Notes in Computer Science, 1317:263– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cambouropoulos:1997:MRF

[764] E. Cambouropoulos. Musical rhythm: A formal model for determining local boundaries, accents and metre in a melodic surface. Lecture Notes in Computer Science, 1317:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krumhansl:1997:EPO

[765] C. Krumhansl. Effects of perceptual organization and musical form on melodic expectancies. Lecture Notes in Computer Science, 1317:294–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larson:1997:CCS

[766] S. Larson. Continuations as completions: Studying melodic, expectation in the creative microdomain seek well. Lecture Notes in Computer Science, 1317:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toiviainen:1997:OSO

[767] P. Toiviainen. Optimizing selforganizing timbre maps: Two approaches. Lecture Notes in Computer Science, 1317:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rusko:1997:TMG

[768] M. Rusko. Towards a more general understanding of the nasality phenomenon. Lecture Notes in Computer Science, 1317:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1997:KES

[769] C. Reuter. Karl Erich Schumann’s principles of timbre as a helpful tool in stream segregation research. Lecture Notes in Computer Science, 1317: 362–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rochebois:1997:CSU

[770] T. Rochebois and G. Charbonneau. Cross-synthesis using interverted principal harmonic sub-spaces. Lecture Notes in Computer Science, 1317:375– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1997:GPM

[771] D. Cohen and S. Dubnov. Gestalt phenomena in musical texture. Lecture Notes in Computer Science, 1317:386– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanguiane:1997:TIE

[772] A. Tanguiane. Technology of interpretation and expressive pulses. Lecture Notes in Computer Science, 1317:409– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fyk:1997:IPP

[773] J. Fyk. Intonational protention in the performance of melodic octaves on the violin. Lecture Notes in Computer Science, 1317:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davida:1997:ACC

[774] G. Davida, Y. Frankel, Y. Tsiounis, and M. Yung. Anonymity control in E-cash systems. Lecture Notes in Computer Science, 1318:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabber:1997:HMP

[775] E. Gabber, P. B. Gibbons, Y. Matias, and A. Mayer. How to make personalized Web browsing simple, secure, and anonymous. Lecture Notes in Computer Science, 1318:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCoy:1997:ANV

[776] J. McCoy. Anonymous networking and virtual intranets: Tools for anonymous corporations. Lecture Notes in Computer Science, 1318:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Syverson:1997:UST

[777] P. F. Syverson, S. G. Stubblebine, and D. M. Goldschlag. Unlinkable serial transactions. Lecture Notes in Computer Science, 1318:39–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radu:1997:EEC

[778] C. Radu, R. Govaerts, and J. Vandewalle. Efficient electronic cash with restricted privacy. Lecture Notes in Computer Science, 1318:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1997:SC

[779] Y. Zheng. The SPEED cipher. Lecture Notes in Computer Science, 1318: 71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lelieveldt:1997:ESE

[780] S. L. Lelieveldt. Evaluating the security of electronic money. Lecture Notes in Computer Science, 1318:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birch:1997:ECT

[781] D. G. W. Birch and N. A. McEvoy. Electronic cash — technology will denationalise money. Lecture Notes in Computer Science, 1318:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maher:1997:FIA

[782] D. P. Maher. Fault induction attacks, tamper resistance, and hostile reverse engineering in perspective. Lecture Notes in Computer Science, 1318:109– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guillou:1997:SCR

[783] L. C. Guillou. Some critical remarks on “dynamic data authentication” as specified in EMV ’96. Lecture Notes in Computer Science, 1318:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1997:SCI

[784] N. Lange. Single-chip implementation of a cryptosystem for financial applications. Lecture Notes in Computer Science, 1318:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivest:1997:PFC

[785] R. L. Rivest. Perspectives on financial cryptography. Lecture Notes in Computer Science, 1318:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1997:AML

[786] M. K. Franklin and D. Malkhi. Auditable metering with lightweight security. Lecture Notes in Computer Science, 1318:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stem:1997:SFM

[787] J. Stem and S. Vaudenay. SVP: A flexible micropayment scheme. Lecture Notes in Computer Science, 1318:161– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarecki:1997:EMS

[788] S. Jarecki and A. Odlyzko. An efficient micropayment system based on probabilistic polling. Lecture Notes in Computer Science, 1318:173–191, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://www.research.
att.com/~amo/doc/polling.pdf;
http://www.research.att.com/~amo/
doc/polling.ps;   http://www.research.
att.com/~amo/doc/polling.tex.

Yacobi:1997:CBL

[789] Y. Yacobi. On the continuum between on-line and off-line E-cash systems — I. Lecture Notes in Computer Science, 1318:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagnia:1997:TMP

[790] H. Pagnia and R. Jansen. Towards multiple-payment schemes for digital money. Lecture Notes in Computer Science, 1318:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobsson:1997:AAT

[791] M. Jakobsson and M. Yung. Applying anti-trust policies to increase trust in a versatile E-money system. Lecture Notes in Computer Science, 1318: 217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Swire:1997:ULF

[792] P. P. Swire. The uses and limits of financial cryptography: A law professor’s perspective. Lecture Notes in Computer Science, 1318:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radlo:1997:LIC

[793] E. J. Radlo. Legal issues in cryptography. Lecture Notes in Computer Science, 1318:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froomkin:1997:DST

[794] A. M. Froomkin. Digital signatures today. Lecture Notes in Computer Science, 1318:287–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merrill:1997:ARD

[795] C. R. Merrill. An attorney’s roadmap to the digital signature guidelines. Lecture Notes in Computer Science, 1318: 291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wright:1997:AVL

[796] B. Wright. Alternative visions for legal signatures and evidence. Lecture Notes in Computer Science, 1318:299– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wayner:1997:MLP

[797] P. C. Wayner. Money laundering: Past, present and future. Lecture Notes in Computer Science, 1318:301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rivest:1997:ELT

[798] R. L. Rivest. Electronic lottery tickets as micropayments. Lecture Notes in Computer Science, 1318:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lampru:1997:STG

[799] P. Lampru. Strategic tasks for government in the Information Age. Lecture Notes in Computer Science, 1318:315– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grigg:1997:UEM

[800] I. Grigg and C. C. Petro. Using electronic markets to achieve efficient task distribution. Lecture Notes in Computer Science, 1318:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldstein:1997:GSM

[801] T. Goldstein. The gateway security model in the Java electronic commerce framework. Lecture Notes in Computer Science, 1318:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kravitz:1997:HSL

[802] D. W. Kravitz. Highly scalable on-line payments via task decoupling. Lecture Notes in Computer Science, 1318:355– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1997:GGU

[803] B. Fox, B. Beckman, and D. Simon. GUMP: Grand unified meta-protocols recipes for simple, standards-based financial cryptography. Lecture Notes in Computer Science, 1318:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esslinger:1997:SNC

[804] B. Esslinger and J. Schneider. Secure network communications and secure store and forward mechanisms with the SAP R/3 system. Lecture Notes in Computer Science, 1318:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beydoun:1997:ASK

[805] G. Beydoun and A. Hoffmann. Acquisition of search knowledge. Lecture Notes in Computer Science, 1319:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beys:1997:SAF

[806] P. Beys and M. Van Someren. A systematic approach to the functionality of problem-solving methods. Lecture Notes in Computer Science, 1319:17– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borst:1997:OAP

[807] P. Borst and H. Akkermans. An ontology approach to product disassembly. Lecture Notes in Computer Science, 1319:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boswell:1997:KRD

[808] R. Boswell, S. Craw, and R. Rowe. Knowledge refinement for a design system. Lecture Notes in Computer Science, 1319:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cornelissen:1997:CVK

[809] F. Cornelissen, C. M. Jonker, and J. Treur. Compositional verification of knowledge-based systems: A case study for diagnostic reasoning. Lecture Notes in Computer Science, 1319: 65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1997:ERS

[810] S. Decker, M. Daniel, M. Erdmann, and R. Studer. An enterprise reference scheme for integrating model based knowledge engineering and enterprise modelling. Lecture Notes in Computer Science, 1319:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fensel:1997:TAM

[811] D. Fensel. The tower-of-adapter method for developing and reusing problem-solving methods. Lecture Notes in Computer Science, 1319:97– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fensel:1997:UOD

[812] D. Fensel, E. Motta, S. Decker, and Z. Zdrahal. Using ontologies for defining tasks, problem-solving methods and their mappings. Lecture Notes in Computer Science, 1319:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1997:EES

[813] U. Hahn and K. Schnattinger. An empirical evaluation of a system for text knowledge acquisition. Lecture Notes in Computer Science, 1319:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1997:ACS

[814] D. M. Jones and R. C. Paton. Acquisition of conceptual structure in scientific theories. Lecture Notes in Computer Science, 1319:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martinez-Bejar:1997:DOC

[815] R. Martinez-Bejar, V. R. Benjamins, and F. Martin-Rubio. Designing operators for constructing domain knowledge ontologies. Lecture Notes in Computer Science, 1319:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perkuhn:1997:RPS

[816] R. Perkuhn. Reuse of problem-solving methods and family resemblances. Lecture Notes in Computer Science, 1319: 174–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierret-Golbreich:1997:SFK

[817] C. Pierret-Golbreich and X. Talon. Specification of flexible knowledgebased systems. Lecture Notes in Computer Science, 1319:190–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pos:1997:PSR

[818] A. Pos, H. Akkermans, and R. Straatman. Problem solving for redesign. Lecture Notes in Computer Science, 1319: 205–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynaud:1997:NRC

[819] C. Reynaud, N. Aussenac-Gilles, P. Tchounikine, and F. Trichet. The notion of role in conceptual modeling. Lecture Notes in Computer Science, 1319:221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richards:1997:KAF

[820] D. Richards and P. Compton. Knowledge acquisition first, modelling later. Lecture Notes in Computer Science, 1319:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:KAO

[821] B. D. Smith, K. Rajan, and N. Muscettola. Knowledge acquisition for the onboard planner of an autonomous spacecraft. Lecture Notes in Computer Science, 1319:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschaitschian:1997:ITK

[822] B. Tschaitschian, A. Abecker, and F. Schmalhofer. Information tuning with KARAT: Capitalizing on existing documents. Lecture Notes in Computer Science, 1319:269–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallis:1997:SPK

[823] S. Wallis and G. Nelson. Syntactic parsing as a knowledge acquisition problem. Lecture Notes in Computer Science, 1319:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zarri:1997:BMU

[824] G. P. Zarri and S. Azzam. Building up and making use of corporate knowledge repositories. Lecture Notes in Computer Science, 1319:301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazier:1997:MCC

[825] F. Brazier, P. Van Eck, and J. Treur. Modelling competitive co-operation of agents in a compositional multi-agent framework. Lecture Notes in Computer Science, 1319:317–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazier:1997:IPD

[826] F. M. T. Brazier and N. J. E. Wijngaards. An instrument for a purpose driven comparison of modelling frameworks. Lecture Notes in Computer Science, 1319:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breidenstein:1997:KDR

[827] T. Breidenstein, I. Bournaud, and F. Wolinski. Knowledge discovery in rule bases. Lecture Notes in Computer Science, 1319:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coupe:1997:SPE

[828] V. M. H. Coupe and L. C. Van der Gaag. Supporting probability elicitation by sensitivity analysis. Lecture Notes in Computer Science, 1319:335– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabriel:1997:TMC

[829] J.-M. Gabriel. Towards more collaboration between machine learning systems and their users. Lecture Notes in Computer Science, 1319:341–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia:1997:CNS

[830] D. Garcia. COATIS, an NLP system to locate expressions of actions connected by causality links. Lecture Notes in Computer Science, 1319:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaspar:1997:UKA

[831] S. Kaspar and A. Hoffmann. Using knowledge acquisition to build spoken language systems. Lecture Notes in Computer Science, 1319:353–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maillet-Contoz:1997:ADP

[832] L. Maillet-Contoz and J. Sallantin. An assistant for a design project: Application to the design of a mixed hardware/ software architecture. Lecture Notes in Computer Science, 1319:359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poittevin:1997:RIR

[833] L. Poittevin. REVINOS: An interactive revision tool based on the concept of situation. Lecture Notes in Computer Science, 1319:365–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Straatman:1997:KK

[834] R. Straatman. KIDS for KADS. Lecture Notes in Computer Science, 1319: 371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talavera:1997:EIB

[835] L. Talavera and U. Cortes. Exploiting inductive bias shift in knowledge acquisition from ill-structured domains. Lecture Notes in Computer Science, 1319:377–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kingston:1997:DHP

[836] J. Kingston. The DARPA high performance knowledge bases programme. Lecture Notes in Computer Science, 1319:383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1997:TFT

[837] F. B. Schneider. Towards fault-tolerant and secure agentry. Lecture Notes in Computer Science, 1320:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reuter:1997:WME

[838] A. Reuter. Workflow management: An exercise in distributed computing. Lecture Notes in Computer Science, 1320: 15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pogosyants:1997:VRC

[839] A. Pogosyants, R. Segala, and N. Lynch. Verification of the randomized consensus algorithm of aspnes and herlihy: A case study. Lecture Notes in Computer Science, 1320:22–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eilam:1997:SDB

[840] T. Eilam, S. Moran, and S. Zaks. A simple DFS-Based algorithm for linear interval routing. Lecture Notes in Computer Science, 1320:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1997:ALB

[841] M. Flammini, E. Nardelli, and G. Proietti. ATM layouts with bounded hop count and congestion. Lecture Notes in Computer Science, 1320:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lui:1997:SSN

[842] K.-S. Lui and S. Zaks. Scheduling in synchronous networks and the greedy algorithm. Lecture Notes in Computer Science, 1320:66–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehrke:1997:RCL

[843] J. E. Gehrke, C. G. Plaxton, and R. Rajaraman. Rapid convergence of a local load balancing algorithm for asynchronous rings. Lecture Notes in Computer Science, 1320:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chlebus:1997:PTR

[844] B. S. Chlebus, R. De Prisco, and A. A. Shvartsman. Performing tasks on restartable message-passing processors. Lecture Notes in Computer Science, 1320:96–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DePrisco:1997:RPA

[845] R. De Prisco, B. Lampson, and N. Lynch. Revisiting the Paxos algorithm. Lecture Notes in Computer Science, 1320:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aguilera:1997:HTF

[846] M. K. Aguilera, W. Chen, and S. Toueg. Heartbeat: A timeoutfree failure detector for quiescent reliable communication. Lecture Notes in Computer Science, 1320:126–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1997:GAM

[847] R. Guerraoui and A. Schiper. Genuine atomic multicast. Lecture Notes in Computer Science, 1320:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:1997:LOT

[848] S. Katz, P. Lincoln, and J. Rushby. Low-overhead time-triggered group membership. Lecture Notes in Computer Science, 1320:155–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helary:1997:VPA

[849] J.-M. Helary, A. Mostefaoui, and M. Raynal. Virtual precedence in asynchronous systems: Concept and applications. Lecture Notes in Computer Science, 1320:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoller:1997:DGP

[850] S. D. Stoller. Detecting global predicates in distributed systems with clocks. Lecture Notes in Computer Science, 1320:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

James:1997:FTB

[851] J. James and A. K. Singh. Fault tolerance bounds for memory consistency. Lecture Notes in Computer Science, 1320:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luchangco:1997:PBM

[852] V. Luchangco. Precedence-based memory models. Lecture Notes in Computer Science, 1320:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joung:1997:SIF

[853] Y.-J. Joung and J.-Y. Liao. Strong interaction fairness in a fully distributed system with unbounded speed variability. Lecture Notes in Computer Science, 1320:230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beimel:1997:RCP

[854] A. Beimel and M. Franklin. Reliable communication over partially authenticated networks. Lecture Notes in Computer Science, 1320:245–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnen:1997:SSD

[855] C. Johnen, G. Alari, J. Beauquier, and A. K. Datta. Self-stabilizing depth-first token passing on rooted networks. Lecture Notes in Computer Science, 1320:260–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garay:1997:SDS

[856] J. A. Garay, R. Gennaro, C. Jutla, and T. Rabin. Secure distributed storage and retrieval. Lecture Notes in Computer Science, 1320:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inoue:1997:OWF

[857] M. Inoue, S. Moriya, T. Masuzawa, and H. Fujiwara. Optimal waitfree clock synchronization protocol on a shared-memory multi-processor system. Lecture Notes in Computer Science, 1320:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moir:1997:TSW

[858] M. Moir. Transparent support for waitfree transactions. Lecture Notes in Computer Science, 1320:305–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jayanti:1997:PMO

[859] P. Jayanti and S. Khanna. On the power of multi-objects. Lecture Notes in Computer Science, 1320:320–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neri:1997:MCC

[860] F. Neri and L. Saitta. Modeling conceptual change: An interdisciplinary approach. Lecture Notes in Computer Science, 1321:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botta:1997:RNT

[861] M. Botta, A. Giordana, and R. Piola. Refining numerical terms in Horn clauses. Lecture Notes in Computer Science, 1321:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malerba:1997:HCD

[862] D. Malerba, F. Esposito, G. Semeraro, and S. Caggese. Handling continuous data in top-down induction of first-order rules. Lecture Notes in Computer Science, 1321:24–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frasconi:1997:IIT

[863] P. Frasconi, M. Gori, M. Maggini, and E. Martinelli. Inductive inference of tree automata by recursive neural networks. Lecture Notes in Computer Science, 1321:36–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ardissono:1997:CMM

[864] L. Ardissono, G. Boella, and R. Damiano. A computational model of misunderstandings in agent communication. Lecture Notes in Computer Science, 1321:48–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbero:1997:WCL

[865] G. Barbero and V. Lombardo. Widecoverage lexicalized grammars. Lecture Notes in Computer Science, 1321:60– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barboni:1997:FRC

[866] P. Barboni and D. Sestero. Flexible response choice using problem-solving plans and rethorical relations. Lecture Notes in Computer Science, 1321:72– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nederhof:1997:VEP

[867] M.-J. Nederhof and G. Satta. A variant of Earley parsing. Lecture Notes in Computer Science, 1321:84–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piaggio:1997:ARN

[868] M. Piaggio, A. Sgorbissa, G. Vercelli, and R. Zaccaria. Autonomous robot navigation using a reactive agent. Lecture Notes in Computer Science, 1321: 96–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chella:1997:HAA

[869] A. Chella, S. Gaglio, G. Sajeva, and F. Torterolo. A hybrid architecture for autonomous agents. Lecture Notes in Computer Science, 1321:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zagrouba:1997:DFC

[870] E. Zagrouba. 3-D facets construction for stereovision. Lecture Notes in Computer Science, 1321:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baldoni:1997:EDF

[871] M. Baldoni, U. Baroglio, D. Cavagnino, and G. Lo Bello. Extraction of discriminant features from image fractal encoding. Lecture Notes in Computer Science, 1321:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armano:1997:LRC

[872] G. Armano and G. Fumera. Learning relational concepts at different levels of granularity. Lecture Notes in Computer Science, 1321:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpineto:1997:IMR

[873] C. Carpineto and G. Romano. Inferring minimal rule covers from relations. Lecture Notes in Computer Science, 1321:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basili:1997:CDU

[874] R. Basili, M. T. Pazienza, and M. Vindigni. Corpus-driven unsupervised learning of verb subcategorization frames. Lecture Notes in Computer Science, 1321:159–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokai:1997:LSS

[875] G. Kokai, J. Csirik, and T. Gyimothy. Learning the syntax and semantic rules of an ECG grammar. Lecture Notes in Computer Science, 1321:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamma:1997:IAE

[876] E. Lamma, P. Mello, M. Milano, and F. Riguzzi. Introducing abduction into (extensional) inductive logic programming system. Lecture Notes in Computer Science, 1321:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusoni:1997:EAT

[877] V. Brusoni, L. Console, P. Terenziani, and D. Theseider Dupre. An efficient algorithm for temporal abduction. Lecture Notes in Computer Science, 1321:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadoli:1997:EAC

[878] M. Cadoli, A. Giovanardi, and M. Schaerf. Experimental analysis of the computational cost of evaluating quantified Boolean formulae. Lecture Notes in Computer Science, 1321:207– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massacci:1997:PTT

[879] F. Massacci. A proof theory for tractable approximations of propositional reasoning. Lecture Notes in Computer Science, 1321:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosati:1997:EMK

[880] R. Rosati. Embedding minimal knowledge into autoepistemic logic. Lecture Notes in Computer Science, 1321:231– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asnicar:1997:UMB

[881] F. A. Asnicar, M. Di Fant, and C. Tasso. User model-based information filtering. Lecture Notes in Computer Science, 1321:242–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portinale:1997:CAH

[882] L. Portinale and P. Torasso. A comparative analysis of Horn models and Bayesian networks for diagnosis. Lecture Notes in Computer Science, 1321: 254–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangineto:1997:MAN

[883] E. Sangineto. Multi-agent negotiation and planning through knowledge contextualization. Lecture Notes in Computer Science, 1321:266–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1997:TDR

[884] C. Castelfranchi and R. Falcone. From task delegation to role delegation. Lecture Notes in Computer Science, 1321: 278–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monekosso:1997:ARB

[885] N. D. Monekosso and P. Remagnino. Automated reasoning on-board autonomous spacecrafts. Lecture Notes in Computer Science, 1321:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oddi:1997:WBS

[886] A. Oddi. A weakly backjumping strategy to solve hard scheduling problems. Lecture Notes in Computer Science, 1321:302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baioletti:1997:CTN

[887] M. Baioletti, S. Marcugini, and A. Milani. Compiling task networks into partial order planning domains. Lecture Notes in Computer Science, 1321:311– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cancedda:1997:HAH

[888] N. Cancedda, G. Kamstrup, E. Pianta, and E. Pietrosanti. A hybrid approach to hypertext generation. Lecture Notes in Computer Science, 1321:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCarolis:1997:GUA

[889] B. De Carolis, F. De Rosis, and S. Pizzutillo. Generating user-adapted hypermedia from discourse plans. Lecture Notes in Computer Science, 1321: 334–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Artale:1997:WIU

[890] A. Artale, B. Magnini, and C. Strapparava. WordNet for Italian and its use for lexical discrimination. Lecture Notes in Computer Science, 1321:346– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaspari:1997:ESR

[891] M. Gaspari. Efficient support for reactive rules in prolog. Lecture Notes in Computer Science, 1321:357–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compatangelo:1997:RBK

[892] E. Compatangelo, F. M. Donini, and G. Rumolo. Reasoning with behavioural knowledge in application domain models. Lecture Notes in Computer Science, 1321:369–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madalinska-Bugaj:1997:HSQ

[893] E. Madalinska-Bugaj. How to solve qualification and ramification using Dijkstra’s semantics for programming languages. Lecture Notes in Computer Science, 1321:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholezyk:1997:TQR

[894] D. Pacholezyk. Towards a qualitative representation of linguistic negation of nuanced properties. Lecture Notes in Computer Science, 1321:393–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avesani:1997:CCB

[895] P. Avesani, A. Perini, and F. Ricci. CBET: A case base exploration tool. Lecture Notes in Computer Science, 1321:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonzano:1997:LFW

[896] A. Bonzano, P. Cunningham, and B. Smyth. Learning feature weights for CBR: Global versus local. Lecture Notes in Computer Science, 1321:417– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Attardi:1997:CSP

[897] G. Attardi, A. Cisternino, and M. Simi. CompAss: a system for plans of study compilation. Lecture Notes in Computer Science, 1321:427–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balderi:1997:PSD

[898] A. Balderi and M. Solimano. A prototypal system for data-validation. Lecture Notes in Computer Science, 1321: 431–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergamaschi:1997:OTD

[899] S. Bergamaschi, C. Sartori, D. Beneventano, and M. Vincini. ODB-Tools: a description logics based tool for schema validation and semantic query optimization in object oriented databases. Lecture Notes in Computer Science, 1321:435–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malerba:1997:PPD

[900] D. Malerba, F. Esposito, G. Semeraro, and L. De Filippis. Processing paper documents with WISDOM. Lecture Notes in Computer Science, 1321: 439–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neri:1997:RFC

[901] F. Neri. REGAL3.2: FOL concept learning by cooperative genetic algorithms. Lecture Notes in Computer Science, 1321:443–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semeraro:1997:IPI

[902] G. Semeraro, F. Esposito, D. Malerba, and N. Fanizzi. IDL: A prototypical intelligent digital library service. Lecture Notes in Computer Science, 1321:447–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazzica:1997:OOA

[903] P. Bazzica, G. Casonato, and A. Cesta. An object-oriented architecture for the DRS scheduling problem. Lecture Notes in Computer Science, 1321:451– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brancaleoni:1997:MPA

[904] R. Brancaleoni, A. Cesta, and D. D’Aloisi. MASMA: A personal assistant for meetings management. Lecture Notes in Computer Science, 1321: 455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:FPR

[905] M. Fuchs. Flexible proof-replay with heuristics. Lecture Notes in Computer Science, 1323:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:FRE

[906] M. Fuchs. Flexible re-enactment of proofs. Lecture Notes in Computer Science, 1323:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1997:IRC

[907] D. Fuchs. Inference rights for controlling search in generating theorem provers. Lecture Notes in Computer Science, 1323:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomes:1997:RME

[908] P. Gomes and C. Bento. A retrieval method for exploration of a case memory. Lecture Notes in Computer Science, 1323:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hahn:1997:ICE

[909] U. Hahn and M. Klenner. Incremental concept evolution based on adaptive feature weighting. Lecture Notes in Computer Science, 1323:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ettaouil:1997:QKP

[910] M. Ettaouil. A 0 − 1 quadratic knapsack problem for modelizing and solving the constraint satisfaction problems. Lecture Notes in Computer Science, 1323:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomas:1997:ASS

[911] A. P. Tomas and M. Filgueiras. An algorithm for solving systems of linear Diophantine equations in naturals. Lecture Notes in Computer Science, 1323:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smirnov:1997:GNA

[912] Y. Smirnov and M. M. Veloso. Gen-SAT: A navigational approach. Lecture Notes in Computer Science, 1323: 85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soares:1997:TUD

[913] P. Soares and N. J. Mamede. Timetabling using demand profiles. Lecture Notes in Computer Science, 1323:97–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1997:IVT

[914] R. Lu, K. Han, Y. Ma, and W. Zhang. Intelligent VR training. Lecture Notes in Computer Science, 1323:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choquet:1997:TSK

[915] C. Choquet, P. Tchounikine, and F. Trichet. Training strategies and knowledge acquisition: Using the same reflective tools for different purposes. Lecture Notes in Computer Science, 1323:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1997:AIM

[916] D. Pacholczyk. About the intended meaning of a linguistic negation. Lecture Notes in Computer Science, 1323: 131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinto:1997:IIS

[917] H. S. Pinto. Integration of inheritance in SNePS. Lecture Notes in Computer Science, 1323:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vejnarova:1997:MUI

[918] J. Vejnarova. Measures of uncertainty and independence concept in different calculi. Lecture Notes in Computer Science, 1323:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bittencourt:1997:MAA

[919] G. Bittencourt and I. Tonin. A multiagent approach to first-order logic. Lecture Notes in Computer Science, 1323: 167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaCostaMora:1997:MDA

[920] M. Da Costa Mora, G. Pereira Lopes, H. Coelho, and R. M. Viccari. Modelling dynamic aspects of intentions. Lecture Notes in Computer Science, 1323:179–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fonseca:1997:MAN

[921] J. M. Fonseca, E. De Oliveira, and A. Steiger-Garcao. Multi-agent negotiation algorithms for resources cost estimation: A case study. Lecture Notes in Computer Science, 1323:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesar:1997:PCL

[922] A. Cesar, P. L. Da Costa, and G. Bittencourt. Parla: A cooperation language for cognitive multi-agent systems. Lecture Notes in Computer Science, 1323:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder:1997:VAA

[923] M. Schroeder, I. Mora, and J. J. Alferes. Vivid agents arguing about distributed extended logic programs. Lecture Notes in Computer Science, 1323:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prendinger:1997:AR

[924] H. Prendinger. Approximate reasoning. Lecture Notes in Computer Science, 1323:229–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farreny:1997:NRA

[925] H. Farreny. New results about subadmissibility for general families of heuristic search algorithms. Lecture Notes in Computer Science, 1323:241– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takac:1997:FPC

[926] M. Takac. Fixed point classification method for qualitative simulation. Lecture Notes in Computer Science, 1323:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matos:1997:CLC

[927] P. A. Matos and J. P. Martins. Contextual logic of change and the ramification problem. Lecture Notes in Computer Science, 1323:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schacht:1997:TRA

[928] S. Schacht and U. Hahn. Temporal reasoning about actor systems. Lecture Notes in Computer Science, 1323: 279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaFonsecaBastosGomes:1997:CMJ

[929] N. F. Da Fonseca Bastos Gomes. A CLP model to the job sequencing problem. Lecture Notes in Computer Science, 1323:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castellanos:1997:NAE

[930] A. L. Castellanos, J. Castellanos, D. Manrique, and A. Martinez. A new approach for extracting rules from a trained neural network. Lecture Notes in Computer Science, 1323:297– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larranaga:1997:BNR

[931] P. Larrañaga, M. J. Gallego, B. Sierra, and L. Urkola. Bayesian networks, rule induction and logistic regression in the prediction of the survival of women suffering from breast cancer. Lecture Notes in Computer Science, 1323:303– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cox:1997:CUG

[932] M. T. Cox and M. M. Veloso. Controlling for unexpected goals when planning in a mixed-initiative setting. Lecture Notes in Computer Science, 1323: 309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BhanuPrasad:1997:CMS

[933] P. V. S. R. Bhanu Prasad and D. Khemani. Cooperative memory structures and commonsense knowledge for planning. Lecture Notes in Computer Science, 1323:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pellegrini:1997:DTR

[934] J. Pellegrini and J. Wainer. Diagonalization and type rewriting in clam. Lecture Notes in Computer Science, 1323:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poria:1997:GE

[935] S. Poria and R. Garigliano. Granularity for explanation. Lecture Notes in Computer Science, 1323:331–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jerinic:1997:OMI

[936] L. Jerinic, V. Devedzic, M. Lomic, and D. Radovic. Object model of intelligent tutoring shell. Lecture Notes in Computer Science, 1323:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cunha:1997:RAA

[937] A. Cunha and O. Belo. Resource allocation on agent meta-societies. Lecture Notes in Computer Science, 1323: 343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dressier:1997:DIY

[938] O. Dressier. Diagnostic information at your fingertips! Lecture Notes in Computer Science, 1323:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1997:RAA

[939] L. M. Pereira and R. Li. Reasoning about actions with abductive logic programming. Lecture Notes in Computer Science, 1323:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varela:1997:DEP

[940] F. J. Varela. Dimensions of embodiments: Possible futures for cognitive science. Lecture Notes in Computer Science, 1323:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1997:MLM

[941] T. Mitchell. Machine learning meets natural language. Lecture Notes in Computer Science, 1323:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stefik:1997:LDP

[942] M. Stefik and G. Lavendel. Libraries and digital property rights. Lecture Notes in Computer Science, 1324:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1997:ODS

[943] S. Abiteboul. Object database support for digital libraries. Lecture Notes in Computer Science, 1324:11–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Procter:1997:ECC

[944] R. Procter, A. McKinlay, A. Goldenberg, and E. Davenport. Enhancing community and collaboration in the virtual library. Lecture Notes in Computer Science, 1324:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slaughter:1997:COR

[945] L. Slaughter, B. Shneiderman, and G. Marchionini. Comprehension and object recognition capabilities for presentations of simultaneous video key frame surrogates. Lecture Notes in Computer Science, 1324:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

French:1997:ACA

[946] J. C. French, A. L. Powell, J. L. Pfaltz, and E. Schulman. Automating the construction of authority files in digital libraries: A case study. Lecture Notes in Computer Science, 1324:55–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tochtermann:1997:USG

[947] K. Tochtermann, W.-F. Riekert, and G. Wiest. Using semantic, geographical, and temporal relationships to enhance search and retrieval in digital catalogs. Lecture Notes in Computer Science, 1324:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iannella:1997:MRU

[948] R. Iannella. Metadata repositories using PICS. Lecture Notes in Computer Science, 1324:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smeaton:1997:RFQ

[949] A. F. Smeaton and F. Crimmins. Relevance feedback and query expansion for searching the Web: A model for searching a digital library. Lecture Notes in Computer Science, 1324:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponte:1997:TST

[950] J. M. Ponte and W. B. Croft. Text segmentation by topic. Lecture Notes in Computer Science, 1324:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawking:1997:STR

[951] D. Hawking. Scalable text retrieval for large digital libraries. Lecture Notes in Computer Science, 1324:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crespo:1997:ASD

[952] A. Crespo and H. Garcia-Molina. Awareness services for digital libraries. Lecture Notes in Computer Science, 1324:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolaou:1997:TCI

[953] C. Nikolaou, M. Marazakis, D. Papadakis, and Y. Yeorgiannakis. Towards a common infrastructure for large-scale distributed applications. Lecture Notes in Computer Science, 1324:173–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Semeraro:1997:MLL

[954] G. Semeraro, F. Esposito, D. Malerba, and N. Fanizzi. Machine learning + online libraries = IDL. Lecture Notes in Computer Science, 1324:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dartois:1997:BME

[955] M. Dartois, T. Sakaguchi, S. Sugimoto, and K. Tabata. Building a multilingual electronic text collection of folk tales as a set of encapsulated document objects: An approach for casual users to browse multilingual documents on the fly. Lecture Notes in Computer Science, 1324:215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferber:1997:AIT

[956] R. Ferber. Automated indexing with thesaurus descriptors: A co-occurrence based approach to multilingual retrieval. Lecture Notes in Computer Science, 1324:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sheridan:1997:CLI

[957] P. Sheridan, M. Braschler, and P. Schaeuble. Cross-language information retrieval in a multilingual legal domain. Lecture Notes in Computer Science, 1324:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozzi:1997:DLC

[958] A. Bozzi and S. Calabretto. The digital library and computational philology: The BAMBI project. Lecture Notes in Computer Science, 1324:269–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Phelps:1997:MA

[959] T. A. Phelps and R. Wilensky. Multivalent annotations. Lecture Notes in Computer Science, 1324:287–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christophides:1997:SNA

[960] V. Christophides, M. Doerr, and I. Fundulaki. A semantic network approach to semi-structured documents repositories. Lecture Notes in Computer Science, 1324:305–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meghini:1997:MRS

[961] C. Meghini, F. Sebastiani, and U. Straccia. Modelling the retrieval of structured documents containing texts and images. Lecture Notes in Computer Science, 1324:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harding:1997:PRO

[962] S. M. Harding, W. B. Croft, and C. Weir. Probabilistic retrieval of OCR degraded text using N-grams. Lecture Notes in Computer Science, 1324:345– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noordermeer:1997:DDE

[963] T. C. Noordermeer. Deposit for Dutch electronic publications: Research and practice in the netherlands. Lecture Notes in Computer Science, 1324:361– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breu:1997:CDL

[964] M. Breu and R. Weber. Charging for a digital library — the business model and the cost models of the MeDoc digital library. Lecture Notes in Computer Science, 1324:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mony:1997:BNF

[965] S. Mony. Bibliothèque Nationale de France’s audiovisual system: Digital audio, video, and photo consultation in a library. Lecture Notes in Computer Science, 1324:387–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bern:1997:ECC

[966] J. Bern, C. Damm, and C. Meinel. The electronic colloquium on computational complexity (ECCC): A digital library in use. Lecture Notes in Computer Science, 1324:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biermann:1997:GOM

[967] A. W. Biermann, C. A. Guinn, M. Fulkerson, and G. A. Keim. Goaloriented multimedia dialogue with variable initiative. Lecture Notes in Computer Science, 1325:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1997:KBI

[968] W. W. Chu, A. F. Cardenas, and R. K. Taira. Knowledge-based image retrieval with spatial and temporal constructs. Lecture Notes in Computer Science, 1325:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lowry:1997:VVO

[969] M. Lowry, K. Havelund, and J. Penix. Verification and validation oral systems that control deep-space spacecraft. Lecture Notes in Computer Science, 1325:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piatetsky-Shapiro:1997:DMK

[970] G. Piatetsky-Shapiro. Data mining and knowledge discovery: The third generation. Lecture Notes in Computer Science, 1325:48–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:EPC

[971] J. Chen. Embedding prioritized circumscription in logic programs. Lecture Notes in Computer Science, 1325: 50–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conrad:1997:ETL

[972] S. Conrad and G. Saake. Extending temporal logic for capturing evolving behaviour. Lecture Notes in Computer Science, 1325:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambrix:1997:DLM

[973] P. Lambrix and L. Padgham. A description logic model for querying knowledge bases for structured documents. Lecture Notes in Computer Science, 1325:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Botta:1997:RFO

[974] M. Botta, A. Giordana, and R. Piola. Refining first order theories with neural networks. Lecture Notes in Computer Science, 1325:84–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhong:1997:WIB

[975] N. Zhong, C. Liu, and S. Ohsuga. A way of increasing both autonomy and versatility of a KDD system. Lecture Notes in Computer Science, 1325:94– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barber:1997:CAS

[976] B. Barber and H. J. Hamilton. A comparison of attribute selection strategies for attribute-oriented generalization. Lecture Notes in Computer Science, 1325:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1997:BRF

[977] H. S. Nguyen and A. Skowron. Boolean reasoning for feature extraction problems. Lecture Notes in Computer Science, 1325:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saitta:1997:WMC

[978] L. Saitta, F. Neri, and A. Tiberghien. Worm model construction in children during physics learning. Lecture Notes in Computer Science, 1325:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feldman:1997:DED

[979] R. Feldman, W. Kloesgen, and A. Zilberstein. Document explorer: discovering knowledge in document collections. Lecture Notes in Computer Science, 1325:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Elomaa:1997:WBE

[980] T. Elomaa and J. Rousu. Well-behaved evaluation functions for numerical attributes. Lecture Notes in Computer Science, 1325:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacprzyk:1997:IIL

[981] J. Kacprzyk and G. Szkatuta. An improved inductive learning algorithm with a preanalysis of data. Lecture Notes in Computer Science, 1325:157– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:EIN

[982] L. Martin and C. Vrain. Efficient induction of numerical constraints. Lecture Notes in Computer Science, 1325: 167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:LES

[983] J. Zhang and H. J. Hamilton. Learning English syllabification for words. Lecture Notes in Computer Science, 1325: 177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoon:1997:TCQ

[984] S. C. Yoon. Towards conceptual query answering. Lecture Notes in Computer Science, 1325:187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaub:1997:IPQ

[985] T. Schaub and P. Nicolas. An implementation platform for queryanswering in default logics: theoretical underpinnings. Lecture Notes in Computer Science, 1325:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Singh:1997:ASD

[986] M. Singh, J. Barnett, and M. P. Singh. Automating spoken dialogue systems. Lecture Notes in Computer Science, 1325:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skarbek:1997:HDR

[987] W. Skarbek and K. Ignasiak. Handwritten digit recognition by local principal components analysis. Lecture Notes in Computer Science, 1325:217– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1997:CMI

[988] J. Debenham. From conceptual model to internal model. Lecture Notes in Computer Science, 1325:227–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:SGB

[989] Q. Zhang. Speeding GA-based attribute selection for image interpretation. Lecture Notes in Computer Science, 1325:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadad:1997:URS

[990] B. S. Hadad and C. F. Eick. Using recurrent selection to improve GA performance. Lecture Notes in Computer Science, 1325:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giordana:1997:CAC

[991] A. Giordana, L. Saitta, and G. Lo Bello. A coevolutionary approach to concept learning. Lecture Notes in Computer Science, 1325:257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Federman:1997:RML

[992] F. Federman and S. F. Dorchak. Representation of music in a learning classifier system. Lecture Notes in Computer Science, 1325:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLaer:1997:MCP

[993] W. Van Laer, L. De Raedt, and S. Dzeroski. On multi-class problems and discretization in inductive logic programming. Lecture Notes in Computer Science, 1325:277–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:DEE

[994] K.-M. Huang and J. M. Zytkow. Discovering empirical equations from robot-collected data. Lecture Notes in Computer Science, 1325:287–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1997:IPN

[995] S. Tsumoto. Induction of positive and negative deterministic rules based on rough set model. Lecture Notes in Computer Science, 1325:298–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mazlack:1997:ADM

[996] L. J. Mazlack. Autonomous database mining and disorder measures. Lecture Notes in Computer Science, 1325:308– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calmet:1997:KSI

[997] J. Calmet, S. Jekutsch, P. Kullmann, and J. Schue. KOMET — a system for the integration of heterogeneous information sources. Lecture Notes in Computer Science, 1325:318–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nissan:1997:TAF

[998] E. Nissan and D. Rousseau. Towards A1 formalisms for legal evidence. Lecture Notes in Computer Science, 1325:328–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Capponi:1997:TEK

[999] C. Capponi. Type extensibility of a knowledge representation system with powersets. Lecture Notes in Computer Science, 1325:338–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1997:PBV

[1000] D. Costal and A. Olive. Planning based on view updating in deductive databases. Lecture Notes in Computer Science, 1325:348–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choubey:1997:GFA

[1001] S. K. Choubey and V. V. Raghavan. Generic and fully automatic content-based image retrieval architecture. Lecture Notes in Computer Science, 1325:360–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eusterbrock:1997:PSE

[1002] J. Eusterbrock. Program synthesis from examples by theory formation. Lecture Notes in Computer Science, 1325:370–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1997:IAU

[1003] Y. Y. Yao and S. K. M. Wong. Interval approaches for uncertain reasoning. Lecture Notes in Computer Science, 1325:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klopotek:1997:QVQ

[1004] M. A. Klopotek and S. T. Wierzchon.

Qualitative versus quantitative interpretation of the mathematical theory of evidence. Lecture Notes in Computer Science, 1325:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1997:KDD

[1005] Q. Zhu and Z. Chen. Knowledge discovery from databases with the guidance of a causal network. Lecture Notes in Computer Science, 1325:401–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Donahoo:1997:DKM

[1006] M. J. Donahoo, J. W. Murdock, A. K. Goel, and S. Navathe. From data to knowledge: method-specific transformations. Lecture Notes in Computer Science, 1325:411–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kryszkiewicz:1997:GRI

[1007] M. Kryszkiewicz. Generalized rules in incomplete information systems. Lecture Notes in Computer Science, 1325: 421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cucchiara:1997:ICB

[1008] R. Cucchiara, E. Lamma, P. Mello, and M. Milano. An interactive constraintbased system for selective attention in visual search. Lecture Notes in Computer Science, 1325:431–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maluf:1997:ARI

[1009] D. A. Maluf and G. Wiederhold. Abstraction of representation for interoperation. Lecture Notes in Computer Science, 1325:441–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carberry:1997:TCE

[1010] S. Carberry and J. R. Clarke. Trauma CASE: exploiting the knowledge base of an existing decision support system to automatically construct medical cases. Lecture Notes in Computer Science, 1325:456–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholczyk:1997:ISD

[1011] D. Pacholczyk. An intelligent system dealing with negative information. Lecture Notes in Computer Science, 1325:467–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bosc:1997:RBQ

[1012] P. Bosc and O. Pivert. On representation-based querying of databases containing ill-known values. Lecture Notes in Computer Science, 1325:477– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slezak:1997:DVO

[1013] D. Slezak. Decision value oriented decomposition of data tables. Lecture Notes in Computer Science, 1325:487– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grzymala-Busse:1997:MLE

[1014] J. W. Grzymala-Busse and L. J. Old.

A machine learning experiment to determine part of speech from wordendings. Lecture Notes in Computer Science, 1325:497–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quafafou:1997:LFC

[1015] M. Quafafou. Learning flexible concepts from uncertain data. Lecture Notes in Computer Science, 1325:507– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hadzikadic:1997:DAR

[1016] M. Hadzikadic and B. F. Bohren. Determining attribute relevance in decision trees. Lecture Notes in Computer Science, 1325:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chai:1997:WBR

[1017] J. Y. Chai and A. W. Biermann. A WordNet based rule generalization engine in meaning extraction system. Lecture Notes in Computer Science, 1325: 529–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calvanese:1997:RRS

[1018] D. Calvanese, G. De Giacomo, and M. Lenzerini. Representing and reasoning on SGML documents. Lecture Notes in Computer Science, 1325:540– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zarri:1997:CMM

[1019] G. P. Zarri. Conceptual modelling of the “meaning” of textual narrative documents. Lecture Notes in Computer Science, 1325:550–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oldham:1997:ICP

[1020] J. D. Oldham, V. W. Marek, and M. Truszczynski. Intelligent computation of presentation documents. Lecture Notes in Computer Science, 1325: 560–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreasen:1997:FDQ

[1021] T. Andreasen. Flexible database querying based on associations of domain values. Lecture Notes in Computer Science, 1325:570–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuppens:1997:MLF

[1022] F. Cuppens and R. Demolombe. A modal logical framework for security policies. Lecture Notes in Computer Science, 1325:579–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haehnle:1997:CLR

[1023] R. Haehnle, N. V. Murray, and E. Rosenthal. Completeness for linear regular negation normal form inference systems. Lecture Notes in Computer Science, 1325:590–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nie:1997:RSN

[1024] X. Nie and Q. Guo. Renaming a set of non-Horn clauses. Lecture Notes in Computer Science, 1325:600–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:CCR

[1025] Y. Zhang. Constraints, causal rules and minimal change in model-based update. Lecture Notes in Computer Science, 1325:609–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1997:KRD

[1026] F. Esposito, D. Malerba, G. Semeraro, and S. Ferilli. Knowledge revision for document understanding. Lecture Notes in Computer Science, 1325:619– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bieber:1997:EIC

[1027] M. Bieber. Enhancing information comprehension through hypertext. Lecture Notes in Computer Science, 1326: 1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brusilovsky:1997:ETA

[1028] P. Brusilovsky. Efficient techniques for adaptive hypermedia. Lecture Notes in Computer Science, 1326:12–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gobbetti:1997:EAE

[1029] E. Gobbetti and R. Turner. Exploring annotated 3D environments on the World Wide Web. Lecture Notes in Computer Science, 1326:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kay:1997:UMC

[1030] J. Kay and B. Kummerfeld. User models for customized hypertext. Lecture Notes in Computer Science, 1326:47– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kent:1997:CAH

[1031] R. E. Kent and C. Neuss. Conceptual analysis of hypertext. Lecture Notes in Computer Science, 1326:70–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayfield:1997:TLM

[1032] J. Mayfield. Two-level models of hypertext. Lecture Notes in Computer Science, 1326:90–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1997:TDH

[1033] C. Pearce and E. Miller. The TELLTALE dynamic hypertext environment: Approaches to scalability. Lecture Notes in Computer Science, 1326:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:DMB

[1034] W. Wang, C. Ghaoui, and R. Rada.

Domain model based hypertext for collaborative authoring. Lecture Notes in Computer Science, 1326:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rus:1997:IRI

[1035] D. Rus and D. Subramanian. Information retrieval, information structure, and information agents. Lecture Notes in Computer Science, 1326:145– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schultz:1997:RRD

[1036] W. Schultz. Reward responses of dopamine neurons: A biological reinforcement signal. Lecture Notes in Computer Science, 1327:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markram:1997:ICA

[1037] H. Markram and M. Tsodyks. The information content of action potential trains — A synaptic basis. Lecture Notes in Computer Science, 1327:13– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1997:CCA

[1038] R. Miller. Cortical cell assemblies, laminar interaction, and thalamocortical interplay. Lecture Notes in Computer Science, 1327:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brunel:1997:CCS

[1039] N. Brunel. Cross-correlations in sparsely connected recurrent networks of spiking neurons. Lecture Notes in Computer Science, 1327:31–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tetko:1997:CSP

[1040] I. V. Tetko and A. E. P. Villa. A comparative study of pattern detection algorithm and dynamical system approach using simulated spike trains. Lecture Notes in Computer Science, 1327:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rouat:1997:STP

[1041] J. Rouat. Spatio-temporal pattern recognition with neural networks: Application to speech. Lecture Notes in Computer Science, 1327:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lansky:1997:NIF

[1042] P. Lansky and V. Lanska. Noise in integrate-and-fire models of neuronal dynamics. Lecture Notes in Computer Science, 1327:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eurich:1997:CCA

[1043] C. W. Eurich, H. R. Dinse, U. Dicke, and B. Godde. Coarse coding accounts for improvement of spatial discrimination after plastic reorganization in rats and humans. Lecture Notes in Computer Science, 1327:55–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schultz:1997:ARM

[1044] S. Schultz, S. Panzeri, A. Treves, and E. T. Rolls. Analogue resolution in a model of the schaffer collaterals. Lecture Notes in Computer Science, 1327:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattia:1997:MNL

[1045] M. Mattia and S. Fusi. Modeling networks with linear (VLSI) integrateand-fire neurons. Lecture Notes in Computer Science, 1327:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deco:1997:ITA

[1046] G. Deco and B. Schuermann. An information-theoretic analysis of temporal coding strategies by spiking central neurons. Lecture Notes in Computer Science, 1327:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritz:1997:CCS

[1047] R. Ritz and T. J. Sejnowski. Correlation coding in stochastic neural networks. Lecture Notes in Computer Science, 1327:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirose:1997:TDH

[1048] A. Hirose. Two-dimensional Hodgkin– Huxley equations for investigating a basis of pulse-processing neural networks. Lecture Notes in Computer Science, 1327:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koerner:1997:CPS

[1049] E. Koerner and U. Koerner. Concurrent parallel-sequential processing in gamma controlled cortical-type networks of spiking neurones. Lecture Notes in Computer Science, 1327:91– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:NRA

[1050] L. S. Smith. A noise-robust auditory modeling front end for voiced speech. Lecture Notes in Computer Science, 1327:97–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1997:NDU

[1051] T. V. Ho and J. Rouat. A novelty detector using a network of integrate and fire neurons. Lecture Notes in Computer Science, 1327:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eggert:1997:DPD

[1052] J. Eggert and J. L. Van Hemmen.

Derivation of pool dynamics from microscopic neuronal models. Lecture Notes in Computer Science, 1327:109– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steuber:1997:HSP

[1053] V. Steuber and D. J. Willshaw. How a single purkinje cell could learn the adaptive timing of the classically conditioned eye-blink response. Lecture Notes in Computer Science, 1327:115– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senn:1997:ASM

[1054] W. Senn, M. Tsodyks, and H. Markram.

An algorithm for synaptic modification based on exact timing of pre- and post-synaptic action potentials. Lecture Notes in Computer Science, 1327: 121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benuskova:1997:MPR

[1055] L. Benuskova. Modeling plasticity in rat barrel cortex induced by one spared whisker. Lecture Notes in Computer Science, 1327:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohli:1997:MAC

[1056] R. Kohli and P. G. H. Clarke. Mathematical analysis of competition between sensory ganglion cells for nerve growth factor in the skin. Lecture Notes in Computer Science, 1327:133– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanOoyen:1997:CAN

[1057] A. Van Ooyen and D. J. Willshaw. Competition amongst neurons for neurotrophins. Lecture Notes in Computer Science, 1327:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Samuelides:1997:IHL

[1058] M. Samuelides, S. Thorpe, and E. Veneau. Implementing Hebbian learning in a rank-based neural network. Lecture Notes in Computer Science, 1327: 145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graham:1997:MCH

[1059] B. Graham and D. J. Willshaw. A model of clipped Hebbian learning in a neocortical pyramidal cell. Lecture Notes in Computer Science, 1327:151– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eurich:1997:HDA

[1060] C. W. Eurich, J. D. Cowan, and J. G. Milton. Hebbian delay adaptation in a network of integrate-and-fire neurons. Lecture Notes in Computer Science, 1327:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loerincz:1997:HFT

[1061] A. Loerincz. Hippocampal formation trains independent components via forcing input reconstruction. Lecture Notes in Computer Science, 1327:163– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loewel:1997:NVN

[1062] S. Loewel, K. E. Schmidt, and W. Singer. Nature vs. nurture in the development of tangential connections and functional maps in the visual cortex. Lecture Notes in Computer Science, 1327:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huebener:1997:GRB

[1063] M. Huebener, D. Shoham, S. Schulze, and G. Braendle. Geometric relationships between feature maps in cat visual cortex. Lecture Notes in Computer Science, 1327:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wimbauer:1997:LHM

[1064] S. Wimbauer, O. Wenisch, and J. L.

Van Hemmen. A linear Hebbian model for the development of spatiotemporal receptive fields of simple cells. Lecture Notes in Computer Science, 1327:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stetter:1997:SCC

[1065] M. Stetter, E. W. Lang, and K. Obermayer. Synapse clustering can drive simultaneous ON-OFF and oculardominance segregation in a model of area 17. Lecture Notes in Computer Science, 1327:189–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1997:MPM

[1066] F. Wolf and T. Geisel. Must pinwheels move during visual development? Lecture Notes in Computer Science, 1327:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frisone:1997:ETM

[1067] F. Frisone, L. Perico, and P. G. Morasso. Extending the TRN model in a biologically plausible way. Lecture Notes in Computer Science, 1327: 201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brockmann:1997:SMD

[1068] D. Brockmann, H. U. Bauer, M. Riesenhuber, and T. Geisel. SOM-Model for the development of oriented receptive fields and orientation maps from non-oriented ON-center OFF-center inputs. Lecture Notes in Computer Science, 1327:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1997:ABF

[1069] U. Bauer, P. Adorjan, M. Scholz, and J. B. Levitt. On the anatomical basis of field size, contrast sensitivity, and orientation selectivity in macaque striate cortex: A model study. Lecture Notes in Computer Science, 1327:213– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziegaus:1997:SNU

[1070] C. Ziegaus and E. W. Lang. Statistics of natural and urban images. Lecture Notes in Computer Science, 1327: 219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burger:1997:CNM

[1071] T. Burger and E. W. Lang. A CBL network model with intracortical plasticity and natural image stimuli. Lecture Notes in Computer Science, 1327:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ernst:1997:GOP

[1072] U. Ernst, K. Pawelzik, F. Wolf, and T. Geisel. Geometry of orientation preference map determines nonclassical receptive field properties. Lecture Notes in Computer Science, 1327:231– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartsch:1997:MOT

[1073] H. Bartsch, M. Stetter, and K. Obermayer. A model for orientation tuning and contextual effects of orientation selective receptive fields. Lecture Notes in Computer Science, 1327:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiskott:1997:OFN

[1074] L. Wiskott and T. J. Sejnowski. Objective functions for neural map formation. Lecture Notes in Computer Science, 1327:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopecz:1997:RTS

[1075] K. Kopecz and K. Mohraz. Relative time scales in the self-organization of pattern classification: From one-shot to statistical learning. Lecture Notes in Computer Science, 1327:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1997:RGI

[1076] J. Chao, Y. Miyata, and S. Yoshida. Realization of geometric illusions and geometry of visual space with neural networks. Lecture Notes in Computer Science, 1327:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vapnik:1997:SVM

[1077] V. N. Vapnik. The support vector method. Lecture Notes in Computer Science, 1327:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutton:1997:SMD

[1078] R. S. Sutton. On the significance of Markov decision processes. Lecture Notes in Computer Science, 1327:273– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatenet:1997:ERL

[1079] N. Chatenet and H. Bersini. Economical reinforcement learning for non stationary problems. Lecture Notes in Computer Science, 1327:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czernichow:1997:DGA

[1080] T. Czernichow. A double gradient algorithm to optimize regularization. Lecture Notes in Computer Science, 1327:289–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradshaw:1997:GLS

[1081] N. P. Bradshaw, A. Duchateau, and H. Bersini. Global least-squares vs. EM training for the Gaussian mixture of experts. Lecture Notes in Computer Science, 1327:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kappen:1997:ALB

[1082] H. J. Kappen and F. B. Rodriguez. Accelerated learning in Boltzmann machines using mean field theory. Lecture Notes in Computer Science, 1327: 301–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boes:1997:AOL

[1083] S. Boes. Adaptive online learning for nonstationary problems. Lecture Notes in Computer Science, 1327:307– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aboukassem:1997:WDD

[1084] M. Aboukassem, S. Schwember, S. Noehte, and R. Maenner. Weight discretization due to optical constraints and its influence on the generalization abilities of a simple perceptron. Lecture Notes in Computer Science, 1327: 313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soltani:1997:WFB

[1085] S. Soltani, S. Canu, D. Boichu, and Y. Grandvalet. Wavelet frames based estimator. Lecture Notes in Computer Science, 1327:319–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mozayyani:1997:STP

[1086] N. Mozayyani and G. Vaucher. A spatio-temporal perceptron for online handwritten character recognition. Lecture Notes in Computer Science, 1327:325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weiss:1997:LOU

[1087] M. G. Weiss. Learning oscillations using adaptive control. Lecture Notes in Computer Science, 1327:331–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eggenberger:1997:CNN

[1088] P. Eggenberger. Creation of neural networks based on developmental and evolutionary principles. Lecture Notes in Computer Science, 1327:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettoni:1997:BAR

[1089] A. Bettoni, P. Campadelli, and M. Parodi. A boosting algorithm for regression. Lecture Notes in Computer Science, 1327:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taniguchi:1997:CRN

[1090] M. Taniguchi and V. Tresp. Combining regularized neural networks. Lecture Notes in Computer Science, 1327: 349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueger:1997:MSN

[1091] S. M. Rueger. Making stochastic networks deterministic. Lecture Notes in Computer Science, 1327:355–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruf:1997:ULN

[1092] B. Ruf and M. Schmitt. Unsupervised learning in networks of spiking neurons using temporal coding. Lecture Notes in Computer Science, 1327:361– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selonen:1997:ERM

[1093] A. Selonen and J. Lampinen. Experiments on regularizing MLP models with background knowledge. Lecture Notes in Computer Science, 1327: 367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutjahr:1997:EBF

[1094] S. Gutjahr and J. Feist. Elliptical basis function networks for classification tasks. Lecture Notes in Computer Science, 1327:373–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galleske:1997:PNN

[1095] I. Galleske and J. Castellanos. Probabilistic neural networks with rotated kernel functions. Lecture Notes in Computer Science, 1327:379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jankowski:1997:SCR

[1096] N. Jankowski and V. Kadirkamanathan. Statistical control of RBFlike networks for classification. Lecture Notes in Computer Science, 1327:385– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scherf:1997:IRN

[1097] M. Scherf and W. Brauer. Improving RBF networks by the feature selection approach EUBAFES. Lecture Notes in Computer Science, 1327:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1997:PCS

[1098] L. Graf, U. Kressel, and J. Franke. Polynomial classifiers and support vector machines. Lecture Notes in Computer Science, 1327:397–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimura:1997:URD

[1099] M. Kimura and R. Nakano. Unique representations of dynamical systems produced by recurrent neural networks. Lecture Notes in Computer Science, 1327:403–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hammer:1997:GEN

[1100] B. Hammer. Generalization of elman networks. Lecture Notes in Computer Science, 1327:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishikawa:1997:DNN

[1101] M. Ishikawa and K. Nishino. Designing neural networks by a combination of structural learning and genetic algorithms. Lecture Notes in Computer Science, 1327:415–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varsta:1997:RSO

[1102] M. Varsta, J. del R. Millan, and J. Heikkonen. A recurrent selforganizing map for temporal sequence processing. Lecture Notes in Computer Science, 1327:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stagge:1997:EEN

[1103] P. Stagge and B. Sendhoff. An extended elman net for modeling time series. Lecture Notes in Computer Science, 1327:427–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuzmina:1997:RAM

[1104] M. Kuzmina, E. Manykin, and I. Surina. Recurrent associative memory network of nonlinear coupled oscillators. Lecture Notes in Computer Science, 1327:433–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wersing:1997:LRN

[1105] H. Wersing, J. J. Steil, and H. Ritter. A layered recurrent neural network for feature grouping. Lecture Notes in Computer Science, 1327:439–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meert:1997:MRT

[1106] K. Meert and J. Ludik. A multilayer real-time recurrent learning algorithm for improved convergence. Lecture Notes in Computer Science, 1327: 445–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storkey:1997:ICH

[1107] A. Storkey. Increasing the capacity of a Hopfield network without sacrificing functionality. Lecture Notes in Computer Science, 1327:451–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:NAN

[1108] H. Wang and D. Bell. A novel associative network accommodating pattern deformation. Lecture Notes in Computer Science, 1327:457–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandvalet:1997:ANI

[1109] Y. Grandvalet and S. Canu. Adaptive noise injection for input variables relevance determination. Lecture Notes in Computer Science, 1327:463–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeLaar:1997:ISP

[1110] P. Van De Laar, S. Gielen, and T. Heskes. Input selection with partial retraining. Lecture Notes in Computer Science, 1327:469–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayoraz:1997:CRI

[1111] E. Mayoraz. On the complexity of recognizing iterated differences of polyhedra. Lecture Notes in Computer Science, 1327:475–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guermeur:1997:OLR

[1112] Y. Guermeur, F. D’Alche-Buc, and P. Gallinari. Optimal linear regression on classifier outputs. Lecture Notes in Computer Science, 1327:481–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quelavoine:1997:LVM

[1113] R. Quelavoine and P. Nocera. Learning verification in multilayer neural networks. Lecture Notes in Computer Science, 1327:487–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwon:1997:DFT

[1114] O. J. Kwon and S. Y. Bang. Design of a fault tolerant multilayer perceptron with a desired level of robustness. Lecture Notes in Computer Science, 1327:493–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moerland:1997:MEE

[1115] P. Moerland. Mixtures of experts estimate A posteriori probabilities. Lecture Notes in Computer Science, 1327: 499–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doering:1997:AOC

[1116] A. Doering, M. Galicki, and H. Witte. Admissibility and optimality of the cascade-correlation algorithm. Lecture Notes in Computer Science, 1327:505– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradshaw:1997:EVD

[1117] N. P. Bradshaw. The effective VC dimension of the n-tuple classifier. Lecture Notes in Computer Science, 1327: 511–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oja:1997:NPC

[1118] E. Oja, J. Karhunen, and A. Hyvaerinen. From neural principal components to neural independent components. Lecture Notes in Computer Science, 1327:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taleb:1997:EOA

[1119] A. Taleb and C. Jutten. Entropy optimization — application to blind source separation. Lecture Notes in Computer Science, 1327:529–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehler:1997:IPI

[1120] B.-U. Koehler, T.-W. Lee, and R. Orglmeister. Improving the performance of Infomax using statistical signal processing techniques. Lecture Notes in Computer Science, 1327:535–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pajunen:1997:MLA

[1121] P. Pajunen and J. Karhunen. A maximum likelihood approach to nonlinear blind source separation. Lecture Notes in Computer Science, 1327:541– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattavelli:1997:PBA

[1122] M. Mattavelli, E. Amaldi, and J.-M.

Vesin. A perceptron-based approach to piecewise linear modeling with an application to time series. Lecture Notes in Computer Science, 1327:547– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oja:1997:LIC

[1123] E. Oja and K. Valkealahti. Local independent component analysis by the self-organizing map. Lecture Notes in Computer Science, 1327:553–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linares:1997:MBD

[1124] G. Linares, P. Nocera, and H. Meloni. Model breaking detection using independent component classifier. Lecture Notes in Computer Science, 1327:559– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paraschiv-Ionescu:1997:NNB

[1125] A. Paraschiv-Ionescu, C. Jutten, and G. Bouvier. Neural network based processing for smart sensors arrays. Lecture Notes in Computer Science, 1327: 565–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiori:1997:AMN

[1126] S. Fiori, A. Uncini, and F. Piazza. Application of the MEC network to principal component analysis and source separation. Lecture Notes in Computer Science, 1327:571–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joutsensalo:1997:SBS

[1127] J. Joutsensalo. Semi-blind source parameter separation. Lecture Notes in Computer Science, 1327:577–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoelkopf:1997:KPC

[1128] B. Schoelkopf, A. J. Smola, and K.-R Mueller. Kernel principal component analysis. Lecture Notes in Computer Science, 1327:583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balachander:1997:ECD

[1129] T. Balachander, R. Kothari, and H. Cualing. An empirical comparison of dimensionality reduction techniques for pattern classification. Lecture Notes in Computer Science, 1327:589–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruske:1997:TRN

[1130] J. Bruske and G. Sommer. Topology representing networks for intrinsic dimensionality estimation. Lecture Notes in Computer Science, 1327:595– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haekkinen:1997:SBV

[1131] E. Haekkinen and P. Koikkalainen.

SOM based visualization in data analysis. Lecture Notes in Computer Science, 1327:601–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carpenter:1997:ADP

[1132] G. A. Carpenter and F. D. M. Wilson. ARTMAP-DS: Pattern discrimination by discounting similarities. Lecture Notes in Computer Science, 1327: 607–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fritzke:1997:SON

[1133] B. Fritzke. A self-organizing network that can follow non-stationary distributions. Lecture Notes in Computer Science, 1327:613–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burger:1997:PTS

[1134] M. Burger, T. Graepel, and K. Obermayer. Phase transitions in soft topographic vector quantization. Lecture Notes in Computer Science, 1327:619– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herrmann:1997:VQO

[1135] M. Herrmann and T. Villmann. Vector quantization by optimal neural gas. Lecture Notes in Computer Science, 1327:625–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fort:1997:CKM

[1136] J.-C. Fort and G. Pages. Convergences of the Kohonen maps: A dynamical system approach. Lecture Notes in Computer Science, 1327:631–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laaksonen:1997:LSC

[1137] J. Laaksonen. Local subspace classifier. Lecture Notes in Computer Science, 1327:637–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delmas:1997:ADA

[1138] J. P. Delmas. Asymptotic distributions associated to unsupervised Oja’s learning equation. Lecture Notes in Computer Science, 1327:643–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlassis:1997:PGC

[1139] N. A. Vlassis, A. Dimopoulos, and G. Papakonstantinou. The probabilistic growing cell structures algorithm. Lecture Notes in Computer Science, 1327:649–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hochreiter:1997:UCL

[1140] S. Hochreiter and J. Schmidhuber. Unsupervised coding with LOCOCODE. Lecture Notes in Computer Science, 1327:655–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobrzewski:1997:WPS

[1141] B. Dobrzewski, D. Ruwish, and M. Bode. Wave propagation in selforganizing feature maps as a means for the representation of temporal sequences. Lecture Notes in Computer Science, 1327:661–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pican:1997:CKS

[1142] N. Pican. Contextual Kohonen SOM with orthogonal weight estimator principle. Lecture Notes in Computer Science, 1327:667–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1997:SOM

[1143] H. Ritter. Self-organizing maps for robot control. Lecture Notes in Computer Science, 1327:675–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harvey:1997:CCE

[1144] I. Harvey. Cognition is not computation; evolution is not optimisation. Lecture Notes in Computer Science, 1327:685–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheier:1997:ITI

[1145] C. Scheier and R. Pfeifer. Information theoretic implications of embodiment for neural network learning. Lecture Notes in Computer Science, 1327:691–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tani:1997:VAL

[1146] J. Tani. Visual attention and learning of a cognitive robot. Lecture Notes in Computer Science, 1327:697–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Egner:1997:FBT

[1147] S. Egner and C. Scheier. Feature binding through temporally correlated neural activity in a robot model of visual perception. Lecture Notes in Computer Science, 1327:703–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huber:1997:MOA

[1148] S. A. Huber and H. H. Buelthoff. Modeling obstacle avoidance behavior of flies using an adaptive autonomous agent. Lecture Notes in Computer Science, 1327:709–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1997:MAO

[1149] T. R. Neumann, S. A. Huber, and H. H.

Buelthoff. Minimalistic approach to 3D obstacle avoidance behavior from simulated evolution. Lecture Notes in Computer Science, 1327:715–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaario:1997:SDE

[1150] J. Vaario and K. Shimohara. Synthesis of developmental and evolutionary modeling of adaptive autonomous agents. Lecture Notes in Computer Science, 1327:721–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buehlmeier:1997:HMN

[1151] A. Buehlmeier, P. Steiner, M. Rossmann, and K. Goser. Hebbian multilayer network in a wheelchair robot. Lecture Notes in Computer Science, 1327:727–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolfi:1997:NNA

[1152] S. Nolfi and D. Parisi. Neural networks in an artificial life perspective. Lecture Notes in Computer Science, 1327:733–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1997:IAL

[1153] J. del R. Millan. Incremental acquisition of local networks for the control of autonomous robots. Lecture Notes in Computer Science, 1327:739–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lund:1997:RAI

[1154] H. H. Lund. Robot-animal interaction. Lecture Notes in Computer Science, 1327:745–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mallot:1997:VGA

[1155] H. A. Mallot, M. Franz, B. Schoelkopf, and H. H. Buelthoff. The viewgraph approach to visual navigation and spatial memory. Lecture Notes in Computer Science, 1327:751–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trullier:1997:PSL

[1156] O. Trullier and J.-A Meyer. Place sequence learning for navigation. Lecture Notes in Computer Science, 1327:757– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billard:1997:LCT

[1157] A. Billard and G. Hayes. Learning to communicate through imitation in autonomous robots. Lecture Notes in Computer Science, 1327:763–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salustowicz:1997:LSS

[1158] R. Salustowicz, M. Wiering, and J. Schmidhuber. On learning soccer strategies. Lecture Notes in Computer Science, 1327:769–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizutani:1997:MLL

[1159] K. Mizutani and T. Omori. A model of logic like inference by memory model PATON. Lecture Notes in Computer Science, 1327:775–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saadia:1997:FFC

[1160] N. Saadia, Y. Amirat, J. Pontnau, and A. Ramdane-Cherif. Force feedback control of an assembly robot by neural networks. Lecture Notes in Computer Science, 1327:781–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dapper:1997:NFC

[1161] M. Dapper, R. Maass, V. Zahn, and R. Eckmiller. Neural force control (NFC) for complex manipulator tasks. Lecture Notes in Computer Science, 1327:787–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frontzek:1997:HPP

[1162] T. Frontzek, N. Goerke, and R. Eckmiller. A hybrid path planning system combining the A-method and RBFnetworks. Lecture Notes in Computer Science, 1327:793–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chella:1997:ANN

[1163] A. Chella, S. Gaglio, V. Mulia, and G. Sajeva. An ASSOM neural network to represent actions performed by an autonomous agent. Lecture Notes in Computer Science, 1327:799–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomon:1997:ARB

[1164] R. Salomon. The application of radial basis function networks with implicit continuity constraints. Lecture Notes in Computer Science, 1327:805–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Indiveri:1997:AVG

[1165] G. Indiveri and P. Vetschure. Autonomous vehicle guidance using analog VLSI neuromorphic sensors. Lecture Notes in Computer Science, 1327: 811–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ortmann:1997:NNV

[1166] V. Ortmann and R. Eckmiller. Neural network visual tracking system. Lecture Notes in Computer Science, 1327: 817–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pasemann:1997:PBD

[1167] F. Pasemann. Pole-balancing with different evolved neurocontrollers. Lecture Notes in Computer Science, 1327:823– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kokcharov:1997:CPR

[1168] I. Kokcharov. Calibration of parallel robots by evolutionary algorithm. Lecture Notes in Computer Science, 1327: 831–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duleba:1997:UAM

[1169] I. Duleba and R. Muszynski. On use of ANNs to model and to control robot manipulators. Lecture Notes in Computer Science, 1327:837–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ledoux:1997:IEA

[1170] C. Ledoux and F. Bonnard. Identification of the electric arc of a furnace. Lecture Notes in Computer Science, 1327:843–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breining:1997:UMS

[1171] C. Breining and G. Alt. On using MLPs for step size control in echo cancellation for hands-free telephone sets. Lecture Notes in Computer Science, 1327:849–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Habtom:1997:NND

[1172] R. Habtom and L. Litz. Neurocontrol of nonlinear dynamic systems subject to unmeasured disturbance inputs. Lecture Notes in Computer Science, 1327:855–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blackburn:1997:LAC

[1173] Patrick Blackburn, Marc Dymetman, Alain Lecomte, Aarne Ranta, Christian Retoré, and Eric Villemonte de la Clergerie. Logical aspects of computational linguistics: An introduction. Lecture Notes in Computer Science, 1328:1– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280001.
pdf.

Joshi:1997:PPT

[1174] Aravind K. Joshi and Seth Kulick. Partial proof trees, resource sensitive logics, and syntactic constraints. Lecture Notes in Computer Science, 1328:21– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280021.
pdf.

Kracht:1997:IF

[1175] Marcus Kracht. Inessential features. Lecture Notes in Computer Science, 1328:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1328/13280043.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1328/13280043.
pdf.

Miller:1997:LLL

[1176] Dale Miller. Linear logic as logic programming: An abstract. Lecture Notes in Computer Science, 1328:63– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/1328/13280063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280063.
pdf.

Stabler:1997:DM

[1177] Edward Stabler. Derivational minimalism. Lecture Notes in Computer Science, 1328:68–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1328/13280068.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1328/13280068.pdf.

Abrusci:1997:TAG

[1178] V. Michele Abrusci, Christophe Fouqueré, and Jacqueline Vauzeilles. Tree adjoining grammars in noncommutative linear logic. Lecture Notes in Computer Science, 1328:96– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280096.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280096.
pdf.

Bechet:1997:CDP

[1179] Denis Bechet and Philippe de Groote.

Constructing different phonological bracketings from a proof net. Lecture Notes in Computer Science, 1328:118– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280118.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280118.
                                        pdf.

Boldini:1997:VTT

[1180] Pascal Boldini. Vagueness and type theory. Lecture Notes in Computer Science, 1328:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1328/13280134.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1328/13280134.pdf.

Coscoy:1997:NLE

[1181] Yann Coscoy. A natural language explanation for formal proofs. Lecture Notes in Computer Science, 1328:149– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1328/13280149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280149.
                                        pdf.

Emms:1997:MPL

[1182] Martin Emms. Models for polymorphic lambek calculus. Lecture Notes in Computer Science, 1328:168– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1328/13280168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280168.
                                        pdf.

Gardent:1997:SI

[1183] Claire Gardent. Sloppy identity. Lecture Notes in Computer Science, 1328: 188–207, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1328/13280188.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1328/13280188.
pdf.

Hegner:1997:FDF

[1184] Stephen J. Hegner. A family of decidable feature logics which support HPSG-style set and list constructions. Lecture Notes in Computer Science, 1328:208–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1328/13280208.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1328/13280208.
pdf.

Kempson:1997:LUP

[1185] Ruth Kempson, Wilfried Meyer Viol, and Dov Gabbay. Language understanding: A procedural perspective. Lecture Notes in Computer Science, 1328:228–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280228.
pdf.

King:1997:ADC

[1186] Paul John King and Kiril Ivanov Simov. The automatic deduction of classificatory systems from linguistic theories (abridged). Lecture Notes in Computer Science, 1328:248–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280248.
pdf.

Manara:1997:BCT

[1187] Lucia H. B. Manara and Anne De Roeck. A belief-centered treatment of pragmatic presupposition. Lecture Notes in Computer Science, 1328:274– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280274.
pdf.

Marciniec:1997:CST

[1188] Jacek Marciniec. Connected sets of types and categorial consequence. Lecture Notes in Computer Science, 1328: 292–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280292.
pdf.

Merenciano:1997:GDL

[1189] Josep M. Merenciano and Glyn Morrill. Generation as deduction on labelled proof nets. Lecture Notes in Computer Science, 1328:310–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280310.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280310.
pdf.

Michaelis:1997:SSI

[1190] Jens Michaelis and Marcus Kracht.

Semilinearity as a syntactic invariant. Lecture Notes in Computer Science, 1328:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1328/13280329.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1328/13280329.
pdf.

Riezler:1997:QCL

[1191] Stefan Riezler. Quantitative constraint logic programming for weighted grammar applications. Lecture Notes in Computer Science, 1328:346–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280346.
pdf.

Rogers:1997:SRL

[1192] James Rogers. Strict LT_{2}: Regular :: Local : Recognizable. Lecture Notes in Computer Science, 1328:366– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1328/13280366.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280366.
                                        pdf.

Schena:1997:PLV

[1193] Irene Schena. Pomset logic and variants in natural languages. Lecture Notes in Computer Science, 1328:386– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1328/13280386.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280386.
                                        pdf.

Stolzenburg:1997:CLP

[1194] Frieder Stolzenburg, Stephan Höhne, Ulrich Koch, and Martin Volk. Constraint logic programming for computational linguistics. Lecture Notes in Computer Science, 1328:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1328/13280406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1328/13280406.
                                        pdf.

Szczerba:1997:RTR

[1195] Marek Szczerba. Representation theorems for residuated groupoids. Lecture Notes in Computer Science, 1328: 426–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1328/13280426.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1328/13280426.
pdf.

Galton:1997:CCS

[1196] A. Galton. Continuous change in spatial regions. Lecture Notes in Computer Science, 1329:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hornsby:1997:QRC

[1197] K. Hornsby and M. J. Egenhofer. Qualitative representation of change. Lecture Notes in Computer Science, 1329: 15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rodriguez:1997:ISB

[1198] M. A. Rodriguez and M. J. Egenhofer. Image-schemata-based spatial inferences: The container-surface algebra. Lecture Notes in Computer Science, 1329:35–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dieberger:1997:CMS

[1199] A. Dieberger. A city metaphor to support navigation in complex information spaces. Lecture Notes in Computer Science, 1329:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Timpf:1997:UHS

[1200] S. Timpf and A. U. Frank. Using hierarchical spatial data structures for hierarchical spatial reasoning. Lecture Notes in Computer Science, 1329:69– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raubal:1997:SSI

[1201] M. Raubal, M. J. Egenhofer, D. Pfoser, and N. Tryfona. Structuring space with image schemata: Wayfinding in airports as a case study. Lecture Notes in Computer Science, 1329:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:FBF

[1202] B. Smith and A. C. Varzi. Fiat and bona fide boundaries: Towards an ontology of spatially extended objects. Lecture Notes in Computer Science, 1329:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plewe:1997:ROT

[1203] B. Plewe. A representation-oriented taxonomy of gradation. Lecture Notes in Computer Science, 1329:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1997:CIR

[1204] M. V. Gray. Classification as an impediment to the reliable and valid use of spatial information: A disaggregate approach. Lecture Notes in Computer Science, 1329:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Couclelis:1997:WMM

[1205] H. Couclelis and J. Gottsegen. What maps mean to people: Denotation, connotation, and geographic visualization in land-use debates. Lecture Notes in Computer Science, 1329:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stell:1997:ASS

[1206] J. G. Stell and M. F. Worboys. The algebraic structure of sets of regions. Lecture Notes in Computer Science, 1329: 163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1997:CRT

[1207] V. H. Nguyen, C. Parent, and S. Spaccapietra. Complex regions in topological queries. Lecture Notes in Computer Science, 1329:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knauff:1997:CAT

[1208] M. Knauff, R. Rauh, and J. Renz. A cognitive assessment of topological spatial relations: Results from an empirical investigation. Lecture Notes in Computer Science, 1329:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hangouet:1997:VDL

[1209] J.-F. Hangouet and R. Djadri. Voronoi diagrams on line segments: Measurements for contextual generalization purposes. Lecture Notes in Computer Science, 1329:207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bittner:1997:QCL

[1210] T. Bittner. A qualitative coordinate language of location of figures within the ground. Lecture Notes in Computer Science, 1329:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tao:1997:IFO

[1211] C. Tao, M. Molenaar, and T. Bouloucos. Identification of fuzzy objects from field observation data. Lecture Notes in Computer Science, 1329:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Federico:1997:LTS

[1212] T. Federico and N. Franklin. Long-term spatial representations from pictorial and textual input. Lecture Notes in Computer Science, 1329:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berendt:1997:FAR

[1213] B. Berendt and P. Jansen-Osmann. Feature accumulation and route structuring in distance estimations — an interdisciplinary approach. Lecture Notes in Computer Science, 1329:279–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montello:1997:PCE

[1214] D. R. Montello. The perception and cognition of environmental distance: Direct source of information. Lecture Notes in Computer Science, 1329:297– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harvey:1997:IMP

[1215] F. Harvey. Improving multi-purpose GIS design: Participative design. Lecture Notes in Computer Science, 1329: 313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Portugali:1997:SOC

[1216] J. Portugali. Self-organization, cities, cognitive maps and information systems. Lecture Notes in Computer Science, 1329:329–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barkowsky:1997:CRM

[1217] T. Barkowsky and C. Freksa. Cognitive requirements on making and interpreting maps. Lecture Notes in Computer Science, 1329:347–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1997:KWW

[1218] G. L. Allen. From knowledge to words to wayfinding: Issues in the production and comprehension of route directions. Lecture Notes in Computer Science, 1329:363–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Epstein:1997:SRP

[1219] S. L. Epstein. Spatial representation for pragmatic navigation. Lecture Notes in Computer Science, 1329:373–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erwig:1997:PC

[1220] M. Erwig and M. Schneider. Partition and conquer. Lecture Notes in Computer Science, 1329:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chase:1997:SES

[1221] S. C. Chase. Supporting emergence in spatial reasoning with shape algebras and formal logic. Lecture Notes in Computer Science, 1329:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haunold:1997:LCG

[1222] P. Haunold, S. Grumbach, G. Kuper, and Z. Lacroix. Linear constraints: Geometric objects represented by inequalities. Lecture Notes in Computer Science, 1329:429–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almeida:1997:EBA

[1223] M. J. Almeida. An event-based approach to spatial information. Lecture Notes in Computer Science, 1329:441– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1997:GNF

[1224] G. Edwards. Geocognostics — A new framework for spatial information theory. Lecture Notes in Computer Science, 1329:455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheylan:1997:GMG

[1225] J.-P. Cheylan, T. Libourel, and C. Mende. Graphical modelling for geographical explanation. Lecture Notes in Computer Science, 1329:473–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gahegan:1997:EUC

[1226] M. Gahegan. Experiments using context and significance to enhance the reporting capabilities of GIS. Lecture Notes in Computer Science, 1329:485– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abella:1997:ASR

[1227] A. Abella and J. R. Kender. Automatic summarization of radiographic imagery. Lecture Notes in Computer Science, 1329:497–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrault:1997:ASN

[1228] M. Barrault. An automated system for name placement which complies with cartographic quality criteria: The hydrographic network. Lecture Notes in Computer Science, 1329:499–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benenson:1997:ABS

[1229] I. Benenson and J. Portugali. Agentbased simulations of a city dynamics in a GIS environment. Lecture Notes in Computer Science, 1329:501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bennett:1997:LAI

[1230] B. Bennett, A. G. Cohn, and A. Isli. A logical approach to incorporating qualitative spatial reasoning into GIS. Lecture Notes in Computer Science, 1329: 503–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaser:1997:UIS

[1231] A. D. Blaser. User interaction in a sketch-based GIS user interface. Lecture Notes in Computer Science, 1329: 505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rupp:1997:MRT

[1232] U. Rupp. Metrical refinement of topological relations. Lecture Notes in Computer Science, 1329:507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zahn:1997:ATR

[1233] F. B. Zahn. Approximation of topological relations between fuzzy regions satisfying a linguistically described query. Lecture Notes in Computer Science, 1329:509–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHentenryck:1997:VSM

[1234] P. Van Hentenryck. Visual solver, a modeling language for constraint programming. Lecture Notes in Computer Science, 1330:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guernalec:1997:NBS

[1235] N. B. Guernalec and A. Colmerauer.

Narrowing a 2n-block of sortings in O(n log n). Lecture Notes in Computer Science, 1330:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1997:SVW

[1236] Y. Caseau and F. Laburthe. Solving various weighted matching problems with constraints. Lecture Notes in Computer Science, 1330:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Regin:1997:FAG

[1237] J.-C. Regin and J.-F. Puget. A filtering algorithm for global sequencing constraints. Lecture Notes in Computer Science, 1330:32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1997:SEP

[1238] F. De Boer, M. Gabbrielli, and M. C.

Meo. Semantics and expressive power of a timed concurrent constraint language. Lecture Notes in Computer Science, 1330:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Devienne:1997:SCS

[1239] P. Devienne, J. M. Talbot, and S. Tison. Solving classes of set constraints with tree automata. Lecture Notes in Computer Science, 1330:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gent:1997:LSA

[1240] I. P. Gent and J. L. Underwood. The logic of search algorithms: Theory and applications. Lecture Notes in Computer Science, 1330:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melzer:1997:VPS

[1241] S. Melzer. Verification of parallel systems using constraint programming. Lecture Notes in Computer Science, 1330:92–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achlioptas:1997:RCS

[1242] D. Achlioptas, L. M. Kirousis, E. Kranakis, and D. Krizanc. Random constraint satisfaction: A more accurate picture. Lecture Notes in Computer Science, 1330:107–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomes:1997:HTD

[1243] C. P. Gomes, B. Selman, and N. Crato. Heavy-tailed distributions in combinatorial search. Lecture Notes in Computer Science, 1330:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Istrate:1997:CSI

[1244] G. Istrate. Counting, structure identification and maximum consistency for binary constraint satisfaction problems. Lecture Notes in Computer Science, 1330:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rish:1997:SAB

[1245] I. Rish and D. Frost. Statistical analysis of backtracking on inconsistent CSPs. Lecture Notes in Computer Science, 1330:150–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caseau:1997:UCP

[1246] Y. Caseau. Using constraint propagation for complex scheduling problems: Managing size, complex resources and travel. Lecture Notes in Computer Science, 1330:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabin:1997:UIM

[1247] D. Sabin and E. C. Freuder. Understanding and improving the MAC algorithm. Lecture Notes in Computer Science, 1330:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1997:MEH

[1248] B. M. Smith and S. A. Grant. Modelling exceptionally hard constraint satisfaction problems. Lecture Notes in Computer Science, 1330:182–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galinier:1997:TSM

[1249] P. Galinier and J.-K. Hao. Tabu search for maximal constraint satisfaction problems. Lecture Notes in Computer Science, 1330:196–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamadi:1997:RAN

[1250] Y. Hamadi and D. Merceron. Reconfigurable architectures: A new vision for optimization problems. Lecture Notes in Computer Science, 1330: 209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirayama:1997:DPC

[1251] K. Hirayama and M. Yokoo. Distributed partial constraint satisfaction problem. Lecture Notes in Computer Science, 1330:222–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1997:LML

[1252] L. Michel and P. Van Hentenryck. Localizer: A modeling language for local search. Lecture Notes in Computer Science, 1330:237–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdennadher:1997:OSC

[1253] S. Abdennadher. Operational semantics and confluence of constraint propagation rules. Lecture Notes in Computer Science, 1330:252–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudet:1997:AUH

[1254] A. Boudet and E. Contejean. AC-Unification of higher-order patterns. Lecture Notes in Computer Science, 1330:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermann:1997:CUD

[1255] M. Hermann and P. G. Kolaitis. On the complexity of unification and disunification in commutative idempotent semigroups. Lecture Notes in Computer Science, 1330:282–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1997:OCF

[1256] M. Mueller, J. Niehren, and A. Podelski. Ordering constraints over feature trees. Lecture Notes in Computer Science, 1330:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debruyne:1997:RPC

[1257] R. Debruyne and C. Bessiere. From restricted path consistency to maxrestricted path consistency. Lecture Notes in Computer Science, 1330:312– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gent:1997:CAC

[1258] I. P. Gent, E. MacIntyre, P. Prosser, and P. Shaw. The constrainedness of arc consistency. Lecture Notes in Computer Science, 1330:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:LAV

[1259] Chu Min Li and Anbulagan. Lookahead versus look-back for satisfiability problems. Lecture Notes in Computer Science, 1330:341–355, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokoo:1997:WAM

[1260] M. Yokoo. Why adding more constraints makes a problem easier for hillclimbing algorithms: Analyzing landscapes of CSPs. Lecture Notes in Computer Science, 1330:356–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michel:1997:IMN

[1261] L. Michel and J.-F Puget. Interval methods for non-linear constraints. Lecture Notes in Computer Science, 1330:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minton:1997:CST

[1262] S. Minton. Configurable solvers: Tailoring general methods to specific applications. Lecture Notes in Computer Science, 1330:372–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baptiste:1997:CPD

[1263] P. Baptiste and C. Le Pape. Constraint propagation and decomposition techniques for highly disjunctive and highly cumulative project scheduling problems. Lecture Notes in Computer Science, 1330:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beck:1997:FPE

[1264] J. C. Beck, A. J. Davenport, and M. S. Fox. Five pitfalls of empirical scheduling research. Lecture Notes in Computer Science, 1330:390–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGivry:1997:BOC

[1265] S. De Givry, G. Verfaillie, and T. Schiex. Bounding the optimum of constraint optimization problems. Lecture Notes in Computer Science, 1330: 405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pesant:1997:GCG

[1266] G. Pesant, M. Gendreau, and J.-M Rousseau. GENIUS-CP: A generic single-vehicle routing algorithm. Lecture Notes in Computer Science, 1330: 420–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bettini:1997:SQT

[1267] C. Bettini, X. Sean Wang, and S. Jajodia. Satisfiability of quantitative temporal constraints with multiple granularities. Lecture Notes in Computer Science, 1330:435–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crumbach:1997:TRG

[1268] S. Crumbach and G. Kuper. Tractable recursion over geometric data. Lecture Notes in Computer Science, 1330:450– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoffmann:1997:FSS

[1269] C. M. Hoffmann, A. Lomonosov, and M. Sitharam. Finding solvable subsets of constraint graphs. Lecture Notes in Computer Science, 1330:463–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1997:TDC

[1270] D. Cohen, P. Jeavons, and M. Koubarakis. Tractable disjunctive constraints. Lecture Notes in Computer Science, 1330: 478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harvey:1997:CCS

[1271] W. Harvey, P. J. Stuckey, and A. Borning. Compiling constraint solving using projection. Lecture Notes in Computer Science, 1330:491–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Labat:1997:OHO

[1272] J.-M. Labat and L. Mynard. Oscillation, heuristic ordering and pruning in neighborhood search. Lecture Notes in Computer Science, 1330:506–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulte:1997:PCI

[1273] C. Schulte. Programming constraint inference engines. Lecture Notes in Computer Science, 1330:519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shvetsov:1997:NOO

[1274] I. Shvetsov, V. Telerman, and D. Ushakov. NeMo+: Object-oriented constraint programming environment based on subdefinite models. Lecture Notes in Computer Science, 1330:534– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pacholski:1997:SCP

[1275] L. Pacholski and A. Podelski. Set constraints: A pearl in research on constraints. Lecture Notes in Computer Science, 1330:549–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merten:1997:CMU

[1276] A. G. Merten. From conceptual modeler to University President. Lecture Notes in Computer Science, 1331:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storey:1997:ODD

[1277] V. C. Storey, H. Ullrich, and S. Sundaresan. An ontology for database design automation. Lecture Notes in Computer Science, 1331:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schewe:1997:WBR

[1278] K. D. Schewe. Well-behaving rule systems for entity-relationship and objectoriented models. Lecture Notes in Computer Science, 1331:14–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lloyd-Williams:1997:EDK

[1279] M. Lloyd-Williams. Exploiting domain knowledge during the automated design of object-oriented databases. Lecture Notes in Computer Science, 1331: 16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purao:1997:ISR

[1280] S. Purao and V. C. Storey. Intelligent support for retrieval and synthesis of patterns for object-oriented design. Lecture Notes in Computer Science, 1331:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petrounias:1997:CDF

[1281] I. Petrounias. A conceptual development framework for temporal information systems. Lecture Notes in Computer Science, 1331:43–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1997:TFC

[1282] D. Costal, A. Olive, and M. R. Sancho. Temporal features of class populations and attributes in conceptual models. Lecture Notes in Computer Science, 1331:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goralwalla:1997:MSE

[1283] I. A. Goralwalla, D. Szafron, M. T. Oezsu, and R. J. Peters. Managing schema evolution using a temporal object model. Lecture Notes in Computer Science, 1331:71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iyengar:1997:DOR

[1284] S. Iyengar. Distributed object repositories: Concepts and standards. Lecture Notes in Computer Science, 1331: 85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dey:1997:ESS

[1285] D. Dey and S. Sarkar. Extended SQL support for uncertain data. Lecture Notes in Computer Science, 1331:102– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloesch:1997:CQU

[1286] A. C. Bloesch and T. A. Halpin. Conceptual queries using ConQuer-II. Lecture Notes in Computer Science, 1331: 113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baekgaard:1997:TBS

[1287] L. Baekgaard. Transaction-based specification of database evolution. Lecture Notes in Computer Science, 1331:127– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreft:1997:BCR

[1288] M. Schreft and M. Stumptner. Behavior consistent refinement of object life cycles. Lecture Notes in Computer Science, 1331:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:ATI

[1289] L. Liu and C. Pu. ActivityFlow:

Towards incremental specification and flexible coordination of workflow activities. Lecture Notes in Computer Science, 1331:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brodie:1997:SBS

[1290] M. L. Brodie. Silver bullet shy on legacy mountain: When neat technology just doesn’t work — or — miracles to save the realm: Faustian bargains or noble pursuits. Lecture Notes in Computer Science, 1331:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moody:1997:MLA

[1291] D. Moody. A multi-level architecture for representing enterprise data models. Lecture Notes in Computer Science, 1331:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sekine:1997:DMC

[1292] J. Sekine, A. Kitai, Y. Ooshima, and Y. Oohara. Data model for customizing DB schemas based on business policies. Lecture Notes in Computer Science, 1331:198–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalianis:1997:ECM

[1293] H. Dalianis and P. Johannesson. Explaining conceptual models — an architecture and design principles. Lecture Notes in Computer Science, 1331: 215–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruska:1997:EOO

[1294] T. Hruska and P. Kolencik. Extending an object-oriented model: Multiple class objects. Lecture Notes in Computer Science, 1331:229–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Misic:1997:FAM

[1295] V. B. Misic and S. Moser. Formal approach to metamodeling: A generic object-oriented perspective. Lecture Notes in Computer Science, 1331:243– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chu:1997:ARO

[1296] W. W. Chu and G. Zhang. Associations and roles in object-oriented modeling. Lecture Notes in Computer Science, 1331:257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Analyti:1997:PCP

[1297] A. Analyti, N. Spyratos, and P. Constantopoulos. Property covering: A powerful construct for schema derivations. Lecture Notes in Computer Science, 1331:271–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gammoudi:1997:IGH

[1298] M. M. Gammoudi, J. D. Mendes, and W. S. Pinto. Inheritance graph hierarchy construction using rectangular decomposition of a binary relation and designer feedback. Lecture Notes in Computer Science, 1331:285–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waiyamai:1997:TOD

[1299] K. Waiyamai, R. Taouil, and L. Lakhal.

Towards an object database approach for managing concept lattices. Lecture Notes in Computer Science, 1331:299– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Batini:1997:EIC

[1300] C. Batini, G. Longobardi, and S. Fornasiero. An experience of integration of conceptual schemas in the Italian public administration. Lecture Notes in Computer Science, 1331:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albrecht:1997:AOD

[1301] M. Albrecht, M. Altus, and M. Steeg. Application-oriented design of behavior: A transformational approach using RADD. Lecture Notes in Computer Science, 1331:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wu:1997:JBF

[1302] D. Wu, D. Agrawai, A. El Abbadi, and A. Singh. A Java-based framework for processing distributed objects. Lecture Notes in Computer Science, 1331:333–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malinowski:1997:FTD

[1303] E. Malinowski and S. Chakravarthy. Fragmentation techniques for distributing object-oriented databases. Lecture Notes in Computer Science, 1331:347– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pissinou:1997:ABM

[1304] N. Pissinou, K. Makki, M. Hong, and L. Ji. An agent based mobile system. Lecture Notes in Computer Science, 1331:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Navathe:1997:FCM

[1305] S. B. Navathe. Is the future of conceptual modeling bleak or bright? Lecture Notes in Computer Science, 1331: 375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1997:SPD

[1306] P. Thompson and J. W. Sweitzer. Successful practices in developing a complex information model. Lecture Notes in Computer Science, 1331:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:RCC

[1307] M. L. Lee and T. W. Ling. Resolving constraint conflicts in the integration of entity-relationship schemas. Lecture Notes in Computer Science, 1331:394– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McBrien:1997:FFS

[1308] P. McBrien and A. Poulovassilis. A formal framework for ER schema transformation. Lecture Notes in Computer Science, 1331:408–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hohenstein:1997:GAD

[1309] U. Hohenstein and V. Plesser. A generative approach to database federation. Lecture Notes in Computer Science, 1331:422–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yee:1997:VRI

[1310] L. Yee and C. Hsu. A virtual reality interface to an enterprise metadatabase. Lecture Notes in Computer Science, 1331:436–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahanayake:1997:FFC

[1311] A. N. W. Dahanayake, H. G. Sol, and J. L. G. Dietz. A fully flexible CAME in a CASE environment. Lecture Notes in Computer Science, 1331:450–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gong:1997:RDM

[1312] M. Gong, L. Scott, Y. Xiao, and R. Offen. A rapid development model for Meta-CASE tool design. Lecture Notes in Computer Science, 1331:464– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropp:1997:WPM

[1313] W. Gropp and E. Lusk. Why are PVM and MPI so different? Lecture Notes in Computer Science, 1332:3–10, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitowski:1997:CPM

[1314] J. Kitowski, K. Boryczko, and J. Moscinski. Comparison of PVM and MPI performance in short-range molecular dynamics simulation. Lecture Notes in Computer Science, 1332: 11–16, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piernas:1997:APM

[1315] J. Piernas, A. Flores, and J. M. Garcia. Analyzing the performance of MPI in a cluster of workstations based on fast Ethernet. Lecture Notes in Computer Science, 1332:17–24, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Resch:1997:CMP

[1316] M. Resch, H. Berger, and T. Boenisch.

A comparison of MPI performance on different MPPs. Lecture Notes in Computer Science, 1332:25–32, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roda:1997:PPI

[1317] J. L. Roda, C. Rodriguez, F. Almeida, and D. Gonzalez-Morales. Predicting the performance of injection communication patterns on PVM. Lecture Notes in Computer Science, 1332:33– 40, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1997:ECP

[1318] L. P. Santos, V. Castro, and A. Proenca. Evaluation of the communication performance on a parallel processing system. Lecture Notes in Computer Science, 1332:41–48, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Souza:1997:EPH

[1319] P. S. Souza, L. J. Senger, M. J. Santana, and R. C. Santana. Evaluating personal high performance computing with PVM on Windows and LINUX environments. Lecture Notes in Computer Science, 1332:49–56, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uminski:1997:EEP

[1320] P. W. Uminski, M. R. Matuszek, and H. Krawczyk. Experimental evaluation of PVM group communication. Lecture Notes in Computer Science, 1332: 57–66, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1997:AOO

[1321] J.-M. Adamo. ARCH, an object oriented MPI-Based library for asynchronous and loosely synchronous parallel system programming. Lecture Notes in Computer Science, 1332:67– 74, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beisel:1997:EMD

[1322] T. Beisel, E. Gabriel, and M. Resch. An extension to MPI for distributed computing on MPPs. Lecture Notes in Computer Science, 1332:75–82, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiNapoli:1997:DCA

[1323] C. Di Napoli, M. Giordano, and M. M. Furnari. Distributed and cooperative applications in PVM. Lecture Notes in Computer Science, 1332:83–90, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagg:1997:HMA

[1324] G. E. Fagg, J. J. Dongarra, and A. Geist. Heterogeneous MPI application interoperation and process management under PVMPI. Lecture Notes in Computer Science, 1332:91– 98, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galibert:1997:YCL

[1325] O. Galibert. YLC, A C++ Linda system on top of PVM. Lecture Notes in Computer Science, 1332:99–106, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1997:ACP

[1326] G. A. Geist. Advanced capabilities in PVM 3.4. Lecture Notes in Computer Science, 1332:107–115, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geist:1997:BPW

[1327] G. A. Geist, J. A. Kohl, P. M. Papadopoulos, and S. L. Scott. Beyond PVM 3.4: What we’ve learned, what’s next, and why. Lecture Notes in Computer Science, 1332:116–126, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoyos-Rivera:1997:UPB

[1328] G. J. Hoyos-Rivera and V. G. Sanchez-Arias. Using PVM to build an interface to support cooperative work in a distributed systems environment. Lecture Notes in Computer Science, 1332:127– 134, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mintchev:1997:TPM

[1329] S. Mintchev and V. Getov. Towards portable message passing in Java: Binding MPI. Lecture Notes in Computer Science, 1332:135–142, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mysliwiec:1997:CAM

[1330] G. Mysliwiec, J. Sipowicz, and R. Schaefer. Control activities in message passing environment. Lecture Notes in Computer Science, 1332:143– 150, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoropoulos:1997:GSP

[1331] P. Theodoropoulos, P. Tsanakas, and G. Papakonstantinou. Global semaphores in a parallel programming environment. Lecture Notes in Computer Science, 1332:151–158, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlassov:1997:SSM

[1332] V. Vlassov and L.-E. Thorelli. A synchronizing shared memory: Model and programming implementation. Lecture Notes in Computer Science, 1332:159– 166, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1997:CMP

[1333] K. Wolf, E. Brakkee, and D. P. Ho. Communication in multi-physics applications. Lecture Notes in Computer Science, 1332:167–176, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1997:ESP

[1334] M. Fischer and J. Simon. Embedding SCI into PVM. Lecture Notes in Computer Science, 1332:177–184, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hempel:1997:IMN

[1335] R. Hempel, H. Ritzdorf, and F. Zimmermann. Implementation of MPI on NEC’s SX-4 multi-node architecture. Lecture Notes in Computer Science, 1332:185–193, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overeinder:1997:BCD

[1336] B. J. Overeinder and P. M. A. Sloot.

Breaking the curse of dynamics by task migration: Pilot experiments in the Polder Metacomputer. Lecture Notes in Computer Science, 1332:194–207, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rough:1997:PRD

[1337] J. Rough, A. Goscinski, and D. De Paoli. PVM on the RHODOS distributed operating system. Lecture Notes in Computer Science, 1332:208– 218, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alfaro:1997:FDW

[1338] F. J. Alfaro, J. A. Gallud, and J. L.

Sanchez. A function to dynamic workload allocation in distributed applications. Lecture Notes in Computer Science, 1332:219–225, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1997:IPW

[1339] R. Baraglia, M. Cosso, D. Laforenza, and M. Nicosia. Integrating PVaniM into WAMM for monitoring metaapplications. Lecture Notes in Computer Science, 1332:226–233, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1997:EPA

[1340] M. Bubak, W. Funika, and J. Moscinski. Evaluation of parallel application’s behavior in message passing environment. Lecture Notes in Computer Science, 1332:234–241, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cotronis:1997:MPP

[1341] J. Y. Cotronis. Message-passing program development by ensemble. Lecture Notes in Computer Science, 1332: 242–249, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1997:IPD

[1342] B. Di Martino, A. Mazzeo, N. Mazzocca, and U. Villano. Interaction patterns detection in PVM programs to support simulation. Lecture Notes in Computer Science, 1332:250–256, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krotz-Vogel:1997:PPP

[1343] W. Krotz-Vogel and H.-C. Hoppe. The PALLAS parallel programming environment. Lecture Notes in Computer Science, 1332:257–266, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ludwig:1997:OUI

[1344] T. Ludwig and R. Wismueller. OMIS 2.0 — A universal interface for monitoring systems. Lecture Notes in Computer Science, 1332:267–276, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manegold:1997:QBM

[1345] S. Manegold, F. Waas, and D. Gudlat.

In quest of the bottleneck — monitoring parallel database systems. Lecture Notes in Computer Science, 1332:277– 284, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunderam:1997:TAS

[1346] V. Sunderam, B. Topol, S. Moyer, and A. Krantz. Tools and auxiliary subsystems in PVM. Lecture Notes in Computer Science, 1332:285–294, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wismueller:1997:DMP

[1347] R. Wismueller. Debugging message passing programs using invisible message tags. Lecture Notes in Computer Science, 1332:295–304, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandrov:1997:PMC

[1348] V. Alexandrov, K. Chan, A. Gibbons, and W. Rytter. On the PVM/ MPI computations of dynamic programming recurrences. Lecture Notes in Computer Science, 1332:305–312, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1997:PBB

[1349] J. L. Alonso, H. Schmidt, and V. N. Alexandrov. Parallel branch and bound algorithms for integer and mixed integer linear programming problems under PVM. Lecture Notes in Computer Science, 1332:313–320, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaheta:1997:PIP

[1350] R. Blaheta, O. Jakl, and J. Stary. PVM-implementation of the PCG method with displacement decomposition. Lecture Notes in Computer Science, 1332:321–328, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciegis:1997:NID

[1351] R. Ciegis, R. Sablinskas, and J. Wasniewski. Numerical integration on distributed-memory parallel systems. Lecture Notes in Computer Science, 1332:329–336, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derakhshan:1997:PEP

[1352] M. Derakhshan, S. Hammarling, and A. Krommer. PINEAPL: A European project on Parallel Industrial Numerical Applications and Portable Libraries. Lecture Notes in Computer Science, 1332:337–342, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:BCA

[1353] J. J. Dongarra, F. Desprez, A. Petitet, and C. Randriamaro. Block-cyclic array redistribution on networks of workstations. Lecture Notes in Computer Science, 1332:343–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mazzariol:1997:PCS

[1354] M. Mazzariol, B. A. Gennart, V. Messerli, and R. D. Hersch. Performance of CAP-Specified linear algebra algorithms. Lecture Notes in Computer Science, 1332:351–358, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mysliwiec:1997:IPS

[1355] G. Mysliwiec, J. Sipowicz, and H. Burkhart. Implementing parallel SBS-Type linear solvers using AL-WAN. Lecture Notes in Computer Science, 1332:359–366, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peinado:1997:HPC

[1356] M. Peinado and R. Venkatesan. Highly parallel cryptographic attacks. Lecture Notes in Computer Science, 1332:367– 374, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petcu:1997:ISM

[1357] D. Petcu. Implementation of some multiprocessor algorithms for ODEs using PVM. Lecture Notes in Computer Science, 1332:375–382, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soch:1997:PGP

[1358] M. Soch, P. Tvrdik, and M. Volf. Parallel graph-partitioning using the mob heuristic. Lecture Notes in Computer Science, 1332:383–389, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zilli:1997:TBN

[1359] G. Zilli and L. Bergamaschi. Truncated block Newton and quasi-Newton methods for sparse systems of nonlinear equations. experiments on parallel platforms. Lecture Notes in Computer Science, 1332:390–400, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aversa:1997:MDP

[1360] R. Aversa, G. Iannello, and N. Mazzocca. An MPI driven parallelization strategy for different computing platforms: A case study. Lecture Notes in Computer Science, 1332:401–408, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bala:1997:PVQ

[1361] P. Bala, T. Clark, P. Grochowski, and B. Lesyng. Parallel version of a quantum classical molecular dynamics code for complex molecular and biomolecular systems. Lecture Notes in Computer Science, 1332:409–416, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bendtsen:1997:RLS

[1362] C. Bendtsen and Z. Zlatev. Running large-scale air pollution models on message passing machines. Lecture Notes in Computer Science, 1332:417–426, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozas:1997:PED

[1363] G. Bozas, M. Fleischhauer, and S. Zimmermann. PVM experiences in developing the MIDAS parallel database system. Lecture Notes in Computer Science, 1332:427–434, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clematis:1997:DNL

[1364] A. Clematis, A. Coda, M. Spagnuolo, and M. Mineter. Developing non-local iterative parallel algorithms for GIS on Cray T3D using MPI. Lecture Notes in Computer Science, 1332:435–442, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demuynck:1997:DOD

[1365] K. Demuynck, J. Broeckhove, and F. Arickx. Dynamic optimization of a distributed VR system by networkbalancing. Lecture Notes in Computer Science, 1332:443–450, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiMartino:1997:MDH

[1366] V. Di Martino and G. Ruocco. Molecular dynamics on hybrid memory machines. Lecture Notes in Computer Science, 1332:451–456, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Exbrayat:1997:OPS

[1367] M. Exbrayat and H. Kosch. Offering parallelism to a sequential database management system on a network of workstations using PVM. Lecture Notes in Computer Science, 1332:457– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galaktionov:1997:MST

[1368] A. S. Galaktionov, P. D. Anderson, and G. W. M. Peters. Mixing simulations: Tracking strongly deforming fluid volumes in 3D flows. Lecture Notes in Computer Science, 1332:463–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goumopoulos:1997:PCS

[1369] C. Goumopoulos, E. Housos, and O. Liljenzin. Parallel crew scheduling on workstation networks using PVM. Lecture Notes in Computer Science, 1332:470–477, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grecki:1997:MPE

[1370] M. Grecki, G. Jablonski, and A. Napieralski. MOPS — parallel environment for simulation of electronic circuits using physical models of semiconductor devices. Lecture Notes in Computer Science, 1332:478–485, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reinhard:1997:MHP

[1371] E. Reinhard and A. Chalmers. Message handling in parallel radiance. Lecture Notes in Computer Science, 1332:486– 493, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roy:1997:PNT

[1372] R. Roy and Z. Stankovski. Parallelization of neutron transport solvers. Lecture Notes in Computer Science, 1332: 494–501, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

GarciaSalcines:1997:PRR

[1373] E. Garcia Salcines, G. Cerruela Garcia, J. I. Benavides Benitez, and F. Muñoz Garcia. Parallel rendering of radiance on distributed memory system by PVM. Lecture Notes in Computer Science, 1332:502–507, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strietzel:1997:PTS

[1374] M. Strietzel. Parallel turbulence simulation: Resolving the inertial subrange of Kolmogorov’s spectra. Lecture Notes in Computer Science, 1332:508–516, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merker:1997:SDP

[1375] R. Merker, U. Eckhardt, D. Fimmel, and H. Schreiber. A system for designing parallel processor arrays. Lecture Notes in Computer Science, 1333:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaffer:1997:RAS

[1376] C. Schaffer. On the role of architectures in systems engineering. Lecture Notes in Computer Science, 1333:13– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sveda:1997:ASC

[1377] M. Sveda. An approach to safety critical systems design. Lecture Notes in Computer Science, 1333:34–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muller-Wipperfurth:1997:ESC

[1378] T. Müller-Wipperfürth. Extensions and semantics of controller state diagrams for practical applications. Lecture Notes in Computer Science, 1333: 50–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auer:1997:DIE

[1379] D. Auer, H. Dobler, and G. Weishaar. Design and implementation of an energy management system for an integrated iron and steel plant. Lecture Notes in Computer Science, 1333:58– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anaya:1997:RMM

[1380] A. Anaya, M. J. Rodriguez, and J. Parets. Representation and management of memory and decision in evolving software systems. Lecture Notes in Computer Science, 1333:71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1997:GGE

[1381] H. Mayr. GEM — A generic engineering framework for mechanical engineering based upon meta models. Lecture Notes in Computer Science, 1333:83– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larnac:1997:FRP

[1382] M. Larnac, V. Chapurlat, J. Magnier, and B. Chenot. Formal representation and proof of the interpreted sequential machine model. Lecture Notes in Computer Science, 1333:95–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glaesser:1997:MCE

[1383] U. Glaesser. Modelling of concurrent and embedded systems. Lecture Notes in Computer Science, 1333:108– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cau:1997:DTT

[1384] A. Cau and W.-P. De Roever. A densetime temporal logic with nice compositionality properties. Lecture Notes in Computer Science, 1333:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabalar:1997:IPD

[1385] P. Cabalar, R. P. Otero, M. Cabarcos, and A. Barreiro. Introducing planning in discrete event systems. Lecture Notes in Computer Science, 1333:146– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruenbacher:1997:MTA

[1386] P. Gruenbacher. Moving towards an agent-based engineering environment. Lecture Notes in Computer Science, 1333:314–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preucil:1997:EMD

[1387] L. Preucil and P. Stepan. Experiencing modeling and development of an intelligent autonomous robot. Lecture Notes in Computer Science, 1333:324– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Probst:1997:LDR

[1388] R. Probst and P. Kopacek. Latest developments in robotics. Lecture Notes in Computer Science, 1333:345– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno-Diaz:1997:ACB

[1389] R. Moreno-Diaz and J. Mira Mira. Architectures for complex behaviour inspired in the nervous system. Lecture Notes in Computer Science, 1333:353– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neto:1997:TUN

[1390] J. P. Neto, H. T. Siegelmann, J. F. Costa, and C. P. Suarez Araujo. Turing Universality of neural nets (revisited). Lecture Notes in Computer Science, 1333:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacak:1997:LLA

[1391] W. Jacak and S. Dreiseitl. Lifelong learning approach to intelligent agents modeling. Lecture Notes in Computer Science, 1333:367–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marik:1997:RQR

[1392] V. Marik, O. Stepankova, and J. Lazansky. Role of qualitative reasoning in a multi-agent system. Lecture Notes in Computer Science, 1333:380–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lhotska:1997:LMA

[1393] L. Lhotska. Learning in multi-agent systems: Theoretical issues. Lecture Notes in Computer Science, 1333:394– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleman-Flores:1997:CMV

[1394] M. Aleman-Flores, K. N. Leibovic, and R. Moreno-Diaz. A computational model for visual size, location and movement. Lecture Notes in Computer Science, 1333:406–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubnicki:1997:LAM

[1395] Z. Bubnicki. Logic-algebraic method for a class of knowledge based systems. Lecture Notes in Computer Science, 1333:420–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scharaenger:1997:ICM

[1396] J. Scharaenger. Image compression by multilevel polynomial interpolation and wavelet texture coding. Lecture Notes in Computer Science, 1333:431– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarez:1997:IEU

[1397] L. Alvarez, J. Esclarin, E. Gonzalez, and L. Mazorra. Image enhancement using and optimum quantizer. Lecture Notes in Computer Science, 1333:444–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nueckel:1997:PAD

[1398] A. Nueckel and A. Klappenecker. On the parametrization of algebraic discrete Fourier transforms. Lecture Notes in Computer Science, 1333:453–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burge:1997:LVIb

[1399] M. Burge and W. Burger. Learning visual ideals. Lecture Notes in Computer Science, 1333:464–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Medina-Rodriguez:1997:MRC

[1400] P. Medina-Rodriguez and E. Fernandez-Garcia. Multiple representation of complex intensity changes for image segmentation. Lecture Notes in Computer Science, 1333:473–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1997:DAA

[1401] J. Santos and R. J. Duro. Design of ANN architectures for handling the temporal dimension in signal processing. Lecture Notes in Computer Science, 1333:486–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gulias:1997:SDF

[1402] V. M. Gulias, J. S. Jorge, J. J. Quintela, and J. L. Freire. From sequential to distributed formal proofs. Lecture Notes in Computer Science, 1333: 501–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vila:1997:ICE

[1403] J. A. Vila, J. Presedo, M. F. Delgado, and R. Iglesias. Improvement in the computational efficiency in the analysis of signals by way of adaptive time frequency distributions. Lecture Notes in Computer Science, 1333:514–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mauerkirchner:1997:EBS

[1404] M. Mauerkirchner. Event based simulation of software development project planning. Lecture Notes in Computer Science, 1333:527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanabu:1997:AIW

[1405] M. Tanabu, J.-I. Iijima, and N. Shiba.

Application of intelligent worldwide DSS to optimization problems. Lecture Notes in Computer Science, 1333: 541–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacak:1997:STA

[1406] W. Jacak, G. Jahn, and J. Rozenblit. System theoretical approach to control and synchronization of teleeducation in open distributed environment. Lecture Notes in Computer Science, 1333:549–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czarczynski:1997:MAT

[1407] W. Czarczynski, R. Klempous, and J. Nikodem. Multilevel approach to traffic control system. Lecture Notes in Computer Science, 1333:561–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kueng:1997:CDT

[1408] J. Kueng, R. Wagner, and W. Woess. A comparison of 4-GL development tools in the area of metadata and dynamic user interfaces. Lecture Notes in Computer Science, 1333:567–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuett:1997:CPB

[1409] D. Schuett, F. Pichler, and J. Scharinger. Cryptographic permutations based on BOOT decompositions of Walsh matrices. Lecture Notes in Computer Science, 1333:580–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceska:1997:PCT

[1410] M. Ceska, V. Janousek, and T. Vojnar. PNtalk — a computerized tool for object oriented Petri nets modelling. Lecture Notes in Computer Science, 1333:591–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaffer:1997:SLD

[1411] C. Schaffer and J. Zeindl. System level design of microcontroller applications. Lecture Notes in Computer Science, 1333:611–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellare:1997:MUR

[1412] M. Bellare and P. Rogaway. Minimizing the use of random oracles in authenticated encryption schemes. Lecture Notes in Computer Science, 1334:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCrescenzo:1997:ZKP

[1413] G. Di Crescenzo, K. Sakurai, and M. Yung. Zero-knowledge proofs of decision power: new protocols and optimal round-complexity. Lecture Notes in Computer Science, 1334:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshiba:1997:CLT

[1414] T. Koshiba. Computational learning theoretic cryptanalysis of language theoretic cryptosystems. Lecture Notes in Computer Science, 1334:28–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1997:LSS

[1415] Y. Bai and V. Varadharajan. A language for specifying sequences of authorization transformations and its applications. Lecture Notes in Computer Science, 1334:39–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sun:1997:DCP

[1416] H.-M. Sun and B.-L. Chen. On the decomposition constructions for perfect secret sharing schemes. Lecture Notes in Computer Science, 1334:50– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biehl:1997:TVC

[1417] I. Biehl and S. Wetzel. Traceable visual cryptography. Lecture Notes in Computer Science, 1334:61–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghodosi:1997:RMA

[1418] H. Ghodosi, J. Pieprzyk, and R. Safavi-Naini. Remarks on the multiple assignment secret sharing scheme. Lecture Notes in Computer Science, 1334: 72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charnes:1997:SSH

[1419] C. Charnes, K. Martin, J. Pieprzyk, and R. Safavi-Naini. Secret sharing in hierarchical groups. Lecture Notes in Computer Science, 1334:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aura:1997:SC

[1420] T. Aura and P. Nikander. Stateless connections. Lecture Notes in Computer Science, 1334:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1997:DSP

[1421] R. Choi, J. Na, K. Lee, and E. Kim. Design of a security platform for CORBA based application. Lecture Notes in Computer Science, 1334:98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lioy:1997:SDM

[1422] A. Lioy, F. Maino, and M. Mezzalama. Secure document management and distribution in an open network environment. Lecture Notes in Computer Science, 1334:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obana:1997:ACA

[1423] S. Obana and K. Kurosawa. A02 — code = affine resolvable + BIBD. Lecture Notes in Computer Science, 1334: 118–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:MAS

[1424] K. M. Martin and R. Safavi-Naini.

Multisender authentication systems with unconditional security. Lecture Notes in Computer Science, 1334:130– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1997:PUI

[1425] K. Hayashi, E. Okamoto, and M. Mambo. Proposal of user identification scheme using mouse. Lecture Notes in Computer Science, 1334:144– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Millan:1997:EGA

[1426] W. Millan, A. Ckark, and E. Dawson. An effective genetic algorithm for finding highly nonlinear Boolean functions. Lecture Notes in Computer Science, 1334:149–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:DBF

[1427] X.-M. Zhang, Y. Zheng, and H. Imai.

Duality of Boolean functions and its cryptographic significance. Lecture Notes in Computer Science, 1334:159– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1997:CCI

[1428] E. Dawson and C.-K Wu. Construction of correlation immune Boolean functions. Lecture Notes in Computer Science, 1334:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihaljevic:1997:IKS

[1429] M. J. Mihaljevic. An improved key stream generator based on the programmable cellular automata. Lecture Notes in Computer Science, 1334:181– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joesang:1997:TPF

[1430] A. Joesang. A trust policy framework. Lecture Notes in Computer Science, 1334:192–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:CAS

[1431] L. Chang and I. S. Moskowitz. Critical analysis of security in voice hiding techniques. Lecture Notes in Computer Science, 1334:203–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1997:TER

[1432] Sangjoon Park, Sangwoo Park, Kwangjo Kim, and Dongho Won. Two efficient RSA multisignature schemes. Lecture Notes in Computer Science, 1334:217– 222, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:PSR

[1433] S. Kim, S. Park, and D. Won. Proxy signatures, revisited. Lecture Notes in Computer Science, 1334:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelsey:1997:RKC

[1434] J. Kelsey, B. Schneier, and D. Wagner.

Related-key cryptanalysis of 3-WAY, Biham-DES, CAST, DES-X NewDES, RC2, and TEA. Lecture Notes in Computer Science, 1334:233–246, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.counterpane.com/
related-key_cryptanalysis.html.

Zhu:1997:MAS

[1435] F. Zhu and B.-A. Guo. A multiplication-addition structure against differential attack. Lecture Notes in Computer Science, 1334:247–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneko:1997:SEM

[1436] Y. Kaneko, S. Moriai, and K. Ohta.

On strict estimation method of provable security against differential. Lecture Notes in Computer Science, 1334: 258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimoyama:1997:IFS

[1437] T. Shimoyama, S. Amada, and S. Moriai. Improved fast software implementation of block ciphers. Lecture Notes in Computer Science, 1334:269– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alabbadi:1997:SCH

[1438] M. M. Alabbadi. Security comments on the Hwang–Chen algebraic-code cryptosystem. Lecture Notes in Computer Science, 1334:274–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miyaji:1997:EEC

[1439] A. Miyaji, T. Ono, and H. Cohen. Efficient elliptic curve exponentiation. Lecture Notes in Computer Science, 1334: 282–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chao:1997:ECS

[1440] J. Chao, N. Matsuda, and S. Tsujii. Efficient construction of secure hyperelliptic discrete logarithm problems. Lecture Notes in Computer Science, 1334:292–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bleichenbacher:1997:NOC

[1441] D. Bleichenbacher, M. Joye, and J.-J. Quisquater. A new and optimal chosenmessage attack on RSA-type cryptosystems. Lecture Notes in Computer Science, 1334:302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakai:1997:WRK

[1442] Y. Sakai, K. Sakurai, and I. Ishizuka. On weak RSA-keys produced from Pretty Good Privacy. Lecture Notes in Computer Science, 1334:314–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobara:1997:SSM

[1443] K. Kobara and H. Imai. Selfsynchronized message randomization methods for subliminal channels. Lecture Notes in Computer Science, 1334: 325–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chapman:1997:HHS

[1444] M. Chapman and D. Davida. Hiding the hidden: A software system for concealing ciphertext as innocuous text. Lecture Notes in Computer Science, 1334:335–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1997:DSP

[1445] C. Boyd. Digital signature and public key cryptosystem in a prime order subgroup of Zn0*. Lecture Notes in Computer Science, 1334:346–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:1997:TOW

[1446] J. Patarin and L. Goubin. Trapdoor one-way permutations and multivariate polynomials. Lecture Notes in Computer Science, 1334:356–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:1997:ACB

[1447] J. Patarin and L. Goubin. Asymmetric cryptography with S-boxes. Lecture Notes in Computer Science, 1334:369– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camion:1997:PS

[1448] P. Camion and H. Chabanne. On the powerline system. Lecture Notes in Computer Science, 1334:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traore:1997:MUF

[1449] J. Traore. Making unfair a “fair” blind signature scheme. Lecture Notes in Computer Science, 1334:386–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1997:ETS

[1450] C. Boyd. Enforcing traceability in software. Lecture Notes in Computer Science, 1334:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1997:PVP

[1451] W. Mao. Publicly verifiable partial key escrow. Lecture Notes in Computer Science, 1334:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1997:SCR

[1452] H. Watanabe and T. Kasami. A secure code for recipient watermarking against conspiracy attacks by all users. Lecture Notes in Computer Science, 1334:414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1997:PIP

[1453] J. W. Gray and K. F. Epsilon. Protocols for issuing public-key certificates over the Internet. Lecture Notes in Computer Science, 1334:424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Postma:1997:DCF

[1454] A. Postma, T. Krol, and E. Molenkamp.

Distributed cryptographic function application protocols. Lecture Notes in Computer Science, 1334:435–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogata:1997:FTA

[1455] W. Ogata, K. Kurosawa, K. Sako, and K. Takatani. Fault tolerant anonymous channel. Lecture Notes in Computer Science, 1334:440–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrer:1997:ISS

[1456] J. D. Ferrer and R. X. Sanchez del Castillo. An implementable scheme for secure delegation of computing and data. Lecture Notes in Computer Science, 1334:445–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanderMerwe:1997:ECS

[1457] J. Van der Merwe and S. H. Von Solms.

Electronic commerce with secure intelligent trade agent. Lecture Notes in Computer Science, 1334:452–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peterson:1997:ESF

[1458] H. Peterson and G. Poupard. Efficient scalable fair cash with off-line extortion prevention. Lecture Notes in Computer Science, 1334:463–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:AUP

[1459] L. Chen and C. J. Mitchell. An anonymous and undeniable payment scheme. Lecture Notes in Computer Science, 1334:478–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williamson:1997:GAL

[1460] D. P. Williamson. Gadgets, approximation, and linear programming: Improving hardness results for cut and satisfiability problems. Lecture Notes in Computer Science, 1335:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alimonti:1997:NOL

[1461] P. Alimonti. Non-oblivious local search for MAX 2-CCSP with application to MAX DICUT. Lecture Notes in Computer Science, 1335:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alt:1997:NSC

[1462] H. Alt, U. Fuchs, and K. Kriegel. On the number of simple cycles in planar graphs. Lecture Notes in Computer Science, 1335:15–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babel:1997:SHD

[1463] L. Babel and S. Olariu. On the separable-homogeneous decomposition of graphs. Lecture Notes in Computer Science, 1335:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babel:1997:PHG

[1464] L. Babel and G. J. Woeginger. Pseudo-Hamiltonian graphs. Lecture Notes in Computer Science, 1335:38–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bermond:1997:AOD

[1465] J.-C. Bermond, M. Di Ianni, M. Flammini, and S. Perennes. Acyclic orientations for deadlock prevention in interconnection networks. Lecture Notes in Computer Science, 1335:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertet:1997:WOE

[1466] K. Bertet, J. Gustedt, and M. Morvan. Weak-order extensions of an order. Lecture Notes in Computer Science, 1335:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoni:1997:UBM

[1467] A. Bertoni, P. Campadelli, and R. Posenato. An upper bound for the maximum cut mean value. Lecture Notes in Computer Science, 1335:78– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1997:NCR

[1468] U. Brandes and D. Handke. NP-Completeness results for minimum planar spanners. Lecture Notes in Computer Science, 1335:85–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandt:1997:CIN

[1469] S. Brandt. Computing the independence number of dense triangle-free graphs. Lecture Notes in Computer Science, 1335:100–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broersma:1997:ATM

[1470] H. J. Broersma, E. Dahlhaus, and T. Kloks. Algorithms for the treewidth and minimum fill-in of HHD-Free graphs. Lecture Notes in Computer Science, 1335:109–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1997:MEO

[1471] E. Dahlhaus. Minimal elimination ordering inside a given chordal graph. Lecture Notes in Computer Science, 1335:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAmore:1997:LAN

[1472] F. D’Amore and F. Iacobini. Online algorithms for networks of temporal constraints. Lecture Notes in Computer Science, 1335:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFluiter:1997:PAT

[1473] B. De Fluiter and H. L. Bodlaender. Parallel algorithms for treewidth two. Lecture Notes in Computer Science, 1335:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinitz:1997:OGE

[1474] Y. Dinitz, M. Feighelstein, and S. Zaks.

On optimal graphs embedded into paths and rings, with analysis using l1-Spheres. Lecture Notes in Computer Science, 1335:171–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dragan:1997:GMO

[1475] F. F. Dragan. On greedy matching ordering and greedy matchable graphs. Lecture Notes in Computer Science, 1335:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erlebach:1997:LLC

[1476] T. Erlebach and K. Jansen. Offline and on-line call-scheduling in stars and trees. Lecture Notes in Computer Science, 1335:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlineny:1997:CCK

[1477] P. Hlineny and J. Kratochvil. Computational complexity of the krausz dimension of graphs. Lecture Notes in Computer Science, 1335:214–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1997:ASG

[1478] T. Kloks, D. Kratsch, and H. Mueller.

Asteroidal sets in graphs. Lecture Notes in Computer Science, 1335:229– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kratochvil:1997:CCG

[1479] J. Kratochvil, A. Proskurowski, and J. A. Telle. Complexity of colored graph covers I. colored directed multigraphs. Lecture Notes in Computer Science, 1335:242–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mosbah:1997:SAR

[1480] M. Mosbah and N. Saheb. A syntactic approach to random walks on graphs. Lecture Notes in Computer Science, 1335:258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prisner:1997:BGI

[1481] E. Prisner. Bicliques in graphs II: Recognizing k-path graphs and underlying graphs of line digraphs. Lecture Notes in Computer Science, 1335:273– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sampels:1997:LNS

[1482] M. Sampels. Large networks with small diameter. Lecture Notes in Computer Science, 1335:288–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skodinis:1997:BTW

[1483] K. Skodinis. The bounded tree-width problem of context-free graph languages. Lecture Notes in Computer Science, 1335:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thorup:1997:SPS

[1484] M. Thorup. Structured programs have small tree-width and good register allocation. Lecture Notes in Computer Science, 1335:318–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uehara:1997:MPL

[1485] R. Uehara. A measure of parallelization for the lexicographically first maximal subgraph problems. Lecture Notes in Computer Science, 1335:333–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vos:1997:MYE

[1486] T. Vos and D. Swierstra. Make your enemies transparent. Lecture Notes in Computer Science, 1335:342–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wada:1997:OFT

[1487] K. Wada, W. Chen, Y. Luo, and K. Kawaguchi. Optimal fault-tolerant ATM-Routings for biconnected graphs. Lecture Notes in Computer Science, 1335:354–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsolf:1997:GOC

[1488] B. A. Marsolf, A. J. C. Bik, K. A.

Gallivan, and H. A. G. Wijshoff. The generation of optimized codes using nonzero structure analysis. Lecture Notes in Computer Science, 1336:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1997:IEE

[1489] G. R. Gao and V. Sarkar. On the importance of an end-to-end view of memory consistency in future computer systems. Lecture Notes in Computer Science, 1336:30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gannon:1997:HPD

[1490] D. Gannon. High performance distributed object systems. Lecture Notes in Computer Science, 1336:42– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veidenbaum:1997:ICP

[1491] A. V. Veidenbaum. Instruction cache prefetching using multilevel branch prediction. Lecture Notes in Computer Science, 1336:51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cybenko:1997:HPW

[1492] G. Cybenko. High performance wireless computing. Lecture Notes in Computer Science, 1336:71–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arabnia:1997:HPC

[1493] H. R. Arabnia. High-performance computing and applications in image processing and computer vision. Lecture Notes in Computer Science, 1336:72– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watanabe:1997:PFH

[1494] T. Watanabe. Present and future of HPC technologies. Lecture Notes in Computer Science, 1336:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eickemeyer:1997:EMP

[1495] R. J. Eickemeyer, R. E. Johnson, S. R.

Kunkel, and B. H. Lim. Evaluation of multithreaded processors and thread-switch policies. Lecture Notes in Computer Science, 1336:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kacsuk:1997:MIC

[1496] P. Kacsuk and M. Amamiya. A multithreaded implementation concept of Prolog on Datarol-II machine. Lecture Notes in Computer Science, 1336:91– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Evripidou:1997:TSU

[1497] S. Evripidou. Thread synchronization unit (TSU): A building block for high performance computers. Lecture Notes in Computer Science, 1336:107– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:DDP

[1498] T. Sato. Data dependence path reduction with tunneling load instructions. Lecture Notes in Computer Science, 1336:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imlig:1997:PEE

[1499] N. Imlig and A. Tsutui. Performance estimation of embedded software with pipeline and cache hazard modeling. Lecture Notes in Computer Science, 1336:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajo:1997:IED

[1500] H. Nakajo, A. Ichikawa, and Y. Kaneda. An implementation and evaluation of a distributed shared-memory system on workstation clusters using fast serial links. Lecture Notes in Computer Science, 1336:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ma:1997:DOC

[1501] J. Ma, R. Huang, and E. Tsuboi. Designing and optimizing 3-connectivity communication networks using a distributed genetic algorithm. Lecture Notes in Computer Science, 1336:159– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funahashi:1997:ARR

[1502] A. Funahashi, T. Hanawa, H. Amano, and T. Kudoh. Adaptive routing on the recursive diagonal torus. Lecture Notes in Computer Science, 1336:171– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brownhill:1997:AML

[1503] C. J. Brownhill, A. Nicolau, S. Novack, and C. D. Polychronopoulos. Achieving multi-level parallelization. Lecture Notes in Computer Science, 1336:183– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubota:1997:TER

[1504] A. Kubota, S. Tatsumi, T. Tanaka, and M. Goshima. A technique to eliminate redundant inter-processor communication on parallelizing compiler TINPAR. Lecture Notes in Computer Science, 1336:195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uehara:1997:AVP

[1505] T. Uehara, Y. Kunieda, and T. Tsuda.

An automatic vectorizing/parallelizing Pascal compiler V-Pascal V.3. Lecture Notes in Computer Science, 1336:205– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishizaki:1997:AAD

[1506] K. Ishizaki, H. Komatsu, and T. Nakatani. An algorithm for automatic detection of loop indices for communication overlapping. Lecture Notes in Computer Science, 1336:217– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sasakura:1997:NVI

[1507] M. Sasakura, K. Joe, and K. Araki.

Nara view: An interactive 3D visualization system for parallelization of programs. Lecture Notes in Computer Science, 1336:231–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inenaga:1997:HAN

[1508] K. Inenaga, S. Kusakabe, T. Morimoto, and M. Amamiya. Hybrid approach for non-strict dataflow program on commodity machine. Lecture Notes in Computer Science, 1336:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nobukuni:1997:RMM

[1509] Y. Nobukuni, T. Matsumoto, and K. Hiraki. Resource management methods for general purpose massively parallel OS SSS-Core. Lecture Notes in Computer Science, 1336:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uchihira:1997:SBH

[1510] N. Uchihira, H. Kawata, and F. Tamura. Scenario-based hypersequential programming: Formulation of parallelization. Lecture Notes in Computer Science, 1336:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akiyama:1997:PSP

[1511] Y. Akiyama, M. Saito, T. Noguchi, and K. Onizuka. Parallelization of space plasma particle simulation. Lecture Notes in Computer Science, 1336:281– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSturler:1997:IIS

[1512] E. De Sturler and D. Loher. Implementing iterative solvers for irregular sparse matrix problems in High Performance Fortran. Lecture Notes in Computer Science, 1336:293–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mutenda:1997:PNN

[1513] L. Mutenda, M. Hiyama, T. Yoshinaga, and T. Baba. Parallel navigation in an A-NETL based parallel OODBMS. Lecture Notes in Computer Science, 1336:305–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimizu:1997:HPP

[1514] N. Shimizu and T. Watanabe. High performance parallel FFT on distributed memory parallel computers. Lecture Notes in Computer Science, 1336:317–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tooyama:1997:PCM

[1515] T. Tooyama and S. Horiguchi. Parallel computation model LogPQ. Lecture Notes in Computer Science, 1336: 327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nanri:1997:CEC

[1516] T. Nanri, H. Sato, and M. Shimasaki.

Cost estimation of coherence protocols of software managed cache on distributed shared memory system. Lecture Notes in Computer Science, 1336: 335–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:PDS

[1517] H. Sato, T. Nanri, and M. Shimasaki.

A portable distributed shared memory system on the cluster environment:

Design and implementation fully in software. Lecture Notes in Computer Science, 1336:343–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Omori:1997:OOF

[1518] Y. Omori, A. Fukuda, and K. Joe. An object-oriented framework for loop parallelization. Lecture Notes in Computer Science, 1336:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ogasawara:1997:MRR

[1519] T. Ogasawara and H. Komatsu. A method for runtime recognition of collective communication on distributedmemory multiprocessors. Lecture Notes in Computer Science, 1336:361–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishi:1997:IPA

[1520] K. Nishi, J. Cheng, and K. Ushijima. Improving the performance of automated forward deduction system EnCal. Lecture Notes in Computer Science, 1336:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuseki:1997:EPM

[1521] H. Mizuseki, K. Esfarjani, Z.-Q. Li, and K. Ohno. Efficiency of parallel machine for large-scale simulation in computational physics. Lecture Notes in Computer Science, 1336:381–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onizuka:1997:PPD

[1522] K. Onizuka, T. Noguchi, M. Saito, and Y. Akiyama. Parallel PDB data retriever “PDB diving booster”. Lecture Notes in Computer Science, 1336:389– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cristea:1997:PMN

[1523] A. I. Cristea and T. Okamoto. A parallelization method for neural networks with weak connection design. Lecture Notes in Computer Science, 1336:397– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torresen:1997:EPC

[1524] J. Torresen, O. Landsverk, S. Mori, and H. Nakashima. Exploiting parallel computers to reduce neural network training time of real applications. Lecture Notes in Computer Science, 1336:405–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauer:1997:MFT

[1525] F. L. Bauer. The might of formulas and their limits. Lecture Notes in Computer Science, 1337:1– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370001.
pdf.

Zemanek:1997:HSE

[1526] Heinz Zemanek. Hardware — software: An equivalence and a contradiction. Lecture Notes in Computer Science, 1337:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1337/13370009.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1337/13370009.
pdf.

Coy:1997:DD

[1527] Wolfgang Coy. Defining discipline. Lecture Notes in Computer Science, 1337: 21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370021.
pdf.

Siefkes:1997:CSC

[1528] Dirk Siefkes. Computer science as cultural development — toward a broader theory. Lecture Notes in Computer Science, 1337:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370037.
pdf.

Gruska:1997:TAI

[1529] Jozef Gruska and Roland Vollmar. Towards adjusting informatics education to information era. Lecture Notes in Computer Science, 1337:49– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370049.
pdf.

Klaeren:1997:ISC

[1530] Herbert Klaeren, Christiane Floyd, and Friedrich Diestelmeier. Informatics and society: A curriculum for distance education. Lecture Notes in Computer Science, 1337:69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370069.
pdf.

Mateescu:1997:SSA

[1531] Alexandru Mateescu, Grzegorz Rozenberg, and Arto Salomaa. Syntactic and semantic aspects of parallelism. Lecture Notes in Computer Science, 1337: 79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370079.
pdf.

Kupka:1997:UFC

[1532] Ingbert Kupka. Unique fixpoints in complete lattices with applications to formal languages and semantics. Lecture Notes in Computer Science, 1337: 107–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370107.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370107.
                                        pdf.

Karner:1997:AFL

[1533] Georg Karner and Werner Kuich. On abstract families of languages, power series, and elements. Lecture Notes in Computer Science, 1337:117–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370117.
                                        pdf.

Lange:1997:TFL

[1534] Klaus-Jörn Lange. Are there formal languages complete for SymSPACE(log n)? Lecture Notes in Computer Science, 1337:125–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1337/13370125.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370125.
                                        pdf.

Jantzen:1997:TCT

[1535] Matthias Jantzen. On twist-closed trios: A new morphic characterization of r.e. sets. Lecture Notes in Computer Science, 1337:135–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370135.
pdf.

Esparza:1997:AAS

[1536] Javier Esparza and Peter Rossmanith.

An automata approach to some problems on context-free grammars. Lecture Notes in Computer Science, 1337: 143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370143.
pdf.

Culik:1997:ASW

[1537] Karel Culik and Jarkko Kari. On aperiodic sets of Wang tiles. Lecture Notes in Computer Science, 1337:153– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370153.
pdf.

Jenner:1997:CUC

[1538] Birgit Jenner. Closure under complementation of logspace complexity classes — A survey. Lecture Notes in Computer Science, 1337:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370163.
                                        pdf.

Kirsig:1997:RBS

[1539] Bernd Kirsig. A relation between sparse and printable sets in NSPACE(log n). Lecture Notes in Computer Science, 1337:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370177.
                                        pdf.

Weihrauch:1997:FCAa

[1540] Klaus Weihrauch. A foundation for computable analysis. Lecture Notes in Computer Science, 1337:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370185.
                                        pdf.

Schmidhuber:1997:CSV

[1541] Jürgen Schmidhuber. A computer scientist’s view of life, the universe, and everything. Lecture Notes in Computer Science, 1337:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370201.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/1337/13370201.
pdf.

Kudlek:1997:CC

[1542] Manfred Kudlek. Calendars and chronologies. Lecture Notes in Computer Science, 1337:209–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/1337/13370209.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1337/13370209.pdf.

Ehrig:1997:UAP

[1543] Hartmut Ehrig and Julia Padberg. A uniform approach to Petri nets. Lecture Notes in Computer Science, 1337: 219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370219.
pdf.

Kiehn:1997:OPO

[1544] Astrid Kiehn. Observing partial order runs of Petri nets. Lecture Notes in Computer Science, 1337:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370233.
pdf.

Meseguer:1997:RTP

[1545] José Meseguer, Ugo Montanari, and Vladimir Sassone. Representation theorems for Petri nets. Lecture Notes in Computer Science, 1337:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370239.
                                        pdf.

Diekert:1997:RTE

[1546] Volker Diekert. A remark on trace equations. Lecture Notes in Computer Science, 1337:251–260, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370251.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370251.
                                        pdf.

Kindler:1997:VDA

[1547] Ekkart Kindler and Wolfgang Reisig. Verification of distributed algorithms with algebraic Petri nets. Lecture Notes in Computer Science, 1337:261– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370261.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370261.
                                        pdf.

Vogler:1997:SSA

[1548] Walter Vogler. A short story on action refinement. Lecture Notes in Computer Science, 1337:271–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370271.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370271.
pdf.

Broy:1997:IRS

[1549] Manfred Broy. Interactive and reactive systems: States, observations, experiments, input, output, nondeterminism, compositionality, and all that. Lecture Notes in Computer Science, 1337:279– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370279.
pdf.

Daduna:1997:DTA

[1550] Hans Daduna. Discrete time analysis of a state dependent tandem with different customer types. Lecture Notes in Computer Science, 1337:287– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370287.
pdf.

Desel:1997:HDA

[1551] Jörg Desel. How distributed algorithms play the token game. Lecture Notes in Computer Science, 1337:297– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370297.
                                        pdf.

Walter:1997:ASR

[1552] Rolf Walter. The asynchronous stack revisited: Rounds set the twilight reeling. Lecture Notes in Computer Science, 1337:307–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370307.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370307.
                                        pdf.

Monien:1997:OSC

[1553] B. Monien, P. Berenbrink, R. Lüling, and M. Riedel. Online scheduling of continuous media streams. Lecture Notes in Computer Science, 1337:313– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370313.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370313.
                                        pdf.

Langmaack:1997:CGG

[1554] Hans Langmaack. Contribution to Goodenough’s and Gerhart’s theory of software testing and verification: relation between strong compiler test and compiler implementation verification. Lecture Notes in Computer Science, 1337:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370321.
pdf.

Hotz:1997:ACU

[1555] Günter Hotz and Hongzhong Wu.

On the arrangement complexity of uniform trees. Lecture Notes in Computer Science, 1337:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370337.
pdf.

Boley:1997:RFI

[1556] Harold Boley. A relational-functional integration for declarative programming. Lecture Notes in Computer Science, 1337:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1337/13370351.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1337/13370351.pdf.

Kanngiesser:1997:IOC

[1557] Siegfried Kanngießer. Inside and outside the Chinese room. Lecture Notes in Computer Science, 1337:359– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370359.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370359.
                                        pdf.

Habel:1997:ASS

[1558] Christopher Habel and Carola Eschenbach. Abstract structures in spatial cognition. Lecture Notes in Computer Science, 1337:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370369.
                                        pdf.

Freksa:1997:STS

[1559] Christian Freksa. Spatial and temporal structures in cognitive processes. Lecture Notes in Computer Science, 1337: 379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370379.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370379.
                                        pdf.

Hernandes:1997:QVF

[1560] Daniel Hernándes. Qualitative vs. fuzzy representations of spatial distance. Lecture Notes in Computer Science, 1337:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1337/13370389.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1337/13370389.pdf.

Hernandez:1997:QVF

[1561] D. Hernandez. Qualitative vs. fuzzy representations of spatial distance. Lecture Notes in Computer Science, 1337: 389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pribbenow:1997:WPF

[1562] Simone Pribbenow. What’s a part? — on formalizing partwhole relations. Lecture Notes in Computer Science, 1337:399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370399.
pdf.

Rollinger:1997:SFS

[1563] Claus R. Rollinger. SPOCK: A feasibility study on the completeness of parts lists. Lecture Notes in Computer Science, 1337:407–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370407.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/1337/13370407.
                                        pdf.

Schill:1997:DSS

[1564] Kerstin Schill. Decision support systems with adaptive reasoning strategies. Lecture Notes in Computer Science, 1337:417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1337/13370417.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1337/13370417.pdf.

Morik:1997:KDD

[1565] Katharina Morik. Knowledge discovery in databases — an inductive logic programming approach. Lecture Notes in Computer Science, 1337:429– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1337/13370429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1337/13370429.
                                        pdf.

Ammon:1997:CH

[1566] Kurt Ammon. The composition heuristic. Lecture Notes in Computer Science, 1337:437–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1337/13370437.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1337/13370437.pdf.

Weiss:1997:JAP

[1567] Gerhard Weiß. The job assignment problem: A study in parallel and distributed machine learning. Lecture Notes in Computer Science, 1337:445– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370445.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370445.
pdf.

Hamdi:1997:SIB

[1568] Mohamed Salah Hamdi and Karl Kaiser. Self-improving behavior arbitration. Lecture Notes in Computer Science, 1337:455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1337/13370455.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1337/13370455.pdf.

Sturm:1997:NNM

[1569] Margit Sturm. Neural networks for manipulator path planning. Lecture Notes in Computer Science, 1337:465– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370465.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370465.
pdf.

Eldracher:1997:PPU

[1570] Martin Eldracher and Thomas Pic.

Path planning using a subgoal graph.

Lecture Notes in Computer Science, 1337:473–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1337/13370473.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1337/13370473.
pdf.

Schittenkopf:1997:NMC

[1571] Christian Schittenkopf and Gustavo Deco. A nonlinear Markovian characterization of time series using neural networks. Lecture Notes in Computer Science, 1337:481–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370481.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370481.
pdf.

Scheler:1997:FBP

[1572] Gabriele Scheler. Feature-based perception of semantic concepts. Lecture Notes in Computer Science, 1337:489– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370489.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370489.
pdf.

Ruge:1997:ADT

[1573] Gerda Ruge. Automatic detection of thesaurus relations for information retrieval applications. Lecture Notes in Computer Science, 1337:499– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370499.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370499.
pdf.

Pfefferer:1997:IVN

[1574] Leo Pfefferer and Dieter Schütt.

InfoSphere^{T M}-V: A new approach to 3D-visualization of information. Lecture Notes in Computer Science, 1337: 507–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1337/13370507.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1337/13370507.
pdf.

Feldmann:1997:CCA

[1575] R. Feldmann. Computer chess: Algorithms and heuristics for a deep look into the future. Lecture Notes in Computer Science, 1338:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKreveld:1997:ATT

[1576] M. Van Kreveld. Algorithms for triangulated terrains. Lecture Notes in Computer Science, 1338:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1997:DRP

[1577] K.-J Lange. On the distributed realization of parallel algorithms. Lecture Notes in Computer Science, 1338:37– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Makowsky:1997:FPD

[1578] J. A. Makowsky and E. V. Ravve. The fundamental problem of database design. Lecture Notes in Computer Science, 1338:53–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nesetril:1997:SAC

[1579] J. Nesetril and D. Turzik. Solving and approximating combinatorial optimization problems (towards MAX CUT and TSP). Lecture Notes in Computer Science, 1338:70–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orponen:1997:CPC

[1580] P. Orponen. The computational power of continuous time neural networks. Lecture Notes in Computer Science, 1338:86–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weihrauch:1997:FCAb

[1581] K. Weihrauch. A foundation for computable analysis. Lecture Notes in Computer Science, 1338:104–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wiedermann:1997:TMC

[1582] J. Wiedermann. Towards machines that can think. Lecture Notes in Computer Science, 1338:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wozniakowski:1997:CCC

[1583] H. Wozniakowski. Computational complexity of continuous problems. Lecture Notes in Computer Science, 1338: 142–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaks:1997:PLA

[1584] S. Zaks. Path layout in ATM networks. Lecture Notes in Computer Science, 1338:144–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Covaci:1997:MAT

[1585] S. Covaci and T. Magedanz. The mobile agent technology. Lecture Notes in Computer Science, 1338:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duce:1997:TPI

[1586] D. A. Duce. Theory and practice in interactionally rich distributed systems. Lecture Notes in Computer Science, 1338:163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Issarny:1997:CBP

[1587] V. Issarny. Configuration-based programming systems. Lecture Notes in Computer Science, 1338:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muthukrishnan:1997:AGP

[1588] P. K. Muthukrishnan and B. R. Bryant.

Automatic generation of parallelizing compilers for object-oriented programming languages from denotational semantics specifications. Lecture Notes in Computer Science, 1338:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bjoerner:1997:FSE

[1589] D. Bjoerner. A formal software engineering paradigm: From domains via requirements to software — formal specification and design calculi. Lecture Notes in Computer Science, 1338: 219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuntzmann-Combelles:1997:WPS

[1590] A. Kuntzmann-Combelles. The whole picture to software process improvement. Lecture Notes in Computer Science, 1338:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pree:1997:OOD

[1591] W. Pree. Object-oriented design patterns. Lecture Notes in Computer Science, 1338:266–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dittrich:1997:OOD

[1592] K. R. Dittrich and A. Geppert. Objectoriented DBMS and beyond. Lecture Notes in Computer Science, 1338:275– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1997:IRO

[1593] M. E. Orlowska, H. Li, and C. Liu. On integration of relational and objectoriented database systems. Lecture Notes in Computer Science, 1338:295– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:OTD

[1594] M. H. Williams and N. W. Paton.

From OO through deduction to active databases — ROCK, ROLL and RAP. Lecture Notes in Computer Science, 1338:313–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zara:1997:IVR

[1595] J. Zara. An introduction to virtual reality modeling language. Lecture Notes in Computer Science, 1338:331–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zlatuska:1997:SSI

[1596] J. Zlatuska. Stepping stones to an information society. Lecture Notes in Computer Science, 1338:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becchetti:1997:LBV

[1597] L. Becchetti and C. Gaibisso. Lower bounds for the virtual path layout problem in ATM networks. Lecture Notes in Computer Science, 1338:375– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dai:1997:QPT

[1598] B. Dai, D. A. Bell, and J. G. Hughes.

Query processing in temporal evidential databases. Lecture Notes in Computer Science, 1338:383–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCastro:1997:FAT

[1599] C. De Castro. A first approach to temporal predicate locking for concurrency detection in temporal relational databases supporting schema versioning. Lecture Notes in Computer Science, 1338:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaibisso:1997:EIA

[1600] C. Gaibisso and G. Proietti. Efficient insertion of approximately sorted sequences of items into a dictionary. Lecture Notes in Computer Science, 1338: 399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galdamez:1997:HAS

[1601] P. Galdamez, F. D. Muñoz-Escoi, and J. M. Bernabeu-Auban. High availability support in CORBA environments. Lecture Notes in Computer Science, 1338:407–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glasnak:1997:SSN

[1602] V. Glasnak. On f-sparse sets in NP– P. Lecture Notes in Computer Science, 1338:415–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hof:1997:ZOE

[1603] M. Hof, H. Moessenboeck, and P. Pirkelbauer. Zero-overhead exception handling using metaprogramming. Lecture Notes in Computer Science, 1338:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janousek:1997:OSF

[1604] J. Janousek and B. Melichar. The output-store formal translator directed by LR parsing. Lecture Notes in Computer Science, 1338:432–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kadlec:1997:PPA

[1605] J. Kadlec. Parallel processing on Alphas under MATLAB 5. Lecture Notes in Computer Science, 1338:440–447, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolman:1997:PLB

[1606] P. Kolman. PRAM lower bound for element distinctness revisited. Lecture Notes in Computer Science, 1338:448– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kopecek:1997:OTS

[1607] I. Kopecek. Optimal trees for searching in codebook. Lecture Notes in Computer Science, 1338:456–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kralovic:1997:TOS

[1608] R. Kralovic. Time optimal selfstabilizing algorithms. Lecture Notes in Computer Science, 1338:464–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kroha:1997:RSI

[1609] P. Kroha and M. Strauss. Requirements specification iteratively combined with reverse engineering. Lecture Notes in Computer Science, 1338: 473–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucera:1997:FRI

[1610] A. Kucera. On finite representations of infinite-state behaviours. Lecture Notes in Computer Science, 1338:481– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lucas:1997:ESS

[1611] S. Lucas. Efficient strong sequentiality using replacement restrictions. Lecture Notes in Computer Science, 1338: 489–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1997:OE

[1612] M. Marchiori. Optimal encodings. Lecture Notes in Computer Science, 1338: 497–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mraz:1997:MRA

[1613] F. Mraz, M. Platek, P. Jancar, and J. Vogel. Monotonic rewriting automata with a restart operation. Lecture Notes in Computer Science, 1338: 505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nienhuys-Cheng:1997:KFP

[1614] S.-H. Nienhuys-Cheng and A. De Bruin. Kahn’s fixed-point characterization for linear dynamic networks. Lecture Notes in Computer Science, 1338: 513–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pala:1997:DAC

[1615] K. Pala, P. Rychly, and P. Smrz. DE-SAM — annotated corpus for Czech. Lecture Notes in Computer Science, 1338:523–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peter:1997:MMC

[1616] Y. Peter. Mobility management in CORBA: A generic implementation of the LifeCycle service. Lecture Notes in Computer Science, 1338:531–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pijls:1997:TGT

[1617] W. Pijls and A. De Bruin. A theory of game trees, based on solution trees. Lecture Notes in Computer Science, 1338:539–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder:1997:AAV

[1618] H. Schroeder, A. E. May, I. Vrt’o, and O. Sykora. Approximation algorithms for the vertex bipartization problem. Lecture Notes in Computer Science, 1338:547–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schroeder:1997:OAA

[1619] H. Schroeder, O. Sykora, and I. Vrt’o. Optical all-to-all communication for some product graphs. Lecture Notes in Computer Science, 1338:555–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strupl:1997:PSO

[1620] D. Strupl and R. Neruda. Parallelizing self-organizing maps. Lecture Notes in Computer Science, 1338:563–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimauro:1997:HRS

[1621] G. Dimauro, S. Impedovo, G. Pirlo, and A. Salzo. Handwriting recognition: State of the art and future trends. Lecture Notes in Computer Science, 1339:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suen:1997:CCR

[1622] C. Y. Suen and Z. C. Li. Crucial combinations for the recognition of handwritten letters. Lecture Notes in Computer Science, 1339:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amin:1997:RPH

[1623] A. Amin. Recognition of printed and handwritten Arabic characters. Lecture Notes in Computer Science, 1339: 40–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

El-Yacoubi:1997:OED

[1624] A. El-Yacoubi, M. Gilloux, R. Sabourin, and C. Y. Suen. Objective evaluation of the discriminant power of features in an HMM-based word recognition system. Lecture Notes in Computer Science, 1339:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerfali:1997:EVL

[1625] W. Guerfali and R. Plamondon. Effect of variability on letters generation with the vectorial delta-lognormal model. Lecture Notes in Computer Science, 1339:74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sabourin:1997:LSV

[1626] R. Sabourin. Off-line signature verification: Recent advances and perspectives. Lecture Notes in Computer Science, 1339:84–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kasturi:1997:RAG

[1627] R. Kasturi and H. Luo. Research advances in graphics recognition: An update. Lecture Notes in Computer Science, 1339:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belaid:1997:FTR

[1628] A. Belaid. Future trends in retrospective document conversion. Lecture Notes in Computer Science, 1339:111– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgefors:1997:LLP

[1629] G. Borgefors, G. Ramella, and G. Sanniti di Baja. A. low level processing using binary pyramids to create multiresolution shape descriptors. Lecture Notes in Computer Science, 1339:129– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frucci:1997:CPC

[1630] M. Frucci and A. Marcelli. Contour pixel classification for character skeletonization. Lecture Notes in Computer Science, 1339:141–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Belaid:1997:CPV

[1631] A. Belaid and Y. Chenevoy. Constraint propagation vs. syntactical analysis for the logical structure recognition of library references. Lecture Notes in Computer Science, 1339:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tayeb-Bey:1997:GFD

[1632] S. Tayeb-Bey and A. S. Saidi. Grammatical formalism for document understanding system: From documents towards HTML text. Lecture Notes in Computer Science, 1339:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diana:1997:DMF

[1633] S. Diana, E. Trupin, Y. Lecourtier, and J. Labiche. Document modeling for form class identification. Lecture Notes in Computer Science, 1339:176– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duffy:1997:CDA

[1634] L. Duffy, F. Lebourgeois, and H. Emptoz. Composite document analysis by means of typographic characteristics. Lecture Notes in Computer Science, 1339:188–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sauvola:1997:TAT

[1635] J. Sauvola, H. Kauniskangas, D. Doermann, and M. Pietikaeinen. Techniques for the automated testing of document analysis algorithms. Lecture Notes in Computer Science, 1339:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sauvola:1997:GTT

[1636] J. Sauvola, H. Kauniskangas, D. Doermann, and C. Shin. Graphical tools and techniques for querying document image databases. Lecture Notes in Computer Science, 1339:213–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RodriguesGomes:1997:HNR

[1637] N. Rodrigues Gomes and L. L. Lee.

Handwritten numeral recognition via fuzzy logic and local discriminating features. Lecture Notes in Computer Science, 1339:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leroy:1997:ULC

[1638] A. Leroy. Unsupervised learning of character prototypes. Lecture Notes in Computer Science, 1339:237–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wimmer:1997:LCW

[1639] Z. Wimmer, S. Garcia-Salicetli, B. Dorizzi, and P. Gallinari. Off-line cursive word recognition with a hybrid Neural-HMM system. Lecture Notes in Computer Science, 1339:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1997:FPL

[1640] C. Simon, E. Levrat, R. Sabourin, and J. Bremont. A fuzzy perception for off-line handwritten signature verification. Lecture Notes in Computer Science, 1339:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martens:1997:EDH

[1641] R. Martens and L. Claesen. An evaluation of different handwriting observation techniques from a signature verification point of view. Lecture Notes in Computer Science, 1339:273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeOliveira:1997:GSD

[1642] C. De Oliveira, C. A. Kaestner, F. Bortolozzi, and R. Sabourin. Generation of signatures by deformations. Lecture Notes in Computer Science, 1339:283– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshimura:1997:ASD

[1643] M. Yoshimura and I. Yoshimura. An application of the sequential dynamic programming matching method to offline signature verification. Lecture Notes in Computer Science, 1339:299– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1997:SAF

[1644] T. W. Rauber, V. B. De Souza, and S. Rossetto. A system for automatic form reading. Lecture Notes in Computer Science, 1339:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koerich:1997:AEF

[1645] A. L. Koerich and L. L. Lee. Automatic extraction of filled-in information from bankchecks based on prior knowledge about layout structure. Lecture Notes in Computer Science, 1339:322– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1997:SMB

[1646] J. E. B. Santos, F. Bortolozzi, and R. Sabourin. A simple methodology to bankcheck segmentation. Lecture Notes in Computer Science, 1339:334– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nievergelt:1997:IGC

[1647] J. Nievergelt. Introduction to geometric computing: From algorithms to software. Lecture Notes in Computer Science, 1340:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gold:1997:VMG

[1648] C. M. Gold, P. R. Remmele, and T. Roos. Voronoi methods in GIS. Lecture Notes in Computer Science, 1340: 21–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKreveld:1997:DEM

[1649] M. Van Kreveld. Digital elevation models and TIN algorithms. Lecture Notes in Computer Science, 1340:37– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBerg:1997:VT

[1650] M. De Berg. Visualization of TINs. Lecture Notes in Computer Science, 1340:79–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weibel:1997:GSD

[1651] R. Weibel. Generalization of spatial data: Principles and selected algorithms. Lecture Notes in Computer Science, 1340:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nievergelt:1997:SDS

[1652] J. Nievergelt and P. Widmayer. Spatial data structures: Concepts and design choices. Lecture Notes in Computer Science, 1340:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bugnion:1997:SFC

[1653] E. Bugnion, T. Roos, R. Wattenhofer, and P. Widmayer. Space filling curves versus random walks. Lecture Notes in Computer Science, 1340:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arge:1997:EMA

[1654] L. Arge. External-memory algorithms with applications in GIS. Lecture Notes in Computer Science, 1340:213– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schirra:1997:PRG

[1655] S. Schirra. Precision and robustness in geometric computations. Lecture Notes in Computer Science, 1340:255–287, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlbach:1997:EDS

[1656] H. J. Ohlbach. Efficient data structures for deduction systems. Lecture Notes in Computer Science, 1341:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flach:1997:ILD

[1657] P. A. Flach. Inductive logic databases:

From extensional to intensional knowledge. Lecture Notes in Computer Science, 1341:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceri:1997:SST

[1658] S. Ceri, P. Fraternali, S. Paraboschi, and G. Psaila. Sharing software tools on the Web: The IDEA Web lab. Lecture Notes in Computer Science, 1341: 4–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vossen:1997:TW

[1659] G. Vossen. Transactional workflows. Lecture Notes in Computer Science, 1341:20–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1997:ATU

[1660] C. Baral and J. Lobo. From action theories to updates, transactions, triggers and agents. Lecture Notes in Computer Science, 1341:26–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1997:DWO

[1661] S. Chaudhuri and U. Dayal. Data warehousing and OLAP for decision support. Lecture Notes in Computer Science, 1341:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1997:MCT

[1662] G. Dong and K. Ramamohanarao. Maintaining constrained transitive closure by conjunctive queries. Lecture Notes in Computer Science, 1341:35– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gluche:1997:IUM

[1663] D. Gluche, T. Grust, C. Mainberger, and M. H. Scholl. Incremental updates for materialized OQL views. Lecture Notes in Computer Science, 1341:52– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shu:1997:VMU

[1664] H. Shu. View maintenance using conditional tables. Lecture Notes in Computer Science, 1341:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgida:1997:AUC

[1665] A. Borgida and G. Weddell. Adding uniqueness constraints to description logics. Lecture Notes in Computer Science, 1341:85–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toman:1997:PBT

[1666] D. Toman. A point-based temporal extension of SQL. Lecture Notes in Computer Science, 1341:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mumick:1997:AOV

[1667] I. S. Mumick and K. A. Ross.

Attribute-oriented view definitions in relational and deductive databases. Lecture Notes in Computer Science, 1341:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahoui-Guerni:1997:EST

[1668] M. Mahoui-Guerni. Exploiting semantic of typed objects under an optimistic control in the transactional deferred update model: A promising approach. Lecture Notes in Computer Science, 1341:140–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Theodoratos:1997:LUO

[1669] D. Theodoratos. Logical updating of object class structures. Lecture Notes in Computer Science, 1341:158–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costal:1997:AOI

[1670] D. Costal, E. Teniente, and T. Urpi. An approach to obtain intensional translations for consistent view updating. Lecture Notes in Computer Science, 1341:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertino:1997:TIO

[1671] E. Bertino, G. Guerrini, and I. Merlo. Trigger inheritance and overriding in active object database systems. Lecture Notes in Computer Science, 1341: 193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sundermier:1997:ODA

[1672] A. Sundermier, T. Ben Abdellatif, S. W. Dietrich, and S. D. Urban. Object deltas in an active database development environment. Lecture Notes in Computer Science, 1341:211–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lautemann:1997:CPD

[1673] S.-E. Lautemann, P. Eigner, and C. Woehrle. The COAST project: Design and implementation. Lecture Notes in Computer Science, 1341:229– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baral:1997:FCA

[1674] C. Baral, J. Lobo, and G. Trajcevski. Formal characterizations of active databases: Part II. Lecture Notes in Computer Science, 1341:247– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaniolo:1997:NSA

[1675] C. Zaniolo. The nonmonotonic semantics of active rules in deductive databases. Lecture Notes in Computer Science, 1341:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giannotti:1997:DBA

[1676] F. Giannotti, G. Manco, M. Nanni, and D. Pedreschi. Datalog++: A basis for active object-oriented databases. Lecture Notes in Computer Science, 1341: 283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruska:1997:CCF

[1677] T. Hruska and P. Kolencik. Comparison of categorical foundations of object-oriented database model. Lecture Notes in Computer Science, 1341: 302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1997:WFS

[1678] W. May, B. Ludaescher, and C. Lausen.

Well-founded semantics for deductive object-oriented database languages. Lecture Notes in Computer Science, 1341:320–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1997:DSD

[1679] S. Greco and D. Sacca. Deterministic semantics for datalog: Complexity and expressive power. Lecture Notes in Computer Science, 1341:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neven:1997:ISD

[1680] F. Neven and J. Van den Bussche. On implementing structured document query facilities on top of a DOOD. Lecture Notes in Computer Science, 1341: 351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jim:1997:DMA

[1681] K. Jim, J. Lai, and B. Wuethrich. A data mining algorithm optimal for single rules. Lecture Notes in Computer Science, 1341:368–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Himmeroeder:1997:DSW

[1682] R. Himmeroeder, G. Lausen, B. Ludaescher, and C. Schlepphorst. On a declarative semantics for Web queries. Lecture Notes in Computer Science, 1341:386–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aberer:1997:RBG

[1683] K. Aberer, D. Che, and K. Boehm. Rule-based generation of logical query plans with controlled complexity. Lecture Notes in Computer Science, 1341: 399–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hacid:1997:RBD

[1684] M.-S. Hacid, P. Marcel, and C. Rigotti. A rule-based data manipulation language for OLAP systems. Lecture Notes in Computer Science, 1341:417– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zimmer:1997:GME

[1685] D. Zimmer, A. Meckenstock, and R. Unland. A general model for event specification in active database management systems. Lecture Notes in Computer Science, 1341:419–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fegaras:1997:OLO

[1686] L. Fegaras. Optimizing large OODB queries. Lecture Notes in Computer Science, 1341:421–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nantajeewarawat:1997:NIT

[1687] E. Nantajeewarawat and V. Wuwongse. Nonmonotonic inheritance through specialization. Lecture Notes in Computer Science, 1341:423–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godfrey:1997:ODQ

[1688] P. Godfrey and J. Gryz. Overview of dynamic query evaluation in intensional query optimization. Lecture Notes in Computer Science, 1341:425– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kandzia:1997:TEE

[1689] P.-T Kandzia. Towards efficient evaluation of methods by reduction. Lecture Notes in Computer Science, 1341: 427–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tam:1997:CAR

[1690] K. Tam, J. Lloyd, Y. Lesperance, and H. Levesque. Controlling autonomous robots with GOLOG. Lecture Notes in Computer Science, 1342:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolle:1997:VQR

[1691] R. M. Bolle, B.-L. Yeo, and M. M. Yeung. Video query and retrieval. Lecture Notes in Computer Science, 1342: 13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sterling:1997:FNW

[1692] L. Sterling. On finding needles in WWW haystacks. Lecture Notes in Computer Science, 1342:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Havens:1997:EDB

[1693] W. S. Havens. Extending dynamic backtracking for distributed constraint satisfaction problems. Lecture Notes in Computer Science, 1342:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pang:1997:CDB

[1694] W. Pang and S. D. Goodwin. Constraint-directed backtracking. Lecture Notes in Computer Science, 1342: 47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thornton:1997:APC

[1695] J. Thornton and A. Sattar. Applied partial constraint satisfaction using weighted iterative repair. Lecture Notes in Computer Science, 1342:57– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1997:AMI

[1696] N. Ritter, R. Owens, K. Yogesan, and P. Van Saarloos. The application of mutual information to the registration of stereo and temporal images of the retina. Lecture Notes in Computer Science, 1342:67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:1997:KFP

[1697] J. R. Cooper and R. O. Hastings.

Kalman filtering from a phase based optical flow operator. Lecture Notes in Computer Science, 1342:77–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perry:1997:ACR

[1698] S. W. Perry and L. Guan. Adaptive constraint restoration and error analysis using a neural network. Lecture Notes in Computer Science, 1342:87– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grinton:1997:EAC

[1699] C. Grinton, L. Sonenberg, and L. Sterling. Exploring agent cooperation: Studies with a simple pursuit game. Lecture Notes in Computer Science, 1342:96–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abtroun:1997:SPP

[1700] M. A. Abtroun and S. Hassas. The serializability problem in a parallel rule-based system: A solution by distributed coordination. Lecture Notes in Computer Science, 1342:106–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bui:1997:OCA

[1701] H. H. Bui, D. Kieronska, and S. Venkatesh. Optimal communication among team members. Lecture Notes in Computer Science, 1342:116–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnier:1997:DRG

[1702] T. Schnier and J. Gero. Dominant and recessive genes in evolutionary systems applied to spatial reasoning. Lecture Notes in Computer Science, 1342:127–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierrot:1997:UMC

[1703] H. J. Pierrot and R. Hinterding. Using multi-chromosomes to solve a simple mixed integer problem. Lecture Notes in Computer Science, 1342:137– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohammadian:1997:HFF

[1704] M. Mohammadian and M. Kingham.

Hierarchical and feed-forward fuzzy logic for financial modelling and prediction. Lecture Notes in Computer Science, 1342:147–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debenham:1997:KBS

[1705] J. Debenham. A knowledge-based system for workflow management using the World Wide Web. Lecture Notes in Computer Science, 1342:157–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colomb:1997:UPF

[1706] R. M. Colomb and P. Y.-P. Chen.

Use of partial functional dependencies to make practical approximate translations among forms of propositional expert systems. Lecture Notes in Computer Science, 1342:167–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beydoun:1997:NAS

[1707] G. Beydoun and A. Hoffmann. NRDR for the acquisition of search knowledge. Lecture Notes in Computer Science, 1342:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lawson:1997:RPE

[1708] G. Lawson and D. Lukose. Rapid prototyping of executable problem solving methods using MODEL-ECS. Lecture Notes in Computer Science, 1342:187– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peppas:1997:PSC

[1709] P. Peppas, M. Pagnucco, M. Prokopenko, and N. Foo. Preferential semantics for causal fixpoints. Lecture Notes in Computer Science, 1342:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ONeill:1997:BUM

[1710] T. O’Neill and N. Foo. Belief update, the Markovian situation calculus, and discrete event systems. Lecture Notes in Computer Science, 1342: 207–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kwok:1997:NCT

[1711] R. B. H. Kwok, A. Nayak, and N. Foo. A notion of correctness with theories containing theoretical terms. Lecture Notes in Computer Science, 1342:217– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foo:1997:IPS

[1712] N. Foo, P. Peppas, and Y. Zhang. Inductive properties of states. Lecture Notes in Computer Science, 1342:227– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jachowicz:1997:DPR

[1713] P. Jachowicz and R. Goebel. Describing plan recognition as nonmonotonic reasoning and belief revision. Lecture Notes in Computer Science, 1342:236– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacNish:1997:SMN

[1714] C. K. MacNish and G. Antoniou. Specification morphisms for nonmonotonic knowledge systems. Lecture Notes in Computer Science, 1342:246–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antoniou:1997:RRD

[1715] G. Antoniou. Representation results for default logics. Lecture Notes in Computer Science, 1342:255–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billington:1997:PQL

[1716] D. Billington. Proving quantified literals in defeasible logic. Lecture Notes in Computer Science, 1342:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arai:1997:FHP

[1717] Y. Arai and K. Hirota. Fuzzy hierarchical pattern recognition for robotics applications. Lecture Notes in Computer Science, 1342:274–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1997:ACB

[1718] A. R. Pearce, T. Caelli, and S. Goss.

Adaptive curvature-based topography for learning symbolic descriptions of terrain maps. Lecture Notes in Computer Science, 1342:282–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:CAI

[1719] R. Chang, L. Guan, and J. A. Burne.

A computer assisted image analysis system for diagnosing movement disorders. Lecture Notes in Computer Science, 1342:290–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yin:1997:VRS

[1720] B. Yin and T. Yu. Vowel recognition for speaker independent Chinese speech recognition. Lecture Notes in Computer Science, 1342:302–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dixon:1997:BIR

[1721] S. Dixon. Beat induction and rhythm recognition. Lecture Notes in Computer Science, 1342:311–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1997:BNN

[1722] H. He and Z. Huang. Boosting neural networks in real world applications: An empirical study. Lecture Notes in Computer Science, 1342:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamilton:1997:MLC

[1723] H. J. Hamilton, N. Shan, and W. Ziarko. Machine learning of credible classifications. Lecture Notes in Computer Science, 1342:330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:1997:MKM

[1724] G. J. Williams and Z. Huang. Mining the knowledge mine: The hot spots methodology for mining large real world databases. Lecture Notes in Computer Science, 1342:340–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiu:1997:UDT

[1725] B. C. Chiu, C. I. Webb, and Z. Zheng.

Using decision trees for agent modelling: A study on resolving confliction predictions. Lecture Notes in Computer Science, 1342:349–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raskutti:1997:SSA

[1726] B. Raskutti and A. Beitz. Sample set assessment for providing personalised recommendations. Lecture Notes in Computer Science, 1342:359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Min:1997:ICI

[1727] K. Min and W. H. Wilson. Integrated correction of ill-formed sentences. Lecture Notes in Computer Science, 1342:369–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hirashima:1997:IFC

[1728] T. Hirashima, N. Matsuda, T. Nomoto, and J. Toyoda. Information filtering for context-sensitive browsing. Lecture Notes in Computer Science, 1342:379– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sauvage:1997:STS

[1729] V. Sauvage. The T-SOM (Tree-SOM). Lecture Notes in Computer Science, 1342:389–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Treadgold:1997:EBC

[1730] N. K. Treadgold and T. D. Gedeon. Extending and benchmarking the CasPer algorithm. Lecture Notes in Computer Science, 1342:398–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thapa:1997:DFC

[1731] R. B. Thapa and R. Sadananda. Disconcepts and fuzzy cognitive maps. Lecture Notes in Computer Science, 1342:407–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peh:1997:WSB

[1732] L. Peh and C. P. Tsang. Weight sensitive Boolean extraction produces compact expressions. Lecture Notes in Computer Science, 1342:417–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briscoe:1997:LTS

[1733] G. Briscoe and T. Caelli. Learning temporal sequences in recurrent selforganising neural nets. Lecture Notes in Computer Science, 1342:427–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofman:1997:TDS

[1734] I. Hofman and R. Jarvis. Threedimensional scene analysis using multiple range finders — data capture, coordinate transformations and initial segmentation. Lecture Notes in Computer Science, 1342:436–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jarvis:1997:SBR

[1735] R. Jarvis. Sensor-based robotic automation of mushroom farming — preliminary considerations. Lecture Notes in Computer Science, 1342:446–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russell:1997:CAM

[1736] R. A. Russell. Creating adroit mechanisms by using large numbers of simple homogeneous structures. Lecture Notes in Computer Science, 1342:456– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russell:1997:NTS

[1737] R. A. Russell. A nephelometric tactile sensor. Lecture Notes in Computer Science, 1342:466–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OHagan:1997:FTR

[1738] R. O’Hagan and A. Zelinsky. Finger track — A robust and real-time gesture interface. Lecture Notes in Computer Science, 1342:475–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:CMK

[1739] C. Liu and M. A. Orgun. A constraint mechanism for knowledge specification of simulation systems based on temporal logic. Lecture Notes in Computer Science, 1342:485–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pisan:1997:CEP

[1740] Y. Pisan. Controlling engineering problem solving. Lecture Notes in Computer Science, 1342:496–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waugh:1997:EQR

[1741] S. Waugh, T. Menzies, and S. Goss. Evaluating a qualitative reasoner. Lecture Notes in Computer Science, 1342: 505–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fink:1997:RTS

[1742] S. J. Fink and S. B. Baden. Run-time support for multi-tier programming of block-structured applications on SMP clusters. Lecture Notes in Computer Science, 1343:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujinami:1997:ART

[1743] N. Fujinami. Automatic run-time code generation in C++. Lecture Notes in Computer Science, 1343:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bassetti:1997:CPE

[1744] F. Bassetti, K. Davis, and D. Quinlan. A comparison of performanceenhancing strategies for parallel numerical object-oriented frameworks. Lecture Notes in Computer Science, 1343: 17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weidmann:1997:DPI

[1745] M. Weidmann. Design and performance improvement of a real-world, object-oriented C++ solver with STL. Lecture Notes in Computer Science, 1343:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chien:1997:EHL

[1746] A. A. Chien, J. Dolby, B. Ganguly, and V. Karamcheti. Evaluating high level parallel programming support for irregular applications in ICC++. Lecture Notes in Computer Science, 1343:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaervi:1997:PSV

[1747] J. Jaervi. Processing sparse vectors during compile time in C++. Lecture Notes in Computer Science, 1343:41– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veldhuizen:1997:WCF

[1748] T. L. Veldhuizen and M. E. Jernigan. Will C++ be faster than Fortran? Lecture Notes in Computer Science, 1343:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Squyres:1997:DEM

[1749] J. M. Squyres, B. Saphir, and A. Lumsdaine. The design and evolution of the MPI-2 C++ interface. Lecture Notes in Computer Science, 1343:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quittek:1997:EES

[1750] J. W. Quittek and B. Weissman. Efficient extensible synchronization in sather. Lecture Notes in Computer Science, 1343:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joosen:1997:EOO

[1751] W. Joosen, B. Robben, H. Van Wulpen, and P. Verbaeten. Experiences with an object-oriented parallel language: The CORRELATE project. Lecture Notes in Computer Science, 1343:73– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuoka:1997:TPC

[1752] S. Matsuoka, A. Nikami, H. Ogawa, and Y. Ishikawa. Towards a parallel C++ programming language based on commodity object-oriented technologies. Lecture Notes in Computer Science, 1343:81–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takahashi:1997:CTM

[1753] T. Takahashi, Y. Ishikawa, M. Sato, and A. Yonezawa. A compiletime meta-level architecture supporting class specific optimization. Lecture Notes in Computer Science, 1343:89– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besch:1997:OOA

[1754] M. Besch, H. Bi, G. Heber, and M. Kessler. An object-oriented approach to the implementation of a highlevel data parallel language. Lecture Notes in Computer Science, 1343:97– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerlach:1997:FPA

[1755] J. Gerlach, M. Sato, and Y. Ishikawa.

A framework for parallel adaptive finite element methods and its template based implementation in CC++. Lecture Notes in Computer Science, 1343: 105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuda:1997:PAC

[1756] M. Matsuda, M. Sato, and Y. Ishikawa.

Parallel array class implementation using C++ STL adaptors. Lecture Notes in Computer Science, 1343:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loeffler:1997:MJF

[1757] G. Loeffler. A multithreaded Java framework for solving linear elliptic partial differential equations in 3D. Lecture Notes in Computer Science, 1343:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mintchev:1997:ABN

[1758] S. Mintchev and V. Getov. Automatic binding of native scientific libraries to Java. Lecture Notes in Computer Science, 1343:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imasaki:1997:JJP

[1759] K. Imasaki. JAPE: The Java parallel environment. Lecture Notes in Computer Science, 1343:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ang:1997:AJM

[1760] J.-P. Ang and Y.-T Tan. An architecture in Java for mobile computation. Lecture Notes in Computer Science, 1343:145–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ichisugi:1997:EJP

[1761] Y. Ichisugi and Y. Roudier. The extensible Java preprocessor kit and a tiny data-parallel Java. Lecture Notes in Computer Science, 1343:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruaset:1997:NSP

[1762] A. Magnus Bruaset, X. Cai, H. P. Langtangen, and A. Tveito. Numerical solution of PDEs on parallel computers utilizing sequential simulators. Lecture Notes in Computer Science, 1343:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calvin:1997:TUP

[1763] C. Calvin and P. Emonot. The TR/ O-unitaire project: A parallel CFD 3-Dimensional code. Lecture Notes in Computer Science, 1343:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1997:OOO

[1764] D. L. Brown, W. D. Henshaw, and D. J. Quinlan. Overture: An objectoriented framework for solving partial differential equations. Lecture Notes in Computer Science, 1343:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1997:ODP

[1765] W. Humphrey, S. Karmesin, F. Bassetti, and J. Reynders. Optimization of data-parallel field expressions in the POOMA framework. Lecture Notes in Computer Science, 1343:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:MTP

[1766] S. R. Lee, J. C. Cummings, S. D.

Nolen, and N. D. Keen. MC++ and a transport physics framework. Lecture Notes in Computer Science, 1343: 195–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCandless:1997:RAH

[1767] B. C. McCandless and A. Lumsdaine. The role of abstraction in high-performance computing. Lecture Notes in Computer Science, 1343:203– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1997:DDC

[1768] T. Ohta. Design of a data class for parallel scientific computing. Lecture Notes in Computer Science, 1343:211– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeng:1997:DOP

[1769] W.-M. Jeng and C. Lin. Describing objects in parallel ECEM image reconstruction. Lecture Notes in Computer Science, 1343:218–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henderson:1997:FPM

[1770] M. E. Henderson and S. E. Lyons. Flow in porous media using NAO finite difference classes. Lecture Notes in Computer Science, 1343:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bashford:1997:OOP

[1771] D. Bashford. An object-oriented programming suite for electrostatic effects in biological molecules. Lecture Notes in Computer Science, 1343:233– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeLuca:1997:POB

[1772] C. DeLuca, C. W. Heisey, R. A. Bond, and J. M. Daly. A portable, object-based parallel library and layered framework for real-time radar signal processing. Lecture Notes in Computer Science, 1343:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Irwin:1997:AOP

[1773] J. Irwin, J.-M. Loingtier, J. R. Gilbert, and G. Kiczales. Aspect-oriented programming of sparse matrix code. Lecture Notes in Computer Science, 1343: 249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddleton:1997:CSA

[1774] R. F. Haddleton and J. L. Pfaltz. Client/server architecture in the ADAMS parallel object-oriented database system. Lecture Notes in Computer Science, 1343:257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MacDonald:1997:PBO

[1775] S. MacDonald, J. Schaeffer, and D. Szafron. Pattern-based objectoriented parallel programming. Lecture Notes in Computer Science, 1343:267– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gray:1997:IEP

[1776] P. A. Gray and V. S. Sunderam. The IceT environment for parallel and distributed computing. Lecture Notes in Computer Science, 1343:275–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramamoorthi:1997:GRR

[1777] R. Ramamoorthi, A. Rifkin, B. Dimitrov, and K. M. Chandy. A general resource reservation framework for scientific computing. Lecture Notes in Computer Science, 1343:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rabin:1997:CPP

[1778] M. O. Rabin. Correctness of programs and protocols through randomization. Lecture Notes in Computer Science, 1345:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayache:1997:MIA

[1779] N. Ayache. Medical image analysis and simulation. Lecture Notes in Computer Science, 1345:4–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1997:VPM

[1780] R. E. Bryant and M. N. Velev. Verification of pipelined microprocessors by comparing memory execution sequences in symbolic simulation. Lecture Notes in Computer Science, 1345: 18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merz:1997:RA

[1781] S. Merz. Rules for abstraction. Lecture Notes in Computer Science, 1345: 32–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buessow:1997:CZT

[1782] R. Buessow and W. Grieskamp. Combining Z and temporal interval logics for the formalization of properties and behaviors of embedded systems. Lecture Notes in Computer Science, 1345:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramanujam:1997:RTC

[1783] R. Ramanujam. Rules for trace consistent reasoning. Lecture Notes in Computer Science, 1345:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Punnekkat:1997:SAR

[1784] S. Punnekkat, R. Davis, and A. Burns. Sensitivity analysis of real-time task sets. Lecture Notes in Computer Science, 1345:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tseng:1997:DMS

[1785] S.-M. Tseng, Y. H. Chin, and W.- P. Yang. Dynamic multiprocessor scheduling for supporting real-time constraints. Lecture Notes in Computer Science, 1345:83–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:HDA

[1786] Z. Huang, W.-J. Lei, C. Sun, and A. Sattar. Heuristic diff acquiring in lazy release consistency model. Lecture Notes in Computer Science, 1345: 98–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizutani:1997:RDC

[1787] T. Mizutani, S. Igarashi, K. Tomita, and M. Shio. Representation of discretely controlled continuous systems in software-oriented formal analysis. Lecture Notes in Computer Science, 1345:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wetprasit:1997:GFR

[1788] R. Wetprasit, A. Sattar, and L. Khatib.

A generalized framework for reasoning with multi-point events. Lecture Notes in Computer Science, 1345:121– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georget:1997:ICR

[1789] Y. Georget, P. Codognet, and F. Rossi.

Implementing constraint retraction for finite domains. Lecture Notes in Computer Science, 1345:136–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plagemann:1997:IIS

[1790] T. Plagemann and V. Goebel. IN-STANCE: The intermediate storage node concept. Lecture Notes in Computer Science, 1345:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:CHA

[1791] X. Li, Dang Van Hung, and Z. Tao. Checking hybrid automata for linear duration invariants. Lecture Notes in Computer Science, 1345:166–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikk:1997:HAM

[1792] E. Mikk, Y. Lakhnech, and M. Siegel. Hierarchical automata as model for statecharts. Lecture Notes in Computer Science, 1345:181–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harao:1997:PDL

[1793] M. Harao. Proof discovery in LK system by analogy. Lecture Notes in Computer Science, 1345:197–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Inuzuka:1997:EIE

[1794] N. Inuzuka, H. Seki, and H. Itoh. Efficient induction of executable logic programs from examples. Lecture Notes in Computer Science, 1345:212–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeCharlier:1997:AVB

[1795] B. Le Charlier, C. Leclere, S. Rossi, and A. Cortesi. Automated verification of behavioural properties of Prolog programs. Lecture Notes in Computer Science, 1345:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boudol:1997:TUR

[1796] G. Boudol. Typing the use of resources in a concurrent calculus. Lecture Notes in Computer Science, 1345: 239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roe:1997:ILR

[1797] P. Roe. An imperative language with read/write type modes. Lecture Notes in Computer Science, 1345:254–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohno:1997:EGS

[1798] K. Ohno, M. Ikawa, M. Goshima, and S.-I. Mori. Efficient goal scheduling in a concurrent logic language using type-based dependency analysis. Lecture Notes in Computer Science, 1345:268–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1997:ADO

[1799] R. Hasegawa. An analysis of divisibility orderings and recursive path orderings. Lecture Notes in Computer Science, 1345:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dinesh:1997:SWM

[1800] T. B. Dinesh and S. M. Ueskuedarh. Share-where maintenance in visual algebraic specifications. Lecture Notes in Computer Science, 1345:297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1997:FTB

[1801] S. Fujita. A fault tolerant broadcast scheme in star graphs. Lecture Notes in Computer Science, 1345:312–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fujita:1997:CCP

[1802] K.-E. Fujita. Calculus of classical proofs I. Lecture Notes in Computer Science, 1345:321–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1997:TEL

[1803] R. Watson and E. Salzman. Tracing the evaluation of lazy functional languages: A model and its implementation. Lecture Notes in Computer Science, 1345:336–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Asher:1997:BRA

[1804] Y. Ben-Asher and E. Stein. Basic results in automatic transformations of shared memory parallel programs into sequential programs. Lecture Notes in Computer Science, 1345:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaipainen:1997:ROS

[1805] M. Kaipainen, P. Papadopoulos, and P. Karhu. Recurrent oscillatory selforganizing map: Adapting to complex environmental periodicities. Lecture Notes in Computer Science, 1345:367– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lursinsap:1997:BBD

[1806] C. Lursinsap, K. Kanchanasut, and T. Siriboon. Basic binary decision diagram operations for image processing. Lecture Notes in Computer Science, 1345:368–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakajima:1997:AOS

[1807] T. Nakajima. Adaptive object storage system for mobile computing environments. Lecture Notes in Computer Science, 1345:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tomohiko:1997:SUI

[1808] M. Tomohiko. Structure of user interface module for practical Internet messages. Lecture Notes in Computer Science, 1345:373–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badonski:1997:SSU

[1809] M. Badonski, M. Ivanovic, and Z. Budimac. Software specification using LASS. Lecture Notes in Computer Science, 1345:375–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horita:1997:NTL

[1810] E. Horita and K. Mano. Nepi^{2}: A twolevel calculus for network programming based on the Π-calculus. Lecture Notes in Computer Science, 1345:377–378, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:SRR

[1811] M. Lin, J. Malec, and S. Nadjm-Tehrani. On semantics of reactive rule-based systems. Lecture Notes in Computer Science, 1345:379–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gagne:1997:NSS

[1812] J.-R. Gagne and J. Plaice. The nonstandard semantics of esterel. Lecture Notes in Computer Science, 1345:381– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1997:HSL

[1813] S. Kusakabe, K. Inenaga, and M. Amamiya. Hybrid support for lenient implementation of arraycomprehension. Lecture Notes in Computer Science, 1345:383–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouzoubaa:1997:SHC

[1814] M. Bouzoubaa. Solver for hierarchical CSP containing several constraint types and multi-output constraints. Lecture Notes in Computer Science, 1345:385–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramachandran:1997:QGP

[1815] Vijaya Ramachandran. QSM: A general purpose shared-memory model for parallel computation. Lecture Notes in Computer Science, 1346:1– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460001.
pdf.

Dey:1997:AGD

[1816] Tamal K. Dey, Arunabha Roy, and Nimish R. Shah. Approximating geometric domains through topological triangulations. Lecture Notes in Computer Science, 1346:6–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460006.
pdf.

Mahajan:1997:SSD

[1817] Sanjeev Mahajan, Edgar A. Ramos, and K. V. Subrahmanyam. Solving some discrepancy problems in NC. Lecture Notes in Computer Science, 1346: 22–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460022.
pdf.

Cirino:1997:GEB

[1818] K. Cirino, S. Muthukrishnan, N. S. Narayanaswamy, and H. Ramesh. Graph editing to bipartite interval graphs: Exact and asymptotic bounds. Lecture Notes in Computer Science, 1346:37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1346/13460037.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460037.
                                        pdf.

Clarke:1997:MC

[1819] Edmund M. Clarke. Model checking. Lecture Notes in Computer Science, 1346:54–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1346/13460054.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460054.
                                        pdf.

Kfoury:1997:RVI

[1820] A. J. Kfoury. Recursion versus iteration at higher-orders. Lecture Notes in Computer Science, 1346:57– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1346/13460057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460057.
                                        pdf.

Gordon:1997:CEI

[1821] A. D. Gordon, P. D. Hankin, and S. B. Lassen. Compilation and equivalence of imperative objects. Lecture Notes in Computer Science, 1346:74– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460074.
pdf.

Marchiori:1997:EPR

[1822] Massimo Marchiori. On the expressive power of rewriting. Lecture Notes in Computer Science, 1346:88– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460088.
pdf.

Kapur:1997:MVA

[1823] Deepak Kapur and M. Subramaniam.

Mechanizing verification of arithmetic circuits: SRT division. Lecture Notes in Computer Science, 1346:103–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460103.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460103.
pdf.

Pontelli:1997:CPI

[1824] E. Pontelli, D. Ranjan, and G. Gupta.

On the complexity of parallel implementation of logic programs (extended abstract). Lecture Notes in Computer Science, 1346:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1346/13460123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460123.
                                        pdf.

You:1997:ASD

[1825] Jia-Huai You, Li Yan Yuan, and Randy Goebel. An abductive semantics for disjunctive logic programs and its proof procedure. Lecture Notes in Computer Science, 1346:138– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1346/13460138.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460138.
                                        pdf.

Mohalik:1997:ACA

[1826] Swarup Mohalik and R. Ramanujam. Assumption-commitment in automata. Lecture Notes in Computer Science, 1346:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/1346/13460153.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460153.
                                        pdf.

Kulkarni:1997:CDM

[1827] Sandeep S. Kulkarni and Anish Arora. Compositional design of multitolerant repetitive Byzantine agreement. Lecture Notes in Computer Science, 1346:

169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460169.
pdf.

Sudan:1997:AIC

[1828] Madhu Sudan. Algorithmic issues in coding theory. Lecture Notes in Computer Science, 1346:184–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460184.
pdf.

Seth:1997:SRE

[1829] Anil Seth. Sharper results on the expressive power of generalized quantifiers. Lecture Notes in Computer Science, 1346:200–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460200.
pdf.

Vinodchandran:1997:ILR

[1830] N. V. Vinodchandran. Improved lowness results for solvable blackbox group problems. Lecture Notes in Computer Science, 1346:220–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1346/13460220.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460220.
                                        pdf.

Arvind:1997:RBM

[1831] V. Arvind and J. Köbler. On resourcebounded measure and pseudorandomness. Lecture Notes in Computer Science, 1346:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/1346/13460235.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        1346/13460235.pdf.

Vardi:1997:VOS

[1832] Moshe Y. Vardi. Verification of open systems. Lecture Notes in Computer Science, 1346:250–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/1346/13460250.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/1346/13460250.
                                        pdf.

Boer:1997:HSC

[1833] F. S. de Boer, U. Hannemann, and W.-P. de Roever. Hoare-style compositional proof systems for reactive shared variable concurrency. Lecture Notes in Computer Science, 1346:267– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/1346/13460267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460267.
pdf.

DeBoer:1997:HSC

[1834] F. S. De Boer, U. Hannemann, and W.-P. De Roever. Hoare-style compositional proof systems for reactive shared variable concurrency. Lecture Notes in Computer Science, 1346:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Namjoshi:1997:SCS

[1835] Kedar S. Namjoshi. A simple characterization of stuttering bisimulation. Lecture Notes in Computer Science, 1346:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1346/13460284.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1346/13460284.
pdf.

Devillers:1997:GRH

[1836] Raymond Devillers, Hanna Klaudel, and Robert-C. Riemann. General refinement for high level Petri nets. Lecture Notes in Computer Science, 1346: 297–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460297.
pdf.

Dufourd:1997:PTM

[1837] Catherine Dufourd and Alain Finkel. Polynomial-time many-one reductions for Petri nets. Lecture Notes in Computer Science, 1346:312–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460312.
pdf.

Graves:1997:CRP

[1838] Burkhard Graves. Computing reachability properties hidden in finite net unfoldings. Lecture Notes in Computer Science, 1346:327–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1346/13460327.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1346/13460327.
pdf.

Kong:1997:TPD

[1839] T. Yung Kong. Topology-preserving deletion of 1’s from 2-, 3- and 4dimensional binary images. Lecture Notes in Computer Science, 1347:3–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gritzmann:1997:RFL

[1840] P. Gritzmann. On the reconstruction of finite lattice sets from their X- rays. Lecture Notes in Computer Science, 1347:19–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lienhardt:1997:ATB

[1841] P. Lienhardt. Aspects in topologybased geometric modeling. Lecture Notes in Computer Science, 1347:33– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kovalevsky:1997:ADS

[1842] V. Kovalevsky. Applications of digital straight segments to economical image encoding. Lecture Notes in Computer Science, 1347:51–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAndrea:1997:MSG

[1843] G. D’Andrea and C. Fiorio. Maximal superpositions of grids and an application. Lecture Notes in Computer Science, 1347:63–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borgefors:1997:MRS

[1844] G. Borgefors, G. Sanniti di Baja, and S. Svensson. Multiresolution representation of shapes in binary images II: Volume images. Lecture Notes in Computer Science, 1347:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schramm:1997:CT

[1845] J. M. Schramm. Coplanar tricubes.

Lecture Notes in Computer Science, 1347:87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vitonne:1997:CTD

[1846] J. Vitonne and J. M. Chassery. Coexistence of tricubes in digital naive plane. Lecture Notes in Computer Science, 1347:99–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertrand:1997:SSP

[1847] G. Bertrand and M. Couprie. Some structural properties of discrete surfacess. Lecture Notes in Computer Science, 1347:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miguet:1997:LAC

[1848] S. Miguet, J. M. Nicod, and D. Sarrut. A linear algorithm for constructing the polygon adjacency relation in isosurfaces of 3D images. Lecture Notes in Computer Science, 1347:125–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayala:1997:DLF

[1849] R. Ayala, E. Dominguez, A. R. Frances, and A. Quintero. Digital lighting functions. Lecture Notes in Computer Science, 1347:139–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ptak:1997:DTR

[1850] P. Ptak, H. Kofler, and W. Kropatsch. Digital topologies revisited: An approach based on the topological pointneighbourhood. Lecture Notes in Computer Science, 1347:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imiya:1997:ECD

[1851] A. Imiya and U. Eckhardt. The Euler characteristic of discrete object. Lecture Notes in Computer Science, 1347: 163–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenoir:1997:FEM

[1852] A. Lenoir. Fast estimation of mean curvature on the surface of a 3D discrete object. Lecture Notes in Computer Science, 1347:175–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sladoje:1997:EET

[1853] N. Sladoje and J. Zunic. Ellipses estimation from their digitization. Lecture Notes in Computer Science, 1347: 187–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchand-Maillet:1997:TEO

[1854] S. Marchand-Maillet and Y. M.

Sharaiha. Topological errors and optimal chamfer distance coefficients. Lecture Notes in Computer Science, 1347: 199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malgouyres:1997:HDD

[1855] R. Malgouyres. Homotopy in 2dimensional digital images. Lecture Notes in Computer Science, 1347:213– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thollot:1997:SMF

[1856] J. Thollot. Set manipulations of fractal objects using matrices of IFS. Lecture Notes in Computer Science, 1347: 223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibanez:1997:RTO

[1857] L. Ibañez, C. Hamitouche, and C. Roux. Ray-tracing and 3D objects representation in the BCC and FCC grids. Lecture Notes in Computer Science, 1347:235–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andres:1997:SSL

[1858] E. Andres, P. Nehlig, and J. Francon. Supercover of straight lines, planes and triangles. Lecture Notes in Computer Science, 1347:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albers:1997:CDE

[1859] Patrick Albers and Malik Ghallab. Context dependent effects in temporal planning. Lecture Notes in Computer Science, 1348:1–12, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashish:1997:IGP

[1860] Naveen Ashish, Craig A. Knoblock, and Alon Y. Levy. Information gathering plans with sensing actions. Lecture Notes in Computer Science, 1348: 13–25, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aylett:1997:WDP

[1861] Ruth Aylett, Alex M. Coddington, David P. Barnes, and Robert A. Ghanea-Hercock. What does a planner need to know about execution? Lecture Notes in Computer Science, 1348: 26–38, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aylett:1997:PCP

[1862] Ruth Aylett, Gary J. Petley, Paul W. H. Chung, James K. Soutter, and Andrew G. Rushton. Planning and chemical plant operating procedure synthesis. Lecture Notes in Computer Science, 1348:39–51, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baioletti:1997:TPP

[1863] Marco Baioletti, Stefano Marcugini, and Alfredo Milani. Task planning and partial order planning: a domain transformation approach. Lecture Notes in Computer Science, 1348:52–63, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beetz:1997:ETS

[1864] Michael Beetz and Drew V. McDermott. Expressing transformations of structured reactive plans. Lecture Notes in Computer Science, 1348:64– 76, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beetz:1997:FPP

[1865] Michael Beetz and Drew V. McDermott. Fast probabilistic plan debugging. Lecture Notes in Computer Science, 1348:77–90, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burgess:1997:QGR

[1866] Andrew Burgess and Sam Steel. Quantification in generative refinement planning. Lecture Notes in Computer Science, 1348:91–103, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castilho:1997:MTR

[1867] Marcos A. Castilho, Olivier Gasquet, and Andreas Herzig. Modal tableaux for reasoning about actions and plans. Lecture Notes in Computer Science, 1348:104–116, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cesta:1997:TRP

[1868] Amedeo Cesta and Cristiano Stella. A time and resource problem for planning architectures. Lecture Notes in Computer Science, 1348:117–129, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cimatti:1997:PMC

[1869] A. Cimatti, E. Giunchiglia, F. Giunchiglia, and P. Traverso. Planning via model checking: a decision procedure for AR. Lecture Notes in Computer Science, 1348:130–142, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pereira:1997:PPR

[1870] Célia Da Costa Pereira, Frédérick Garcia, Jérôme Lang, and Roger Martin-Clouaire. Possibilistic planning: representation and complexity. Lecture Notes in Computer Science, 1348:143– 155, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacomo:1997:PSM

[1871] Giuseppe De Giacomo, Luca Iocchi, Daniele Nardi, and Riccardo Rosati. Planning with sensing for a mobile robot. Lecture Notes in Computer Science, 1348:156–168, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:1997:EPP

[1872] Yannis Dimopoulos, Bernhard Nebel, and Jana Koehler. Encoding planning problems in non-monotonic logic programs. Lecture Notes in Computer Science, 1348:169–181, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estlin:1997:AHH

[1873] Tara A. Estlin, Steve Chien, and Xuemei Wang. An argument for a hybrid HTN/ operator-based approach to planning. Lecture Notes in Computer Science, 1348:182–194, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1997:NHP

[1874] Maria Fox. Natural hierarchical planning using operator decomposition. Lecture Notes in Computer Science, 1348:195–207, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garcia-Martinez:1997:PLE

[1875] Ramón García-Martínez and Daniel Borrajo. Planning, learning, and executing in autonomous systems. Lecture Notes in Computer Science, 1348:208– 220, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gazen:1997:CEU

[1876] B. Cenk Gazen and Craig A. Knoblock. Combining the expressivity of UCPOP with the efficiency of Graphplan. Lecture Notes in Computer Science, 1348: 221–233, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Givan:1997:BPM

[1877] Robert Givan, Sonia M. Leach, and Thomas Dean. Bounded parameter Markov decision processes. Lecture Notes in Computer Science, 1348:234– 246, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hepplewhite:1997:PST

[1878] Richard T. Hepplewhite and Jeremy W. Baxter. Planning and search techniques for intelligent behaviour of battlefield entities. Lecture Notes in Computer Science, 1348:247–259, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kambhampati:1997:UEG

[1879] Subbarao Kambhampati, Eric Parker, and Eric Lambrecht. Understanding and extending Graphplan. Lecture Notes in Computer Science, 1348:260– 272, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehler:1997:EPG

[1880] Jana Koehler, Bernhard Nebel, Jörg Hoffmann, and Yannis Dimopoulos. Extending planning graphs to an ADL subset. Lecture Notes in Computer Science, 1348:273–285, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Massacci:1997:BSP

[1881] Fabio Massacci. Breaking security protocols as an AI planning problem. Lecture Notes in Computer Science, 1348:286–298, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miksch:1997:TOS

[1882] Silvia Miksch, Yuval Shahar, Werner Horn, Christian Popow, Franz Paky, and Peter Johnson. Time-oriented skeletal plans: support to design and execution. Lecture Notes in Computer Science, 1348:299–311, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muslea:1997:SLP

[1883] Ion Muslea. SINERGY: a linear planner based on genetic programming. Lecture Notes in Computer Science, 1348:312–324, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Munoz-Avila:1997:CSM

[1884] Hector Muñoz-Avila and Frank Weberskirch. A case study on the mergeability of cases with a partial-order planner. Lecture Notes in Computer Science, 1348:325–337, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nebel:1997:IIF

[1885] Bernhard Nebel, Yannis Dimopoulos, and Jana Koehler. Ignoring irrelevant facts and operators in plan generation. Lecture Notes in Computer Science, 1348:338–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oddi:1997:TSS

[1886] Angelo Oddi and Amedeo Cesta. A tabu search strategy to solve scheduling problems with deadlines and complex metric constraints. Lecture Notes in Computer Science, 1348:351–363, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Onder:1997:CSP

[1887] Nilufer Onder and Martha E. Pollack. Contingency selection in plan generation. Lecture Notes in Computer Science, 1348:364–376, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1997:APP

[1888] Marcus V. Santos and Marcio Rillo. Approaching the plans are programs paradigm using transaction logic. Lecture Notes in Computer Science, 1348: 377–389, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shanahan:1997:ECP

[1889] Murray Shanahan. Event calculus planning revisited. Lecture Notes in Computer Science, 1348:390–402, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spragg:1997:CBR

[1890] J. E. Spragg, G. Fozzard, and D. J.

Tyler. Constraint-based reactive rescheduling in a stochastic environment. Lecture Notes in Computer Science, 1348:403–413, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsuneto:1997:PRS

[1891] Reiko Tsuneto, Dana S. Nau, and James A. Hendler. Plan-refinement strategies and search-space size. Lecture Notes in Computer Science, 1348: 414–426, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:RUH

[1892] Xuemei Wang and Steve Chien. Replanning using hierarchical task network and operator-based planning. Lecture Notes in Computer Science, 1348:427–439, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Washington:1997:BPB

[1893] Richard Washington. BI-POMDP:

Bounded, incremental, partiallyobservable Markov-model planning. Lecture Notes in Computer Science, 1348:440–451, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lamine:1997:SCI

[1894] S. B. Lamine and J. Plaice. Software configuration with information systems. Lecture Notes in Computer Science, 1349:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertot:1997:HTS

[1895] Y. Bertot. Head-tactics simplification. Lecture Notes in Computer Science, 1349:16–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1997:IT

[1896] S. L. Bloom, Z. Esik, A. Labella, and E. G. Manes. Iteration 2-theories. Lecture Notes in Computer Science, 1349: 30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruns:1997:MCF

[1897] G. Bruns and I. Sutherland. Model checking and fault tolerance. Lecture Notes in Computer Science, 1349:45– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buth:1997:DAF

[1898] B. Buth, M. Kouvaras, J. Peleska, and H. Shi. Deadlock analysis for a fault-tolerant system. Lecture Notes in Computer Science, 1349:60–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cenciarelli:1997:SMT

[1899] P. Cenciarelli, A. Knapp, B. Reus, and M. Wirsing. From sequential to multithreaded Java: An event-based operational semantics. Lecture Notes in Computer Science, 1349:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioli:1997:PSP

[1900] M. Cerioli, A. Haxthausen, B. Krieg-Brueckner, and T. Mossakowski. Permissive subsorted partial logic in CASL. Lecture Notes in Computer Science, 1349:91–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antonio:1997:STC

[1901] C. Antonio and G. J. Milne. Specification of timing constraints within the circal process algebra. Lecture Notes in Computer Science, 1349:108–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:SVP

[1902] Xiao Jun Chen and F. Corradini. On the specification and verification of performance properties for a timed process algebra. Lecture Notes in Computer Science, 1349:123–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousot:1997:AIA

[1903] P. Cousot and R. Cousot. Abstract interpretation of algebraic polynomial systems. Lecture Notes in Computer Science, 1349:138–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroot:1997:MRM

[1904] M. De Groot and K. Robinson. Modular refinement and model building. Lecture Notes in Computer Science, 1349: 155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denker:1997:LTL

[1905] G. Denker, J. Ramos, C. Caleiro, and A. Sernadas. A linear temporal logic approach to objects with transactions. Lecture Notes in Computer Science, 1349:170–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Du:1997:SDS

[1906] X. Du, K. T. McDonnell, E. Nanos, and Y. S. Ramakrishna. Software design, specification, and verification: Lessons learned from the rether case study. Lecture Notes in Computer Science, 1349:185–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidge:1997:RRR

[1907] C. J. Fidge. Refinement rules for realtime multi-tasking programs. Lecture Notes in Computer Science, 1349:199– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

France:1997:ROO

[1908] R. B. France, J.-M. Bruel, M. M. Larrondo-Petrie, and E. Grant. Rigorous object oriented-modeling: Integrating formal and informal notations. Lecture Notes in Computer Science, 1349:216–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giacobazzi:1997:CAI

[1909] R. Giacobazzi and F. Ranzato. Completeness in abstract interpretation: A domain perspective. Lecture Notes in Computer Science, 1349:231–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1997:FPV

[1910] J. Harrison. Floating point verification in HOL light: The exponential function. Lecture Notes in Computer Science, 1349:246–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hooman:1997:VDR

[1911] J. Hooman. Verification of distributed real-time and fault-tolerant protocols. Lecture Notes in Computer Science, 1349:261–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1997:IBC

[1912] B. Jacobs. Invariants, bisimulations and the correctness of coalgebraic refinements. Lecture Notes in Computer Science, 1349:276–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Janowski:1997:BFM

[1913] T. Janowski. On bisimulation, fault-monotonicity and provable faulttolerance. Lecture Notes in Computer Science, 1349:292–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katis:1997:SGC

[1914] P. Katis, N. Sabadini, and R. F. C.

Walters. Span(graph): A categorical algebra of transition systems. Lecture Notes in Computer Science, 1349:307– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katis:1997:RPT

[1915] P. Katis, N. Sabadini, and R. F. C. Walters. Representing place/transition nets in span(graph). Lecture Notes in Computer Science, 1349:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesens:1997:IPB

[1916] D. Lesens. Invariants of parameterized binary tree networks as greatest fixpoints. Lecture Notes in Computer Science, 1349:337–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy:1997:MSC

[1917] N. Levy and J. Souquieres. Modelling specification construction by successive approximations. Lecture Notes in Computer Science, 1349:351–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Limet:1997:PVL

[1918] S. Limet and F. Saubion. On partial validation of logic programs. Lecture Notes in Computer Science, 1349:365– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopes:1997:PRS

[1919] A. Lopes and J. L. Fiadeiro. Preservation and reflection in specification. Lecture Notes in Computer Science, 1349: 380–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paige:1997:CSU

[1920] R. F. Paige. Case studies in using a meta-method for formal method integration. Lecture Notes in Computer Science, 1349:395–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrow:1997:UC

[1921] J. Parrow and B. Victor. The update calculus. Lecture Notes in Computer Science, 1349:409–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piessens:1997:SAE

[1922] F. Piessens and E. Steegmans. Selective attribute elimination for categorical data specifications. Lecture Notes in Computer Science, 1349:424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajan:1997:ASD

[1923] S. P. Rajan and M. Fujita. ATM switch design: Parametric high-level modeling and formal verification. Lecture Notes in Computer Science, 1349:437– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenegge:1997:HFQ

[1924] A. Schoenegge. The hidden function question revisited. Lecture Notes in Computer Science, 1349:451–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sernadas:1997:SLM

[1925] A. Sernadas, C. Sernadas, and C. Caleiro. Synchronization of logics with mixed rules: Completeness preservation. Lecture Notes in Computer Science, 1349:465–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankland:1997:SBF

[1926] C. Shankland and M. Thomas. Symbolic bisimulation for full LOTOS. Lecture Notes in Computer Science, 1349: 479–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simons:1997:ACR

[1927] M. Simons and M. Sintzoff. Algebraic composition and refinement of proofs. Lecture Notes in Computer Science, 1349:494–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telford:1997:ESF

[1928] A. Telford and D. Turner. Ensuring streams flow. Lecture Notes in Computer Science, 1349:509–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ulidowski:1997:EPL

[1929] I. Ulidowski and S. Yuen. Extending process languages with time. Lecture Notes in Computer Science, 1349:524– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:PAC

[1930] F. Wang and P.-A Hsiung. Parametric analysis of computer systems. Lecture Notes in Computer Science, 1349: 539–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almeida:1997:CPR

[1931] J. J. Almeida, L. S. Barbosa, F. L.

Neves, and J. N. Oliveira. CAMILA: Prototyping and refinement of constructive specifications. Lecture Notes in Computer Science, 1349:554–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buth:1997:P

[1932] B. Buth. Pamela+PVS. Lecture Notes in Computer Science, 1349:560– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerone:1997:CS

[1933] A. Cerone, A. J. Cowie, and G. J.

Milne. The circal system. Lecture Notes in Computer Science, 1349:563– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davies:1997:RTC

[1934] R. Davies. A refinement-type checker for standard ML. Lecture Notes in Computer Science, 1349:565–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grundy:1997:RHP

[1935] J. Grundy and T. Laangbacka. Recording HOL proofs in a structured browsable format. Lecture Notes in Computer Science, 1349:567–571, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Low:1997:AMA

[1936] C. K. Low. Analysing multi-agent system traces with IDAF. Lecture Notes in Computer Science, 1349:572–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ozols:1997:DTD

[1937] M. A. Ozols, K. A. Eastaughffe, and A. Cant. DOVE: A tool for design oriented verification and evaluation. Lecture Notes in Computer Science, 1349:574–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1997:BMB

[1938] K. Robinson. The B method and the B toolkit. Lecture Notes in Computer Science, 1349:576–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rus:1997:ALP

[1939] T. Rus, T. Halverson, E. Van Wyk, and R. Kooima. An algebraic language processing environment. Lecture Notes in Computer Science, 1349:581–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Traynor:1997:CDS

[1940] O. Traynor, D. Hazel, P. Kearney, and A. Martin. The cogito development system. Lecture Notes in Computer Science, 1349:586–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibaraki:1997:SNH

[1941] T. Ibaraki. Solving NP-hard combinatorial problems in the practical sense. Lecture Notes in Computer Science, 1350:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tajima:1997:ACS

[1942] A. Tajima and S. Misono. Airline crewscheduling problem with many irregular flights. Lecture Notes in Computer Science, 1350:2–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hidaka:1997:PAF

[1943] K. Hidaka and H. Okano. Practical approach to a facility location problem for large-scale logistics. Lecture Notes in Computer Science, 1350:12–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horie:1997:HIG

[1944] S. Horie and O. Watanabe. Hard instance generation for SAT. Lecture Notes in Computer Science, 1350:22– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kutylowski:1997:PTM

[1945] M. Kutylowski and R. Wanka. Playing tetris on meshes and multi-dimensional SHEARSORT. Lecture Notes in Computer Science, 1350:32–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsuura:1997:FAS

[1946] A. Matsuura and A. Nagoya. Formulation of the addition-shift-sequence problem and its complexity. Lecture Notes in Computer Science, 1350:42– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1997:WUS

[1947] T. Hayashi, K. Nakano, and S. Olariu. Weighted and unweighted selection algorithms for k sorted sequences. Lecture Notes in Computer Science, 1350: 52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1997:ADF

[1948] L. Bai, H. Ebara, H. Nakano, and H. Maeda. An adaptive distributed fault-tolerant routing algorithm for the star graph. Lecture Notes in Computer Science, 1350:62–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1997:MCR

[1949] Q.-P. Gu and H. Tamaki. Multicolor routing in the undirected hypercube. Lecture Notes in Computer Science, 1350:72–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yeh:1997:CSR

[1950] T.-H. Yeh, C.-M. Kuo, C.-L. Lei, and H.-C Yen. Competitive source routing on tori and meshes. Lecture Notes in Computer Science, 1350:82–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uno:1997:AEA

[1951] T. Uno. Algorithms for enumerating all perfect, maximum and maximal matchings in bipartite graphs. Lecture Notes in Computer Science, 1350:92– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishii:1997:AEV

[1952] T. Ishii, H. Nagamochi, and T. Ibaraki.

Augmenting edge and vertex connectivities simultaneously. Lecture Notes in Computer Science, 1350:102–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1997:TFH

[1953] T. Eiter, T. Ibaraki, and K. Makino.

Two-face Horn extensions. Lecture Notes in Computer Science, 1350:112– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ausiello:1997:DMR

[1954] G. Ausiello, P. G. Franciosa, D. Frigioni, and R. Giaccio. Decremental maintenance of reachability in hypergraphs and minimum models of Horn formulae. Lecture Notes in Computer Science, 1350:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leiserson:1997:AAM

[1955] C. E. Leiserson. Algorithmic analysis of multithreaded algorithms. Lecture Notes in Computer Science, 1350:132– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tamaki:1997:CPG

[1956] H. Tamaki and T. Tokuyama. A characterization of planar graphs by pseudo-line arrangements. Lecture Notes in Computer Science, 1350:133– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Panaite:1997:OFT

[1957] P. Panaite and A. Pelc. Optimal faulttolerant broadcasting in trees. Lecture Notes in Computer Science, 1350:143– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1997:TFH

[1958] T. Asano, K. Hori, T. Ono, and T. Hirata. A theoretical framework of hybrid approaches to MAX SAT. Lecture Notes in Computer Science, 1350: 153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dannoura:1997:LVL

[1959] E. Dannoura and K. Sakurai. Online versus off-line in money-making strategies with BROKERAGE. Lecture Notes in Computer Science, 1350:173– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1997:DMH

[1960] X. Deng and C. Papadimitriou. Decision-making by hierarchies of discordant agents. Lecture Notes in Computer Science, 1350:183–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1997:NEL

[1961] Khanh Quoc Nguyen, V. Varadharajan, and Y. Mu. A new efficient offline anonymous cash scheme. Lecture Notes in Computer Science, 1350:193– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:AUC

[1962] Z.-Z. Chen. Approximating unweighted connectivity problems in parallel. Lecture Notes in Computer Science, 1350: 202–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poon:1997:RLW

[1963] Chung Keung Poon and V. Ramachandran. A randomized linear work EREW PRAM algorithm to find a minimum spanning forest. Lecture Notes in Computer Science, 1350:212–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atallah:1997:EPA

[1964] M. J. Atallah, D. Z. Chen, and O. Daescu. Efficient parallel algorithms for planar st-graphs. Lecture Notes in Computer Science, 1350:223–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravikumar:1997:PSS

[1965] B. Ravikumar. Peg-solitaire, string rewriting systems and finite automata. Lecture Notes in Computer Science, 1350:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klauck:1997:SPF

[1966] H. Klauck. On the size of probabilistic formulae. Lecture Notes in Computer Science, 1350:243–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ryabko:1997:HCL

[1967] B. Ryabko and A. Fionov. Homophonic coding with logarithmic memory size. Lecture Notes in Computer Science, 1350:253–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohring:1997:CMA

[1968] R. H. Möhring and M. Müller-Hannemann. Complexity and modeling aspects of mesh refinement into quadrilaterals. Lecture Notes in Computer Science, 1350:263–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minakawa:1997:TOV

[1969] T. Minakawa and K. Sugihara. Topology oriented vs. exact arithmetic — experience in implementing the threedimensional convex hull algorithm. Lecture Notes in Computer Science, 1350:273–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ott:1997:CLB

[1970] M. Ott and F. Stephan. The complexity of learning branches and strategies from queries. Lecture Notes in Computer Science, 1350:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neyer:1997:SMS

[1971] G. Neyer and P. Widmayer. Singularities make spatial join scheduling hard. Lecture Notes in Computer Science, 1350:293–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:FOD

[1972] H.-F. S. Chen and D. T. Lee.

A faster one-dimensional topological compaction algorithm. Lecture Notes in Computer Science, 1350:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chan:1997:AFO

[1973] W.-T. Chan and F. Y. L. Chin. Algorithms for finding optimal disjoint paths around a rectangle. Lecture Notes in Computer Science, 1350:314– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakari:1997:AFR

[1974] Y. Kusakari and T. Nishizeki. An algorithm for finding a region with the minimum total L1 distance from prescribed terminals. Lecture Notes in Computer Science, 1350:324–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haxell:1997:DSB

[1975] P. E. Haxell and M. Loebl. On defect sets in bipartite graphs. Lecture Notes in Computer Science, 1350:334– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1997:DPD

[1976] M.-S. Chang, S.-Y. Hsieh, and G.- H Chen. Dynamic programming on distance-hereditary graphs. Lecture Notes in Computer Science, 1350:344– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicerone:1997:ECA

[1977] S. Cicerone and G. Di Stefano. On the equivalence in complexity among basic problems on bipartite and parity graphs. Lecture Notes in Computer Science, 1350:354–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kao:1997:ACM

[1978] M.-Y. Kao, Tak Wah Lam, Wing Kin Sung, and Hing Fung Ting. All-cavity maximum matchings. Lecture Notes in Computer Science, 1350:364–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rao:1997:FAC

[1979] S. V. Rao and A. Mukhopadhyay. Fast algorithms for computing β-skeletons and their relatives. Lecture Notes in Computer Science, 1350:374–383, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kyoda:1997:BCA

[1980] Y. Kyoda, K. Imai, F. Takeuchi, and A. Tajima. A branch-and-cut approach for minimum weight triangulation. Lecture Notes in Computer Science, 1350:384–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kellerer:1997:EAS

[1981] H. Kellerer, U. Pferschy, and M. G. Speranza. An efficient approximation scheme for the subset-sum problem. Lecture Notes in Computer Science, 1350:394–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pantziou:1997:CCC

[1982] G. E. Pantziou, G. P. Pentaris, and P. G. Spirakis. Competitive call control in mobile networks. Lecture Notes in Computer Science, 1350:404–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1997:GSP

[1983] J. Oommen and J. Dong. Generalized swap-with-parent schemes for selforganizing sequential linear lists. Lecture Notes in Computer Science, 1350: 414–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdallah:1997:ORO

[1984] S. M. Abdallah, E. M. Nebot, and D. C.

Rye. Object recognition and orientation via Zernike moments. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdellatif:1997:ICB

[1985] M. Abdellatif, Y. Tanaka, A. Gofuku, and I. Nagai. Illumination color from the blurred inter-reflection of a reference nose. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adan:1997:FMD

[1986] A. Adan, C. Cerrada, and V. Feliu. A fast mesh deformation method to build spherical representation models of 3D objects. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Akutsu:1997:AVC

[1987] T. Akutsu, K. Arakawa, and H. Murase. Active viewpoint control for shape from occluding contours. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandrov:1997:DAD

[1988] V. V. Alexandrov, E. V. Laikov, and B. E. Frenkel. Dynamic adaptive data structures for semantic analysis and synthesis of video information. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bab-Hadiashar:1997:RTL

[1989] A. Bab-Hadiashar and D. Suter. Robust total least squares based optic flow computation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bai:1997:ISC

[1990] X. Bai and G. Xu. Improved supervised color constancy for color inspection. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolle:1997:RIB

[1991] R. M. Bolle, N. K. Ratha, and S. Pankanti. Research issues in biometrics. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1997:FSM

[1992] M. S. Brown and W. B. Seales.

Fast stereo matching in compressed video. Lecture Notes in Computer Science, 1351:II–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burschka:1997:IRS

[1993] D. Burschka and S. A. Blum. Identification of 3D reference structures for video-based localization. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Califano:1997:HDI

[1994] A. Califano, B. Germain, and S. Colville. A high-dimensional indexing scheme for scalable fingerprintbased identification. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cha:1997:FIT

[1995] S.-H Cha. Fast image template and dictionary matching algorithms. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaen:1997:ATD

[1996] A. Chaen, K. Yamazawa, N. Yokoya, and H. Takemura. Acquisition of threedimensional information using omnidirectional stereo vision. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chai:1997:REG

[1997] J. Chai and S. Ma. Robust epipolar geometry estimation using genetic algorithm. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:DES

[1998] D. Chen, C.-M. Chang, and A. Laine. Detection and enhancement of small masses via precision multiscale analysis. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:NNS

[1999] Y. Chen, X. Ding, Y. Wu, and M. Chen. A new nonlinear shape normalization method for off-line handwritten Chinese character recognition. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chow:1997:PIF

[2000] S. K. Chow, M. Gillies, and S. L. Chan. Parallel implementation of fractal image compression using multiple digital signal processors. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chung:1997:RMH

[2001] J.-M. Chung and N. Ohnishi. Robust matching and hierarchical recognition of 2-D shapes using “chain of circles”. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collina:1997:STQ

[2002] C. Collina, M. Ferri, P. Frosini, and E. Porcellini. Sketch up: Towards qualitative shape data management. Lecture Notes in Computer Science, 1351: I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cumani:1997:ECE

[2003] A. Cumani. Efficient contour extraction in color images. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daum:1997:DUS

[2004] M. Daum and G. Dudek. Out of the dark: Using shadows to reconstruct 3D surfaces. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deriche:1997:FPL

[2005] R. Deriche, C. Bouvin, and O. Faugeras.

Front propagation and level-set approach for geodesic active stereovision. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandes:1997:EDT

[2006] J. L. Fernandes and J. R. A. Torreao.

Estimating depth through the fusion of photometric stereo images. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferri:1997:PSN

[2007] M. Ferri, P. Frosini, A. Lovato, and C. Zambelli. Point selection: A new comparison scheme for size functions (with an application to monogram recognition). Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funahashi:1997:GMD

[2008] T. Funahashi, T. Miyazaki, M. Yamada, and H. Seki. A generating method for 3-dimensional knitting cloth shapes. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimson:1997:IGS

[2009] E. Grimson. Image guided surgical systems. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gross:1997:DGM

[2010] A. Gross and L. Latecki. Digital geometric methods in image analysis and compression. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanmandlu:1997:CMT

[2011] M. Hanmandlu and V. Shantaram. Contour matching technique for 3D object recognition using Kalman filter. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanmandlu:1997:KFB

[2012] M. Hanmandlu and V. Shantaram.

Kalman filter based matching technique for 3D object recognition. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauta-Kasari:1997:UFM

[2013] M. Hauta-Kasari, W. Wang, S. Toyooka, and J. Parkkinen. Unsupervised filtering of munsell spectra. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1997:IFF

[2014] L. Hong and A. Jain. Integrating faces and fingerprints for personal identification. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howell:1997:LIB

[2015] A. J. Howell and H. Buxton. Learning identity and behaviour with neural networks. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hu:1997:RWC

[2016] W. Hu and W. B. Seales. Recognition in wavelet-compressed imagery. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1997:ACC

[2017] Q. Huang, Y. Cui, S. Samarasekera, and M. Greiffenhagen. Auto cameraman via collaborative sensing agents. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imagawa:1997:RTT

[2018] K. Imagawa, S. Lu, and S. Igi. Realtime tracking of human hands from a sign-language image sequence. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ipson:1997:NTP

[2019] S. S. Ipson, M. Melhi, and W. Booth. A novel triangulation procedure for thinning cursive text. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwahori:1997:SSC

[2020] Y. Iwahori, M. Murakami, R. J. Woodham, and N. Ishii. Sign of surface curvature from shading images using neural network. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jia:1997:LRO

[2021] J. Jia and K. Abe. Learning and recognizing 3D objects by using partial planar curve matching method. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katafuchi:1997:MIP

[2022] N. Katafuchi, M. Sano, S. Ohara, and M. Okudaira. A method of industrial parts surface inspection based on an optics model. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kato:1997:MCC

[2023] Z. Kato, T.-C. Pong, and J. C.- M. Lee. Motion compensated color video classification using Markov random fields. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaveti:1997:RPS

[2024] S. Kaveti, Eam Khwang Teoh, and H. Wang. Recognition of planar shapes using algebraic invariants from higher degree implicit polynomials. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:FCR

[2025] Hyoung Seop Kim, N. Motomura, and S. Ishikawa. Finding the center of rotational symmetry from noisy forms. Lecture Notes in Computer Science, 1351: I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:SMB

[2026] Jin Wook Kim, Eun Yi Kim, Se Hyun Park, and Hang Joon Kim. Segmentation of MRF based image using hierarchical genetic algorithm. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimmel:1997:IPB

[2027] R. Kimmel, R. Malladi, and N. Sochen. Image processing via the Beltrami operator. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimmel:1997:NNC

[2028] R. Kimmel. A natural norm for color processing. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnamoorthi:1997:CED

[2029] R. Krishnamoorthi and P. Bhattacharyya. Color edge detection using orthogonal polynomials. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lai:1997:QER

[2030] K. F. Lai, H. Zhou, and S. Chan. Query expansion by raw image features and text annotations in image retrieval. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lan:1997:PMR

[2031] Z.-D. Lan, R. Mohr, and L. Quan. Precise matching by robust estimation of deformation and local coherence. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lang:1997:ABM

[2032] J. Lang and M. R. M. Jenkin. Actively building models with VIRTUE. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1997:MID

[2033] Tak Kan Lau and I. King. Montage: An image database for the fashion, textile, and clothing industry in Hong Kong. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:DCA

[2034] F. Li and M. Brady. Dynamic calibration of an active vision system to compute the ground plane transformation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:EPS

[2035] S. Z. Li, Y. H. Huang, J. S. Fu, and K. L. Chan. Edge-preserving smoothing by convex minimization. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1997:SZM

[2036] S. X. Liao and M. Pawlak. A study of Zernike moment computing. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lien:1997:MBD

[2037] C.-C. Lien and C.-L. Huang. The model-based dynamic hand posture identification using genetic algorithm. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:EAR

[2038] X. Lin, X. Ding, Y. Chen, and J. Liu. Evaluation and application of recognition confidence in OCR. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maes:1997:OSS

[2039] S. H. Maes and H. S. M. Beigi. Open sesame! speech, password or key to secure your door? Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchand:1997:BNB

[2040] E. Marchand and F. Chaumette. A Bayes nets-based prediction/verification scheme for active visual reconstruction. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McIvor:1997:RSC

[2041] A. M. McIvor and P. T. Waltenberg. Recognition of simple curved surfaces from 3D surface data. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKenna:1997:FRS

[2042] S. J. McKenna and S. Gong. Face recognition from sequences using models of identity. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKenna:1997:OTU

[2043] S. J. McKenna, Y. Raja, and S. Gong.

Object tracking using adaptive color mixture models. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Menard:1997:SCS

[2044] C. Menard. Stereo correspondences in scale space. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukaigawa:1997:FSA

[2045] Y. Mukaigawa, Y. Nakamura, and Y. Ohta. Face synthesis with arbitrary pose and expression from several images — an integration of image-based and model-based approaches. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nalwa:1997:ALS

[2046] V. S. Nalwa. Automatic on-line signature verification. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nathan:1997:PCO

[2047] K. Nathan. Pen computing — an overview. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nozaki:1997:LCS

[2048] K. Nozaki, M. Niimi, R. O. Eason, and E. Kawaguchi. A large capacity steganography using color BMP images. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okatani:1997:CSP

[2049] T. Okatani and K. Deguchi. On the classification of singular points for the global shape from shading problem: A study of the constraints imposed by isophotes. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okatani:1997:DSG

[2050] T. Okatani and K. Deguchi. Determination of sign of Gaussian curvature of surface having general reflectance property. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oshiro:1997:FVS

[2051] N. Oshiro, A. Nishikawa, N. Maru, and F. Miyazaki. Foveated vision for scene exploration. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pagliari:1997:EMA

[2052] C. I. Pagliari and T. J. Dennis. Evolutionary methods applied to binocular disparity estimation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1997:CNA

[2053] In Kyu Park, Il Dong Yun, and Sang Uk Lee. A color normalization algorithm for image indexing. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Psarrou:1997:URN

[2054] A. Psarrou and J. Tanner. Using RBF networks to map GWT ridge images to pose. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qi:1997:CCL

[2055] Z. Qi and W. Gu. Combining camera and laser radar for ALV navigation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qi:1997:CMF

[2056] H. Qi, W. E. Snyder, and G. L. Bilbro. Comparison of mean field annealing and multiresolution analysis in missing data estimation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qi:1997:DTS

[2057] Z. Qi, Z. Duric, and R. S. Michalski. Detecting targets in SAR images: A machine learning approach. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raja:1997:STU

[2058] Y. Raja, S. J. McKenna, and S. Gong. Segmentation and tracking using color mixture models. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramakrishna:1997:EAS

[2059] R. S. Ramakrishna and B. Vaidyanathan. Error analysis in stereo vision. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rehrmann:1997:FRS

[2060] V. Rehrmann and L. Priese. Fast and robust segmentation of natural color scenes. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubner:1997:ACI

[2061] Y. Rubner, C. Tomasi, and L. J. Guibas. Adaptive color-image embeddings for database navigation. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuetz:1997:SAO

[2062] C. Schuetz, T. Jost, and H. Huegli. Semi-automatic 3D object digitizing system using range images. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Se:1997:SVB

[2063] S. Se and M. Brady. Stereo vision-based obstacle detection for partially sighted people. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shen:1997:AFP

[2064] W. Shen. Automated fingerprint pattern classification error analysis. Lecture Notes in Computer Science, 1351: I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shimada:1997:DPE

[2065] N. Shimada, Y. Shirai, Y. Kuno, and J. Miura. 3-D pose estimation and model refinement of an articulated object from a monocular image sequence. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasa:1997:LAF

[2066] N. Srinivasa and N. Ahuja. A learning approach to fixating on 3D targets with active cameras. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanaka:1997:LFE

[2067] H. T. Tanaka, A. Ishizawa, and H. Adachi. Live facial expression generation based on mixed reality. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tang:1997:ADT

[2068] C.-Y. Tang, Y.-P. Hung, and Z. Chen. Automatic detection and tracking of human heads using an active stereo vision system. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tian:1997:SCL

[2069] Y.-L. Tian, H. T. Tsui, and S. Y. Yeung. Spherical and cylindrical light source models for shape recovery. Lecture Notes in Computer Science, 1351: I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tian:1997:SRO

[2070] Y.-L. Tian, H. T. Tsui, and S. Y. Yeung. Shape recovery from one image under multiple light sources. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tominaga:1997:ERP

[2071] S. Tominaga. Estimation of reflection parameters from a color image. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tong:1997:DRV

[2072] W. B. Tong, S. K. Tso, S. Lang, and G. Z. Lu. Directing robots with visual primitives for navigation and micro-manipulation. Lecture Notes in Computer Science, 1351:I–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:UFI

[2073] Y.-F. Wang and R.-B. O. Alferez. A unified framework for image-derived invariants. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirtz:1997:TEB

[2074] B. Wirtz. Technical evaluation of biometric systems. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xie:1997:NDS

[2075] M. Xie. New development of stereo vision: A solution of motion stereo correspondence. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:PSR

[2076] J. Yang, D. Zhang, N. Ohnishi, and N. Sugie. Polyhedral shape recovery based on interreflections. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yow:1997:EHF

[2077] Kin Choong Yow and R. Cipolla. Enhancing human face detection using motion and active contours. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zha:1997:RFS

[2078] H. Zha, T. Hoshide, and T. Hasegawa. A recursive fitting-and-splitting algorithm for 3-D object modeling based on superquadrics. Lecture Notes in Computer Science, 1351:I–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hayashi:1997:TIH

[2079] J.-I. Hayashi, K. Kato, T. Endoh, and K. Murakami. On typical implementations of Hough transform for improving its performances. Lecture Notes in Computer Science, 1352:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:HSR

[2080] Y. Chen and S. Ma. Hierarchical segmentation and representation with dynamic link architecture neural network. Lecture Notes in Computer Science, 1352:9–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:PCS

[2081] N. Zhang and Wee Kheng Leow. Perceptually consistent segmentation of texture using multiple channel filter. Lecture Notes in Computer Science, 1352:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhuiyan:1997:OED

[2082] M. Shoaib Bhuiyan, Y. Iwahori, and A. Iwata. Optimal edge detection under difficult imaging conditions. Lecture Notes in Computer Science, 1352: 25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ratakonda:1997:RIQ

[2083] K. Ratakonda and N. Ahuja. Restoring image quality through structure preserving de-noising. Lecture Notes in Computer Science, 1352:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jenkinson:1997:FSN

[2084] M. Jenkinson and M. Brady. Feature saliency from noise variations in invariants. Lecture Notes in Computer Science, 1352:41–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1997:MIR

[2085] F. Chen and D. Surer. Multiscale image representation and edge detection. Lecture Notes in Computer Science, 1352:49–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fountain:1997:RIT

[2086] S. R. Fountain and T. N. Tan. Rotation invariant texture features from Gabor filters. Lecture Notes in Computer Science, 1352:57–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salden:1997:EIL

[2087] A. Salden. Euclidean invariants of linear scale-spaces. Lecture Notes in Computer Science, 1352:65–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mokhtarian:1997:SOM

[2088] F. Mokhtarian. Segmenting objects at multiple scales: A robust approach. Lecture Notes in Computer Science, 1352:73–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Qiu:1997:MGE

[2089] G. Qiu. Multi-grid edge models for magnifying digital images. Lecture Notes in Computer Science, 1352:81– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurita:1997:SRI

[2090] T. Kurita, K. Hotta, and T. Mishima. Scale and rotation invariant recognition method using higher-order local autocorrelation features of log-polar image. Lecture Notes in Computer Science, 1352:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peake:1997:SLI

[2091] G. S. Peake and T. N. Tan. Script and language identification from document images. Lecture Notes in Computer Science, 1352:97–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masai:1997:DCD

[2092] H. Masai and T. Watanabe. Document categorization for document image understanding. Lecture Notes in Computer Science, 1352:105–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yokokura:1997:RVB

[2093] N. Yokokura and T. Watanabe. Recognition of various bar-graph structures based on layout model. Lecture Notes in Computer Science, 1352:113–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1997:WCB

[2094] P.-K. Wong and C. Chan. Word-class bigram statistics language model for a hand-written Chinese character recognizer. Lecture Notes in Computer Science, 1352:121–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:LCS

[2095] X. Wang. Log classification by single X-ray scans using texture features from growth rings. Lecture Notes in Computer Science, 1352:129–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Takebe:1997:PFF

[2096] H. Takebe, Y. Katsuyama, and S. Naoi.

Precise and fast form identification method by using adaptive base lines for matching. Lecture Notes in Computer Science, 1352:137–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1997:CCC

[2097] J. Zheng, X. Ding, Y. Wu, and F. Guo. Combinatorial coarse classification method for OLCCR. Lecture Notes in Computer Science, 1352:145– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:DCG

[2098] Y. Liu, T. Yamamura, N. Ohnishi, and N. Sugie. Detecting characters in greyscale scene images. Lecture Notes in Computer Science, 1352:153–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugimoto:1997:CBI

[2099] A. Sugimoto. Conic based image transfer for 2-D objects: A linear algorithm. Lecture Notes in Computer Science, 1352:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heyden:1997:MCI

[2100] A. Heyden and K. Aastroem. Minimal conditions on intrinsic paramenters for Euclidean reconstruction. Lecture Notes in Computer Science, 1352:169– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Byne:1997:SBH

[2101] J. H. M. Byne and J. A. D. W. Anderson. Surface based hypothesis verification in intensity images using geometric and appearance data. Lecture Notes in Computer Science, 1352:177–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zha:1997:NBV

[2102] H. Zha, K. Morooka, and T. Masegawa. Next best viewpoint (NBV) planning for active object modeling based on a learning-by-showing approach. Lecture Notes in Computer Science, 1352:185– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lourens:1997:ORM

[2103] T. Lourens and R. P. Wuertz. Object recognition by matching symbolic edge graphs. Lecture Notes in Computer Science, 1352:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westling:1997:ICS

[2104] M. E. Westling and L. S. Davis. Interpretation of complex scenes using Bayesian networks. Lecture Notes in Computer Science, 1352:201–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:RUS

[2105] P. Liu, W. Wu, K. Ikeuchi, and M. Sakauchi. Recognition of urban scene using silhouette of buildings and city map database. Lecture Notes in Computer Science, 1352:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishijima:1997:CIM

[2106] M. Nishijima and T. Watanabe. A cooperative inference mechanism for extracting road information automatically. Lecture Notes in Computer Science, 1352:217–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:MBA

[2107] T. Liu, Kap Luk Chan, and Stan Ziqing Li. Model-based active object recognition using MRF matching and sensor planning. Lecture Notes in Computer Science, 1352:225–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seales:1997:IIC

[2108] W. B. Seales and Cheng Jiun Yuan. Improved image classification using morphing. Lecture Notes in Computer Science, 1352:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:RNM

[2109] C.-H. Kim and T.-J. Suh. Reconstruction of non-manifold objects from two orthographic views. Lecture Notes in Computer Science, 1352:241–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sumi:1997:ORU

[2110] Y. Sumi and F. Tomita. 3D object recognition using segment-based stereo vision. Lecture Notes in Computer Science, 1352:249–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Funt:1997:SCV

[2111] B. Funt. The state of color vision research. Lecture Notes in Computer Science, 1352:257–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tominaga:1997:CVC

[2112] S. Tominaga. Color vision and color media processing research in asia. Lecture Notes in Computer Science, 1352: 258–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noronha:1997:RAD

[2113] S. J. Noronha and R. Nevatia. Recent advances in detection and description of buildings from multiple aerial images. Lecture Notes in Computer Science, 1352:259–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Davis:1997:VSH

[2114] L. Davis, S. Fejes, D. Harwood, and Y. Yacoob. Visual surveillance of human activity. Lecture Notes in Computer Science, 1352:267–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aggarwal:1997:BPR

[2115] J. K. Aggarwal and S. Shah. Bayesian paradigm for recognition of objects — innovative applications. Lecture Notes in Computer Science, 1352:275– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bolle:1997:TMP

[2116] R. Bolle, Y. Aloimonos, and C. Fermueller. Toward motion picture grammars. Lecture Notes in Computer Science, 1352:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bajcsy:1997:HTS

[2117] P. Bajcsy and N. Ahuja. Hierarchical texture segmentation. Lecture Notes in Computer Science, 1352:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1997:RIS

[2118] X. Jiang and H. Bunke. Range image segmentation: Adaptive grouping of edges into regions. Lecture Notes in Computer Science, 1352:299–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mirmehdi:1997:OCI

[2119] M. Mirmehdi, P. L. Palmer, and J. Kittler. Optimising the complete image feature extraction chain. Lecture Notes in Computer Science, 1352:307– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1997:UFS

[2120] M.-S. Lee and G. Medioni. A unified framework for salient curves, regions, and junctions inference. Lecture Notes in Computer Science, 1352:315– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrin:1997:LMI

[2121] B. Perrin, N. Ahuja, and N. Srinivasa. Learning multiscale image models of 2D object classes. Lecture Notes in Computer Science, 1352:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minoh:1997:MCF

[2122] M. Minoh. 3D model centered framework for CV and VR. Lecture Notes in Computer Science, 1352:332–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:IBG

[2123] Z. Zhang. Image-based geometricallycorrect photorealistic scene/object modeling (IBPhM). Lecture Notes in Computer Science, 1352:340–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1997:MOS

[2124] Y. Sato, M. D. Wheeler, and K. Ikeuchi. Measuring object surface shape and reflectance properties. Lecture Notes in Computer Science, 1352: 350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berger:1997:RIC

[2125] M.-O. Berger and G. Simon. Robust image composition algorithms for augmented reality. Lecture Notes in Computer Science, 1352:360–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jo:1997:CBR

[2126] K.-H. Jo, Y. Kuno, and Y. Shirai.

Context-based recognition of manipulative hand gestures for human computer interaction. Lecture Notes in Computer Science, 1352:368–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trajkovic:1997:ARS

[2127] M. Trajkovic and M. Hedley. An algorithm for recursive structure and motion recovery under affine projection. Lecture Notes in Computer Science, 1352:376–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:RAD

[2128] Z.-Y. Zhang and H.-T. Tsui. Relative affine depth: Structure from motion by an uncalibrated camera. Lecture Notes in Computer Science, 1352: 384–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trajkovic:1997:EMR

[2129] M. Trajkovic and M. Hedley. The eigenspace method for rigid motion recovery from less than eight point correspondences. Lecture Notes in Computer Science, 1352:392–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:SMA

[2130] Y.-F. Wang and P. Liang. 3D shape and motion analysis from image blur and smear: A unified approach. Lecture Notes in Computer Science, 1352: 400–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1997:LES

[2131] P. Wang, K. Ikeuchi, and M. Sakauchi. 3D line’s extraction from 2D spatiotemporal image created by sine slit. Lecture Notes in Computer Science, 1352:408–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bottino:1997:TNI

[2132] A. Bottino, A. Laurentini, and P. Zuccone. Toward non-intrusive motion capture. Lecture Notes in Computer Science, 1352:416–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohba:1997:ABV

[2133] K. Ohba, Y. Sato, and K. Ikeuchi. Appearance based visual learning and object recognition with illumination invariance. Lecture Notes in Computer Science, 1352:424–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taniguchi:1997:EBS

[2134] Y. Taniguchi and Y. Shirai. Evidencebased scene interpretation considering subjective certainty of recognition. Lecture Notes in Computer Science, 1352:432–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jurie:1997:RHV

[2135] F. Jurie. Robust hypothesis verification for model based object recognition using Gaussian error model. Lecture Notes in Computer Science, 1352:440– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirihara:1997:SMM

[2136] S. Kirihara and H. Saito. Shape modeling from multiple view images using GAs. Lecture Notes in Computer Science, 1352:448–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jojic:1997:DRM

[2137] N. Jojic, J. Gu, H. C. Shen, and T. S.

Huang. 3-D reconstruction of multipart self-occluding objects. Lecture Notes in Computer Science, 1352:455–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jojic:1997:ACD

[2138] N. Jojic and T. S. Huang. On analysis of cloth drape range data. Lecture Notes in Computer Science, 1352: 463–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shinya:1997:VME

[2139] M. Shinya, T. Saito, T. Mori, and N. Osumi. VR models from epipolar images: An approach to minimize errors in synthesized images. Lecture Notes in Computer Science, 1352:471– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yonemoto:1997:SPP

[2140] S. Yonemoto, N. Tsuruta, and R.- I. Taniguchi. Shape and pose parameter estimation of 3D multi-part objects. Lecture Notes in Computer Science, 1352:479–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zheng:1997:GMO

[2141] Jiang Yu Zheng, A. Murata, and N. Abe. Generating 3D models of objects using multiple visual cues in image sequences. Lecture Notes in Computer Science, 1352:487–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mita:1997:STP

[2142] T. Mita, S. Hiura, H. Kato, and S. Inokuchi. Strategical tracking of polyhedral objects by reactive change of projection pattern — reactive range finder. Lecture Notes in Computer Science, 1352:495–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hwang:1997:AVG

[2143] W.-S. Hwang and J. Weng. Autonomous vision-guided robot manipulation control. Lecture Notes in Computer Science, 1352:503–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:NAA

[2144] Q. Zhang, W. Gu, and X. Ye. A new adaptive approach on rapid obstacle detection in range image. Lecture Notes in Computer Science, 1352:511– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uchimura:1997:RSM

[2145] K. Uchimura and Z. Hu. Recognition of shape model for general roads. Lecture Notes in Computer Science, 1352: 519–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lourakis:1997:VDO

[2146] M. I. A. Lourakis and S. C. Orphanoudakis. Visual detection of obstacles assuming a locally planar ground. Lecture Notes in Computer Science, 1352:527–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chuang:1997:PBM

[2147] J.-H. Chuang, C.-H. Tsai, W.-H. Tsai, and C.-Y. Yang. Potential-based modeling of 2D regions using non-uniform source distributions. Lecture Notes in Computer Science, 1352:535–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1997:LAM

[2148] G. Xu and N. Sugimoto. A linear algorithm for motion from three weak perspective images using Euler angles. Lecture Notes in Computer Science, 1352:543–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1997:LST

[2149] A. R. Pearce, T. Caelli, and S. Goss. On learning spatio-temporal relational structures in two different domains. Lecture Notes in Computer Science, 1352:551–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Or:1997:EIP

[2150] S. H. Or, W. S. Luk, K. H. Wong, and I. King. An efficient iterative pose estimation algorithm. Lecture Notes in Computer Science, 1352:559–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1997:NMA

[2151] Z. Zhang. A new multistage approach to motion and structure estimation by gradually enforcing geometric constraints. Lecture Notes in Computer Science, 1352:567–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ye:1997:TPP

[2152] Y. Ye, J. K. Tsotsos, K. Bennet, and E. Harley. Tracking a person with prerecorded image database and a pan, tilt, and zoom camera. Lecture Notes in Computer Science, 1352:575–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukai:1997:RMS

[2153] T. Mukai and N. Ohnishi. Recovery of motion and structure from optical flow under perspective projection by solving linear simultaneous equations. Lecture Notes in Computer Science, 1352:583–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quek:1997:VCM

[2154] F. K. H. Quek and R. K. Bryll. Vector coherence mapping: A parallelizable approach to image flow computation. Lecture Notes in Computer Science, 1352:591–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bab-Hadiashar:1997:RMS

[2155] A. Bab-Hadiashar and D. Suter. Robust motion segmentation using rank ordering estimators. Lecture Notes in Computer Science, 1352:599–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:OFS

[2156] Q. Yang and S. Ma. Optical flow in the scale space. Lecture Notes in Computer Science, 1352:607–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ngan:1997:MDT

[2157] P. M. Ngan. Motion detection in temporal clutter. Lecture Notes in Computer Science, 1352:615–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Booth:1997:NFT

[2158] W. Booth, J. M. Noras, and D. Xu.

A novel fast three-step search algorithm for block-matching motion estimation. Lecture Notes in Computer Science, 1352:623–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:1997:MVD

[2159] Y. Lu, J. Miller, and Tie Qi Chen. Moving vehicle detection and tracking in image sequences. Lecture Notes in Computer Science, 1352:631–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iwai:1997:GRI

[2160] Y. Iwai, T. Hata, and M. Yachida. Gesture recognition from image motion based on subspace method and HMM. Lecture Notes in Computer Science, 1352:639–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pramadihanto:1997:IFU

[2161] D. Pramadihanto, Y. Iwai, M. Yachida, and H. Wu. Identifying faces under varying pose using a single example view. Lecture Notes in Computer Science, 1352:647–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Utsumi:1997:MCB

[2162] A. Utsumi, H. Mori, J. Ohya, and M. Yachida. Multiple camera based human motion estimation. Lecture Notes in Computer Science, 1352:655– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murakami:1997:AFC

[2163] K. Murakami, M. Takai, and H. Koshimizu. An autonomous facial caricaturing based on a model of visual illusionexperimental modeling of visual illusion. Lecture Notes in Computer Science, 1352:663–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishikawa:1997:EFM

[2164] T. Ishikawa, H. Sera, S. Morishima, and D. Terzopoulos. 3D estimation of facial muscle parameter from the 2D marker movement using neural network. Lecture Notes in Computer Science, 1352:671–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gong:1997:ABF

[2165] S. Gong, E.-J. Ong, and P. J. Loft.

Appearance-based face recognition under large head rotations in depth. Lecture Notes in Computer Science, 1352: 679–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1997:SCM

[2166] J. Yang, W. Lu, and A. Waibel. Skincolor modeling and adaptation. Lecture Notes in Computer Science, 1352: 687–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ariki:1997:HIR

[2167] Y. Ariki, N. Ishikawa, and Y. Sugiyama.

Human information retrieval by face extraction and recognition on TV news images using subspace method. Lecture Notes in Computer Science, 1352: 695–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otsuka:1997:CFE

[2168] T. Otsuka and J. Ohya. Converting facial expressions using recognitionbased analysis of image sequences. Lecture Notes in Computer Science, 1352: 703–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohta:1997:MBF

[2169] H. Ohta, H. Saji, and H. Nakatani. Muscle-based feature models for analyzing facial expressions. Lecture Notes in Computer Science, 1352:711–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:1997:MMM

[2170] Y. Li, S. Ma, and H. Lu. A morphological method for moving object segmentation and posture recognition. Lecture Notes in Computer Science, 1352:719–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jiang:1997:DGF

[2171] X. Jiang, M. Binkert, B. Achermann, and H. Bunke. Detection of glasses in facial images. Lecture Notes in Computer Science, 1352:726–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishimura:1997:NMC

[2172] T. Nishimura, T. Mukai, and R. Oka. Non-monotonic continuous dynamic programming for spotting recognition of hesitated gestures from time-varying images. Lecture Notes in Computer Science, 1352:734–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liao:1997:FRU

[2173] H.-Y. M. Liao, C.-C. Han, G.-J. Yu, and H.-R. Tyan. Face recognition using a face-only database: A new approach. Lecture Notes in Computer Science, 1352:742–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenhart:1997:DFM

[2174] W. Lenhart and G. Liotta. Drawable and forbidden minimum weight triangulations. Lecture Notes in Computer Science, 1353:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juenger:1997:PAM

[2175] M. Juenger, E. K. Lee, P. Mutzel, and T. Odenthal. A polyhedral approach to the multi-layer crossing minimization problem. Lecture Notes in Computer Science, 1353:13–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1997:EOP

[2176] P. Bose. On embedding an outer-planar graph in a point set. Lecture Notes in Computer Science, 1353:25–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1997:BCN

[2177] F. Shahrokhi, O. Sykora, L. A. Szekely, and I. Vrt’o. Bipartite crossing numbers of meshes and hypercubes. Lecture Notes in Computer Science, 1353: 37–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pach:1997:TDG

[2178] J. Pach, T. Thiele, and G. Toth. Three-dimensional grid drawings of graphs. Lecture Notes in Computer Science, 1353:47–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1997:IOG

[2179] A. Papakostas and I. G. Tollis. Incremental orthogonal graph drawing in three dimensions. Lecture Notes in Computer Science, 1353:52–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calamoneri:1997:TDL

[2180] T. Calamoneri and A. Massini. On three-dimensional layout of interconnection networks. Lecture Notes in Computer Science, 1353:64–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1997:ODG

[2181] T. Biedl, T. Shermer, S. Whitesides, and S. Wismath. Orthogonal 3-D graph drawing. Lecture Notes in Computer Science, 1353:76–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:FBV

[2182] P. Eades, M. E. Houle, and R. Webber. Finding the best viewpoints for three-dimensional graph drawings. Lecture Notes in Computer Science, 1353: 87–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rahman:1997:LAO

[2183] M. Saidur Rahman, S.-I. Nakano, and T. Nishizeki. A linear algorithm for optimal orthogonal drawings of triconnected cubic plane graphs. Lecture Notes in Computer Science, 1353:99– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:IOG

[2184] U. Foessmeier. Interactive orthogonal graph drawing: Algorithms and bounds. Lecture Notes in Computer Science, 1353:111–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garrido:1997:EGG

[2185] M. A. Garrido and A. Marquez. Embedding a graph in the grid of a surface with the minimum number of bends is NP-hard. Lecture Notes in Computer Science, 1353:124–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1997:AAB

[2186] U. Foessmeier and M. Kaufmann. Algorithms and area bounds for nonplanar orthogonal drawings. Lecture Notes in Computer Science, 1353:134–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:DCG

[2187] P. Eades and Q.-W Feng. Drawing clustered graphs on an orthogonal grid. Lecture Notes in Computer Science, 1353:146–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1997:GCC

[2188] F. J. Brandenburg. Graph clustering I: Circles of cliques. Lecture Notes in Computer Science, 1353:158–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kakoulis:1997:ALE

[2189] K. C. Kakoulis and I. G. Tollis. An algorithm for labeling edges of hierarchical drawings. Lecture Notes in Computer Science, 1353:169–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Iturriaga:1997:ELT

[2190] C. Iturriaga and A. Lubiw. Elastic labels: the two-axis case. Lecture Notes in Computer Science, 1353:181– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juenger:1997:PUP

[2191] M. Juenger, S. Leipert, and P. Mutzel. Pitfalls of using PQ-Trees in automatic graph drawing. Lecture Notes in Computer Science, 1353:193–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valtr:1997:GDN

[2192] P. Valtr. Graph drawings with no k pairwise crossing edges. Lecture Notes in Computer Science, 1353:205– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:ARD

[2193] X. Lin and P. Eades. Area requirements for drawing hierarchically planar graphs. Lecture Notes in Computer Science, 1353:219–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFraysseix:1997:SPG

[2194] H. De Fraysseix and P. Ossona de Mendez. A short proof of a Gauss problem. Lecture Notes in Computer Science, 1353:230–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandes:1997:BPD

[2195] U. Brandes and D. Wagner. A Bayesian paradigm for dynamic graph layout. Lecture Notes in Computer Science, 1353:236–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Purchase:1997:WAG

[2196] H. Purchase. Which aesthetic has the greatest effect on human understanding? Lecture Notes in Computer Science, 1353:248–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dobkin:1997:IGP

[2197] D. P. Dobkin, E. R. Gansner, E. Koutsofios, and S. C. North. Implementing a general-purpose edge router. Lecture Notes in Computer Science, 1353: 262–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fekete:1997:WLE

[2198] S. P. Fekete, M. E. Houle, and S. Whitesides. The wobbly logic engine: Proving hardness of non-rigid geometric graph representation problems. Lecture Notes in Computer Science, 1353:272–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patrignani:1997:TTD

[2199] M. Patrignani and F. Vargiu. 3DCube: A tool for three dimensional graph drawing. Lecture Notes in Computer Science, 1353:284–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roxborough:1997:GCU

[2200] T. Roxborough and A. Sen. Graph clustering using multiway ratio cut. Lecture Notes in Computer Science, 1353:291–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hundack:1997:AGD

[2201] C. Hundack, P. Mutzel, I. Pouchkarev, and S. Thome. ArchE: A graph drawing system for archaeology. Lecture Notes in Computer Science, 1353:297– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bridgeman:1997:IAI

[2202] S. S. Bridgeman, J. Fanto, A. Garg, and R. Tamassia. InteractiveGiotto: An algorithm for interactive orthogonal graph drawing. Lecture Notes in Computer Science, 1353:303–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Didimo:1997:GIT

[2203] W. Didimo and A. Leonforte. GRID: An interactive tool for computing orthogonal drawings with the minimum number of bends. Lecture Notes in Computer Science, 1353:309–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tunkelang:1997:LNU

[2204] D. Tunkelang, R. J. Byrd, and J. W.

Cooper. Lexical navigation: Using incremental graph drawing for query refinement. Lecture Notes in Computer Science, 1353:316–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andalman:1997:DGB

[2205] B. Andalman, K. Ryall, W. Ruml, and J. Marks. Design gallery browsers based on 2D and 3D graph drawing. Lecture Notes in Computer Science, 1353:322–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:OAG

[2206] P. Eades, R. F. Cohen, and Mao Lin Huang. Online animated graph drawing for Web navigation. Lecture Notes in Computer Science, 1353:330–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barghouti:1997:GGP

[2207] N. S. Barghouti, J. Mocenigo, and W. Lee. Grappa: A GRAPh PAckage in Java. Lecture Notes in Computer Science, 1353:336–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lauer:1997:GSD

[2208] H. Lauer, M. Ettrich, and K. Soukup.

GraVis — system demonstration. Lecture Notes in Computer Science, 1353: 344–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hlineny:1997:TGU

[2209] P. Hlineny. Touching graphs of unit balls. Lecture Notes in Computer Science, 1353:350–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Czyzowicz:1997:DRC

[2210] J. Czyzowicz, E. Kranakis, D. Krizanc, and J. Urrutia. Discrete realizations of contact and intersection graphs. Lecture Notes in Computer Science, 1353: 359–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1997:MAD

[2211] P. Crescenzi and P. Penna. Minimumarea h-v drawings of complete binary trees. Lecture Notes in Computer Science, 1353:371–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biedl:1997:TPM

[2212] T. C. Biedl, B. P. Madden, and I. G. Tollis. The three-phase method: A unified approach to orthogonal graph drawing. Lecture Notes in Computer Science, 1353:391–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wills:1997:NIV

[2213] G. J. Wills. NicheWorks — interactive visualization of very large graphs. Lecture Notes in Computer Science, 1353:403–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seemann:1997:ESA

[2214] J. Seemann. Extending the sugiyama algorithm for drawing UML class diagrams: Towards automatic layout of object-oriented software diagrams. Lecture Notes in Computer Science, 1353: 415–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berry:1997:GDM

[2215] J. Berry, N. Dean, M. Goldberg, and G. Shannon. Graph drawing and manipulation with LINK. Lecture Notes in Computer Science, 1353:425–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1997:GDC

[2216] P. Eades, J. Marks, and S. North.

Graph-drawing contest report. Lecture Notes in Computer Science, 1353:438– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkart:1997:I

[2217] Olaf Burkart. 1. introduction. Lecture Notes in Computer Science, 1354: 1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1354/13540001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1354/13540001.
pdf.

Burkart:1997:B

[2218] Olaf Burkart. 2. background. Lecture Notes in Computer Science, 1354:9– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1354/13540009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1354/13540009.
pdf.

Burkart:1997:PP

[2219] Olaf Burkart. 3. pushdown processes. Lecture Notes in Computer Science, 1354:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1354/13540033.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1354/13540033.
pdf.

Burkart:1997:MC

[2220] Olaf Burkart. 4. model checking. Lecture Notes in Computer Science, 1354: 67–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1354/13540067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1354/13540067.
pdf.

Burkart:1997:EC

[2221] Olaf Burkart. 5. equivalence checking. Lecture Notes in Computer Science, 1354:115–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/1354/13540115.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/1354/13540115.
pdf.

Burkart:1997:SP

[2222] Olaf Burkart. 6. summary and perspectives. Lecture Notes in Computer Science, 1354:151–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1354/13540151.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1354/13540151.pdf.

Al-Dabbagh:1997:TAR

[2223] A. Al-Dabbagh and M. Darnell. The theory and application of reciprocal pairs of periodic sequences. Lecture Notes in Computer Science, 1355:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AlJabri:1997:ZEC

[2224] A. K. Al Jabri and S. Al-Issa. Zeroerror codes for correlated information sources. Lecture Notes in Computer Science, 1355:17–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benachour:1997:TDT

[2225] P. Benachour, G. Markarian, and B. Honary. Trellis decoding techniques and their performance in the adder channel for synchronous and asynchronous CCMA codes. Lecture Notes in Computer Science, 1355:23– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blake-Wilson:1997:KAP

[2226] S. Blake-Wilson, D. Johnson, and A. Menezes. Key agreement protocols and their security analysis. Lecture Notes in Computer Science, 1355: 30–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bond:1997:LDP

[2227] J. Bond, S. Hui, and H. Schmidt. Low density parity check codes based on sparse matrices with no small cycles. Lecture Notes in Computer Science, 1355:46–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brincat:1997:SC

[2228] K. Brincat and A. Meijer. On the SAFER cryptosystem. Lecture Notes in Computer Science, 1355:59–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1997:AAU

[2229] K. L. Brown and M. Darnell. An adaptive approach to T of N user multiaccess communications channel using an orthogonal coded multiplexer. Lecture Notes in Computer Science, 1355: 69–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1997:BLC

[2230] F. L. Carter. The breaking of the Lorenz cipher: An introduction to the theory behind the operational role of “Colossus” at BP. Lecture Notes in Computer Science, 1355:74–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cocks:1997:SKG

[2231] C. Cocks. Split knowledge generation of RSA parameters. Lecture Notes in Computer Science, 1355:89–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fan:1997:AEC

[2232] P. Fan. Analysis of error control in digital trunked radio systems. Lecture Notes in Computer Science, 1355:96– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filiol:1997:RCE

[2233] E. Filiol. Reconstruction of convolutional encoders over GF(q). Lecture Notes in Computer Science, 1355:101– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harari:1997:HHF

[2234] S. Harari. HCC: A hash function using error correcting codes. Lecture Notes in Computer Science, 1355:110–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hess:1997:PKC

[2235] E. Hess. Public-key cryptosystems based on elliptic curves — an evolutionary approach. Lecture Notes in Computer Science, 1355:118–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honary:1997:NAT

[2236] B. Honary, B. Thomas, P. Coulton, and M. Darnell. Novel application of turbo decoding for radio channels. Lecture Notes in Computer Science, 1355:119–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Howgrave-Graham:1997:FSR

[2237] N. Howgrave-Graham. Finding small roots of univariate modular equations revisited. Lecture Notes in Computer Science, 1355:131–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Husni:1997:RRS

[2238] E. Husni and P. Sweeney. Robust reed solomon coded MPSK modulation. Lecture Notes in Computer Science, 1355:143–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joye:1997:RTS

[2239] M. Joye, J. J. Quisquater, F. Bao, and R. H. Deng. RSA-type signatures in the presence of transient faults. Lecture Notes in Computer Science, 1355:155–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kabatianskii:1997:DSS

[2240] G. Kabatianskii, E. Krouk, and B. Smeets. A digital signature scheme based on random error-correcting codes. Lecture Notes in Computer Science, 1355:161–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1997:VRAa

[2241] V. K. N. Lau and M. D. Macleod. Variable rate adaptive trellis coded QAM for high bandwidth efficiency applications under Rayleigh fading channel. Lecture Notes in Computer Science, 1355:168–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lau:1997:VRAb

[2242] V. K. N. Lau and S. V. Maric. Variable rate adaptive channel coding for coherent and non-coherent Rayleigh fading channel. Lecture Notes in Computer Science, 1355:180–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1997:LNU

[2243] B. Lin and S. Shepherd. LABYRINTH:

a new ultra high speed stream cipher. Lecture Notes in Computer Science, 1355:192–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1997:RBH

[2244] X. Liu, P. G. Farrell, and C. A.

Boyd. Resisting the Bergen-Hogan attack on adaptive arithmetic coding. Lecture Notes in Computer Science, 1355:199–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Markarian:1997:NDT

[2245] G. Markarian, P. Benachour, and B. Honary. Novel decoding technique for the synchronous and quasisynchronous multiple access adder channel. Lecture Notes in Computer Science, 1355:209–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1997:IEI

[2246] K. M. Martin. Increasing efficiency of international key escrow in mutually mistrusting domains. Lecture Notes in Computer Science, 1355:221–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maucher:1997:MDC

[2247] J. Maucher. Multi dimensional compartment schemes. Lecture Notes in Computer Science, 1355:233–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McManus:1997:ESA

[2248] S. McManus and P. G. Farrell. Evaluation of standard approximation to loglikelihood ratio addition in the MAP algorithm, and its application in block code (turbo) iterative decoding algorithms. Lecture Notes in Computer Science, 1355:239–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mow:1997:MCB

[2249] W. H. Mow. Multiuser coding based on detecting matrices for synchronous-CDMA systems. Lecture Notes in Computer Science, 1355:251–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ODonoghue:1997:ECC

[2250] C. P. O’Donoghue and C. J. Burkley. Enumeration of convolutional codes and minimal encoders. Lecture Notes in Computer Science, 1355:258–??, 1997.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pinch:1997:UCN

[2251] R. G. E. Pinch. On using Carmichael numbers for public key encryption systems. Lecture Notes in Computer Science, 1355:265–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Preneel:1997:HFM

[2252] B. Preneel. Hash functions and MAC algorithms based on block ciphers. Lecture Notes in Computer Science, 1355: 270–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radu:1997:WHR

[2253] C. Radu, R. Govaerts, and J. Vandewalle. Witness hiding restrictive blind signature scheme. Lecture Notes in Computer Science, 1355:283–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schneider:1997:NCU

[2254] M. Schneider. A note on the construction and upper bounds of correlationimmune functions. Lecture Notes in Computer Science, 1355:295–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sonander:1997:GCC

[2255] S. Sonander and B. Honary. On generalised concatenated codes. Lecture Notes in Computer Science, 1355:307– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tandon:1997:EPA

[2256] A. Tandon and P. G. Farrell. Error performance analysis of different interleaving strategies applied to eight track digital tape systems. Lecture Notes in Computer Science, 1355:317–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zuquete:1997:EEP

[2257] A. Zuquete and P. Guedes. Efficient error-propagating block chaining. Lecture Notes in Computer Science, 1355:323–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keramane:1997:OBC

[2258] C. Keramane and A. Duda. Operator based composition of structured multimedia presentations. Lecture Notes in Computer Science, 1356:1–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tassel:1997:EEP

[2259] J. Tassel, B. Briscoe, and A. Smith. An end to end price-based QoS control component using reflective Java. Lecture Notes in Computer Science, 1356: 18–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tschammer:1997:SSE

[2260] V. Tschammer, M. Mendes, V. Ouzounis, and M. Tschichholz. SECCO — support environment for electronic commerce. Lecture Notes in Computer Science, 1356:33–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1997:BHC

[2261] D. J. Kim, H. Namgoong, and Y.- C. Lew. A bridge for heterogeneous communication between CORBA and DCE. Lecture Notes in Computer Science, 1356:45–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olivir:1997:DCM

[2262] H. Olivir, M. Banfield, and D. Hutchison. Distributed connection management for real-time multimedia services. Lecture Notes in Computer Science, 1356:59–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunkelmann:1997:EDV

[2263] T. Kunkelmann, T. Blecher, R. Reinema, and R. Steinmetz. Evaluation of different video encryption methods for a secure multimedia conferencing gateway. Lecture Notes in Computer Science, 1356:75–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rizzo:1997:FAD

[2264] L. Rizzo. The FreeBSD audio driver.

Lecture Notes in Computer Science, 1356:90–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mangues-Bafalluy:1997:CDM

[2265] J. Mangues-Bafalluy and J. Domingo-Pascual. Cell discarding mechanisms with minimum flow quality. Lecture Notes in Computer Science, 1356:105– ??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruppel:1997:IDS

[2266] W. Ruppel. Implementing the digital storage media — command and control (DSM-CC) standard in the framework of an on-line service. Lecture Notes in Computer Science, 1356:122–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfeifer:1997:RSH

[2267] T. Pfeifer, S. Arbanowski, and R. Popescu-Zeletin. Resource selection in heterogeneous communication environments using the teleservice descriptor. Lecture Notes in Computer Science, 1356:132–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wittmann:1997:AAM

[2268] R. Wittmann and M. Zitterbart. AMnet: Active multicasting network. Lecture Notes in Computer Science, 1356: 154–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1997:IVG

[2269] M. Hofmann and M. Rohrmueller. Impact of virtual group structure on multicast performance. Lecture Notes in Computer Science, 1356:165–??, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kompe:1997:PSU

[2270] Ralf Kompe. Prosody in speech understanding systems, volume 1307 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63580-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xix + 357 pp. LCCN TK7895.S65 K66 1997.

Hussmann:1997:FFS

[2271] Heinrich Hussmann. Formal foundations for software engineering methods, volume 1322 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CO-DEN LNCSD9. ISBN 3-540-63613-7. ISSN 0302-9743 (print), 1611-3349 (electronic). x + 286 pp. LCCN QA76.758 .H87 1997.

Ausnit-Hood:1997:AQS

[2272] Christine Ausnit-Hood, editor. Ada 95 quality and style: Guidelines for Professional Programmers, volume 1344 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63823-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). xiv + 292 pp. LCCN QA76.73.A35A216 1997.

Burkart:1997:AVS

[2273] Olaf Burkart. Automatic verification of sequential infinite-state processes, volume 1354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63982-9 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 163 pp. LCCN ????